A Complete Bibliography of Publications in SIAM Journal on Applied Dynamical Systems

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

26 August 2024 Version 2.37

Title word cross-reference

(1/3, 1/2) [GAS18]. (N + 1) [Riv13]. 0 [DDvGS07, FGMW07]. 1 [FGMW07, GLNW15, SG11, TW18b]. 1 + N [DHW22, BCH10]. 1: 1 [MO20]. 2 [CMW11, FHKK21, GRS^{+}22, IW21, LDB20, Ste14, UW14]. 2: 3 [RAM15]. 2N [PRAC23]. 3 [BdCT12, LRR08, TXKW17]. 3x + 1 [LP21]. 4 [BM20a, WSB16]. ^{2+} [KVB22]. α [HGT15, YB22]. β [GS09b]. C^{1} [Chi08]. C_{60} [GAKTWW21]. O(ϵ) [Wil22]. d [DNO23b, DNO23a]. D_{n} [CBR05]. ẍ = f(x, t) [AH09]. E [FE12]. E^{3} [Noa08]. G_{0} [PMBM05]. H mod K [FG10]. H ∈ (1/3, 1/2] [GALS16]. I [FE12]. J_{2} [BCPS08]. κ_{1}/κ_{2} = c0 [BB24]. L_{1} [CR12].

L_{2} [Cap12]. R^{2} [Ike23]. R^{3} [Kri20, BM15, KR11, KH15a, LZ23, Wec05]. R^{3} × SO(3) × T [CEvdM24]. N [ACMM20, CT22, MG18, NX22, PW24, DKaK^{+}08, GH05, Rob13, VA21]. N + 1 [ABCV23]. ϕ^{4} [GH05]. π [DDvGS07]. Q [PP12]. R^{2}r [Vil18]. R_{0} [HJL16, HJL17, MJJL12]. Σ_{3} [BCGH08]. ε [ETS21]. T̂ C [TD08].

-Body [ABCV23, CT22, MG18, PW24, PRAC23, Riv13, ACMM20]. -Bounce [GH05]. -Breathers [PP12]. -cells [GS09b]. -Chaos [BCGH08]. -Clusters [FHKK21]. -Concave [DNO23b, DNO23a]. -D [GLNW15, IW21, TW18b, TXKW17]. -Dimensional [DKaK^{+}08, LDB20, SG11]. -Dynamical [LP21]. -Induced [KVB22].

-Kinks [DDvGS07]. -Models [HGT15]. -Parameter [BdCT12]. -Reduction [CEvdM24]. -Spot [NX22]. -stable [YB22]. -Vortex [Rob13].

19 [Chu21].

2 [DDGK13, GHH^{+}21]. 2d [BT10, BHLM21, QCARL21]. 2D-Normal [QCARL21].

84 [CWZ17].

A-Current [ZBN09]. abc [HSS13, LPS13, MXYZ16]. Abel [HTV20]. Abelian [FG10]. Absolute [ACK17, JC23, Rad06, STW23, SRS09]. Absolutely [GBIB06]. Absorber [TKB17]. Absorbing [JS06]. Absorption [FP16]. Abstraction [SA13]. Abundance [Rod21, RC23]. Accelerator [DEV04]. Accumulation [CV14, KSKJ20]. Accuracy [PW21, dWRS18]. Accurately [MR06]. across [Bal17, BHP^{+}21]. Action [BH20, CF12, MR24]. Activator [VSC23]. Activator-Inhibitor [VSC23]. Active [BB16]. Actively [GFB03, SDK20]. Activities [RWZ21]. Activity [CGG^{+}22, JR05, KN14, PR22, PTK09, RT02, ZRP18]. Activity-Driven [ZRP18]. Actuated [OPR24]. Acute [GLS21, PES12]. Adaptation [CE04, JLG21, PE18]. Adapted [Gia15, LKH^{+}22]. Adapting [BTK12, SAS11]. Adaptive [FH14, HNP16, HLLP18, Jam10, KPW17, LMKY23, RBI21, SA13, Wil21, Wil22, WS24, ZLL22]. Adaptively [BSY19]. Addiction [DAFM19]. Adding [DK18, LCDS12]. Additive [GH21, WLW15]. Adjoint [Ni23]. Adjoints [LD18]. Admissible [Gor13, SHA23]. Advection [GHLP22, HL22, HP17, MJM05, YNZ22]. Advective [CLW24, TC21]. Aerosol [SDW15]. Affect [GH04b]. Affected [FRB19, Rob04]. Affine [GINR20, HSSY23]. Afterdepolarizations [KVB20, KVB22]. Age [NC21, SAA^{+}18]. Age-Structured [SAA^{+}18]. Agent [AHS14, JJ20]. Agent-Based [AHS14, JJ20]. Agents [DA12]. Aggregated [FSS19]. Aggregation [BT16, BE14, BFH14, EK16, EFK17, GBCV20, HKLN20]. Aircraft [HKLN13]. Albedo [HAS16, ML12, MW14]. Algae [GM15, HP17, SJLY17]. Algebraic [DLRB19, NB^{+}23, PEH22, VC12]. Algorithm [BH23, BP08, COT19, Hül16, KHG21, Ni23]. Algorithmic [KKV18, MSW15]. Algorithms [DFT08, DTK20, HDL^{+}08, HdlL13]. All-to-All [VM11]. Allelopathic [TR19]. Alleviation [PVVCS21]. Allocation [YCG^{+}22]. Allow [PBK18]. Almost [HS09, OBK18]. Alone [TLRB11]. along [HS10a, Rob04, Ton19]. Alpha [GC20]. Alternate [LS15, LS16]. Alternating [LR19]. Alternative [CF20]. Among [FHKK21, RT02, WIN16]. Amplified [RS13]. Amplitude [Blö03, BN23, KT24, LZH^{+}17, NCA^{+}21, Wil21, Wil22, WS24, YB22]. Anaerobic [FSDAS21]. Analogue [Wul08]. Analysis [ACL23, ABMS15, AY23b, AKK^{+}09, AHS14, BKPS19, BCH14, BB12, BK24a, BRRS02, BHLMW24, BDG^{+}16, BK20b, CDKS19, Chu21, CZSV23, CLBdB09, CJAMV20, DJM04, De 03, De 07, DDvGS07, DLRB19, ESZ04, ES22, EKL07, FvdSG20, Fol11, GVNS09, GL24, GKML09, GC20, GBCV20, GS13, HAS16, HJL16, HJL17, HRYZ19, HS05, HP14, IPS19, KPK23, KVB20, KP23, Lan16, Las18, LR20, LGLC15, LV14, LE10, MSB^{+}14, MW16, MJJL12, MW10, MIK19, PP20, PPK14, SP03, Sie02, SRMPM08, SLF24, SMM21, TKB17, TDL17, TKKCG16, TT20, VF10, WSYTA23, WW20, YCG^{+}22, YNZ22, YvLKL22, YB11, ZSL23]. Analytic [BMSY21, JM13, KKJ18, PE10, dlLL12].

Analytical [BS19, GHS10, RBBG20]. Analyzing [AMBV22, BM18, CW22b]. Anchoring [CSJM18]. Anesthesia [MK12]. Angular [BFH22, BH23, BCHM16, TD12]. Animal [BE14]. Anisotropic [AVV23, CDKS19, GST03, GKCG15]. Annihilation [GW24]. Annular [GB09]. Annulus [DE22]. Anomalous [TWW18]. Anti [HM22, HM24, TLRB11]. Anti-integrability [HM22, HM24]. Anti-Phase [TLRB11]. Antidiffusion [BHV11]. Antigravity [KRW13]. Antigravity-Type [KRW13]. Antikink [GH05]. Antiphase [LT13, LT15]. Antisynchrony [AD21, NSS19]. Antithetic [Bri20, BK20b]. Aperiodic [FKS20]. Appearance [CL09]. Appearing [DNO23b, DNO23a]. Application [ABBC20, BGB05, CWZ17, CBR19, DC16, FGMW07, FGDKC15, FH22, FH23, GL24, GSJ23, HDL^{+}08, IM16, JL10, LFFC^{+}20, SMS18, SL19, TS23, VH08, VCK09]. Applications [BR13, BR23, CP17, DEV04, GS22, GSB^{+}16, KKBK20, Oro14, PLST20, PB10, SW16, YvLKL22]. Approach [Aga18, AJB^{+}16, BvdBV18, Bal17, BMMP20, Bri19, Bri20, CW22b, CDT21, DRC09, DEL14, DKZ17, DIL^{+}24a, DDMG16, DvHX16, FA13, FDS14, GN14, GL24, GMS14, GC20, GHS10, GH09, GS20, HG10, HMN09, KKV18, KDKR13, KBGD^{+}23, LR20, MM22, MSW15, NB^{+}23, PFGV14, PEH22, PA19, RBBG20, SAA^{+}18, ST13, VC12, VCNSD24, WB14, Wil21, YW10]. Approximate [BS18, DIL^{+}24b, GAG^{+}21]. Approximating [GV04, PYGR06, SV09]. Approximation [BT16, CW17, CCL24, Chi08, CH13, CFR04, GN14, GZED20, GMCM19, GMCM21, KKK20, KBS14, MW14, Moh19, QCARL21, Tup09, WM22, dWSLD21]. Approximators [MPC^{+}22]. Arbitrary [LT21, Wil23b]. Architecture [JR05]. Arcs [JL10]. Arctic [HAS16]. Area [DHP23, Jam10].

Argument [FEIvdD12]. Arising

[CM03, GL13, SO09, VC12]. Armbruster [AHARS18]. Arnol’d [SO09, AGMS23]. Arrays [KE08]. Arrhythmias [PK05a]. Arrows [GST05]. Arterial [TW18a]. Arteries [CM03]. Articulated [MRR06]. Aspects [AV19]. Assembly [GVY17]. Assessment [FH23]. Asset [DSC12]. Assimilation [BBJ21, Brö17, COT19, FGHM^{+}20, FMT16, OBK18, OPL21, dLDF^{+}18, dLD22]. Assisted [BCGH08, Cap12, CWZ17, CZ15, CW18, FdlL17, Ipp11, KKJ18, SZ13, Wil05, WZ09, WSB16, WB17, Zgl02, dlLJ16, GLJY23]. Assists [RS07]. Associated [DGG16, VBW13]. Asteroid [GKMS06]. Asymmetric [Fer20, GG18, NWW21, PTK09, Tro08]. Asymmetry [DHW22]. Asymptotic [AEL08, Ada23, BC09, BFH14, CDH23, CHH23, DNDY16, EK10a, Has21, HBB13b, LM19, LTB09, LNOR21, MMNS22, MZ11, NS13]. Asymptotically [Jel24, SPCT12, Sul23]. Asymptotics [AHW21, Bri19, KRW12, Noa08, Van08]. Asynchronous [RCG^{+}23, TW18b, Yak08]. Atom [ESZ04]. Attached [BCH14]. Attenuation [HLvdD22]. Attitude [CDPVY21]. Attracting [CZ15, DA12, Rod21, Zgl02, GFE20]. Attraction [FH22, FH23, GKM21, HS10a, WR13]. Attractive [CH14, HKLN20, JJ20, LTB09]. Attractive-Repulsive [LTB09]. Attractivity [BCKN14]. Attractor [GH21, GMS14, MM22, Wil10, ZYO05]. Attractors [AP16, CW22a, DL21, EMNT15, GZED20, GS09a, HM24, LWK23, PACM22, Rod21, SAR13, SCD07, TFC19, WLW15]. Autocatalytic [HKO17]. Autoencoder [OR19]. Automated [NSS06]. Automatic [GOH20, WV19]. Automation [FT12]. Autonomous [Sul23, SMM21].

Autoregression [HARB21]. Auxin [FY13]. Averaged [CEvdM24, GL24]. Averages [BJ24]. Averaging [CH13, DEV04, LV14, MPY11, RRW15, YPMD08]. Averaging-Extrapolation [LV14]. Axial [CLOS14, Xia08]. Axially [RS21]. Axisymmetric [HHW21].

B [GKO17, HDL^{+}08, ZW24]. B-DNA [HDL^{+}08]. Backscatter [PRY23]. Backward [KDKR13, LKO15]. Backward-Forward [KDKR13]. Backward-Time [LKO15]. Bacterial [DRH19]. Bad [BCJ19]. Balance [BFK^{+}22, IMS15, PR24, TL24, USW05, Wid13]. Balanced [CJ18a, FvdSG20, KC13, MR24, Van06]. Balances [HS03]. Balancing [RS11]. Ballistic [MXYZ16]. Banach [VRS22]. Banana [CS18]. Barriers [HKK20]. Based [AHS14, BT19, BAB13, BM24, BU23, CK15, Chi08, Chi09, CSJM18, CV09, CF12, DS23, DKB^{+}23, FY13, ILM20, JJ20, KGB^{+}17, KVX04, LTPL23, LPH22, MCL^{+}20, SBKS15, Wil21, dLDF^{+}18, dLD22, BT21, FE10, OM10, FJ18]. Basic [WZ12]. Basin [AD15]. Basis [BKPS19, BGOŻ08]. Bayesian [GINR20, KSKJ20, PSV22]. Be [RPY20, GBIB06]. Beam [BFG21, DEV04]. Behavior [BFH14, DNDY16, GYAU24, HK15, KKP16, LZF23, MMNS22]. Behaviors [DHK20, FHP22, Leg11, Leg13, THF12]. Behaviour [Sul23]. Belousov [GS13]. Bénard [FGHM^{+}20]. Bends [Rob04]. Benney [HSS22]. Benthic [HJL16, HJL17]. Benthic-Drift [HJL16, HJL17]. Between [BHM24, DRC09, MKO18, Vel13, BHLM21, Bal17, BK24b, CP12, DEL14, ESK24, IW21, KE08, KT24, RRW15, Tup09, WSB16, vdDZ04]. Beverton [HP14]. Beyond [NVC18, OW20]. Bi [CLOS14]. Bi-Axial [CLOS14]. Bianchi [CH10]. Bias [OPL21]. Biased [CBWS24]. Bichromatic [HMS19].

Bicluster [HKLN20]. Bidimensional [TB09]. Bidomain [MMNS22]. Bifurcation [AAM05, AKO13, ADP08, AHS14, BCH14, BN23, BK24a, BK24b, BDG^{+}16, BTBK14, CGP16, CM07, CLBdB09, CHS12, CL09, DKB^{+}23, DLRB19, ELB15, ES22, EKL07, GS16, GH04b, GKO17, GKO18, GST03, GSJ23, GKML09, GC20, Guc08, GM12, GL15, HRYZ19, Hül05, ILM20, Jac06, Jel24, KLW13, KPR15, KOP07, KPR12b, Kri20, LM16, LLZ17, LGLC15, LS17, LPH22, LNOR21, MPW04, Mak17, MW10, MKO18, NSS24, PFGV14, PRCA^{+}23, PPK14, RAM15, SSS06, SP03, SAV12, Sie02, SJLY17, SG11, Ste14, SO09, Tak16, TC23, TV14, TKB17, TWW18, VSC23, VNSG08, Wec05, WV19, WS09, XCC07, YY19, YLWK16, YB11, ZG11, ZKE15, dWSLD21, vdBKV11]. Bifurcation-Theory [GC20]. Bifurcations [Agu15, AVV23, AdBG^{+}09, APBB22, BR13, BBK23, BEW11, BJK20, BK24a, BE14, BC15, BCF^{+}18, CB18, CD10, CEvdM24, DDGK13, DDCK11, DM09, EG06, EPCL05, FE12, GGP^{+}20, GKO18, GKSV23, GL09, GKM05, GKC14, GHS10, GHW03, GK10, HHW21, HE15, KF14, KPR12a, KGB^{+}17, KH15b, KL21, KM24, LFOG17, LCKO08, LCDS12, MP09, Mat11, NS15, NC16, OW20, OdBS08, PCNL12, PB22, PW07, PRY23, QCARL21, RS13, RS16, SW16, SBB10, SM08, SHK13, SvGKP24, TW23, Ton23, VHS22, Vas23, VSABM23, WG15, WS14, ZR20, ZHKR15, vdBLQ21]. Big [KVB20]. Bilayer [MDW23]. Bilayers [PY14]. Bilinear [Kra21, OPR24]. Billiards [CRR11, CFG21, DKaK^{+}08]. Bimodal [GHH^{+}24]. Binary [AMBV22, GKMS06, LZX22, WIN16, ZLX20]. Binocular [KB10]. Binomial [EK10b]. Binomial-Like [EK10b]. Biochemical [BM20b, MV21, MFVW17]. Biolocomotion [EJ16]. Biological [BT16, CW22b, DRH19, MD18, PA19, SAS11]. Biology [Bri20, VC12]. Biology-Inspired [Bri20].

Biomass [FSS19]. Bioregulatory [SSR18]. Bipartite [VM09]. Bipedal [GO15]. Birds [GLW10]. Birkeland [NPV12]. Birkhoff [BJ24]. Birth [VSC23]. Bistability [CM07, SL12, TF21, WWXR24]. Bistable [CATA20, HJSS23]. Bivirus [AY23b]. BizJet [PVVCS21]. Blinking [HBB13a, HBB13b, MJM05]. Block [CW22b, CP06]. Block-Diagonalization [CP06]. Blocks [SW14]. Blood [CM03, GCKW07]. Blow [BRW05, KP23, Mat18, SS09]. Blow-Up [BRW05, KP23, Mat18, SS09]. Blowing [KS11]. Blowup [DDDGZ16, KH15b, KH15a, Kri21, MZ11]. Bodies [CEvdM24, Ver08]. Body [AV19, ABCV23, BN19, BCPS08, BCH10, CR12, CH03, CPY19, CT22, DV19, HGS15, MG18, MRR06, PW24, PRAC23, Riv13, RS07, RS16, SDR09, ACMM20]. Bogdanov [AHGKM16, BK24a, BK24b]. Boiling [SRMPM08]. Boolean [SAR13, TFC19, WHT13]. Border [Gle14, GKC14, PK05a]. Bose [GKCG15, KLK10, PK05b]. Both [KVB22]. Bottom [PRY23]. Bounce [GH05]. Bound [MRS17]. Boundaries [DE06, FK17, LS17, WGCT21]. Boundary [AEHV05, AIT18, BJL^{+}17, BGT10, DDCK11, ELB15, FH22, FH23, GSDN15, LK15, Law16, LD18, LW02, LBR18, MW16, MS15, SDW15, TC23]. Boundary-Hopf-Fold [ELB15]. Boundary-Value [LD18, SDW15]. Bounded [BCP24, CKMW12, HKB^{+}22a, HKB^{+}22b, IS17, MJB14, VNSG08, YB22]. Bounded-Confidence [BCP24, HKB^{+}22a, HKB^{+}22b]. Bounding [FG20]. Bounds [BF18, DFT08, DF19, FGHC16, Fer20, KKJ18, OBK18]. Boussinesq [CCD^{+}10, HSS13, LL08, WW02]. Box [Hen11]. Bragg [MHC09]. Braided [vdBL08]. Braids [FT07, SD22, vdBKV11].

Branch [NSS06]. Branches [WB17]. Branching [Soa17]. Breakdown [FH12, HdlL07]. Breaking [BC15, BCF^{+}18, CL14, CATA20, FGS^{+}23, MHB07, SG11, Ste14]. Breakup [BC23]. Breathers [GBH11, PP12, Yos17]. Breathing [FB04, Fol11]. Brownian [GALS16]. Brusselator [RRW14]. Bubbles [BM20a, Kur17]. Buck [CLBdB09]. Buckling [HMP02]. Bucy [dWRS18]. Budyko [MW14, Wid13]. Bulk [DDN22, GWW19, GLNW15, PLNW19]. Bulk-Membrane [GWW19]. Bulk-Surface [PLNW19]. Bump [CEK22, GH21]. Bumps [CC06a, KE13]. Bundles [AM23, CL13, Hül16]. Buoyancy [Pat03]. Burgers [BW09, CZ15, KZ21, MW16]. Burglaries [SAA^{+}18]. Burner [GB09]. Burst [aAA10]. Burster [LCDS12]. Bursters [EDKC16, LSB11, aAA10]. Bursting [BTK16, BBR^{+}05, DR10, FDS14, FS16, GH04b, GS09b, KVDC12, MRMM14, RRW15, SG10, VBW13]. Bus [WFM^{+}14].

C. [JLG21]. Ca [KVB22]. Cahn

[BT16, BSW16, CMW11, DEP^{+}11, PY14]. Calcium [JPK^{+}22, TZKS12, WFM^{+}14]. Calculation [HMP02]. Calculus [HMP02]. Cam [AdBG^{+}09, OdBS08]. Cam-Follower [AdBG^{+}09, OdBS08]. Camassa [MZ11]. Can [ACL23, AH19, CHH23, KVB20, SRS09]. Canard [KH18b, KH18a, RWK08, CR11, EDKC16, HKO17, KVB20, Rob16, RCG12]. Canard-Like [RCG12]. Canard-Mediated [EDKC16]. Canards [BEG^{+}03, BBK17, DK18, EJK20, EW09, KVB22, LR22, MW17b, RRW15, Wec05]. Cannot [GBIB06]. Canonical [DK18, FPT12, Moh19, PW07]. Cantilever [BRRS02]. Cantor [STY23]. Capture [RS07]. Captures [BGZ16]. Capturing [Sco13]. Car [BP08, SGW09].

Car-Following [SGW09]. Cardiac [BJSW08, CP17, ES22, KVB22]. Cardinality [CDH23]. Cardiomyocytes [KVB20]. Caricature [Rot22]. Cascades [GKO18, GRSB19]. Case [AH09, AK10, BCH10, BR23, Chu21, CZSV23, GS07, GKO18, KLK10, NWKR15, SMS18, VSABM23, Wec05, ZZ20, GKO17]. Cases [FG19, QCARL21]. Catastrophe [DKB^{+}23, ACJ24]. Catch [LPH22]. Category [GT18]. Causal [STB15]. Causality [CGS15]. Causation [STB15]. Caused [YPTW24]. Causes [RB21]. Cavity [YW10, YC10]. Celestial [CHP17]. Cell [ADR16, ADP08, DH19, DLRB19, DR10, EG06, ES22, GST05, JR05, MvB18, Mor15, PK05a, RPY20, RWK08, Rot22, ACJ24, SAH21, SHA23, SS14, SGP03, VC17, VC19, KC13]. Cell-Type [MvB18]. Cells [CE04, PMBM05, ST13, SG11, Ste14, GS09b]. Center [BHM24, BK24b, CR12, CKMW12, Chi17, FDS12, LHW23, RAM15, WS14]. Centering [HHKB20]. Centers [Pat03]. Central [AV19, ASH18, AH19, CHP17, CL11, GH04a, LZ23, RS16]. Centrality [DS23]. Centrality-Based [DS23]. Certain [GNR18, HGT15, Pat03]. Chain [CBR19, LP18]. Chains [CL08, DLP21, GG18, Ton19, VH08]. Change [GT18, LHW23]. Changes [HACY22, aAA10]. Changing [BN19]. Channel [New14]. Chaos [AAM05, AdBG^{+}09, BCGH08, BTK12, BEW11, BK20a, Bre23, CJ11, EPCL05, FGMW07, GKO18, GLS21, GM09, GO02, HKO13, JL08, Lin06, LRH12, Ni23, WSB16, Wil19, vdBL08]. Chaotic [Bal05, CW22a, CLOS14, DKaK^{+}08, GL24, GHC11, Las18, LW02, MFE05, MJM05, NUY05, PP08, RWZ21, RHT13, SS12, VLS13, WM14b, YNN^{+}23, vdBL08]. Characteristic [SS11, Xia08]. Characteristics [AST07]. Characterization [CL11, HL18, SÜvLM16, WS24]. Characterizes [WE18]. Characterizing [MS15]. Charged [GVY17, LPS13]. Charges [ZL20]. Cheating [DRH19]. Chemical [ACK17, BP16, DIL^{+}24b, DB13, ERT11, FG19, FvdSG20, GD23, GS11, HN14, HKO17, LRK12, Vas23, ZKCS19]. Chemostat [FSS19]. Chemotaxis [GLS21, IS17, OY09, ZSL23]. Chimera [OWK18]. Chimeras [CG24, Lai17]. China [ZZ20]. Choreographies [MG16]. Circadian [APBB22, BXB17, LDB20]. Circle [LSB11]. Circles [DM09, OP08]. Circuits [Bok22, MW10, PCG16, RS11, TFC19]. Circular [AV19, BCH10, CDPVY21, DL21, Rob04]. Circularization [KO24]. Clamping [QT20]. Class [BH20, BK20b, DK03, DB13, DTG^{+}16, HHW21, KGK21, OdBS08, PRK22, RGAB16]. Classes [CKCG19]. Classical [BMSM24, OW20]. Classification [BTBK14, BYK08, DNDY16, GHH^{+}24, GH15b, LPK15, SS04, YBO22]. Climate [NWW21]. Clocks [ACL23]. Closed [BB24, DKB16]. Closing [Chu21]. Closing-Reopening [Chu21]. Closure [DDDGZ16, PD18]. Closures [KM24]. Cloud [SDW15]. Cloud-Formation [SDW15]. Cluster [AOWT07, DA12, Oro14]. Clustered [KE08]. Clustering [DA12, Smi24]. Clusters [EW09, FHKK21, RPY20]. Co [CPY19]. Co-Orbital [CPY19]. Coarse [AHS14, BTK16, OK23]. Coarse-Grained [OK23]. Coarse-Graining [BTK16]. Codes [DMS05, GKKZ05]. Codim [DDGK13]. Codimension [AAM05, BJK20, BE14, CKH21, GHS10, LCDS12, LHW23, MP09, PRCA^{+}23, VA21]. Codimension- [VA21]. Codimension-2 [CKH21, GHS10]. Codimension-Three [PRCA^{+}23]. Codimension-Two [BE14, LCDS12, MP09]. Coefficient [VM09]. Coefficients [BCBD20, TW23, WYM^{+}24, ZR20]. Coexistence [AdBG^{+}09, BMWY18, WSB16, vdBDLJ15]. Coherence [MB14a, MB14b]. Coherent [FJ18, FKS20, GM15, KS07]. Coin [PP12]. Coincident [Pat03]. Cointegration [SØRD24]. Collapse [AY20, Ton23]. Collections [TFTK22]. Collective [AJ14, CL08, DDN22, KM08, Leg11, Leg13]. Collinear [DV19, LRK12, MR18, PYGR06]. Collision [BCPS08, FMOW03, Gle14, GKC14, WIN16]. Collisionless [CKPP19]. Collisions [CDH23, GH05]. Collocation [Gie19]. Combinatorial [CKCG19, CGH^{+}16, CGG^{+}22, DJK^{+}19, DMS22, PACM22]. Combining [LPH22]. Common [LS05, MV21]. Communication [IW21]. Community [SBV23]. Commutative [SAH21]. Compact [Hen11, WL22, dlLJ16]. Compactifications [Mat18]. Compared [AHGKM16]. Comparison [CKCG19]. Compartmental [BM24]. Compartments [IW21]. Competing [VVZ15, vdDZ04]. Competition [CSRR08, DRC09, HL22, SSS06, Sla20, SWR05, TC21, YNZ22]. Competition-Diffusion-Advection [YNZ22]. Competitive [GYAU24, KB10, MDIC24, SBV23]. Complete [CHH24, CH14, VM09]. Completion [CMV23]. Complex [ABMS15, AJB^{+}16, BRW05, CR09, GCY22, HDL^{+}08, IBB^{+}10, KGB^{+}17, LBHM05, MRB^{+}13, MR24, PYVG14, RLZT23, SSR10, VNSG08, Wei03]. Complex-Balanced [MR24]. Complexes [AG23]. Component [Ike23, NX22, TvH21, vHDKP10]. Components [TFC19]. Composite [BHLMW24]. Compound [EKL06]. Compressible [HKK20, HK18, PY14]. Compressive [BTBK14]. Compressor [Xia08]. Compromise [CBWS24]. Computation [AEHV05, AM17, BHM24, BC23, CLN24, CJ17, CW22a, CFG21, DDMG16, FH12, FGLdlL17, FKS20, GH15a, Gie19, GYdlL21, GJNO22, HdlL07, HLJ23, HdlL13, ILM20, Jam10, JL10, JM13, JO09, KC13, Kri15, LKH^{+}22, MIK19, MSW15, OM10, PW21, SÜvLM16, WB06, ZDG19, vdBGW15]. Computational [DDGK13, DIL^{+}24a, FT12, FGH14, GH09, GK18, LW20, PE10]. Computations [AM06, BHLZ18, FdlL17, vdBDLJ15]. Computed [TY07]. Computer [BCGH08, Cap12, CWZ17, CZ15, CW18, FdlL17, GLJY23, Ipp11, JO09, KKJ18, SZ13, Wil05, WZ09, WSB16, WB17, Zgl02, dlLJ16]. Computer-Assisted [BCGH08, CWZ17, CW18, Ipp11, WSB16, WB17, GLJY23]. Computing [BK20a, CL13, DF19, DsMS24, EKO04, EKO05, FKO18, FM16, GK09, Hen05, Hon21, Hül16, KO03, LMNT09, MG16, VM09, Wei03]. Concave [DNO23b, DNO23a]. Concentration [ACK17, JC23]. Concepts [ILM20]. Conceptual [NWW21]. Condensates [GKCG15, KLK10, PK05b]. Condition [USW05, YCMP22]. Conditional [OPL21, RT02]. Conditions [BJL^{+}17, BCDG16, BM24, DA12, GSDN15, LK15, LW02, MW16, MS15, MR24, SBR06, Vas23]. Conductance [FE10, GPTV17]. Conductance-based [FE10]. Conductances [GH04b]. Conduction [BC21, De 07, Lee22, STW23]. Cone [Gia15]. Confidence [BGH^{+}24, BCP24, HKB^{+}22a, HKB^{+}22b, MJB14]. Configurations [LZ23, LBR18, RS16, VBG^{+}09]. Confinement [CP12, Pat03]. Conjecture [CLZ21, GMS14, HL22]. Conjugacy [FM16]. Conjugate [BJ22, GKS03]. Conley [BM16, Bat17, BMMP20, BCHM16, DF19, DMS22, FT12, KGK21, Mat11, MSW15, SW14]. Connected [GOH20, KRW13]. Connecting [HM24, Hül05, JM13, SLF24, WZ16, dlLJ16].

Connection [DsMS24, Kri20]. Connections [BHM24, CW18, Tro08, Wil05, WSB16]. Connectivity [AJB^{+}16, JL08]. Connectomes [NC21]. Consensus [BKPP24, CBWS24, GRSB19, Has21, HDDL21, JJ20, RB21, ZZQ18]. Conservation [DIL^{+}24a, DIL^{+}24b, HLLP18, JZ11, MFVW17, VC17, VC19]. Conservative [vdBKV11]. Conserve [SAR13]. Conserving [MFN^{+}23]. Consistent [AYB19, MvB18]. Constant [LPH22]. Constant-Catch [LPH22]. Constrained [LD18]. Constraints [BKPP24, Wil22]. Construct [MSM17, McC15]. Constructing [SBN09]. Construction [DKB16, LD18, TV14]. Consumer [AHS14]. Contact [IS24]. Containment [CJAMV20]. Content [WHT13]. Continuation [BT10, Chu21, CHS12, DD13, DKB^{+}23, EKO05, KKJ18, LS17, NS15, OM10, PSSJ23, SOV05, TD08, UN21, WS06, vdBLQ21]. Continuation-based [OM10]. Continuations [GS09a]. Continuity [BGO11, FH22]. Continuous [BKPS19, GAS18, GBIB06, PJW05, Rob16, SM08]. Continuum [CP12, HKM21, Lai17, MW17a, SS14]. Contour [Hül16]. Contractile [AAK12]. Contraction [FEIvdD12, Gie19]. Contractive [JMB^{+}13]. Contrast [CB18]. Contrasting [BAB13]. Control [AAC23, BTK12, BR23, Bri20, BK20b, CKPP19, CGG^{+}22, CSRR08, DKB^{+}23, GS16, GL24, HMD^{+}23, ILM20, IMS15, KKK20, KM08, KPW17, KVX04, LCMA05, PSSJ23, POR20, Pos09, PBK16, PBK18, PPK14, SBKS15, TGPP19, Wil19, Wil21, ZB20, ZLL22, ZME19]. Control-Based [DKB^{+}23]. Controllable [WYM^{+}24]. Controlled [BR23]. Controllers [Bri20]. Controlling [ZB20]. Convection [AMBV22, FGHM^{+}20, GSDN15, TBR23, UN21, WIN16]. Convective [GS07]. Convectons [TBR23]. Convergence [BTK12, CCL24, HdlL13, MSB^{+}14, RL24, SMS18, VM11]. Convergent [BHP^{+}21]. Converter [CLBdB09]. Convex [BS18, BKPP24, FG20]. Conveying [BRRS02]. Cook [BSW16]. Cooperative [WWC^{+}18]. Coorbital [CH03, DHW22]. Coordinate [AHW21, GHH^{+}21, TT20, Wil21, WS24]. Coordinate-Free [AHW21]. Coordinate-Invariant [TT20]. Coordinates [BHLZ18, BCGFS13]. Coordinating [RK18]. Coordination [AMDH24, SARTA20]. Copy [MW10]. Core [HG10]. Corner [LRK12, OdBS08]. Corner-Impact [OdBS08]. Correction [DHP23]. Correlations [Ipp11]. Correspondence [SSR18]. Corrigendum [CT19]. Cortex [CB16, GST03, RWK08]. Cortical [PTK09, VCK09]. Cosmological [CH10]. Coupled [ADR16, ADP08, AOWT07, AK06, BSY19, BEW11, BC15, BMWY18, CG24, CJ08, CGL19, CP06, DE22, DKTG12, DK03, EG06, EW09, EDKC16, EVC18, EKL06, FE10, Fol11, GST05, GWW19, GLNW15, HI23, Jac06, JR05, JL08, KC13, KM08, KBGD^{+}23, KT24, KP23, LT13, LW02, LPK15, LE10, Ly14, Lyu18, MV14, MHC09, MP13, Mor15, NWKR15, NSS19, OE21, PFGV14, PLNW19, PW21, PK05a, PB17, QT20, RGAB16, RBI21, RRW15, SAH21, SHA23, ST13, SS12, SÜvLM16, SGP03, VM08, WYM^{+}24, XWC^{+}23, YCL08, ZZ09, aAA10]. Coupled-Mode [CP06]. Coupling [BMWY18, CSKR06, DB11, Fer20, FGDKC15, GST03, JKKL24, JLG21, KE08, MMP16, MV14, PW21, PR22, PW24, SARTA20, Smi24, TLRB11, VM09, VM11, Zha07, ZZ09]. Couplings [CH14, HNP16, HLLP18, HKLN20, LZX22, LA18, ZLX20]. CoV [GHH^{+}21]. COVID [Chu21]. COVID-19 [Chu21]. Cowan [HE15]. CPA [GOH20]. CPG [VH08]. Crime [GW24, SBB10, TW18b]. Criteria [WF13]. Critical [BL08, BC23, De 07, DB11, DNO23b, DNO23a, GMM08, OP08, VM09]. Cross [HSV24]. Crossing [AS18, BN19, MG18, MNG07]. Crossings [LM19]. Crowd [KKC06]. Crystal [BCF^{+}18]. Cube [NSS13]. Cubic [Blö03]. Cucker [CKPP19, Has21]. Curl [EG05]. Current [BM15, ZBN09]. Currents [GH04b, MLTC21]. Curvature [HL18, MB14a, MB14b]. Curve [CK15]. Curved [GL13, HHBS22, HSV24, TT20, VHS22]. Curves [AG05, GH09, KT24, Ly14, MNG07, WG15]. Cusp [KOP07, PRCA^{+}23]. Cusp-Cusp [KOP07]. Cusp-Turing [PRCA^{+}23]. Cusped [KP23]. Cut [Ste23]. Cutting [SUOL18]. Cutting-and-Shuffling [SUOL18]. Cycle [BM18, CFST08, CKK^{+}09, KT24, LFOG17, NSTZ20, RPY20, WGCT21, WE18, XWC^{+}23]. Cycles [BJK20, CFdSL22, Chu21, CD10, DDGK13, GC20, HRR^{+}03, HTV20, HdlL13, Mak17, MO15, NWW21, PCNL12, RCG12, SSS06, STY23, SPCT12, TC23]. Cyclic [DvG09, LGLC15, TFC19]. Cyclically [SBV23]. Cycling [BAB13]. Cylinder [HKM21, MST03]. Cylinders [AD15, BvdBV18]. Cylindrical [AMBV22, YB22].

D [CBR19, BM20a, CMW11, GRS^{+}22, GLNW15, IW21, LRR08, TW18b, TXKW17, UW14, WSB16]. D-Chain [CBR19]. Dafermos [SS09]. Daido [Chi17]. Damage [CRSN07]. Damped [BCGH08, KO24, RK23a, WZ09, ZYO05]. Damped-Driven [RK23a]. Damper [EPCL05]. Damping [GS22, SCD07]. Data [ABBC20, BHP^{+}21, BGH^{+}24, BBJ21, Brö17, COT19, Chu21, CZSV23, DTG^{+}16, FGMW07, FGHM^{+}20, FMT16, GHTW23, HHHY09, HHKB20, KHG21, LTLA21, MCL^{+}20, MIK19, OBK18, OPL21, OK23, PD18, POR20, RABK19, TFTK22, TY07, VCNSD24, ZKCS19, dLDF^{+}18, dLD22]. Data-Driven [ABBC20, BHP^{+}21, DTG^{+}16, LTLA21, OK23, PD18, POR20, RABK19]. Data-Rate [KHG21]. Database [AKK^{+}09]. Databases [BCHM16]. Death [Hsu19]. Debris [CGP16]. Decay [CGG^{+}22, Ipp11]. Decentralized [CKPP19]. Decision [CD17, EWLH11]. Decomposing [GCD^{+}21]. Decomposition [AMBV22, ADR23, AM17, AK18, AZAK22, AYB19, BMSY21, EMKB19, HMD^{+}23, HHKB20, KFB16, LV17, LP23, Oro14, PN20, PBK16, RK23b, SHA23, TFTK22, ZRDC19]. Decompositions [DJK^{+}19]. Deep [CMV23, GS20]. Defect [DvHX16]. Defects [IS24, SS04, SS07]. Deficiency [Des23]. Definition [WE18]. Deformation [TD12]. Deformations [WW20]. Degeneracies [LWK23]. Degenerate [BMSM24, BT16, BSW16, DDCK11, RAM15, VSC23]. Degree [SL19]. Delay [AMDH24, BKPS19, BCGFS13, BR13, BJK20, BK24a, BDG^{+}16, BL08, BC15, CHK17, DHK20, EKL06, GYdlL21, GLW10, Has21, HLJ23, KKBK20, KRW13, McC15, MNG07, NPRW19, QSvdH19, SP03, SS11, SÜvLM16, TR19, VRS22, WLW15, WZ18, WYD22, YW10, YCL08, YCMP22, YB11, Zha07, ZW24, dWSLD21]. Delay-Coupled [BC15, EKL06, SÜvLM16]. Delay-Differential [BL08]. Delay-Induced [SP03]. Delayed [CSKR06, DS23, GLJY23, HMN09, IMS15, KKP15, KKP16, LPR22, LCMA05, Oro14, Pos09, PPK14, PYVG14, SV09, TKB17, WZ17, XCC07, ZKE15]. Delayed-Mutual [CSKR06]. Delays [AH06, AK06, BK20b, CHK17, GMB16, GSB^{+}16, Lee22, LGLC15, LPK15, RBK15, STW23, SvGKP24, Vel13]. Demodulation [CH13]. Dendrite [SL12]. Dengue [CJAMV20]. Densities [BHLM21, Ipp11]. Density [BKS06]. Dependent [BGB05, BS19, Bok22, CHK17, DMCK15, FRB19, GZED20, GYdlL21, HMP02, IS17, JKKL24, Kim20, PR22, SS14, WZM23, GLJY23]. Deployment [GW24]. Depression [Fay13, KB10]. Depth [CMV23, WSWK12]. Derivation [LK15]. Derivative [CSS17, PSV22]. Derivative-Free [PSV22]. Deriving [JR05]. Descending [KN14]. Descent [BKPP24]. Describe [PID21]. Describing [CCD^{+}10, CV14]. Description [BS19, NS13]. Design [GMM08, KVX04, RS09]. Designing [MM17]. Desingularization [vdBLQ21]. Despite [ACL23]. Destabilization [DRdRV18]. Desynchronization [WM14b]. Desynchrony [LSB11]. Detailed [FvdSG20]. Detecting [DKB^{+}23, SS12]. Detection [CHS12, MKM23]. Determinant [WF13]. Determination [GOH20]. Determining [CGS15, DS19, NPRW19, SK08]. Deterministic [CW17, DDDGZ16, ESK24, FGHC16, OK24, SMS18, WLW15]. Detuning [HK05]. Development [YHM^{+}02]. Deviation [Sco13]. Diagonalization [CP06]. Diagram [CMW11]. Diagrams [GH04b, SGW09, TV14, WV19]. Diatomic [DLP21]. Die-Out [AY23a]. Diffeomorphisms [HM24, MM06]. Difference [LA18, NSS19, dlLL12]. Difference-Coupled [NSS19]. Different [BSOM20, CZSV23, SSR18, Sco13]. Differentiable [BSKR16]. Differential [AEHV05, AY20, AK10, BR13, BJK20, BK24a, BK20a, BL08, CHK17, CD20, DLRB19, DNO23b, DNO23a, FG19, GAS18, HL18, HLJ23, HJSS23, KRW13, KL21, LD18, LNOR21, MB14a, MB14b, Mat11, Mat18, McC15, MK24, RABK19, SS11, VA21, VRS22, YM23, YB11, dWSLD21]. Differentially [LL08]. Differentiation [SDT17]. Diffusion [AGMS23, ADF20, ADF21, AJB^{+}16, BvdBV18, BCGFS13, BHLMW24, CL08, CGK08, DK03, FP16, GW24, GHLP22, GLNW15, GSJ23, HH19, HH20, HCW23, HL22, HP17, IS24, IW21, Law16, MRB^{+}13, MP13, MRS14, MFN^{+}23, NUY05, NX22, PLNW19, PID21, Rad13, SW21, SRS09, SvGKP24, TC21, TW18b, TXKW17, TWW18, UW14, VSABM23, WZ12, WR13, WM24, WW20, Wri10, XSW24, YNZ22]. Diffusion-Mapped [BCGFS13]. Diffusion-Mediated [IW21]. Diffusions [GIHLS20]. Diffusive [CJ18b, HKK20, JKKL24, SJLY17, SÜvLM16, TBR23]. Diffusivity [CFG21]. Digestion [FSDAS21]. Digraph [DHK20]. Digraphs [Don24]. Dimension [DvHX16, EWLH11, NRS20, STY23]. Dimensional [Agu15, BJSW08, BXB17, Brö17, BIPK24, CLL12, CW22a, CW11, CLOS14, CZ15, CW18, CL09, DJM04, DKaK^{+}08, DL22, EKO04, EKO05, FMT16, FEIvdD12, GJNO22, GL13, HM22, HM24, KKK20, KZ21, Lai17, Leg11, Leg13, LDB20, MSM17, MP13, MM17, MFE05, MM12, MJM05, NSUW09, NC16, PLNW19, PBB22, RCG12, SSS06, SRMPM08, SØRD24, SS14, SG11, Ste14, TZ22, YY19, ZDG19, dlLJ16, LO10, RLZT23]. Dimensionality [WB14, WSWK12]. Dimensions [BAA^{+}19, DDMG16, JZ11, TWW18, Wri10]. Dipoles [GKCG15]. Direct [Bal05]. Direction [CB18]. Directions [ABBC20]. Disambiguating [ESK24]. Discontinuities [BSKR16]. Discontinuity [CKH21, CD10, KK19, OdBS08, VA21]. Discontinuity-Induced [OdBS08]. Discontinuous [AEG24, BCDG16, FPT12, GH15b, HAS16, JC09, LSB11, LPH22, PB10]. Discovery [CBK19, OK23, PD18, YBO22]. Discrepancy [ESK24]. Discrete [ABG^{+}17, BM16, Bat17, Brö17, COT19, DJM04, ELT22, FMT16, FE10, HQW^{+}20, HS10b, IW23, KM10, Kie20, LS15, LRH12, MCP09, NUY05, RBK15, VVZ15, Yak08, Yos17, YC10, LS16]. Discrete-in-Time [COT19]. Discrete-State [HQW^{+}20]. Discrete-Time [HQW^{+}20, Kie20]. Discretization [BDG^{+}16, MR06]. Discriminant [TFTK22]. Disease [DS24, WZ17, vdDZ04]. Disguised [HSSY23]. Disorder [SBV23]. Dispersal [CLW24, She14]. Dispersals [Yak08]. Dispersion [Rob04, SSR10, SB10]. Dispersion-Managed [SB10]. Dispersive [CCD^{+}10, MHC09, NP15]. Displacement [Bal11]. Dissecting [SG10]. Dissipated [BC09]. Dissipation [BRMR04, EJ16]. Dissipation-Induced [BRMR04]. Dissipative [AGMS23, AK10, BYK08, BMWY18, EMNT15, OBK18, TKKCG16, dlLK19]. Distancing [GHH^{+}21]. Distinct [CEK22, FSS19, FSDAS21]. Distinguishing [NWKR15]. Distributed [AH06, HMD^{+}23, MNG07, NSUW09, RBK15, SvGKP24, YB11]. Distribution [Tup09]. Distributions [ACK17, Bok22, GLW10, GD23, PEH22]. Disturbance [FH22]. Disturbances [GMM08, XWC^{+}23]. Divergence [BU23]. Diversity [MDIC24]. DMD [KGB^{+}17]. DMD-Based [KGB^{+}17]. DNA [CRSN07, HDL^{+}08]. Do [Zha07]. Domain [CR12, CDS10, Daw09, IS17, RS21, SMM21, TW18a]. Domains [CLN24, FK17, GL13, Kra21, Lai17, YB22]. Dominant [BHL16]. Dominated [AY20]. Dominating [BGZ16]. Double [KLW13, QCARL21, UN21, XCC07]. Doubling [DS19, SS07]. Doubly [CJ18b, TBR23]. Downscaling [COT19]. Downstream [BHP^{+}21]. Drag [PRY23]. Drift [CLW24, HJL16, HJL17, KPG19, PE10, WM24]. Drillstring [GVNS09]. Driven [ABBC20, BHP^{+}21, BT16, DTG^{+}16, GALS16, GAS18, GVY17, GKC14, Lin06, LTLA21, MDW23, OK24, OE21, OK23, PD18, POR20, RK23a, RABK19, WZM23, YB22, ZRP18]. Driver [ABC^{+}22]. Driving [WFM^{+}14]. Droplet [SS12]. Droplets [LBR18, RK23a]. Dry [GHS10]. Dry-Friction [GHS10]. Dryland [GLS23]. Dual [GS09b, HKLN13]. Dual-Wheel [HKLN13]. Ducks [KVB20]. Ducts [HHBS22, HSV24, VHS22]. Due [KRW13, PRY23, APBB22]. Dumbbell [DM20]. Duplex [MCL^{+}20]. Dwell [Aga18]. Dynamic [AMBV22, ADR23, AST07, AM17, AK18, AZAK22, AYB19, BBR^{+}05, EMKB19, GO15, GB09, GRSB19, HMD^{+}23, HHW21, HHKB20, JC23, KFB16, LV17, LP23, LLZ17, MPW04, MKM23, NVC18, PN20, PBK16, RB21, RK23b, SBKS15, TFTK22, ZRDC19]. Dynamical [ACFK09, ABMS15, AEG24, ADR23, AM23, BM16, Bat17, BMSY21, BD23, BFH22, BH23, BLDK18, BTBK14, BIPK24, CLL12, CII23, CL11, DJM04, DA12, DKZ17, Des23, DMS22, DTG^{+}16, DJCJC22, DAFM19, FGHC16, GKMS06, GAG^{+}21, GALS16, GCD^{+}21, GS22, GHTW23, GCY22, GH15b, Guo10, GS20, HN14, HQW^{+}20, JS06, Kie20, KH15a, Kur17, LPR22, LM19, LP21, LS15, LS16, LMKY23, LO10, LdST09, LKH^{+}22, LPK15, MM11, NRS20, PVMP17, RK18, SW16, SRS14, SM20, SMM21, TW18a, TD08, VRS22, VC12, WB14, WSB16, Wil23b, YCG^{+}22, YvLKL22, ZDG19]. Dynamics [AJ14, AV19, ABM^{+}04, AS18, AAC23, AKK^{+}09, ABG^{+}17, AOWT07, AD15, BKM22, BKS06, BHP^{+}21, BFK18, BMMP20, BGH^{+}24, BB12, BTK16, BMCGW14, BC14, BCP24, BMWY18, BCHM16, CD17, CR23, CSKR06, CG18, CW11, CRSN07, CLOS14, CP23, CKCG19, CTAA18, CP17, DS24, DHMO05, DFT08, DF19, DSC12, DH19, DDN22, DV19, DEP^{+}11, DJK^{+}19, DGMW12, DvG09, Don24, DNO23b, DNO23a, DEV04, DR10, EWLH11, EGS24, ELT22, EVC18, EPCL05, FG20, FHKK21, FGS^{+}23, FGH14, GIKR20, GW24, GINR20, GRS^{+}22, GFB03, GVNS09, Gia15, GKCG15, GLNW15, GRSB19, GH15b, HKL14, HLLP18, HK19, HP20, HKM21, HHBS22, HKP22, HHHY09, HMD^{+}23, HSV24, HBB13a, HBB13b, HK05, HKB^{+}22a, HKB^{+}22b, HDL^{+}08, HKLN13, IS17, IW21, Jef14, JPK^{+}22, JMB^{+}13, KKV18, KNWH11, Kim20, KL17, KRW13, KTK20, LCMA05, LR20]. Dynamics [LTB09, LDB20, LNOR21, Ly14, MR06, MEvdD13, MSM17, MG18, MSB09, MP09, MvB18, MW14, MJB14, MO15, Moh19, MST03, MDIC24, MW17c, NPV12, NX22, NVC18, NPRW19, OZM11, OR19, OK23, PCNL12, PD20, PLNW19, PSV22, PID21, PRK22, PB10, RB21, RBK15, RCG^{+}23, RAM15, RBI21, RPY20, RLZT23, RHT13, SW21, SHdlL17, SMRB11, SRS09, SJLY17, SWR05, TC21, Ton23, TB09, THF12, TZKS12, Tup09, TXKW17, VHS22, VCK09, VBW13, WZ17, WYM^{+}24, WIN16, Wei03, Wid13, XSW24, YNN^{+}23, ZBN09, ZW24]. Dynamics-Adapted [Gia15].

Early [KVB20, KVB22]. Earth [CG18, ML12]. Ecological [SBV23]. Ecology [PLST20, YNZ22]. Ecosystems [HCW23, OW20]. Edge [YLWK16]. Effect [BCPS08, KSWW06, MP09, MK12, MS15, TPC24]. Effective [CM03, HDDL21]. Effects [AMDH24, AMNB06, BJL^{+}17, ESK24, GMB16, HKL14, LLYZ13, LZH^{+}17, QSvdH19, RWZ21, SK13, TGPP19, ZL20, ZRP18]. Efficacy [CGS15]. Efficient [CLJ15, CW22a, FT12, SD22, XCC07]. Eigendecomposition [DC16]. Eigenfields [EG05]. Eigenfunction [PZ22]. Eigenfunctions [BLDK18, GCY22, MK24]. Eigenvalue [HMN09, NC21, SG11, Ste14]. Eigenvalues [AR12, BL08, CII23]. Einstein [GKCG15, KLK10, PK05b]. Elastic [CM03, HMP02, Mun11, SS14]. Electrical [BJSW08, KVB22, TLRB11]. Electrically [CGL19, DE06, GVY17, LT13]. Electrochemical [Rot22]. Electroencephalographic [SHdlL17]. Electrolysis [DLRB19]. Electrostatic [Guo10]. elegans [JLG21]. Elementary [ADR16, GN14]. Elimination [CW17]. Elite [AY20]. Elite-Dominated [AY20]. Ellipsoid [HGS15]. Ellipsoids [CRR11]. Elliptic [BM20a, BCPS08, CCHZ20, CLZ21, IS17, LSB11, NSS06, NSS13, WB17, aAA10]. Embedding [ZDG19]. Embeddings [CBK19, PD20]. Emergence [BCP24, HKZ18, HKLN20, JJ20]. Emergent [DHK20, FHP22, HLLP18, HK19, HP20, HKM21, HKP22, MSB^{+}14, MDIC24, SDT17]. Enclosures [GJM12]. Encoding [BHP^{+}21, CB18]. Endemic [GYAU24]. Endocrine [EVC18]. Endomorphisms [KOP07]. Endotactic [CD20]. Energy [BGZ16, CF20, LHRK04, MPC^{+}22, PR24, PRY23, Wid13, WM14a]. Energy-Balance [PR24]. Energy-Optimal [WM14a]. Engine [GC20]. Engineering [VC12, VCNSD24]. Enhancement [MXYZ16]. Enlarged [LWK23]. Enrichment [EGS24]. Ensemble [DDSW22, GIHLS20, Kim20, dWRS18]. Ensembles [SKL23]. Entorhinal [RWK08]. Entrainment [BCRR21, CH14, LSB11, LDB20, Lin06, SKL23, Wil21, ZL14]. Entropy [DFT08, FT07, SD22, STB15, VLS13]. Environment [FS09, Law16]. Environmental [BT16, LHW23]. Environments [CLW24, TC21]. Enzyme [ETS21]. Epidemic [AY23b, DS23, DS24, GSJ23, KRW13, SBKS15, WZ12, WWC^{+}18]. epiroticus} [NPV12]. Equal [MR18, RPY20, RS16]. Equation [ALB^{+}10, BM12, BCGH08, BCKN14, BW09, BEG^{+}03, BRW05, BD12, CHK17, CB16, CSS17, CHH23, CR09, CHS12, CDT21, CZ15, DHMO05, DvG09, DLRB19, ELT22, Fay13, FdlL17, FP16, GBK15, GAL20, GH05, GBCV20, GHW03, GK10, Hu23, IBB^{+}10, KZ21, KRW12, LP22, LSAC08, LS17, Llo19, Llo21, LBHM05, MPW04, MSB^{+}14, MS13, MW16, MHB07, NP15, PW07, PY14, PYVG14, RLZT23, SP03, SLS23, SSR10, SMS18, SBN09, TDK18, TKKCG16, VD13, WYM^{+}24, WW02, WZ16, WV19, XSW24, YBO22, YB11, Zgl02, vdBL08]. Equation-Free [CHS12, MSB^{+}14, SMS18]. Equations [AY20, ADF20, ADF21, AK10, BvdBV18, BT10, BT04, BR13, BBJ21, Blö03, BN23, BR23, BJK20, BK24a, BDG^{+}16, BK20a, BK15, BL08, Buz24, CM03, CJN15, CO04, CFR04, CZ16, DK03, DNO23b, DNO23a, EK10a, EK10b, FG19, FGLdlL17, FMT16, GL17, GALS16, GAS18, Guo10, HH19, HH20, HE15, HLJ23, HACY22, HTV20, HMS19, HJSS23, IM16, Jac06, KPK23, KKK20, KL21, Kur17, Lai05, LLZ17, LNOR21, MHC09, Mat11, Mat18, McC15, MM06, MZ11, MK24, NCA^{+}21, OK23, PRY23, RGAB16, RABK19, She14, SS11, SCD07, VA21, VRS22, VF10, Vel13, VNSG08, WLW15, WZM23, WW02, Wri10, YM23, YB22, ZYO05, ZME19, dWSLD21, dlLL12]. Equilateral [TD12]. Equilibria [AY23b, BHL16, BFK18, BT11, BJK20, CJ18a, CG18, CEvdM24, DB13, DGK^{+}21, ESZ04, EFK17, FK17, HGS15, HDDL21, MR18, MR24, NB^{+}23, PYGR06, Pat03, Rob13, Scl23, SLF24, TC23, Ver08, YY19, ZZQ18]. Equilibrium [ABCV23, CKK^{+}09, DDCK11, HS15, LBR18, SL19, Tak16, VBG^{+}09]. Equilibrium-to-Periodic [CKK^{+}09]. Equivalence [DEL14, Des23, HSSY23, KC13]. Equivariant [BT04, FHKK21]. Ergodic [AM17, Bri19, VRS22]. Ergodicity [BK20b, LM19, Sco13, Tup05]. Erratic [RB21]. Erratum [ADF21, HKB^{+}22b, HJL17, KH18a, Leg13, LS16, MB14b, VC19]. Error [KKJ18, OBK18, TS23]. Errors [SK08]. Escape [FS09, RS07]. Escapes [CTAA18, CATA20]. Essential [Rad06]. Estimate [SDW15]. Estimates [ADF20, ADF21, WM24]. Estimation [ACFK09, GPTV17, KHG21, OPL21]. Estimators [PZ22]. Euclidean [DKaK^{+}08]. Euler [BHLZ18, MM06]. Eulerian [DDMG16, MJB14, Ver08]. Evacuation [PSSJ23]. Evans [BHLZ18, BD11, CO04]. Evans-Function [BHLZ18]. Even [BLDK18, WIN16, Yos17]. Even-Symmetric [WIN16]. Event [BSKR16]. Event-Selected [BSKR16]. Events [FG20]. Evidence [KSKJ20, RWZ21]. Evolution [BKS06, BT04, CM16, CLW24, FGLdlL17, GL17, GALS16, GL13, Kul16, KVX04, LLZ17, MB14a, MB14b, PVMP17]. Evolutionary [HS03, PCNL12]. Evolving [AJB^{+}16, RBBG20]. Exact [BIPK24, Lai15, Vil18]. Example [AHS14, ESZ04, KPK08]. Examples [DDGK13]. Exchange [SS09]. Excitability [FDS12, New14]. Excitable [AP16, BGO11, CKK^{+}07, CL09, DZ14, DKTG12, DL22, HS22, HG10, OE21, SMRB11, Ton19, Tro08, WM14a, YNT14]. Excitation [AAK12]. Excitatory [CEK22, CE04, FB04, LT19]. Excited [KLK10]. Existence [BHL16, BC21, BAB13, Brö17, Cap12, CR12, CFST08, CC06a, CJ18b, CV14, CZ15, CZ16, DL22, Dys20, Fay13, GHLP22, GC05a, GC05b, Guo12, HKP22, HvHM^{+}14, Ike23, KTK20, LT17, MR18, MW17c, PJW05, PY14, Wec05, WZ09, Yos17, dlLJ16]. Exogenous [MJB14]. Expansion [YPTW24]. Expansions [AEL08, Bre23, GJNO22]. Expensive [DDSW22]. Experiment [DKB^{+}23]. Experimental [AdBG^{+}09, FGMW07]. Experimentally [GBIB06]. Experiments [CL08]. Explain [KVB20, SRS09]. Explaining [RRW15]. Explanation [MW16, ACJ24]. Explicit [CZ16]. Exploiting [CW22b]. Exploration [KN14, WV19]. Explorations [HdlL07]. Explosion [AG05, DK18, Rob16, RCG12]. Exponent [CG24]. Exponential [Bri20]. Exponentially [DGG16, HH20]. Exponents [BN23, TY07]. Extended [AS18, BCBD20, BD12, EJK20, GKML09, HL18]. Extending [CGG^{+}22]. Extension [ADR23, Chi09]. Extensions [CH13]. External [MRS17, RSTY12, YW10]. Extinction [AY23a]. Extinctions [YPTW24]. Extract [FGH14]. Extraction [FJ18]. Extrapolation [LV14]. Extremal [LFFC^{+}20]. Extreme [FG20]. Extremism [RB21]. Extrinsic [GD23, MV21].

Factors [BT16, Jam10]. Failure [KFB08]. Fallacies [FH14]. Families [AK17, BD11, BdCT12, GNR18, LO08, TS07]. Family [AIT18, BLL12, DJCJC22, HS03, JL10, MCZM18]. Farming [AAC23]. Fast [BB12, Bok22, CJ18b, DD13, DDMG16, DTG^{+}16, DAFM19, EJK20, GMM08, GV04, HLLP18, HdlL13, Jel24, KBS14, KM17, LW20, Ni23, OK24, PVMP17, SBR06, TLRB11, WZ18, RBI21]. Fast-Slow [BB12, DAFM19, EJK20, KM17]. Fat [Hen05]. Faux [MW17b]. Features [Jud20, LR18, TS07]. Feed [RS13]. Feed-Forward [RS13]. Feedback [AH06, BTK12, Bok22, EKL07, Ged10, GKS03, GKML09, HMN09, KKK20, KKP15, KKP16, KN14, KPW17, KVX04, LCMA05, LZF23, ML12, MW14, MHB07, PCG16, PCG18, Pos09, PPK14, PYVG14, RPY20, TKB17]. Feedbacks [XCC07]. Feedforward [SW24]. FEM [FJ18, KZ21]. FEM-Based [FJ18]. Fenichel [CT17, CT19]. Fermi [Yos17]. Few [BKS06]. Fiber [Hül16, MHC09]. Fibers [KBS14]. Field [AV19, BT19, BT21, BSKR16, CCL24, CB16, CB18, CEK22, CO04, De 03, Fay13, GB23, GK22, IM16, IW23, NC16, PE18, PLST20, PZ22, PK17, PbG09, SHdlL17, SL19, THF12, VF10, Vel13, ZME19]. Fields [AH06, BW12, BC14, BK15, BM15, Bri19, BdCT12, Chi08, Dys20, FE12, JC09, Jud20, KE13, KF14, KO03, KS14, Lai15, Lan16, LPS13, MLTC21, NSS19, NCA^{+}21, STW23, SvGKP24, VRS22]. Filament [KTK20]. Filippov [AS18, DDCK11, ELB15, GHH^{+}24, NSTZ20]. Filling [OY09]. Film [CV14, EHLW15, KRW12]. Films [TGPP19]. Filter [dWRS18]. Filtered [EKL07]. Filtering [COT19, HI23, PA19]. Filters [SDT17]. Financial [GRS^{+}22]. Finding [CJAMV20, CKO17, LFFC^{+}20, LR18, LPK15]. Finite [ACK17, BN23, BIPK24, Daw09, DDDGZ16, DKTG12, FJ18, FKS20, GG18, HBB13a, Hu23, LFOG17, MB14a, MB14b, MM17, PYGR06, SLS23, SMS18, VC12, VM08, WSWK12, ZB20]. Finite-Dimensional [BIPK24]. Finite-Size [SLS23]. Finite-Time [DDDGZ16, FJ18, FKS20, MB14a, MB14b, MM17, ZB20, BN23]. Fire [CJ08, CGL19, JMB^{+}13, NC16, SL12, TB09]. Firing [CO04, Dys20, GB23, HE15]. Firing-Rate [GB23]. First [CP06, CKO17, DEV04, IW23, Rad13, Tak16]. First-Order [CP06, DEV04, IW23]. Fish [AAC23]. Fitting [GMY18]. FitzHugh [CS18, CJ18b, CP17, CZ16, GK10, HS10b, Ike23, KP23, QT20, RCG12, TvH21]. Five [TFC19]. Fixed [DRC09, Don24, Hül05, KM10, Zgl02]. Flame [AHW21]. Flames [GB09]. Flip [AKO13, APBB22, GKO17, GKO18, Jac06]. Flip-Flop [APBB22]. Flocking [DR22]. Flocks [Don24, FHP22]. Floer [BvdBV18]. Flop [APBB22]. Floquet [CL13, CLJ15]. Flow [Bal17, BSOM20, Bri19, CM03, CLOS14, CHS12, CV14, CDS10, DSC12, De 07, FMHM24, GCKW07, GHW03, HHW21, HL18, Hen11, Kim20, KGB^{+}17, LFFC^{+}20, LR18, LLYZ13, MJM05, Pro20, Xia08]. Flows [AJ14, Bal05, BU23, BSKR16, CJ11, FKS20, GMCM21, HKK20, HK18, KGB^{+}17, MXYZ16, MFE05, MM12, PBB22, SW14, ZL20]. Fluctuations [GT18]. Fluid [AMBV22, BRRS02, CM03, HHW21, KSG14, LL08, MRR06, Mun11, UN21, WIN16, ZHKR15]. Fluids [Bal17]. Flux [Bal05, BJL^{+}17, FY13]. Flux-Based [FY13]. Fly [TV14]. Focused [TW18b]. Focusing [HSV24, VHS22]. Fold [ELB15, FGGT^{+}12, GKSV23, JC09, KH15b, Mak17, CJ11]. Fold-Fold [Mak17]. Folded [DKO08, DK18, Kri20, MW17b, RRW15, Wec05]. Folded-Saddle [DK18]. Foliation [CV09, CKO17, HdlL23]. Foliation-Based [CV09]. Follower [AdBG^{+}09, OdBS08, ZBN09]. Following [GZED20, NSS06, SGW09]. Forced [ANR14, ANR18, BCGH08, BEG^{+}03, BYK08, GHW03, LDB20, ML12, RS09, RHT13, SM20, VCK09, WZ09, WG15, XSS20, ZG11, ZL14]. Forces [JJ20, WZM23]. Forcing [CR09, CZ15, EMNT15, GBK15, Hu23, KNWH11, KKP15, Kri21, QSvdH19, RSTY12, SK13, WLW15, ZKE15]. Forecast [GHTW23]. Forecasting [ADR23, BGB05]. Form [CLJ15, Gle14, HdRS23, Hon21, PPK14, PEH22]. Formalism [LTLA21, SAH21]. Formation [BB16, CKPP19, CE04, Daw08, DL10, DA12, GLS23, GLS21, KE08, PLNW19, PES12, SDW15, TR19, TKKCG16, WCM08]. Formed [MvB18]. Forms [BK24b, QCARL21]. Formulae [LM19]. Formulas [DDGK13, Sco13]. Formulation [FvdSG20, LS05]. Forward [EGS24, KDKR13, LKO15, RS13]. Forward-Time [LKO15]. Foucault [Moe15]. Foundations [PLST20]. Four [Brö17, ESZ04, LHW23, MR18, RS16].

Four-Body [RS16]. Four-Dimensional [Brö17]. Four-Vortex [MR18]. Fourth [NP15, vdBKV11]. Fourth-Order [NP15]. FPU [HdRS23]. Fractal [NSS06]. Fractional [GALS16]. FRAM [LR19]. Frame [LR18]. Framework [BMSY21, ESK24, FGLdlL17, HCW23, IM16, KS14, LTLA21, VQPVV23, Wil19, Wil22, Wil23b, WS24]. Frameworks [CZSV23]. Free [AHW21, CHS12, GAL20, MSB^{+}14, PSV22, SMS18]. Freezing [BT04, BST08, WSYTA23]. Frequencies [LA18]. Frequency [Ada23, CRR11, CDT21, CSRR08, DGG16, DAFM19, LR19, LV14, PE18, RSTY12, RBI21, XSS20]. Friction [BBK17, GHS10]. Frobenius [CII23, SK22]. Front [BW12, HdRS23, KS14, MXYZ16, vHDKP10]. Fronts [AHW21, Dys20, GS07, GKSV23, Guo12, Llo19, Llo21, MMNS22, PP20]. Frustration [HKL14, HKZ18]. Full [CT22, CEvdM24, GK10]. Fullerene [GAKTWW21]. Fully [GGP^{+}20, MP13, dWRS18]. Function [BHLZ18, BD11, CO04, FG19, PVVCS21]. Functional [AEHV05, CMW11, YB11]. Functionalized [PY14]. Functionals [McC15]. Functions [AY23a, CO04, GH15a, GOH20, MMP16, McC15, PW21, PW24, SHA23, SARTA20]. Fundamental [BCDG16, GRS^{+}22, KGK21, SGW09]. Fundamentals [BFH22].

Gain [SS09]. Gain-of-Stability [SS09]. Gait [ASH18, JLG21, WSYTA23]. Gaits [AH19]. Game [CFdSL22, GRSB19]. Game-Theoretic [GRSB19]. Games [IW23]. Gamma [MNG07]. Gamma-Distributed [MNG07]. Gap [BD11, Coo08, DG05, DP08, Lai15]. Gear [HKLN13, MP09]. Gene [DEL14, MEvdD13, MW10]. General [Bal05, DHK20, GH21, GSJ23, IM16, JL08, SDR09, ST13, WL22]. Generalist [LHW23]. Generalization [LTLA21]. Generalized [BM12, Bre23, CII23, DGMW12, DR22, GKM05, HTV20, JMB^{+}13, KM17, MR24, NSUW09, Riv13, SD17, SS16, SBN09, YCG^{+}22]. Generalizing [PBK18]. Generated [CKCG19, KS07]. Generation [MFN^{+}23]. Generator [ASH18, CF07, CV09, GH04a]. Generators [AH19, OPR24, POR20, YNT14]. Generic [EG05, PVVCS21, QCARL21, WHT13]. Genetic [GMB16, GSB^{+}16, LGLC15]. Geodesic [KO03]. Geodesics [BB24]. Geographically [HCW23]. Geometric [ABI24, BHLMW24, CW22b, DHP23, DvHX16, GMS14, GS11, GH09, HL18, HS05, JPK^{+}22, KPK23, LT17, MPY11, Rad06, VBW13, WB14]. Geometries [DL21]. Geometry [BN19, CHK22, CFG21, CDS10, DKO08, GS09b, HKLN13, KK19, MB14a, MB14b]. Geophysical [HHW21]. Gierer [GWW19, KSWW06, KR11, MW17c, SWR05, VD13]. Ginzburg [BRW05, CR09, DL21, IBB^{+}10, LBHM05, MPW04, PYVG14, RLZT23, SSR10, VNSG08, WZ16, WV19, vdBGW15]. Glacial [NWW21]. Global [AKO13, Agu15, ACMM20, AKK^{+}09, BS18, BCKN14, BW09, CKCG19, CL09, DJM04, DHMO05, DDMG16, DB13, EK10a, EKO05, EMNT15, FHKK21, GKMS06, GAKTWW21, GKO17, GKO18, GHLP22, GMS14, GKR24, HKO13, KO03, LCKO08, LZH^{+}17, Lyu18, MP09, MRMM14, MKO18, OM10, PP08, PPK14, SHdlL17, TC21, VF10, VM08, VSABM23, YCL08, ZHKR15, ZZ09, dlLK19]. Globally [AOWT07, CG24, CJ08, CZ15, ZB20, ZLL22]. GnRH [CF07, CV09, EVC18]. Go [BCJ19]. Good [BCJ19]. Governing [OK23]. Gradient [BFK18, CLL12, GIHLS20, Kim20]. Gradients [CB18]. Grain [LS17]. Grained [OK23]. Graining [BTK16]. Granular [CLOS14, DLP21, WCM08]. Graph [BM24, FA13, NSS19, VM09, XWC^{+}23, YCMP22]. Graph-Based [BM24]. Graph-Theoretic [YCMP22]. Graphic [FvdSG20]. Graphically [CJ18a]. Graphop [GK22]. Graphs [ABI24, DJD19, DMS22, KPG19, SMRB11, Sla20, UE15]. Gratings [MHC09]. Gravitational [AV19, ACMM20, CT22, SL19]. Gravity [EHLW15, RS07]. Gray [SD17, CW11, SWR05]. Grazer [YLWK16]. Grazing [EPCL05, SHK13, SO09]. Grazing-Sliding [SO09]. Greitzer [Xia08]. Grid [GS22]. Gridding [SA13]. Gross [TKKCG16, WYM^{+}24]. Ground [BLL12]. Group [AAM05, CHH24, Chi08, Chi09, Kim20]. Groupoids [SGP03]. Groups [CD17, WL22]. Growing [AGG^{+}19]. Growth [CJN15, PRY23, YCG^{+}22]. Guaranteed [PD20]. Guarantees [ABI24]. Guckenheimer [AHARS18]. Guide [MRS14]. Gust [PVVCS21]. Gyroscopic [USW05]. Gyrostat [Ver08].

Hair [BMCGW14]. Hallucinations [NCA^{+}21]. Hamilton [BC09, KKK20]. Hamiltonian [AH09, AHARS18, AK17, BGZ16, BMWY18, BC23, CFR04, FMHM24, KRK14, LS05, MPY11, NS13, PRK22, Sul23, TAtN09, Tup05, WR02, Wul08, WS09, YY19, YPMD08]. Hamiltonian-Like [BMWY18]. Hamiltonians [But20]. HANDY [Ton23]. Hard [LLYZ13, WGCT21]. Hard-Sphere [LLYZ13]. Harmonic [Smi24]. Harvesting [LPH22]. Hausdorff [FH22, STY23]. Having [NSTZ20]. Hawkes [GP20]. Heart [KVB20]. Heat [BC21, De 07]. Heated [LL08]. Heaviside [CO04]. Heavy [GB09, SDR09]. Hebbian [HK15, LZX22, ZLX20]. Hegselmann [Has21]. Helmholtz [MM06, NSUW09]. Hematopoietic [CM07, DH19, PMBM05].

Hemispheres [NWW21]. Hénon [AIT18, GKM05, JL10, Tak16, WZ09]. Hénon-Like [Tak16]. Hepatitis [ZW24]. Hermitian [HP20]. Heterochaos [STY23]. Heteroclinic [AP16, BHM24, CJ17, CFST08, CL16, CKK^{+}09, CW18, FHKK21, JL10, KPR12a, MO15, QCARL21, SPCT12, Wil05]. Heterodimensional [STY23]. Heterogeneity [DE16, Lai17, YNT14]. Heterogeneous [AH19, GKR24, HCW23, JKKL24, Ly14, OK23, RT02, SK13, SKL23, TC21]. Hexagon [LSAC08, Llo21]. Hexagonal [IR22]. Hexagons [vdBDLJ15]. Hidden [GH15b]. Hierarchical [LDB20, WIN16]. Hierarchy [Jef14]. High [BGH^{+}24, CDT21, DDMG16, GJNO22, GJ17, KKK20, PW21, QCARL21, SØRD24]. High-Dimensional [GJNO22, KKK20, SØRD24]. High-Frequency [CDT21]. High-Order [GJNO22, GJ17, PW21, QCARL21]. Higher [AMBV22, BBK23, BAA^{+}19, LV17, PW24, Smi24, Ste14, WM24, WW02]. Higher-Dimensional [Ste14]. Higher-Order [BBK23, PW24]. Hill [RGAB16]. Hilliard [BSW16, BT16, CMW11, PY14]. Hindmarsh [LCDS12]. Hinged [BFG21]. HIV [WWXR24]. HJB [KVX04]. HJB-POD-Based [KVX04]. HKB [AMDH24]. Hodgkin [GO02, KPK23, Lin06]. Hohenberg [ALB^{+}10, BD12, DHMO05, GBK15, LSAC08, LS17, Llo19, Llo21, MS13, PW07, vdBL08]. Hölder [GAS18]. Holes [BFGTM14]. Holistic [MR06]. Holm [MZ11]. Holt [HP14]. Homeostasis [GS18]. Homeostatic [APBB22]. Homoclinic [AKO13, Agu15, AHGKM16, AAK12, AM06, BK24b, CS18, CKMW12, DMCK15, DvdP02, DRdRV18, GKO17, GKO18, GKM05, GK10, GL15, HLJ23, Jac06, KW08, LCDS12, Llo19, Llo21, Lu16, MO15, SZ13, TBR23, WZ09]. Homoclinics [CWZ17]. Homogeneous [AR12, FY13, GL09, LZ23]. Homogenization [LK15]. Homology [BvdBV18, DJK^{+}19]. Homotopy [Chu21, EGS24]. Hopf [AY23b, ADP08, CKK^{+}07, CHS12, CR11, ELB15, EG06, FG10, GS16, Guc08, GM12, KLW13, LM16, MKO18, PCNL12, PPK14, QCARL21, RAM15, RS13, SP03, SJLY17, SvGKP24, TW23, TWW18, UN21, XCC07, YB11, ZR20, ZG11, dWSLD21, vdBLQ21]. Hopf-Zero [GS16, QCARL21]. Horseshoe [AIT18, Jam10, STY23]. Horseshoes [But20]. Host [YPTW24]. Host-Range [YPTW24]. Hotspot [GW24, RWZ21]. Hotspots [SBB10, TW18b]. Hubei [ZZ20]. Hubs [GFE20]. Hubs-attracting [GFE20]. Human [AMDH24, BXB17, DS24]. Hunting [ETS21]. Hurst [GALS16]. Huxley [GO02, KPK23, Lin06]. Huygens [KLW13]. Hybrid [KGK21, LZF23, RSRT21, TC23, TD08, WL22, YCL08]. Hydraulic [EPCL05]. Hydrocarbon [GB09]. Hydrocarbon-Oxygen [GB09]. Hydrodynamic [EG05]. Hyperbolic [DKaK^{+}08, Ni23, Wil10]. Hyperbolicity [HdlL07]. Hyperchaos [WSB16]. Hyperdissipative [WZM23]. Hypergraphs [HKB^{+}22a, HKB^{+}22b]. Hypernetworks [RBBG20]. Hysteresis [ABG^{+}17, KM17, LL08]. Hysteretic [KNWH11].

Ice [HAS16, ML12, MW14, PR24]. Ice-Albedo [ML12]. Identical [CJN15, HKZ18, ZZQ18]. Identifiability [GD23, SRS14]. Identification [GK18, KGB^{+}17, MH17, RABK19, VCNSD24, Wil23a, YNN^{+}23]. Identifying [BGH^{+}24]. IDM [ABC^{+}22]. II [ANR18, Bat17, BH23, BEG^{+}03, GC05b, HBB13b, LZX22, RWK08]. III [RSRT21]. Illustrated [WV19]. Imitation [GRSB19].

Immune [BR13, WWXR24, ZSL23]. Impact [GS22, GG18, OdBS08, PRK18, PRK22, RHT13, TDL17]. Impacting [PB10, TC23]. Impacts [KRK14]. Impairment [WWXR24]. Impedance [LR19]. Implementation [DDGK13, GM09]. Implications [DRH19]. Implicit [MSB^{+}14]. Importance [SB10]. Improvements [ABC^{+}22]. Improves [HHKB20]. In-Phase [TLRB11]. Inclined [AMBV22, BCPS08]. Include [CGG^{+}22]. Inclusions [CD20]. Incomplete [FJ18]. Incompressible [CM03]. Incorporating [CM16, Lai15]. Increased [NC21]. Incubation [WWZ19]. Indecision [FGS^{+}23]. Index [BM16, Bat17, BMMP20, DF19, FT12, GAS18, HMP02, Mat11, MSW15]. Indices [KR21, SW14]. Indistinguishability [BM24]. Individual [BJL^{+}17, CR23]. Individuals [CP12]. Induced [BEW11, BRMR04, CM16, CDW24, CD10, CTAA18, DE06, HCW23, Jel24, Kie20, KVB22, LT21, LNOR21, NCA^{+}21, OdBS08, RWK08, SP03, THF12, ZSL23, vdDZ04]. Inequalities [CF12]. Inertia [CDH23, HKL14]. Inertial [HHBS22, HGT15, HSV24, Hon21, KDKR13, VHS22]. Infection [SLF24, ZW24]. Infections [KRW13]. Inference [DDSW22, GINR20, STB15]. Infinite [CLL12, CMV23, DJM04, HKM21, LO10, LA13, RC23, WSWK12, YY19, ZDG19, dlLJ16]. Infinite-Dimensional [CLL12, DJM04, YY19, LO10]. Infinitely [Scl23]. Infinitesimal [KT24]. Infinity [LLZ17]. Inflammation [GLS21, PES12]. Influence [HKLN13, Zha07, ZBN09, ZZ20]. Influenza [ABM^{+}04]. Information [AJB^{+}16, GRSB19, JR05, LPR22]. Information-Theoretical [AJB^{+}16]. Informed [PD20]. Inherent [AST07, Don24]. Inhibition [CK15, Dys20, FSDAS21, FA13, Guo12, TLRB11].

Inhibition-Based [CK15]. Inhibitor

[VSC23]. Inhibitory [CEK22, CE04, LT15, LT19, Lee22, LBR18, PR22, SK13, ZBN09]. Inhomogeneities [KS07]. Inhomogeneous [GGP^{+}20, KFB08]. Initialization [AHGKM16]. Initiation [BMCGW14, HS22]. Injectivity [BP16]. Inner [BM12]. Innovation [Kul16, MRB^{+}13]. Inputs [AMNB06, AH19, GH21, GS18, LSB11, MJB14, PBK18, Rot22, RSRT21, Wil21, Wil22]. Insect [ASH18, AH19, GH04a]. Insights [GRS^{+}22]. Insoluble [EHLW15]. Inspired [Bri20, SØRD24, Wil23a]. Instabilities [BS19, CDT21, DS19, SWR05, TW18b]. Instability [BRMR04, DG05, EG05, GS07, HS10a, MM06]. Instantaneous [BKPS19, WGCT21]. Integrability [PRK18, HM22, HM24]. Integral [Bri20, BK20b, CO04]. Integrate [CJ08, JMB^{+}13, NC16, SL12, TB09]. Integrate-and-Fire [CJ08, JMB^{+}13, SL12, TB09]. Integrating [Hen05]. Integration [AK10, BT10, FKS20, GH21, PK17]. Integrators [FMOW03]. Integro [LD18]. Integro-Differential [LD18]. Integrodifferential [She14]. Intelligent [ABC^{+}22]. Intensity [BT19, BT21, MM22]. Intensity-Based [BT19, BT21]. Interacting [AJ14, CD17, CEvdM24, FE12, GINR20, GIHLS20, HKO13, Leg11, Leg13, PZ22, RLZT23, WF13]. Interaction [BS18, BR13, CDKS19, CP12, CR11, EGS24, GM15, HRS04, HdRS23, HP17, KLK10, Rad13, Ver08, Wri10]. Interactions [BBK23, CMW11, DK03, EGF18, KE08, Kur17, LKO15, LR19, PW24, VVZ15, vHDKP10]. Interconnected [WWC^{+}18]. Interface [Bal17, DP08]. Interfaces [FP16, MvB18, MFN^{+}23]. Interior [AR12, ADP08]. Intermediate [CW17]. Interneuron [EW09]. Interplay [BK24b, Vel13]. Interpolant [COT19].

Interpolated [POR20]. Interpolation [GCD^{+}21]. Interpolatory [PVVCS21]. Intersecting [KK19]. Intersection [Jef14]. Intersections [Cap12, STY23]. Interval [ZW24]. Intralayer [RBBG20]. Intraseasonal [TS23]. Intrinsic [CJN15, DR10, GD23, LA18]. Invariance [CGK08, HL18, SAS11]. Invariant [AKO13, Agu15, AG05, BHLM21, BW09, BGT10, BK20a, Bri19, BC23, Cap12, CCL24, CLJ15, CJAMV20, CP17, DJCJC22, DM09, EK10b, FMHM24, FGLdlL17, FM16, GN14, GL17, GINR20, GJNO22, GKO17, GBIB06, GV04, HS10a, HdlL07, HCW23, HL18, Hen05, Hen11, HKO13, Ipp11, JO09, MSM17, MP13, MK24, MKO18, MJM05, NSS20, NSS24, OP08, PYGR06, SV09, SOV05, SAH21, SM20, TT20, VSABM23, WZM23, WB17, YNN^{+}23, dlLL12]. Invasion [HvHM^{+}14, Llo19, Llo21]. Inverse [EKO04, HI23, HACY22, PbG09]. Inversion [BRMR04, PSV22]. Inverted [LCMA05]. Investigating [KKP16]. Investigation [AdBG^{+}09, AST07]. Ion [LLYZ13, New14]. Ionic [ZL20]. Irreducible [GK18]. Irregular [KKP16, SL19]. Irregularities [HRR^{+}03]. Isaacs [KKK20]. Islands [AG05]. Islets [WFM^{+}14]. Isochrons [DDMG16, GYdlL21, HdlL13, LKO15, MRMM14, OM10]. Isolated [CKH21, PP20]. Isolating [SW14]. Isonomy [Jef14]. Isotropic [BM15, Bri19]. Iterative [KBS14, Kri15]. IX [CH10].

Jacobi [KKK20]. Jam [NVC18]. Jordan [Mor15]. Josephson [DDvGS07]. Jump [BN13, YNT14]. Jump-Type [YNT14]. Junction [DDvGS07]. Junctions [Coo08, Lai15]. Jupiter [Cap12].

Kalman [GIHLS20, dWRS18]. KAM [AVV23, IW23, MPY11, PVVY17, Pat03]. Kawahara [CDT21, TDK18]. Kawasaki [CW18]. KdV [LP22]. Kelvin [MM06].

Kepler [KO24]. Keplerian [CF20, DM20]. Kernels [Gia15]. Kim [AHARS18]. Kinetic [PRY23]. Kinetics [ETS21, GLNW15, WF13]. Kink [GH05]. Kink-Antikink [GH05]. Kinks [DDvGS07]. Kirchhoff [NSUW09]. Klausmeier [BCBD20, SD17]. Knife [YLWK16]. Knots [KTK20]. Knudsen [CFG21]. Kobayashi [YW10]. Koopman [AM17, BLDK18, DTK20, GCY22, GMCM19, GMCM21, KKBK20, LP21, LMKY23, LTLA21, MCZM18, MK24, OPR24, PP20, POR20, PBK18, SK22, SMM21, WM22, Wil23a]. KPP [HS09]. Krause [Has21]. Kuramoto [BF18, CHH24, Chi17, CDH23, CDS10, DT13, DJD19, DC16, DHK20, DB11, Fer18, Fer20, FdlL17, GL17, GK22, HKL14, HNP16, HLLP18, HKZ18, HK15, LZX22, LA18, MR06, MW17a, NB^{+}23, Scl23, SØRD24, Ste23, VM09, VM11, WYD22, Zgl02, ZLX20]. Kuramoto-Type [GK22, Ste23]. Kuznetsov [Wil10].

Labeled [TFTK22]. Labyrinthine [YHM^{+}02]. Lag [Chi17]. Lagrange [BHLZ18]. Lagrangian [BU23, FMOW03, HGS15, HK18, Vil18]. Lambda [SRS09]. Lambda-Omega [SRS09]. Laminar [MDW23]. Lamprey [VH08]. Landau [BRW05, CR09, DL21, IBB^{+}10, LBHM05, MPW04, PYVG14, RLZT23, SSR10, VNSG08, WZ16, WV19, vdBGW15]. Landing [HKLN13]. Landscape [GMY18]. Lang [YW10]. Langevin [BK15, EGS24, GINR20, GIHLS20]. Laplace [SMM21]. Laplace-Domain [SMM21]. Laplacian [GFE20]. Large [Daw08, FGHM^{+}20, FHKK21, HKL14, HHHY09, HG10, KVB20, NS15, RCG^{+}23, TS07, Tro08, Wil21, YW10, ZL20]. Large-Scale [Daw08, NS15, RCG^{+}23, TS07, Tro08]. Large-Time [HKL14]. Largest [NC21].

Laser [EKL06, EKL07, GFB03, GKS03, GKML09, SDK20, TKB17]. Lasers [BEW11, BC15, CSKR06, EKL06, Sie02, YvLKL22]. Lasota [BFGTM14]. Lateral [Dys20, FA13, GST03, Guo12]. Lattice [BHV11, CCL24, CL14, DMCK15, DP09, GBH11, HMS19, KC13, KL17, LWK23, SW16]. Lattices [IR22, JL08, KKV18, KR21, KPT13, LW02, MV14, Yos17]. Law [HLLP18, WF13]. Laws [BFK^{+}22, CSKR06, DIL^{+}24a, DIL^{+}24b, JZ11, MFVW17]. Layer [CFR04, DL10, RWK08, SS14, WCM08]. Layered [WWC^{+}18]. Layers [BHP^{+}21]. Leading [GS16]. Leaky [CK15, SL12]. Learned [DGMW12]. Learning [BD23, ESK24, GHTW23, HLLP18, Jud20, LTLA21, LTPL23, OY03, OR19, OPR24, PD20, ZKCS19]. Lecar [NWKR15, New14]. Legacy [GKKZ05]. Lemma [NSS24, SS09, Soa17]. Less [EGS24]. Level [KO03]. Lévy [DHP23, YB22]. Libraries [BTBK14]. Lid [CFR04]. Lie [Noa08, WL22]. Life [FKM23]. Light [MHC09]. Lightwave [SK08]. Like [BMWY18, DsMS24, EK10b, RCG12, Tak16]. Likelihoods [DDSW22]. Limit [BT21, BLL12, BCH10, BM18, CMV23, DDGK13, DRdRV18, GC20, HKM21, HG10, HAS16, HTV20, HdlL13, KRW12, KT24, Lai17, LFOG17, Mak17, PCNL12, PZ22, RCG12, SSS06, SPCT12, TC23, WGCT21, WE18, YW10]. Limit-Cycle [KT24]. Limitations [ABC^{+}22]. Limiters [FH14]. Limits [BT19, GK22, KM17]. Lindstedt [AHGKM16]. Line [BJM20, CDKS19, MW14, PR24]. Linear [BHLM21, BCRR21, BFH22, BH23, BM24, BL08, CFST08, CW22a, CMV23, CF12, Coo08, DEL14, FGDKC15, FPT12, GRS^{+}22, GCY22, GP20, GL09, GWW19, GLNW15, JKKL24, MRS14, MDIC24, MR24, Ni23, PD20, PACM22, PPM14, RCG12, SSR10, SS11, SRS14, TDL17, VRS22, WGCT21]. Linear-in-Parameters [SRS14]. Linearly [OR19, YCL08]. Link [BKPS19]. Link-Delay [BKPS19]. Linkage [BH20]. Linked [DL22]. Linking [LR22, PID21]. Links [BCJ19, KTK20]. Liouville [MM12]. Liouvillean [XSS20]. Lipschitzian [ZLL22]. Liquid [TGPP19]. Lizard [CFdSL22]. Load [PVVCS21]. Local [AAM05, BBJ21, BFH14, CSS17, DB13, GAS18, GHLP22, KKJ18, KE08, LLYZ13, TFC19, VF10, ZHKR15, ZZ09]. Local/Global [VF10]. Localization [GG18]. Localized [ANR14, ANR18, BAA^{+}19, BMCGW14, BYK08, BD12, CW11, Daw08, Daw09, DL10, FE10, GVY17, GL13, HS10a, HRS04, LSAC08, MHC09, PRCA^{+}23, RRW14, TXKW17, TWW18, VSC23, WW20, vdBGW15]. Localizing [LT17]. Locally [Blö03, CH10, JMB^{+}13, ZLL22]. Locating [Bal11]. Lock [AHS14]. Lock-In [AHS14]. Locked [CH13, GFB03, SDK20]. Locking [BF18, CDH23, Fer20, GKS03, HKZ18, HdlL23, RSTY12, TLRB11, VM08, WYD22]. Locomotion [GH04a, Mun11]. Locus [AIT18, HSSY23]. Loewner [VQPVV23]. Logics [WHT13]. Logistic [Bri20, IS17]. Lohe [CH14, HP20, HKP22, Kim20, ZZQ18]. Long [AH06, CM03, CCD^{+}10, CMW11, EGF18, GJ17, HH20, HK15, MG18, NPV12, Pos09, PMBM05, Sul23, dWRS18]. Long-Period [Pos09]. Long-Range [AH06, CMW11, EGF18]. Long-Term [NPV12, Sul23]. Long-Time [dWRS18]. Longitudinal [HACY22]. Loop [Aga18, Pro20]. Loops [CH13, MO15]. LOR [LR20]. Lorentz [RGAB16]. Lorenz [CWZ17, CKO17]. Lorenz-84 [CWZ17]. Loss [EJK20, HAS16, RT02, WZ18]. Lotka [BR23, CJN15, Sla20, TC21]. Low [BE03, BTBK14, CG18, GSDN15, HARB21, MFE05, OZM11, VQPVV23, WSWK12]. Low-Dimensional [MFE05]. Low-Order [VQPVV23]. Low-Rank [BTBK14, HARB21].

Low-Reynolds-Number [OZM11]. Lower [DF19]. Lubrication [BT10]. Lunar [YPMD08]. Lunisolar [CGP16]. Lyapunov [AY23a, BN23, CG24, Cap12, FG19, GH15a, GOH20, McC15, RS11, TY07, WS14, ZR20, dlLK19]. Lyme [WZ17].

Machine [DKB^{+}23]. MacKay [TPC24]. Macro [TW18a]. Macroscopic [FGH14, HDL^{+}08]. Magnetic [De 03, LPS13]. Magnitude [Wil21]. Main [HKLN13]. Making [EWLH11]. Malaria [AAC23]. Managed [SB10]. Maneuvers [AST07]. Manev [ABCV23]. Manifold [AY23b, BFK18, BK24b, CR12, Chi17, CJAMV20, FKO18, GKKZ05, GCD^{+}21, GJM12, HCW23, KPR12b, SBN09, Van08]. Manifolds [AKO13, Agu15, Bal11, BHM24, BJM20, BW09, BGT10, Cap12, CLJ15, CW22a, DKO08, EK10b, EKO04, EKO05, GMM08, GJNO22, GKO17, GJ17, GV04, GK09, HS10a, HGT15, HL18, Hen05, Hen11, JM13, KKJ18, KDKR13, KO03, Kri15, MSM17, MP13, MK24, MKO18, PBB22, SV09, SÜvLM16, SBN09, VSABM23, WZM23, Wri10, ZDG19, dlLL12, dlLK19]. Many [Scl23]. Map [AK06, BM20a, BKS06, BXB17, CRR11, FMHM24, GRS^{+}22, GKM05, GH05, HKO13, JL08, LR19, LW02, LPH22, MV14, WZ18, WZ09, Wil10]. Mapped [BCGFS13]. Mappings [DM12]. Maps [ABMS15, BGB05, BM12, BFGTM14, CG24, DF19, DM09, DEV04, EJK20, EKO04, EKO05, FE10, FM16, GJ17, GKC14, HM22, HdlL07, HdlL23, Hül05, Hül16, Jam10, JM13, LSB11, LCKO08, LDB20, LRR08, MSM17, MM17, MSW15, OP08, PB17, PB10, STY23, SM08, Tak16, Vil18, dlLJ16]. Margins [ILM20]. Market [GRS^{+}22]. Markov [BGB05, BGOŻ08, BN13, CGK08, DDDGZ16]. Markovian [CP23]. Martingale [PZ22]. Mass [ABBC20, BH20, CF12, MFN^{+}23, MR24, Pat03, VC17, VC19].

Mass-Action [BH20, MR24]. Mass-Conserving [MFN^{+}23]. Masses [BN19, RS16]. Master [EK10a, EK10b, PP08]. Master-Slave [PP08]. Matched [CHH23]. Matching [BLDK18]. Mathematical [BKPS19, BMCGW14, CF07, DH19, FSDAS21, GRS^{+}22, KS14, SAA^{+}18, TPC24, YNZ22, ZZ20]. Mathematics [Ren12]. Matrices [BCDG16, BGOŻ08, CC06b, DsMS24, NSS20, SS11]. Matrix [CMV23, FHP22, Kim20, TFBN21]. Matter [CH10, WYM^{+}24]. Matters [Kul16, PB17]. Max [Ste23]. Max-Cut [Ste23]. Maximal [TD12]. Maximally [SD22]. Maxwell [PSW12]. McKean [GPTV17]. Mean [BT19, BT21, GB23, GK22, IW23, MG18, NC16, PLST20, PZ22, SHdlL17, THF12]. Mean-Field [GB23, GK22, PLST20]. Meandering [KL17, Wul08]. Meanders [AST07]. Measure [GMCM19, GMCM21]. Measure-Preserving [GMCM21]. Measurement [dWRS18]. Measurements [FGHM^{+}20, MH17]. Measures [Bri19, CJ18a, CCL24, CHP17, GN14, GBIB06, Tak16, YNN^{+}23]. Mechanical [GBH11, JLG21]. Mechanics [CHP17, FMOW03]. Mechanism [BB16, GBK15]. Mechanisms [CSRR08, CR11, DRdRV18, HdlL07, JLG21, MDW23]. Mechanistic [HACY22]. Media [CCD^{+}10, CL09, DL22, HS22, HG10, HS09, KKC06, KFB08, KS07, LWK23, SS04, WM14a, YC10]. Medial [RWK08]. Mediated [EDKC16, IW21]. Medium [BGO11, FRB19, YNT14]. Meets [CKK^{+}07]. Meinhardt [GWW19, KSWW06, KR11, MW17c, SWR05, VD13]. Melnikov [GHS12, PBB22]. Members [CD17]. Membrane [GWW19, GLNW15]. Memory [BKPS19, PK17, PB17, XSW24, ZRP18]. MEMS [GAL20, Guo10, IPS19]. Mesh [CFR04, KBGD^{+}23]. Meshfree [Gie19].

MESSI [MD18]. Metabolic [LZF23, WFM^{+}14]. Metal [De 03]. Metapopulation [GSJ23]. Metapopulations [DS24, HLvdD22]. Metastability [BW09, BN13, MW16]. Metastable [CDW24, EK16, HDL^{+}08]. Meteorological [MIK19]. Method [AHGKM16, BHM24, BJM20, BD11, BJ24, BFGTM14, BM18, BCHM16, Chi08, Chi09, CFR04, DJM04, DG05, FGH14, GZED20, GHS12, GV04, HdlL07, Ipp11, IW23, KBS14, LMKY23, LW20, LKH^{+}22, NSS13, SAR13, SK08, TS23, Vil18, XCC07, YBO22, dLD22]. Methodology [WM14a]. Methods [AK18, AZAK22, CLN24, Chi09, CGS15, DDSW22, FMHM24, FKM23, Hen11, HI23, Kri15, LT03, LV14, MSB^{+}14, PVVCS21, SV09, SMS18]. Metric [Gie19, MM22, VLS13]. Metrics [LKH^{+}22]. Michelson [CFST08, Wil05]. Micro [MP13]. Micro-scale [MP13]. Microbial [DLRB19, YCG^{+}22]. Microcircuit [VCK09]. Microfluidic [HSV24]. Microscopic [MSB^{+}14, SGW09]. Microtubule [CSJM18]. Microtubule-Based [CSJM18]. Microtus [NPV12]. Microvascular [GCKW07]. Migration [HHBS22, SS14]. Migratory [GLW10]. MIMO [VQPVV23]. Minimal [CJN15, GH04b, GH04a, LPK15]. Minimizers [BS18, CMW11, LP22]. Mirror [BKPP24]. Missing [ESK24]. Mitigate [GHH^{+}21]. Mitigating [TS23]. Mixed [AEHV05, GS11, GL15, HKO17, KP23, KPK08, KVDC12, LGLC15, MDW23, RWK08, VBW13]. Mixed-Mode [GS11, GL15, HKO17, KP23, KPK08, KVDC12, RWK08]. Mixer [MM17]. Mixing [SUOL18, SD22]. Mobile [ABBC20]. Mobility [DS24]. Modal [TDL17]. Mode [AMBV22, ADR23, AM17, AK18, AZAK22, AYB19, BGZ16, CP06, Daw08, EMKB19, GFB03, GS11, GL15, HMD^{+}23, HKO17, HRS04, HHKB20, KP23, KPK08, KVDC12, KFB16, LFFC^{+}20, LV17, LP23, MHC09, PN20, PBK16, RK23b, RWK08, SDK20, TFTK22, TGPP19, VBW13, ZRDC19]. Mode-Locked [GFB03, SDK20]. Model [AY20, AGMS23, ABC^{+}22, ABM^{+}04, AMDH24, AHW21, ASH18, AY23b, APBB22, BSOM20, BCBD20, BKPS19, BCH14, BJSW08, BT16, BBR^{+}05, BSW16, BXB17, BF18, BCP24, BHV11, CBWS24, CWZ17, CDKS19, CB16, CB18, CW11, Chi17, CRSN07, CDH23, CKPP19, CC06a, CV09, CM07, CP12, CHS12, CEvdM24, CGHM18, CR11, CW18, DL10, DSC12, DA12, DH19, DDN22, DT13, DJD19, DK18, DKB16, DNDY16, DR22, DHK20, DLRB19, DAFM19, EWLH11, ES22, EW09, EK16, EFK17, FSS19, FSDAS21, Fer18, Fer20, FDS12, FS16, GLS23, GH21, GM15, GAHK03, GH04a, GB23, GLS21, GS09b, GWW19, GT18, GLNW15, GSJ23, GYAU24, GKM21, GKR24, GO02, GS11, GS13, GPTV17, HKZ18, HK19, HKLN20, HP20, HKM21, HKP22, HvHM^{+}14, HL22, HKB^{+}22a, HKB^{+}22b, HdRS23, HAS16, HKO13, HP17]. Model [HJL16, HJL17, HRYZ19, HS05, HP14, IBB^{+}10, IPS19, IW21, JJ20, KKP15, KKP16, KVB22, KRW13, KPR15, KSWW06, LFFC^{+}20, LT13, LT19, Lee22, LP23, LRK12, LHW23, MR06, MP09, MMNS22, ML12, MW14, MJJL12, MW17a, MCP09, MFE05, MW17c, NWW21, NSUW09, New14, NPV12, OPL21, OY09, PE18, PR22, PB22, PR24, PK05a, POR20, PES12, PK17, PMBM05, QT20, QSvdH19, RB21, RGAB16, RS11, RRW15, RPY20, RWK08, Rot22, SD17, SG10, SHdlL17, Sla20, SJLY17, SS14, SLF24, TPC24, TvH21, TKB17, TR19, Ton23, TS23, TZKS12, TW18b, TXKW17, TT20, Van06, VC17, VC19, VM09, VM11, VBG^{+}09, WZ17, WWZ19, WWXR24, Wid13, WCM08, Xia08, YCG^{+}22, YLWK16, ZZQ18, ZHKR15]. Modeled [ACK17, BN13, GKR24]. Modeling [ABPM22, AAC23, BFK^{+}22, BS19, BMCGW14, CM03, CSJM18, CF07, CZSV23, CGG^{+}22, ESK24, FDS14, GAL20, Guo10, HACY22, KN14, MDW23, SAA^{+}18, TW18a, VQPVV23, WSYTA23, Wil23b, YvLKL22, ZZ20, ZRP18]. Models [ABBC20, AHS14, BAB13, BR13, BEW11, BM24, BHLMW24, BE14, BFH14, CH10, CR23, CCL24, CP23, Coo08, CKCG19, CSRR08, DEL14, DDSW22, EGS24, FY13, FE10, FEIvdD12, GH04b, GHLP22, GK22, GLW10, GKC14, HGT15, Has21, HS09, IS17, KPW17, Kra21, KPR11, LT03, LRH12, MEvdD13, MSB^{+}14, MRB^{+}13, OK24, OPR24, PD18, PLST20, Pro20, RCG12, RSRT21, RL24, SAR13, SSS06, SSR18, SGW09, SBKS15, SARTA20, WZ12, Yak08, ZL20, dLD22]. Modern [SBKS15]. Modes [BS19, EKL06, GAKTWW21, SDK20, YW10]. Modified [ES22, HGS15, LP22, MZ11]. Modular [AST07]. Modulated [Com06, Daw09, WYM^{+}24]. Modulating [HdRS23]. Modulation [BXB17, CGG^{+}22, FDS14, WW02, YB22]. Molecular [Bok22, Tup09]. Molecule [ESZ04, GAKTWW21]. Molecules [CR23, WYM^{+}24]. Moment [BKS06, DDDGZ16, KM24]. Moments [BKS06, GP20, YM23]. Momentum [MM06, TD12]. Monoid [SAH21]. Monomials [AD15]. Monotone [Ged10]. Montreal [Chu21]. Moon [BE03]. Moore [Xia08]. Morales [AH09, AHARS18]. Mori [LTLA21, LTPL23]. Morral [DEP^{+}11]. Morris [NWKR15, New14]. Morse [BCHM16, DJK^{+}19, DMS22]. Moser [BM20a]. Mosquito [HRYZ19]. Mosquitoes [HRYZ19]. Most [SK08, TFC19]. Motifs [HDDL21, JC23, SP21]. Motion [BFK18, CDPVY21, CBR05, CHH23, CPY19, GH21, HACY22, KM08, LT21, LPS13, MG18, MRR06, MFN^{+}23, NSUW09, RS21, USW05, Van06]. Motions [SS12]. Motivated [SARTA20]. Motivations [RK18]. Motoneurons [GH04a]. Motor [AMDH24, SARTA20]. Moving [BFK^{+}22, HdRS23, HS15, PR24]. mRNA [CSJM18]. Multi [RLZT23]. Multi-pulses [RLZT23]. Multiagent [FGS^{+}23]. Multibump [BRW05]. Multichaos [DY17]. Multiclusters [BSY19]. Multidimensional [DA12, LMNT09, XSS20]. Multifronts [BST08]. Multigroup [DSC12]. Multilane [BFK^{+}22]. Multilayer [BCJ19, BMSY21, PK17]. Multilegged [AST07]. Multimodel [TS23]. Multioption [FGS^{+}23]. Multiparameter [AKK^{+}09, BRRS02, MRB^{+}13]. Multiple [AR12, AY23a, BKM22, CT17, CT19, DZ14, EVC18, FGDKC15, FS09, GH04b, GST05, GS18, GS13, Jef14, KPK23, KW08, KVDC12, LS05, LS15, LS16, LPK15, RCLR21, RS07, VSABM23, VBW13, WYD22, YW10, YC10]. Multiple-Timescale [KPK23]. Multiplex [BMSY21, RBBG20]. Multiplicative [GALS16, HH19, LP23, MV21, Sul23, VRS22]. Multiply [BFG21]. Multipulse [SD17]. Multipulses [BST08]. Multiresolution [KFB16]. Multiresonant [CR09]. Multiscale [CBK19, DTG^{+}16, FGH14, Lan16, PSV22]. Multisection [Sie02]. Multispecies [GHLP22]. Multistability [TX21, TZ22]. Multistationarity [BP16, FKdWY23, WF13]. Multistationary [JS17]. Multivalued [ABMS15, BM16, Bat17]. Mussel [GM15, HP17, SJLY17]. Mussel-Algae [GM15]. Mutual [Bal17, CSKR06]. Mutually [EKL06, Lee22]. Myocytes [KVB22].

n [FKdWY23]. n- [FKdWY23]. Nagumo [BHV11, CS18, CJ18b, CP17, CZ16, GK10, HS10b, HMS19, Ike23, KP23, QT20, RCG12, TvH21]. Nanoptera [DLP21, LP18]. Nanowires [RS21]. Natural [TBR23]. Navier [BBJ21, Buz24, FMT16, WZM23].

Near [AKO13, BFK18, CGK08, DRdRV18, HK05, HRS04, MW17c, MKO18, OZM11, PRK18, PRK22, Rod21, SJLY17, Wil23b, WR02, ZKE15, Agu15, AHGKM16, BK24b, DKO08, GS07, GKO18, PYGR06, PRCA^{+}23]. Near-Resonant [HRS04]. Near-Ring [MW17c]. Nearest [HdRS23]. Necessary [BCDG16]. Negative [CF20, Ged10, KKP16, PCG16, PCG18, RPY20, TFC19]. Negligible [Rob04]. Neighbor [HdRS23]. Neighborhood [GKO17]. Neimark [SM08]. Neocortex [SHdlL17]. Nernst [AEL08, BJL^{+}17, LLYZ13, ZL20]. Network [AD15, BHP^{+}21, BBR^{+}05, BXB17, BN13, CR23, CC06a, CKCG19, CTAA18, CGG^{+}22, CE04, DS24, DGMW12, DR10, EWLH11, EGF18, FB04, Fol11, GSJ23, GRSB19, Guo12, HK15, JC23, KC13, KB10, KN14, LT13, LT15, LT19, LR19, LZX22, MEvdD13, MMP16, NSS19, OK23, PCNL12, PB22, PID21, PJW05, Rod21, SAR13, SDT17, SBKS15, SMRB11, SBR06, STB15, TR19, Tro08, VCNSD24, WIN16, WWC^{+}18, ZBN09, ZLX20]. Network-Based [SBKS15]. Network-Organized [GSJ23]. Networked [AY23b, GYAU24]. Networks [AR12, ADR16, AD21, ADP08, AOWT07, AP16, AK06, BT19, BT21, BP16, BKPS19, BAB13, BCRR21, BCJ19, BSY19, BMSY21, BBK23, BP08, BK20b, CW17, CJ18a, CL16, CBR19, CP23, CMV23, CF12, Coo08, CD20, CATA20, CGH^{+}16, DS23, DZ14, DEL14, Des23, DIL^{+}24b, DP09, DKTG12, DB13, DGK^{+}21, FRB19, FG19, FvdSG20, FLWW22, FKdWY23, FE10, GFE20, GGP^{+}20, GP20, GST05, GL09, GMB16, GSB^{+}16, GD23, GC05a, GC05b, GK18, GS20, HHHY09, HSSY23, HDDL21, JKKL24, JS17, KSKJ20, KSG14, KPR12a, LGLC15, LA13, LA18, LE10, MCL^{+}20, MFVW17, MH17, MRB^{+}13, Mor15, MDIC24, MKM23, NB^{+}23, Oro14, OR19, PACM22, PEH22, RCG^{+}23, RS13, SP21, SSR18, Scl23, SAH21, SHA23, SK13, Soa17, SGP03, SG11, Ste14, SW24, TX21, TZ22, TW23, TL24, TFC19]. Networks [Ton10, TF21, TS07, Vas23, WHT13, WF13, YvLKL22, YCL08, Zha07, ZKCS19, ZRP18]. Neural [AMNB06, AH06, BT19, BT21, BW12, BN13, BC14, BK15, CCL24, CB16, CB18, CC06a, CEK22, CO04, CR11, DZ14, Dys20, EWLH11, Fay13, FB04, Fol11, FE12, GC05a, GC05b, Guo12, GS20, IM16, IMS15, JLG21, KE08, KFB08, KB10, KE13, KF14, KS14, Lai15, Lan16, LE10, Ly14, MPC^{+}22, MLTC21, NCA^{+}21, PE18, PB22, PK17, PbG09, RBK15, STW23, SvGKP24, Ton10, THF12, VF10, Vel13, WM14b, YvLKL22, ZME19]. Neurodynamics [Oro14]. Neuroendocrine [FGDKC15]. Neurologically [SARTA20]. Neuromechanical [JLG21]. Neuromodulation [EWLH11]. Neuron [ABPM22, Coo08, FEIvdD12, GPTV17, RRW15, RSRT21, SL12, SG10]. Neuronal [BHP^{+}21, Coo08, CSRR08, FE10, FDS12, FDS14, PR22, PJW05, SLS23, TS07, Tro08, Zha07]. Neurons [BTK16, CGL19, DKTG12, EVC18, GH04b, JMB^{+}13, KN14, LT13, LT15, LT19, Lee22, MRMM14, NC16, TB09]. Newton [BCHM16, NSS13]. Newtonian [Ver08]. Next [HdRS23]. Next-to-Nearest [HdRS23]. NFAT [Ren12]. Nile [WWZ19]. Nilpotent [EG06]. Nine [LFFC^{+}20]. Nine-Mode [LFFC^{+}20]. Ninño [KKP15]. Niño [KKP16]. No [OW20]. Node [Agu15, DKO08, FRB19, Kri20, LNOR21, SW16, Vas23, Wec05, ZKE15]. Node-to-Medium [FRB19]. Node-to-Node [FRB19]. Nodes [CATA20, SMRB11]. Noise [ACL23, AHS14, CCL24, CM16, CDW24, CTAA18, ERT11, EK16, FRB19, GIKR20, GALS16, GD23, HH19, KBGD^{+}23, LM19, LP23, LA13, Lu16, MV21, New14, OBK18, OE21, Sul23, THF12, WLW15, ZYO05, dWRS18].

Noise-Driven [OE21]. Noise-Induced [CM16, CTAA18, THF12]. Noisy [AP16, CL08, DDSW22, LE10, Ly14]. Non [AKO13, Bri19, CP23, GSDN15]. Non-Ergodic [Bri19]. Non-Markovian [CP23]. Non-REM [BAB13]. Non-Slip [GSDN15]. Nonautonomous [AH09, BFH22, BH23, Blö03, BIPK24, CZ15, DNO23b, DNO23a, EK10a, EK10b, HP14, Hül16, LNOR21, MCZM18, SCD07, WLW15, WZ16, YBO22]. Noncentral [Agu15]. Nondeterministic [CJ11]. Nonexistence [LR22]. Nonfeedback [Wil19]. Nongeneric [Pat03]. Nonhomogeneous [BFG21]. Nonhyperbolic [BJK20, CD10, Hül05, MB14a, MB14b]. Nonidentical [MMP16]. Nonintegrability [AHARS18]. Noninteracting [FLWW22]. Noninvasive [PSSJ23]. Noninvertibility [BCDG16]. Noninvertible [Hül16]. Nonisothermal [FvdSG20]. Nonlinear [BKS06, BGH^{+}24, BFG21, BHLMW24, BDG^{+}16, BK15, BTBK14, BFH14, BIPK24, CDPVY21, CBK19, CSS17, DSC12, DLP21, Don24, FG20, Fol11, GAKTWW21, GVNS09, GS07, GCY22, HS03, HDDL21, IBB^{+}10, JZ11, KKK20, KM10, KPT13, KBGD^{+}23, Kur17, LLZ17, LHW23, MV14, MHC09, MRS17, MK24, NP15, NPRW19, Oro14, OPL21, PD20, PSW12, RSRT21, SDW15, SBB10, SK08, SCD07, SLF24, SMM21, TKKCG16, WR13, Wei03, Wil23a, Wil23b, WS24, WW20, XCC07, YB22, Zha07, ZYO05, ZME19, ZL14]. Nonlinearities [Blö03, DGK^{+}21, Rot22]. Nonlinearity [DP08, VD13, ZR20]. Nonlinearly [Jac06]. Nonlocal [BFK^{+}22, BT16, BBK23, GHLP22, LT03, Lai05, She14, VNSG08, XSW24, Zha07]. Nonlocally [DE22]. Nonmonotonic [PCG16, PCG18]. Nonorthogonal [LZX22]. Nonpolar [ELT22]. Nonresonant [QCARL21]. Nonsmooth [CD10, FMOW03, HE15, LM19, LdST09, NWW21, NC16, TDL17]. Nonspherical [Kur17]. Nonstandard [LW20]. Nonuniqueness [AEG24]. Nonwandering [FH23]. Nonzero [ANR18]. Normal [BK24b, CLJ15, De 03, GAKTWW21, GMM08, Gle14, PPK14, QCARL21]. Normalization [DDGK13, Has21]. Normalized [NC21]. Novel [AdBG^{+}09, KM08, MLTC21, OdBS08, TD08]. Novikov [De 03]. Nucleation [BSW16, CL08, DEP^{+}11, GW24]. Nucleation-Diffusion [CL08]. Nucleus [GB23]. Null [Noa08]. Number [CP17, GSDN15, Hsu19, HTV20, LPK15, MW10, OZM11]. Numbers [WZ12]. Numerical [AdBG^{+}09, BD11, BRRS02, BDG^{+}16, BC23, CL08, DJM04, De 03, DD13, DDGK13, DKZ17, DG05, DKB16, FGLdlL17, FdlL17, FKM23, GYdlL21, GJNO22, HLJ23, HG10, KKK20, LWK23, RAM15, SV09, SAV12, Sie02, Tup05, UW14, WV19, WS06, WS09, YNZ22, ZDG19]. Numerics [CL13, CLJ15, DHMO05, vdBL08].

Objects [CG18, FGLdlL17, GL17, GKO17]. Observables [BBJ21, COT19, DKB16, KKBK20]. Observation [GKMS06, OBK18, OY03]. Observed [GBIB06, OPR24, dLD22, dWRS18]. Observing [LO10]. Obstacles [BFK^{+}22]. Octahedral [LM16]. Odd [Yos17]. ODE [CKCG19, GYdlL21, IW21, SZ13]. ODEs [BHLM21, Chi09, MM22]. Off [PPM14, RWZ21]. Off-Hotspot [RWZ21]. Ohta [CW18]. Old [LR18]. Omega [SRS09]. On- [RWZ21]. One [BLL12, BJSW08, BXB17, CW22a, CZ15, CW18, CL09, Des23, DL22, DvHX16, EKO04, EKO05, GL13, Hül05, JZ11, KZ21, LR20, MP13, MO15, RLZT23, SS14, TZ22, TW23, ZLL22]. One-Dimensional [BJSW08, BXB17, CW22a, CZ15, CW18, CL09, DL22, EKO04, EKO05, GL13, KZ21, MP13, SS14, TZ22, RLZT23]. One-Parameter [BLL12]. One-Sided [ZLL22]. One-Stop [LR20]. Online [KPW17, ZRDC19]. Online-Adaptive [KPW17]. Only [Des23, ZLL22]. Onset [GCKW07, GS07]. Open [DKaK^{+}08]. Operational [WE18]. Operations [ADR16]. Operator [AM17, ABG^{+}17, BLDK18, DTK20, GKR24, MCZM18, WM22, Wil23a]. Operators [AM23, BKJ15, CII23, KM17, LP21, LTPL23, SK22]. Opinion [BCP24, CP23, HKB^{+}22a, HKB^{+}22b, MJB14, RB21]. Optical [CM16, EKL07, GKML09, HK05, TKB17, YvLKL22]. Optimal [BE03, BTK12, BR23, BIPK24, DR10, FKS20, GL24, KVX04, MPC^{+}22, SKL23, STB15, TGPP19, TD12, VLS13, WM14a, WM14b, Wil19, Wil21, YNN^{+}23, ZB20, ZL14]. Optimally [BS19]. Optimization [BKPP24, FGHC16, FG20, GMY18, HMN09, KM10, KHG21, LFFC^{+}20, LD18, YCG^{+}22]. Optimized [AK18, LP23, SUOL18]. Optimizing [MM17]. Orbit [Agu15, CDPVY21, CKMW12, Gie19, ML12, MO20, SPCT12, SZ13]. Orbital [CPY19, NP15]. Orbits [AKO13, BCPS08, Cap12, CL13, CLJ15, CGP16, CCHZ20, CH03, DM20, DMS05, FdlL17, FKM23, GKMS06, GL17, GYdlL21, GLJY23, GJ17, GHS12, GK10, HKO17, HLJ23, Hül05, JM13, KW08, LFFC^{+}20, Las18, LO08, Lu16, MXYZ16, MFE05, NS15, NSTZ20, Pos09, SLF24, WZ16, WB17, WR02, WS06, WS09, dlLJ16]. Order [AMBV22, BGZ16, BBK23, BS19, CP06, DEV04, GS22, GJNO22, GJ17, HK19, IW23, KKV18, Kra21, LV17, LZX22, NP15, PW21, PW24, QCARL21, Rad13, STW23, SBV23, VQPVV23, WM24, WW02, Wil23b, WS24, ZLX20, vdBKV11]. Ordinal [MKM23]. Ordinary [DNO23b, DNO23a, KL21, MK24].

Organization [DJCJC22, PRCA^{+}23]. Organized [GSJ23]. Organizing [FDS12, LBR18, LHW23, RAM15]. Orientability [AKO13]. Orientation [CB16, DMCK15]. Orientation-Dependent [DMCK15]. Oriented [DJD19, DKZ17, GZED20, Jam10, JPK^{+}22]. Origin [ETS21]. Oscillating [Kur17, RWZ21]. Oscillation [KKP15, Rob16, Wil21, KKP16]. Oscillations [BCF^{+}18, BYK08, CH10, DE16, ERT11, FY13, FE10, GCKW07, Ged10, GS11, GL15, HKO17, Hsu19, KSG14, KP23, KPK08, KVDC12, LT13, LT15, NWKR15, NPRW19, OE21, PRAC23, PCG16, PMBM05, RWK08, TS23, WFM^{+}14, WS24, vdDZ04]. Oscillator [AMNB06, ASH18, BBK23, BBK17, CK15, GS09b, GHS10, HK05, KNWH11, KM08, KS11, Kri21, LFOG17, Lin06, NB^{+}23, OK24, PW24, RGAB16, SKL23, ZBN09]. Oscillator-Follower [ZBN09]. Oscillators [Ada23, AOWT07, BSY19, BM18, BM20b, BCF^{+}18, BMWY18, CJ08, CHH24, CH14, DE22, DB11, FGDKC15, HKL14, HNP16, HLLP18, HKZ18, HK15, KE08, KT24, KP23, LZX22, LA13, LA18, LE10, Ly14, Lyu18, MV21, PFGV14, PW21, PP08, RBI21, RT02, RHT13, SS12, Ste23, TDL17, VH08, VM08, WE18, XWC^{+}23, ZZQ18, ZLX20, ZL14, ZZ09]. Oscillatory [CTAA18, CATA20, DG05, GSDN15, GLNW15, IW21, KN14, KS07, LT19, NX22, PLNW19, Rot22, SS04, SWR05]. Other [Kri20, LR18]. Outbreaks [GHH^{+}21]. Outlier [GS20]. Oxygen [GB09].

p53 [CRSN07]. Pacemaker [BBR^{+}05, BGO11]. Painlevé [CHK22, ELT22, KH18b, KH18a, NVC18]. Pair [BT21, FE12, Lee22]. Pair-Replica-Mean-Field [BT21]. Pairs [BM16, Bat17, FT12]. Pairwise [BS18]. Paleoclimate [ML12, QSvdH19]. Pancreatic [GS09b, WFM^{+}14]. Pandemic [Chu21]. Paper [CFdSL22]. Parabolic [DV19, IS17, MRS14, WLW15]. Parabolic-Elliptic [IS17]. Paradox [CHK22, NVC18, PE10]. Parallel [JO09]. Parameter [ACFK09, BLL12, BdCT12, CV09, CLBdB09, CGH^{+}16, CGHM18, CZ16, DKZ17, FKdWY23, FH22, GZED20, GKSV23, GKS03, HMP02, YNN^{+}23]. Parameter-Dependent [HMP02]. Parameterization [BHM24, BJM20, BJ24, CLJ15, FMHM24, GJ17, HdlL07, Vil18]. Parameters [GALS16, HMD^{+}23, KM10, KPW17, SRS14, WB14]. Parametric [ADR23, CW22b, HK05, RABK19, SS16, VQPVV23]. Parametrically [RS09]. Parity [Yos17]. Part [BFH22, Bat17, BH23, HBB13a, HBB13b]. Partial [AK10, CHH24, FG19, RABK19, SÜvLM16]. Partially [OPR24, dLD22]. Particle [AV19, CHS12, CFR04, HK19, HHBS22, PZ22]. Particle-Mesh [CFR04]. Particles [AJ14, HSV24, Leg11, Leg13, LPS13, SDW15, VHS22]. Partition [MKM23]. Partitioning [FA13]. Partitions [BGOŻ08]. Passage [GKSV23, LSB11, NS13]. Passive [BB16, Mun11, SL12]. Passivity [RS11]. Passivity-Preserving [RS11]. Past [BKPS19]. Pasta [Yos17]. Patch [KPR15]. Patches [BMCGW14]. Patchy [CLW24, GLW10]. Path [BU23, CL11, GAG^{+}21, GZED20]. Path-Based [BU23]. Paths [BP08, DHP23, GAS18, Gor13]. Pathway [LZF23, Ren12]. Patrol [TW18b]. Pattern [ASH18, AH19, BB16, CE04, Daw08, DL10, GLS23, GH04a, GLS21, LZX22, PLNW19, PES12, TR19, WCM08, ZLX20]. Patterning [MLTC21, PA19]. Patterns [ANR14, ANR18, APBB22, BAA^{+}19, BK15, BCF^{+}18, CDKS19, CB16, CW11, CZSV23, CR09, DL21, DP09, DGMW12, DvdP02, DRdRV18, EVC18, FA13, GW24, GL13, GST05, GWW19, HRS04, IM16, IR22, JR05, LZX22, LSAC08, MDW23, PR22, Rot22, RRW14, RS09, ACJ24, SAV12, SAH21, SD17, SSR10, SBB10, SM11, SGP03, TS07, TXKW17, TWW18, TT20, UE15, UW14, VVZ15, VSC23, YHM^{+}02, ANR18]. PCR3BP [Cap12]. PDE [GL17, IW21, LT03, Law16, NSS06, NSS13, Xia08]. PDE-ODE [IW21]. PDEs [AK17, CLN24, HS03, KKK20, Wei03]. Peak [KRW13]. Peanut [WW20]. Peanut-Shaped [WW20]. Pedestrian [CHS12, PSSJ23]. Penalization [LM19]. Pendulum [BMSM24, BCGH08, CBR05, LCMA05, Moe15, WZ09]. Pentagon [TD12]. Perception [GRS^{+}22]. Perfect [MRR06, Mun11]. Period [DS19, LP18, Pos09, PMBM05, SS07]. Period-2 [LP18]. Period-Doubling [DS19, SS07]. Periodic [AVV23, AK17, AG05, BCPS08, BCRR21, CL13, CLJ15, CCHZ20, CLZ21, CKK^{+}09, CKMW12, CH03, CPY19, CDS10, CZ16, DDGK13, DP09, DM20, DC16, DRdRV18, FMHM24, FdlL17, GNR18, GL17, GBK15, Gie19, GYdlL21, GLJY23, GJ17, HSS22, HdlL07, HLJ23, HRS04, HS09, JZ11, KNWH11, LFFC^{+}20, Las18, LP22, Lee22, LR20, LW02, LO08, LBHM05, LV14, MW16, MPY11, MFE05, NS15, NSTZ20, PVVY17, PSW12, PRAC23, PCG16, PCG18, Pos09, QSvdH19, Riv13, RHT13, SK13, SD17, She14, SS11, TD08, TDK18, VVZ15, WWZ19, WB17, WSWK12, WR02, WS06, WS09, YPMD08, ZKE15, Hu23, SOV05, SS12]. Periodically [ANR14, ANR18, BYK08, CDW24, GKC14, LFOG17, Rod21, VCK09, ZG11, XSS20]. Periodicity [Bal05, DZ14]. Periods [WWZ19]. Permanence [BH20]. Permanent [BR23, HdRS23, ZL20]. Permutation [FHKK21]. Perron [CII23, SK22]. Persistence [DMS22, DsMS24, GLJY23, KPR12b, Lu16, MMP16, MKM23, SS09, dlLK19]. Persistence-Like [DsMS24]. Persistent [DJK^{+}19, JKKL24, PID21]. Perspective [GKMS06, LPR22, MB14a, MB14b]. Perspectives [Law16, PE10, Sco13]. Perturbation [AHGKM16, BHLMW24, CCL24, CT17, CT19, Chi09, CBR05, FLWW22, FKS20, GYdlL21, HS05, IPS19, JPK^{+}22, KPK23, KS11, LT17, LW20, LdST09, Lu16, WGCT21]. Perturbations [BMSM24, CL14, FH23, GLJY23, JS06, KL17, LT21, NSTZ20, dlLK19]. Perturbative [CDT21]. Perturbed [Bal05, Bal11, CDW24, GKKZ05, Rod21, WE18, Wil23b, WW20]. Phantom [EDKC16, KVDC12]. Phase [ASH18, AOWT07, BSY19, BBK23, BM20b, BF18, BMWY18, CK15, CM16, CB16, Chi17, CH13, CDH23, CMW11, DE06, Fer20, GVY17, GKS03, GH09, HKZ18, HdlL23, HL18, KT24, LR18, LA18, Ly14, MRMM14, SPCT12, SG10, SKL23, TLRB11, VM08, WE18, Wil21, Wil22, WS24, WHT13, XWC^{+}23]. Phase-Amplitude [KT24, Wil22, WS24]. Phase-Amplitude-Coordinate-Based [Wil21]. Phase-Conjugate [GKS03]. Phase-Coupled [VM08]. Phase-Lag [Chi17]. Phase-Locked [CH13]. Phase-Locking [BF18, CDH23, Fer20, HKZ18, VM08]. Phase-Selective [SKL23]. Phaseless [SPCT12]. Phases [MS15]. Phenomena [ADF20, ADF21, AEG24, CHK17, HdlL23, KKP15, Kri21, OW20]. Phenomenon [WSYTA23]. Phone [ABBC20]. Phonetic [GT18]. Phosphenes [DE06]. Phosphorylation [FKdWY23]. Photorefractive [KKC06]. Physical [DKB^{+}23]. Physics [ESK24, PD20]. Physics-Informed [PD20]. Physiologically [BAB13]. Piecewise [AG05, BSKR16, BdCT12, CFST08, CJ11, Coo08, DDDGZ16, DEL14, FGDKC15, FPT12, GRS^{+}22, GHH^{+}24, GHS12, KK19, KH15b, KH15a, LPH22, PR22, Rob16, RCG12, SM08, TDL17, ZR20]. Piecewise-Differentiable [BSKR16]. Piecewise-Linear [DEL14, RCG12, TDL17]. Piecewise-Smooth [GHS12, LPH22, Rob16, SM08]. Pinned [GBH11, ACJ24]. Pinning [DMCK15]. Pipe [BCH14, Rob04]. Pipes [Pro20, Rob04]. Pitaevskii [TKKCG16, WYM^{+}24]. Pitchfork [GKSV23, Kri20]. Placement [BKM22]. Planar [AVV23, ALB^{+}10, Bal05, BCH10, Coo08, DD13, DMCK15, DKaK^{+}08, DDCK11, EJK20, EKO04, FDS12, FPT12, GHH^{+}24, HKO13, KOP07, KH15b, LSAC08, Llo19, Llo21, Mak17, MM06, MRR06, MG16, NSTZ20, Rob13, Rob16, SM08, SZ13]. Planck [AEL08, BJL^{+}17, LLYZ13, ZL20]. Plane [BCHM16, BdCT12, GYdlL21, LR18, PYVG14]. Planetary [FKM23]. Plankton [TR19]. Planning [MRR06]. Plant [BMCGW14]. Plasma [ZHKR15]. Plasticity [LT17]. plateau [VBW13]. Play [KM17]. POD [KVX04, TV14]. Poincaré [AHGKM16, AY23b, EKO05, LCKO08, Mat18, MSW15, SDR09, Wil10]. Poincaré-Type [Mat18]. Point [CFST08, CHS12, DRC09, De 07, Hül05, LT21, SL19]. Points [AHGKM16, ABCV23, BJ22, DKB^{+}23, GHH^{+}24, KM10, OW20, SS09, TFBN21, UN21, Zgl02, ZKE15, vdBKV11]. Poisson [AEL08, BJL^{+}17, BC09, CT22, LLYZ13, MRS17, ZL20]. Pol [BEG^{+}03, GHW03, PFGV14]. Polarity [MDW23, ACJ24]. Polarity-Driven [MDW23]. Polarization [BCP24, VC17, VC19]. Police [GW24, TW18b]. Policing [RWZ21]. Policy [AJB^{+}16]. Polychromatic [PSW12]. Polynomial [BK20a, Bre23, DIL^{+}24a, LFFC^{+}20, LT21, YM23]. Pool [SRMPM08].

Pool-Boiling [SRMPM08]. Population [AY20, DH19, DNO23b, DNO23a, GM15, HRR^{+}03, HJL16, HJL17, HRYZ19, LRH12, MJJL12, RPY20, SAA^{+}18, SLS23, SvGKP24, Yak08]. Populations [AY23a, DvG09, EVC18, FH14, VM08, WM14b]. Porous [LWK23]. Portrait [LR18]. Posedness [AHW21, MRS14]. Position [BGB05]. Positive [Bri20, CF20, KKP16]. Positivity [Bri19]. Posteriori [Bre23, FGLdlL17, GL17, JM13]. Potential [ABCV23, KKC06, LLYZ13, LZ23, MRS17, PVVY17, SDW15]. Potentials [BLL12]. Power [GS22, WF13]. Power-Law [WF13]. Practical [BHLZ18, CR12, DRH19]. Prandtl [FGHM^{+}20, GSDN15]. Precession [HL02]. Precipitation [GLS23, GMS11]. Preclusion [WF13]. Predator [Hsu19, LHW23, PVMP17]. Predator-Prey [LHW23]. Predicting [DAFM19, FA13, HRR^{+}03]. Prediction [VLS13]. Predictive [BTK12, POR20]. Predictors [BK24b]. Preference [PPM14]. Preisach [GKR24]. Premixed [GB09]. Presence [GIKR20, GMM08]. Preserving [DM09, DM12, FM16, GMCM19, GMCM21, HdlL23, Jam10, JL10, KM24, LRR08, MSM17, RS11]. Pressure [BCH14]. Prey [Hsu19, LHW23, PPM14, PVMP17]. Primer [BT11]. Principal [DM20]. Principles [DHP23, DEV04]. Prion [ABPM22]. Prioritizing [RK18]. Probabilistic [PD20, PLST20]. Probabilities [LM19]. Probable [SK08]. Problem [AVV23, ACMM20, ABCV23, BCPS08, BCH10, BC21, CR12, CCHZ20, CLZ21, CH03, CPY19, CT22, De 03, DV19, DHW22, GNR18, HGS15, KHG21, KO24, LP22, LO08, MG18, MO20, Moh19, PRAC23, Riv13, Rob13, RS07, RS16, XSS20, YPMD08, vdBGW15]. Problems [AEHV05, BS18, CT17, CT19, HI23, HMP02, HACY22, KS11, Kri20, KVX04, Law16, LD18, PbG09, SDW15]. Procedures [SA13].

Process [BN13, CBR19, DA12, JPK^{+}22]. Process-Oriented [JPK^{+}22]. Processes [DDDGZ16, GMM08, SP21, SA13, YB22, dWRS18]. Producer [YLWK16]. Producer-Grazer [YLWK16]. Product [BC14, PEH22]. Product-form [PEH22]. Production [ABPM22]. Profiles [CSS17, De 07, MZ11]. Projected [LKH^{+}22, dLDF^{+}18]. Projecting [GKKZ05]. Projection [AK18, KN14, LTPL23]. Proliferation [SS14]. Proof [BCGH08, CWZ17, CLZ21, CZ15, CW18, Has21, SZ13, TF21, Wil05, WZ09, WSB16, WB17, Zgl02]. Proofs [Cap12, FdlL17, GLJY23, dlLJ16]. Propagation [ADF20, ADF21, BW12, CJAMV20, GP20, HS09, HJSS23, KFB08, KS14, MHC09, SW24, Ton10, VVZ15, Vel13, WL22, YM23]. Propagations [Ton19]. Propelled [Mun11]. Properties [AM17, AM23, Bat17, Bri19, CK15, DT13, FSS19, HBB13a, HBB13b, OP08]. Propofol [MK12]. Prospects [BDG^{+}16, FH14]. Protein [ABPM22]. Prototypical [KPK08]. Proven [SAR13]. Pseudo [VBW13]. Pseudo-plateau [VBW13]. Pseudogenerators [BKJ15]. Pseudospectral [BDG^{+}16, dWSLD21]. PT [KPT13]. PT-Symmetric [KPT13]. Pullback [AM23]. Pulsatile [EVC18]. Pulse [BCBD20, CF07, CV09, DKTG12, DK03, DRdRV18, GFB03, HdRS23, HS10b, KF14, Lin06, LE10, Lyu18, TvH21, YNT14]. Pulse-Coupled [DKTG12, LE10, Lyu18]. Pulse-Driven [Lin06]. Pulsed [GLS23]. Pulsed-Precipitation [GLS23]. Pulses [Fay13, FB04, Fol11, GC05a, GC05b, HSS13, Jac06, KFB08, MMNS22, NUY05, PJW05, VD13, WIN16, RLZT23]. Pulsing [TKB17]. Pumping [Pro20]. Punctual [AV19]. Pyragas [PPK14].

QR [CC06b]. Quadfurcation [BM20a].

Quadratic [BM20a, DGG16, DM09, GCY22, HM22, HM24, Kra21, LNOR21, MW14]. Quadratic-Bilinear [Kra21]. Quadratically [HKP22]. Quadratics [Noa08]. Quadratization [BIPK24]. Quadristability [LHW23]. Qualitative [HP14, LZF23]. Quantification [Bal05, KL21]. Quantifying [AP16, CZSV23, MM22]. Quantitative [ADF20, ADF21, GNR18, Xia08]. Quartic [AHARS18]. Quasi [BFK18, CPY19, ETS21, FLWW22, FMHM24, HdlL07, HRS04, Hu23, LV14, MRS14, SOV05, SS12, XSS20]. Quasi-Equilibria [BFK18]. Quasi-Linear [MRS14]. Quasi-Periodic [CPY19, FMHM24, HdlL07, LV14, HRS04, Hu23, SOV05, SS12]. Quasi-periodically [XSS20]. Quasi-Steady-State [ETS21, FLWW22]. Quasilinear [NS13]. Quasipatterns [IR22, RS09]. Quasiperiodically [SM20]. Quasiperiodicity [DY17]. Queueing [NPRW19]. Queues [LPR22, NPRW19]. Quintic [RLZT23]. Quiver [NRS20]. Quorum [FRB19].

Rabi [WYM^{+}24]. Rabi-Coupled [WYM^{+}24]. Radial [BAA^{+}19]. Radially [GBCV20, vdBGW15]. Radiation [CM16, Van06]. Ramis [AH09, AHARS18]. Ramp [GKSV23]. Random [ABI24, BGB05, BHLM21, BFH22, BK20a, CLL12, CFG21, ESK24, GAG^{+}21, GALS16, GKM21, HQW^{+}20, JS06, KL21, SMRB11, WLW15, WHT13, ZYO05]. Randomized [EMKB19]. Randomly [Law16]. Range [AH06, BBR^{+}05, CMW11, CZ16, EGF18, YPTW24]. Rank [BTBK14, GMY18, HARB21, MO15]. Rank-1 [GMY18]. Rank-2 [GMY18]. Rapid [BMSM24, PVMP17]. Rarefactions [SS09]. Rate [CJN15, CO04, DRC09, GB23, GKR24, HE15, HCW23, HRYZ19, Jel24, KHG21, Kie20, LNOR21]. Rate-Induced [HCW23, Kie20, LNOR21]. Rates [BU23, CLW24, Dys20, FSS19, FSDAS21, FS09, Hsu19, Ipp11, SDK20]. Ratio [DGG16]. Rayleigh [FGHM^{+}20]. RC [RS11]. Reachable [CC06b]. Reaction [ADF20, ADF21, BvdBV18, BP16, BHLMW24, BK20b, CW17, CJ18a, CR23, DIL^{+}24b, DK03, DB13, FG19, FvdSG20, FLWW22, GW24, GSJ23, GD23, GS13, GK18, HH19, HH20, HN14, HSSY23, HKO17, HCW23, HP17, IS24, JS17, JC23, MFVW17, MP13, MRS14, MFN^{+}23, NUY05, NX22, PLNW19, PEH22, PID21, Rad13, SW21, SRS09, TX21, TZ22, TW23, TF21, TW18b, TXKW17, TWW18, UW14, Vas23, VSABM23, WZ12, WR13, WW20, Wri10, XSW24, YCMP22, ZKCS19]. Reaction-Diffusion [ADF21, HCW23, IS24, XSW24]. Reactions [LRK12]. Reactivity [HLvdD22]. Real [AKO13, FKM23, SG11, Ste14, WZ16]. Real-Life [FKM23]. Reality [OY03]. Realizability [BM15, Bri19]. Realization [BL08]. Realizations [Des23]. Rebel [FHKK21]. Rebound [MK12, MLTC21]. Recollisions [DJCJC22]. Reconstruction [BKM22, BHP^{+}21, Bri19, CGS15, MCL^{+}20]. Recovery [FH22]. Rectangular [VHS22]. Rectifying [BLDK18]. Recurrent [DF19, GMS11, LE10, OR19]. Recursive [HMD^{+}23, LS05]. Reduced [BS19, BEG^{+}03, CHH23, KR21, KPW17, Kra21, Wil23b, WS24]. Reduced-Order [BS19, Kra21]. Reducible [JO09]. Reduction [AHS14, BH23, BK24b, BM20b, Chi17, CF20, CT22, CEvdM24, CGHM18, DIL^{+}24b, DGMW12, DTG^{+}16, EWLH11, FLWW22, IBB^{+}10, KBGD^{+}23, KT24, NRS20, PE18, RS11, SAR13, WB14, Wil22]. Reductions [DsMS24, ETS21]. Reflections [DL22]. Regime [GPTV17, MW17c, TBR23]. Regimes [BTBK14, KGB^{+}17]. Region [BF18, FKdWY23, Fer20, FH22, FH23, GKS03, Llo19, Llo21, NSS06]. Regions [AAC23, KRW13]. Regression [GS20, LTPL23]. Regression-Based [LTPL23]. Regular [AHGKM16, SHK13, Soa17, SG11, Ste14, UE15]. Regularity [OP08]. Regularization [BBK17, CKH21, KK19, SS09]. Regularizations [AEG24, KH15b, KH15a]. Regularized [BHLMW24, IPS19, PB10]. Regulate [CJAMV20]. Regulation [CM07]. Regulator [GCY22]. Regulatory [BAB13, BXB17, CGG^{+}22, DEL14, GSB^{+}16, LGLC15, MW10]. Rejection [CGHM18]. Relapse [DAFM19]. Related [GFE20, KRW13]. Relating [SMRB11]. Relation [KT24, Tup09]. Relations [AEL08, KC13, Rad06]. Relationship [FSDAS21]. Relative [BHL16, CEvdM24, ESZ04, HGS15, JLG21, LBHM05, MR18, Pat03, Rob13, WR02, WS09, WS14, YY19]. Relaxation [BS18, Hsu19, KS11, LFOG17, Rob16, VH08]. Relay [KNWH11]. Release [HRYZ19]. Reliable [FH12]. Relief [BCH14, EPCL05]. REM [BB12, BAB13]. REM/Non [BAB13]. REM/Non-REM [BAB13]. Remote [KHG21]. Removal [FSS19, FSDAS21]. Renormalization [ACMM20, Chi08, Chi09]. Reopening [Chu21]. Repetitions [RCLR21]. Replica [BT19, BT21]. Replica-Mean-Field [BT19]. Replicator [DvG09]. Representation [CGH^{+}16]. Representations [NRS20]. Reproduction [WZ12]. Repulsion [BFH14]. Repulsive [CH14, HKLN20, JJ20, LTB09]. Reservoir [Hon21]. Resetting [BGO11, GH09, Ly14, SPCT12]. Residual [BGZ16]. Residuals [ESK24]. Resistance [HDDL21]. Resolvent [SMM21]. Resonance [CHK17, GBK15, GH05, HK05, KKP15, KPR12a, LR19, LHRK04, MO20, NS13, SS16]. Resonances [CSKR06, CGP16, CG18, DM12, HdlL23, MG18].

Resonant [HRS04, HL02, RAM15, XCC07]. Resonate [CGL19]. Resonate-and-Fire [CGL19]. Resource [YCG^{+}22]. Respiratory [BBR^{+}05]. Response [ABPM22, CK15, CW22a, FH12, KT24, LR19, Rot22, SPCT12, SG10, WE18, WG15]. Responses [Ni23, RSRT21, WGCT21]. Restricted [ABCV23, BCPS08, BCH10, CR12, DV19, MG18, RS07]. Restriction [DS23]. Result [PCG18, VM11]. Results [BP16, MSB^{+}14, MO20, RAM15, UW14]. Retrieve [LZX22, ZLX20]. Return [OW20, WZ18]. Revealed [CK15, LDB20]. Revealing [PSSJ23]. Reversal [HJSS23]. Reverse [VC12, VCNSD24]. Reversible [BH20, CFST08, Des23, KW08, Kri20, NSTZ20, XSS20]. Reversing [FP16]. Revisit [CL11]. Revisited [SGW09]. Reynolds [OZM11]. Rhythms [CK15, LT19]. Ribosome [BSOM20]. Rich [KPR15]. Ricker [HP14]. Rigid [CFR04, CEvdM24, Pro20, SDR09, Ver08]. Rigid-Lid [CFR04]. Rigorous [AM06, CLN24, CL13, Chu21, DJM04, DHMO05, DFT08, GN14, GJM12, JM13, KKJ18, KZ21, KS14, Mat11, MO20, SW14, vdBL08, vdBGW15, vdBDLJ15, vdBLQ21]. Rikitake [TAtN09]. Ring [AV19, BCH10, KR11, MW17c, PR22, ZZ09]. Rings [BC15, BCF^{+}18, BMWY18, NX22]. Rivalry [DGMW12, KB10]. River [YNZ22]. Rivers [LR18, LR22]. Road [BP08, SGW09]. Robin [LK15]. Robot [AST07]. Robust [AZAK22, BKM22, DG05, FH12, FJ18, GMM08, GS20, ILM20, KKK20, KM10, KPR12a, LP23, STY23, WM22, YNT14]. Robustness [ACK17, BAB13, JC23, MM22]. Rock [CFdSL22]. Rocking [CHH23]. Role [BHLZ18, DS24, EJ16, HL22, RCLR21]. Roles [JLG21]. Roll [BAA^{+}19]. Rolls [MJM05, vdBDLJ15]. Root [BMCGW14]. Rose [LCDS12]. Rössler [WSB16]. Rotating [Com06, DE22, GSDN15, LL08, PRY23, UN21, Xia08]. Rotation [BN19, ESZ04, FHP22, TD12]. Rotational [DMS05, RSTY12]. Rotationally [CH10, HGS15]. Rotations [AG05]. Rough [DHP23, GAG^{+}21]. Routes [MLTC21]. Routing [BKPS19]. Rules [BHLMW24]. Rumor [RCLR21]. Running [GAHK03].

Sacker [SM08]. Saddle [AKO13, Agu15, CKMW12, DK18, FKO18, GKO17, GK09, Kri15, LRK12, LNOR21, SW16, Vas23, ZKE15]. Saddle-Center [CKMW12]. Saddle-Node [Agu15, LNOR21, SW16, Vas23]. Saddle-Type [Kri15]. Saddles [Bal11, MW17b]. Sakaguchi [WYD22]. Salerno [MCP09]. Same [PP12]. Sampled [BMMP20, Jud20]. Sampler [GIHLS20]. Sampling [CBK19, SB10]. SARS [GHH^{+}21]. SARS-CoV-2 [GHH^{+}21]. Satellite [CDPVY21]. Satellites [KPR11]. Saturable [TKB17, YC10]. Saturated [HRYZ19, MW17c]. Saturation [KSWW06]. Scalable [AZAK22]. Scalar [BL08, CHK17, DNO23b, DNO23a, PE18]. Scalars [HKK20]. Scale [ABPM22, BCGFS13, CT17, CT19, CH13, Daw08, DK18, FGDKC15, FS16, GS13, HGT15, Kri21, KPK08, KVDC12, NWKR15, NS15, RCG^{+}23, SMS18, TS07, Tro08, YW10, MP13]. Scales [FS09]. Scaling [CSKR06, KS11, TFBN21, TWW18]. Scattered [FJ18]. Scenario [PSSJ23]. Schema [AKK^{+}09]. Scheme [FMT16]. Schemes [PSSJ23]. Schistosomiasis [ZZ20]. Schizophrenia [VCK09]. Schnakenberg [TXKW17, UW14]. Schrödinger [CSS17, Hu23, Jac06, MRS17, NP15]. Schwarzschild [AVV23]. Schwarzschild-Type [AVV23]. Scissors [CFdSL22]. Scott [CW11, SD17, SWR05]. SDE [Law16]. Sea [HAS16]. Search [FKM23]. Seasonal [KKP15, ZZ20].

Seasonality [WZ17]. Second [BGZ16, GS22, HK19, LZX22, NC21, Rad13, STW23, ZLX20]. Second-Order [GS22, HK19, LZX22, STW23, ZLX20]. Secretion [EVC18]. Section [HSV24, SDR09]. Sectional [HL18]. Secular [CGP16]. Seen [BTK16]. Segment [PYGR06]. Selected [BSKR16]. Selection [BLL12, BHLMW24, GMS11, VCNSD24]. Selective [SKL23]. Selectivity [CB16]. SELEX [LS15, LS16]. Self [BRW05, FP16, GVY17, GH21, LBR18, Mun11, WR13, WYD22]. Self-Assembly [GVY17]. Self-Locking [WYD22]. Self-Motion [GH21]. Self-Organizing [LBR18]. Self-Propelled [Mun11]. Self-Similar [BRW05, FP16]. Self-Similarity [WR13]. Selkov [UW14]. Semelparous [DvG09]. Semianalytical [BHM24]. Semianonymous [RCG^{+}23]. Semiconductor [BC15, EKL07, GKS03, Sie02, TKB17]. Semidiscretization [IMS15]. Semiglobal [RAM15]. Semilinear [CLN24, NSS06, NSS13]. Semistrong [DK03, MW17c, Rad13]. Sensing [BTBK14, FRB19, KGB^{+}17]. Sensitivity [BKM22, DLRB19, GL24, GC20, Las18, MRMM14]. Sensor [BKM22]. Sensors [BKM22]. Sensory [BHP^{+}21, NCA^{+}21]. Separable [HKP22]. Separation [BCGFS13, GVY17, JJ20, PN20, SMS18]. Separatrices [BMSM24, DGG16]. Separatrix [BN19, GH05, LRR08]. Sequences [LPH22, Ton10]. Sequential [CTAA18, CATA20, PACM22, SA13]. Series [ABMS15, CGHM18, MIK19, PN20, WSYTA23]. Set [DKZ17, FH23, GZED20, Jam10, PID21]. Set-Oriented [DKZ17, GZED20, Jam10]. Sets [BK20a, CJ17, CKH21, EKO04, FJ18, FKS20, GOH20, HKO13, KK19, KO03, MFE05, NSS20, PYGR06, SM20, VA21, Wil23b].

Shadow [GCD^{+}21]. Shadowing [GL24, Tup09, dLDF^{+}18, dLD22]. Shadowing-Based [dLDF^{+}18, dLD22]. Shallow [CFR04, PRY23]. Shallow-Water [CFR04]. Shape [AH19, HACY22, KT24, MB14a, MB14b, Rot22, WGCT21]. Shaped [WW20]. Shapes [KN14]. Shaping [MPC^{+}22]. Shared [LSB11]. Shear [BEW11, LFFC^{+}20, MFE05, Rob04]. Shear-Induced [BEW11]. Sheets [MM06]. Shell [GBCV20, LL08]. Shells [GSDN15]. Shift [CC06b]. Shifted [KPG19, MNG07]. Shifting [CBWS24, HCW23]. Shifts [DF19]. Shil’nikov [CWZ17, CKK^{+}07, GL15]. Shimmy [HKLN13]. Shock [BHLMW24]. Shop [LR20]. Short [NPV12]. Short-Term [NPV12]. Show [NC21]. Shuffling [SUOL18]. Side [HKLN13]. Side-Stay [HKLN13]. Sided [ZLL22]. Sides [PP12]. Sigmoidal [BCP24, DEL14, DGK^{+}21, Dys20]. Signal [Ton19, YBO22]. Signaling [IW21, MDW23, Ren12]. Signals [GCD^{+}21, SW24, VLS13]. Signatures [CKCG19]. Signed [DJD19, TL24]. Silence [TW18a]. Similar [BRW05, FP16]. Similarity [WR13]. Simple [Aga18, GT18, Has21, KSG14, SG11, Ste14, Van06]. Simplicial [ABMS15, AG23]. Simplification [BTK16]. Simplified [BCH14, TZKS12]. Simply [GAHK03]. Simulating [KDKR13]. Simulation [Tup05]. Simulations [ACL23]. Simultaneous [SSS06]. SINDy [RL24, WM24]. Single [BH20, KT24, SRS14]. Singly [But20]. Singular [BHLMW24, CK15, CT17, CT19, CSS17, Chi09, CKPP19, CR11, FLWW22, GS09a, GHH^{+}24, Guc08, GM12, Guo10, HS05, IPS19, JPK^{+}22, JJ20, KPK23, KRW12, KS11, LT17, LW20, LdST09, MEvdD13, MKO18, RK23b, SZ13, VRS22, WZ16]. Singularities [AV19, BHLM21, GS16, KH15a, KP23, LdST09, MG18, QCARL21, RRW15, Wei03]. Singularity [CH10, CJ11, FGGT^{+}12, FDS14, JC09, Mak17, NVC18]. Singularly [GKKZ05, WW20]. Sinks [SSR10]. SIR [CBR19, DNDY16, GKR24]. SIS [GYAU24]. Site [BSOM20, FKdWY23]. Sitnikov [CCHZ20, CLZ21, GNR18, LO08, Riv13]. Sivashinsky [CDS10, DC16, FdlL17, GL17, MR06, Zgl02]. Size [LLYZ13, NPRW19, RCLR21, RPY20, SLS23]. Sizes [BSOM20]. Skeleton [WIN16]. Skew [BMWY18]. Skew-Symmetric [BMWY18]. Slanted [Daw08]. Slave [PP08]. Sleep [APBB22, BB12, BAB13, BXB17]. Sleep-Wake [APBB22, BXB17]. Sliding [AS18, GHS10, Jef14, SO09]. Slightly [AMBV22]. Slip [GO15, GSDN15, SDK20]. Slow [BB12, Bok22, DD13, DKO08, DTG^{+}16, DAFM19, EJK20, FKO18, FP16, GKKZ05, GKSV23, Guc08, GK09, GJM12, HL18, Jel24, KPR12b, KBS14, Kri15, KM17, LSB11, LW20, MP13, MKO18, PVMP17, RBI21, SWR05, TXKW17, Van08, VD13, WZ18, GHW03]. Slow-Fast [Bok22, DD13, Jel24, KBS14, LW20, WZ18, RBI21]. Slowly [AJB^{+}16, Kri21, OK24]. Smale [CKPP19, Has21, Wil10]. Small [BW09, CM03, CDW24, CZ16, DGG16, DE16, HHBS22, Hsu19, IW21, JS17, Kri21, TX21, dWRS18]. Small-Noise-Induced [CDW24]. Smoke [KR11]. Smoke-Ring [KR11]. Smooth [BdCT12, CJ11, GHH^{+}24, GHS12, Hu23, KK19, KRK14, KH15b, KH15a, LPH22, PR22, PCG16, PCG18, Rob16, SM08, ZR20]. Smoothness [WZ18]. Snake [ALB^{+}10]. Snakes [CKMW12]. Snaking [Daw08, DMCK15, KW08, Llo19, Llo21, TBR23, UW14, YC10]. Social [BT16, GHH^{+}21, KSKJ20, Kul16]. Societal [Ton23]. Societies [AY20]. Sofic [DF19]. Soft [KRK14]. Solenoidal [GS16]. Solid [AV19]. Solitary [DG05, KKC06, PSW12]. Soliton [WYM^{+}24]. Solitons [BD11, CM16, DP08, MCP09, PP20, SB10, WYM^{+}24, YC10]. Solution [AAK12, BCKN14, BCDG16, CHH23, vdBGW15]. Solutions [AHGKM16, AVV23, AK17, BCBD20, BC21, BCRR21, BT04, BBK17, BRW05, CLN24, CW17, CLZ21, CCD^{+}10, CV14, CZ15, CZ16, DvHX16, EK10a, ELT22, FMT16, FE12, FP16, GNR18, GBCV20, Guo10, HvHM^{+}14, HLJ23, HdRS23, HP17, Hu23, HS10b, Ike23, JZ11, KR11, LT17, Lee22, LBHM05, Mat18, MPY11, MW17c, PVVY17, PCG18, PYVG14, Riv13, RHT13, SWR05, TvH21, TDK18, VF10, VSABM23, VNSG08, YPMD08, ZME19, vdBL08]. Solving [MK24]. Some [AK17, BP16, FG19, MO20, PEH22, PW07, PRK18, RS16, UW14, VF10]. Somersault [DT16]. Source [Des23, DS19, PN20]. Source-Only [Des23]. Sources [IS17, SSR10]. Southern [KKP15, KKP16]. Space [BAA^{+}19, CGP16, CGS15, CGH^{+}16, CDS10, DvHX16, Gor13, HL18, HS09, IS17, VQPVV23, XWC^{+}23]. Spaces [BC14, GK18, SSR18, VRS22]. Sparse [BHP^{+}21, BGH^{+}24, FJ18, KGB^{+}17, MH17]. Sparsity [SK22]. Spatial [BKJ15, CDPVY21, CB18, GMS11, PLST20, PA19, PK17, TW18a, XSW24, Zha07]. Spatially [ABCV23, BCBD20, BYK08, DP09, FE10, GVY17, GKML09, PRCA^{+}23, SD17, WYM^{+}24, Yak08]. Spatio [FS16, XSW24]. Spatio-Temporal [FS16, XSW24]. Spatiotemporal [BC14, GLW10, MJJL12, SRS09, SJLY17, TFTK22, ZB20]. SPDEs [BN23, YB22]. Special [FG19, Mor15]. Species [CW17, EFK17, FLWW22, HL22, KLK10, PID21, WF13, vdDZ04]. Spectra [GMCM19, GMCM21, MRS14, Rad06, WB06]. Spectral [AM17, AM23, BJ22, Buz24, CLN24, COT19, CCD^{+}10, HSS13, LMKY23, MH17, MW16, dlLK19].

Spectrally [KPG19]. Spectrum

[CBR19, CFG21, KRW12, MCZM18]. Speed [BTK12, MXYZ16]. Speeds [AH06, HS09]. Sphere [Com06, HK19, HKLN20, HP20, LLYZ13, RRW14, SvGKP24]. Spheres [SM11, UN21]. Spherical [CBR05, CLOS14, DL21, GSDN15, GBCV20, HSV24, LL08, MG16]. Spike [DK18, GIKR20, GWW19, LCDS12, PE18, SWR05, Ton10]. Spike-Adding [DK18]. Spiking [CC06a, EWLH11, FEIvdD12, GP20, GKC14, GPTV17, LA13, MK12, RRW15, TB09]. Spin [MO20]. Spinning [DKaK^{+}08]. Spintronic [RS21]. Spiral [CL14, CL09, DS19, DL22, HS22, HG10, KL17, Lai05, OWK18, SS07, SM11, WB06]. Spirals [DL21]. Splay [DKTG12, ZZ09]. Splitting [BMSM24, DGG16, LRR08, MB14a, MB14b]. Spock [CFdSL22]. Spontaneous [ABI24, FY13, KSG14, New14]. Spot [CW11, Ike23, NX22, RRW14, SW21, TXKW17, TWW18, TT20]. Spots [MS13, WW20]. Spread [HS09, KRW13]. Spreading [TL24, WWC^{+}18]. Spring [HL02]. Spurious [TY07]. Square [AAM05, HHBS22, SAV12]. Squares [FGHC16]. Stability [Aga18, ABCV23, AK06, BHL16, BJ22, BGT10, BFG21, BBK23, BC09, BD11, BRRS02, CDPVY21, CDKS19, CFdSL22, CCHZ20, CLZ21, CCD^{+}10, CW11, CJ18b, DSC12, DDvGS07, DKB^{+}23, DL22, DB13, DGK^{+}21, EK10a, EJK20, Fay13, FGGT^{+}12, FH23, GNR18, GAS18, GS22, GL13, GSB^{+}16, GWW19, GT18, GKR24, GC05a, GC05b, Guo12, HSS13, HSS22, HHHY09, HH19, HH20, HMP02, HK15, Hsu19, Ike23, JZ11, KSWW06, KTK20, Kra21, LMNT09, LT15, LZX22, LGLC15, LZH^{+}17, LT21, LA18, MW17a, MR18, MPY11, MRS14, MNG07, MW17c, MR24, NP15, OWK18, OZM11, PD20, Pat03, PJW05, PCG18, PYVG14, RGAB16, Rob13, RRW14, SS09, STW23, SRS09, SRMPM08, TLRB11, TDK18, TXKW17, TWW18, USW05, WZ18, Wil19, YW10, YLWK16, Yos17, YCMP22, YB11, ZLX20, ZHKR15, ZSL23, dWRS18]. Stabilizability [HMN09]. Stabilization [Pos09]. Stabilized [GAHK03]. Stabilizing [FH14]. Stable [CL13, CJ08, EKO04, FKO18, GJ17, Hül16, JM13, KPG19, KKJ18, LT13, Lee22, Scl23, SW24, WB17, KO03, YB22]. Stable/Unstable [GJ17]. Stacked [LZ23]. Staged [LD18]. Stages [CF20, CT22]. Stall [Xia08]. Standard [BM12, LZX22]. Standing [BJM20, BRRS02, GC05a, GC05b, Ike23, NP15, WSWK12, YHM^{+}02]. Standing-Wave [YHM^{+}02]. Star [KPG19]. State [ACFK09, BJL^{+}17, BT10, CHK17, CGS15, CDS10, ETS21, FRB19, FLWW22, GYdlL21, GLJY23, GSJ23, GK18, HQW^{+}20, KHG21, Kim20, MKM23, VQPVV23]. State-Dependent [CHK17, FRB19, GYdlL21, Kim20, GLJY23]. State-Space [VQPVV23]. State-Spaces [GK18]. States [AOWT07, BLL12, BBK23, BFH14, BD12, Daw09, DL10, DDN22, DKTG12, EK16, FY13, Fer18, GB09, HKP22, JS06, KPG19, KLK10, KE08, KPT13, LA13, MHC09, MRS17, MW17a, OWK18, Oro14, PSSJ23, ZZ09]. Static [GRSB19, JC23, WGCT21]. Stationary [Bok22, BFH14, CJ18a, CCD^{+}10, DvHX16, FE12, GD23, IW23, KPT13, LA13, PEH22, VF10, vdBDLJ15]. Statistics [FMT16, Law16]. Stay [HKLN13]. Steady [BJL^{+}17, BT10, ETS21, FY13, FLWW22, GSJ23, HRS04, WIN16]. Steady-State [BJL^{+}17, BT10]. Steep [DGK^{+}21]. Steering [DR22, KM08]. Stellar [BC21]. Stellate [RWK08]. Stem [DH19, PMBM05]. Stem-Cell [DH19]. Step [CC06b]. Stepwise [HL02]. Sterile [HRYZ19]. Stick [GO15]. Stick-Slip [GO15]. Stiction [BBK17, Kri21]. Stimulated [LFOG17].

Stimuli [BHP^{+}21]. Stirling [GC20].

Stirring [MM17]. Stochastic

[Ada23, AHS14, BGOŻ08, BKPP24, BU23, BW12, BN13, BK15, BM18, BM20b, BK20b, CEK22, DEP^{+}11, DNDY16, DTG^{+}16, FGHC16, FS09, GALS16, Gle14, GSB^{+}16, GK18, HH20, HN14, IM16, KDKR13, KE13, KF14, KS14, Lan16, MV21, OK24, PR24, PP08, SMS18, SHK13, SA13, VRS22, WR13, WL22, WM22, WM24, XWC^{+}23, YM23, ACL23]. Stochastically [ACK17, HBB13a, HBB13b, LK15, PB17, ZLL22]. Stoichiometric [TZ22, TW23, YLWK16]. Stoker [XSS20]. Stokes [BBJ21, Buz24, FMT16, WZM23]. Stop [ABG^{+}17, LR20]. Straight [PYGR06]. Strange [Rod21]. Strategies [CBK19, CP12, CJAMV20, LPH22]. Streakline [Bal17]. Stream [HJL16, HJL17, MJJL12]. Streaming [GHTW23]. Stretch [SS14]. Stretch-Dependent [SS14]. Stripe [DvdP02, KSWW06]. Striped [MS15]. Stripes [AGG^{+}19, MvB18]. Stroboscopic [GKC14]. Strong [De 03, FP16]. Strongly [PW21, QT20, RBI21, Wil23b]. Structural [FGGT^{+}12, LZX22, TL24, Vas23]. Structure [CBR19, CW22b, CSS17, CT22, DS24, EMNT15, GIHLS20, HS03, KPR15, MD18, PID21, SMRB11, WIN16]. Structured [LRH12, SAA^{+}18]. Structures [AJB^{+}16, GM15, HN14, KS07, PRCA^{+}23, SK22]. Strut [HMP02]. Study [AK10, BJL^{+}17, CM03, CLOS14, Chu21, CZSV23, Coo08, DHMO05, DG05, GS07, GKO18, GKS03, KLK10, KH15a, LdST09, NWKR15, NSTZ20, SAV12, VSABM23, YNZ22, ZZ20]. Studying [XCC07]. Subcenter [dlLK19]. Subcritical [PPK14, VNSG08]. Subgradient [KHG21, LKH^{+}22]. Subgrid [HGT15]. Subgrid-Scale [HGT15]. Subharmonic [GHS12, ZL14]. Subject [CFR04, GKS03, MV21, MW10, PPK14]. Subjected [CBR05]. Sublevel [GOH20].

Submanifolds [Buz24, dlLK19]. Subpopulations [DZ14]. Subsonic [HSS13]. Subspaces [KR21, Mor15, NSS20, NSS24, TZ22]. Substrate [FSDAS21]. Substratum [AAK12]. Subvolumes [MSB09]. Sufficient [BCDG16, BM24, MR24, SBR06]. Sum [FGHC16]. Sum-of-Squares [FGHC16]. Sun [Cap12]. Superconducting [SAV12]. Supercritical [TBR23]. Superlattice [CR09]. Superlattices [PK05b]. Superlong [MV14]. Superposition [IR22]. Superslow [DK18]. Suppression [SBB10]. Suprachiasmatic [GB23]. Sure [OBK18]. Surface [BD11, DP08, Gle14, PLNW19, SW21, SD22]. Surfaces [BB24, BC23, GH09, HK18, LHRK04]. Surfactant [EHLW15]. Surge [CF07, CV09, EVC18, Xia08]. Surrogate [OK24, RL24]. Surrounded [AV19]. Surrounding [WB17]. Sustainability [Ton23]. Swamps [Moh19]. Swarm [BT11, FK17, HK19, HKLN20]. Swarmalators [Smi24]. Swarms [DE16, LTB09]. Swift [ALB^{+}10, BD12, DHMO05, GBK15, LSAC08, LS17, Llo19, Llo21, MS13, PW07, vdBL08]. Swimmer [Mun11]. Swimming [OZM11]. Swinging [HL02]. Switched [Aga18, Mak17]. Switching [BGB05, BHLM21, BJK20, CL16, CF12, CGH^{+}16, Don24, Gle14, Jef14, LK15, Law16, PPM14, RC23, SBR06]. Symbolic [CII23, DFT08, DMS05, MSM17, TF21]. Symmetric [AAM05, BL08, BMWY18, CH10, CB18, CCHZ20, CLZ21, CBR05, CP06, EJK20, GWW19, GBCV20, HGS15, KPT13, RS21, SM11, WIN16, Wil05, WS06, vdBGW15]. Symmetries [AR12, ADP08]. Symmetry [AAM05, BHL16, BE14, BC15, BCF^{+}18, CL14, DHW22, EJ16, KLW13, LM16, MHB07, NSS06, NSS13, PB22, RSTY12, SAS11, SS12, SGP03]. Symmetry-Breaking [BC15, BCF^{+}18, CL14, MHB07]. Symplectic [FKM23, MSB09, Vil18]. Synaptic [Fay13, GPTV17, KB10, KN14, PTK09, Vel13, Zha07]. Synaptically [FE10, Fol11]. Sync [LZH^{+}17]. Synchronism [TR19]. Synchronizability [NC21]. Synchronization [ABI24, AG23, BCJ19, CGL19, CHH24, DT13, EW09, EDKC16, EGF18, FRB19, GFE20, HNP16, HQW^{+}20, JKKL24, LW02, Lyu18, MV21, MMP16, PFGV14, PP08, PB17, QT20, RBBG20, ST13, STW23, SÜvLM16, SBR06, SBN09, VH08, WM14a, XWC^{+}23, YCL08, ZLL22]. Synchronized [GLNW15, Lee22, RT02]. Synchronous [FE10, JL08, NWW21, SW24, Yak08]. Synchrony [ADR16, AD21, CJ08, CATA20, DZ14, DP09, GST05, GL09, IW21, KR21, Mor15, NSS19, NSS20, RT02, SAH21, Soa17, SGP03, SG11, Ste14, Tro08, aAA10]. Synchrony-Breaking [SG11, Ste14]. Synergetic [SP03]. Syntrophic [FSDAS21]. System [AEL08, AGMS23, AdBG^{+}09, AM23, BR13, BCDG16, BEG^{+}03, CFST08, CS18, CCD^{+}10, CL11, CP06, CJ18b, CKO17, CP17, DM20, DC16, DAFM19, FGDKC15, GW24, GHS12, GK10, GS11, HSS13, HSS22, HMD^{+}23, HS10b, Ike23, KR11, KVDC12, LP21, LBR18, MRS17, MFN^{+}23, NWKR15, NX22, PLNW19, PSW12, PVMP17, RK23a, RK18, SW21, SP03, SPCT12, TC21, TAtN09, VVZ15, VSABM23, Wil05, Wil10, WSB16, Wil23a, WYD22, YNZ22, YW10, YvLKL22, ZSL23, vHDKP10]. Systematic [ESK24, WB17]. Systems [AH09, Aga18, ANR14, ANR18, AJB^{+}16, ACK17, AEG24, ADR23, AS18, AKK^{+}09, ABG^{+}17, AG05, BJL^{+}17, BM16, Bat17, BGZ16, BD23, BFH22, BH23, BC09, BGOŻ08, BLDK18, BH20, Bri20, BTBK14, BYK08, BC23, BIPK24, CLL12, CBK19, CKK^{+}07, CDW24, CW22b, CII23, CD10, CGK08, CF20, DJM04, DD13, DKZ17, DDCK11, DEP^{+}11, DMS22, DTG^{+}16, ELB15, EG06, EMNT15, FvdSG20, FGHC16, FMHM24, FA13, FHKK21, FKM23, FPT12, FGH14, GAG^{+}21, GALS16, GKKZ05, Ged10, GCD^{+}21, GS22, GHTW23, GCY22, GL24, GHH^{+}24, GSB^{+}16, GSJ23, Guc08, GH15b, GS20, HN14, HCW23, HBB13a, HBB13b, HDL^{+}08, Hsu19, HQW^{+}20, IS24, JS06, Jel24, JR05, KK19, KM10, Kie20, KBGD^{+}23, KRK14, KW08, KPW17, KBS14, KH15b, KH15a, KPK08, KM17, KGK21, LPR22, Las18, LM19]. Systems [Leg11, Leg13, LS15, LS16, LMKY23, LWK23, LDB20, LO10, LLYZ13, LW20, LdST09, LKH^{+}22, LPK15, MCZM18, Mak17, MPY11, MRS14, MD18, MNG07, MM11, MR24, NS13, NS15, NSS19, NC16, NRS20, NUY05, NSTZ20, OBK18, OE21, OdBS08, PZ22, PID21, PTK09, PRK18, PRK22, PCG18, PB10, Rad13, RBK15, RSTY12, Rob16, SV09, SW16, SAS11, SMS18, SRS09, SRMPM08, SK08, SØRD24, SRS14, SÜvLM16, Sul23, SWR05, SM20, SMM21, TC23, TD08, Tup05, TWW18, UW14, VC12, VSC23, VQPVV23, WR13, WZ18, WL22, WB14, Wil23b, WW20, WR02, XCC07, XSS20, YY19, YPMD08, YCMP22, ZR20, ZRDC19, ZLL22, ZDG19, vdBKV11].

T [CFST08]. T-Point [CFST08]. Tails [NX22]. Takens [AHGKM16, BK24a, BK24b]. Tangencies [AM06, CKO17, GKM05, MKO18]. Tangency [LCKO08, WZ09]. Tangent [CKK^{+}09, PRK22]. Tangential [Bal11]. Target [LS15, LS16]. TASEP [GKM21]. Task [SDT17]. TC [TD08]. TC-HAT [TD08]. TCP [HHHY09]. Team [AY23a]. Techniques [DD13, GL24, WV19, ZDG19]. Temperature [BLL12, GMB16, Tak16]. Temporal [AG23, CP23, FS16, Kul16, Zha07, XSW24]. Temporally [WYM^{+}24]. Tend [RPY20]. Tendency [HS15]. Tensor [GMY18, HP20, HKP22, Moh19]. Tensors [HARB21]. Term [HK15, MG18, NPV12, Sul23]. Terms [GS16]. Terrain [TT20]. Terrestrial [LR22]. Territorial [VBG^{+}09]. Test [FGMW07, GM09]. Tethered [KPR11]. Tetrahedral [ESZ04, LM16]. Their [AEG24, Bal11, Bal17, BCRR21, BD11, FMT16, FGH14, GKO17, HdlL07, HdlL13, KC13, KR21, LKO15, NC21, OP08, PZ22, SW14, DGK^{+}21]. Theorem [FG10, KGK21, VRS22]. Theoretic [GRSB19, YCMP22]. Theoretical [AJB^{+}16]. Theory [AH09, AHARS18, AY23b, AM17, Bal11, BD23, BH23, BN23, CT17, CT19, CL13, CHP17, DF19, FDS14, FT12, GAG^{+}21, GL09, GC20, HDL^{+}08, JPK^{+}22, KKBK20, LT17, NPRW19, PbG09, PBB22, PBK18, ACJ24, SL19, VRS22, Mat11]. Thermal [FGHM^{+}20, UN21]. Thermodynamic [Gor13]. Thermostated [But20]. Thermostats [LS05]. Thin [CV14, EHLW15, KRW12, TGPP19]. Thin-Film [CV14]. Three [BCPS08, BEW11, CR12, CHH24, CLOS14, CH03, CPY19, DV19, DK18, DR10, FS16, HM22, HM24, HGS15, Ike23, KKC06, KPK08, MSM17, MR18, MM12, MJM05, NWKR15, NX22, PRCA^{+}23, PBB22, RS07, SSS06, SS12, TvH21, vHDKP10]. Three-Body [BCPS08, CR12, CPY19, DV19, HGS15, RS07]. Three-Cell [DR10]. Three-Component [Ike23, NX22, TvH21, vHDKP10]. Three-Dimensional [CLOS14, HM22, HM24, MSM17, MM12, MJM05, PBB22]. Three-Group [CHH24]. Three-Scale [FS16]. Three-Time-Scale [DK18, NWKR15]. Threshold [ADF20, ADF21, BCRR21, LM19, LPH22, MDIC24, PACM22]. Threshold-Based [LPH22]. Threshold-Linear [BCRR21, MDIC24, PACM22]. Thresholds [DS24, TWW18]. Thrust [BE03]. Tick [ACL23]. Tiling [Hen11]. Time [Aga18, ABMS15, ACK17, ACMM20, ABG^{+}17, Bal05, BKPS19, BT10, BCGFS13, BS19, Bok22, Brö17, CT17, CT19, COT19, CH13, CGHM18, DDDGZ16, DK18, DHK20, FGDKC15, FJ18, FKS20, GBK15, GL24, GMB16, GS13, HKL14, HKM21, HARB21, HBB13a, HS09, HQW^{+}20, HMN09, IMS15, IS17, JKKL24, KKBK20, KM10, Kie20, Kri20, KPK08, KVDC12, LKO15, LGLC15, LA18, MB14a, MB14b, MM17, MIK19, NWKR15, Pos09, PN20, PPK14, RBK15, SMS18, SBR06, WZ17, WZM23, WSYTA23, WSWK12, YW10, YBO22, YCL08, ZRDC19, ZB20, dWRS18, BN23]. Time-Continuous [BKPS19]. Time-Delay [KKBK20]. Time-Delayed [HMN09, IMS15, Pos09, PPK14, WZ17]. Time-Dependent [BS19, Bok22, JKKL24, WZM23]. Time-Periodic [GBK15, WSWK12]. Time-Periodicity [Bal05]. Time-Renormalization [ACMM20]. Time-Reversible [Kri20]. Time-Scale [BCGFS13, CH13, KPK08]. Time-Series [MIK19]. Time-Varying [GMB16, HARB21, LA18, SBR06, YCL08, ZRDC19]. Times [KRW13, LFOG17]. Timescale [EVC18, KPK23, LS05, PR22]. Timescales [GH04b, HH20]. Timing [AMNB06, WGCT21, Wil21]. Tinkerbell [GHC11]. Tippe [BRMR04, USW05]. Tipping [HCW23, Kie20, LNOR21, OW20, TFBN21, ZKE15]. Tissue [BJSW08]. Tissues [MDW23]. Toda [LP18]. Together [ACL23]. Tokamak [ZHKR15]. Tomography [CBR19]. Tongues [SO09]. Toolbox [TD08]. Top [BRMR04, USW05]. Topological [CZSV23, DDN22, Fer18, FT07, Jam10, MIK19, MSW15, SD22, SL19].

Topologies [GRSB19, SBR06]. Topology [JL08]. Tori [AVV23, BC23, DGG16, DJCJC22, FMHM24, FH12, FM16, GJNO22, HdlL07, JO09, MPY11, PVVY17, SOV05, Vil18, WB17]. Toric [CD20, HSSY23]. Toroidal [KTK20]. Torsion [Rob04]. Torus [EMNT15, FT07, HdlL23, OWK18, RRW15, SW21]. Tracking [BP08, HMD^{+}23]. Trade [PPM14]. Trade-Off [PPM14]. Traffic [BKPS19, BFK^{+}22, DS23, MSB^{+}14]. Trains [Rad06]. Trajectories [BE03, FJ18, Hen05, KM08, LA18, OPR24, TD08, ZB20]. Trajectory [FKS20, SRS14, Wil23b, ZKCS19]. Transfer [ABBC20, BKJ15]. Transformations [GMCM19]. Transient [BS19, GMM08, RSRT21, WE18]. Transients [DRC09, HLvdD22, LHW23, MV14]. Transition [BGOŻ08, CDW24, Com06, GLS21, Wul08]. Transitions [ASH18, CL08, DNO23b, DNO23a, GO15, HHW21, Jel24, MV14, RRW15, WHT13]. Transitory [MM11, MM12]. Transmission [ABM^{+}04, AH06, GKR24, WWZ19]. Transonic [BC21, De 07]. Transport [Bal17, BB16, CLOS14, CSJM18, FY13, HKK20, HK18, MM11, MM12, YNN^{+}23]. Transversal [Cap12]. Transversality [DGG16]. Transverse [CG24, HLJ23, TGPP19]. Trap [SSR18]. Trapezoidal [HSV24]. Travel [KRW13]. Travel-Related [KRW13]. Traveling [BvdBV18, BHV11, CJ18b, Dys20, EHLW15, Fay13, Guo12, HSS13, HH19, HH20, HvHM^{+}14, HS10b, IM16, JZ11, KFB08, Lan16, LMNT09, MCP09, OY09, PJW05, She14, SS14, TvH21, TDK18, TS07, Tro08, TZKS12, VH08, Zha07]. Traveling-Wave [JZ11]. Travelling [HMS19, MRS14]. Tree [Gor13]. Trees [DT13, HDDL21, HJSS23, Lyu18]. Tri [GYAU24]. Tri-virus [GYAU24]. Triangular [CW22b]. Triatomic [LRK12]. Triaxial [PVVY17]. Triggered [CRSN07]. Tristability [LHW23]. Tropical [TS23]. Truncated [QCARL21]. Truncation [IS24]. Truth [WHT13]. Truth-Content [WHT13]. Tube [CM03]. Tumbler [CLOS14]. Tumor [BR13, HvHM^{+}14]. Tuning [CV09]. Turbulence [HGT15]. Turing [BB16, GS07, PRCA^{+}23, SJLY17, VSC23]. Turing-Unstable [GS07]. Turning [AST07, SS09]. Twin [HKO17]. Twist [DM12, Moe15, OP08]. Twisted [BBK23, HMP02, MW17a]. Twisting [DT16]. Two [AAM05, Agu15, BHLM21, Bal17, BJK20, BE14, CHK17, CW11, CHH24, CLBdB09, CJ11, CFR04, CKCG19, CEvdM24, DDCK11, DGMW12, EFK17, FGGT^{+}12, FMT16, FEIvdD12, GKS03, Guc08, HL22, IR22, JC09, JZ11, KLK10, KPR15, KH15b, Lai17, Leg11, Leg13, LCDS12, MP09, MM17, NSUW09, NC16, PLNW19, PR22, PP12, PBB22, RCG12, SRMPM08, SvGKP24, SWR05, TWW18, Ver08, WWC^{+}18, Wri10, ZHKR15, vdDZ04]. Two-Dimensional [Agu15, CW11, FMT16, FEIvdD12, Lai17, Leg11, Leg13, MM17, NSUW09, NC16, PLNW19, PBB22, RCG12, SRMPM08]. Two-Fluid [ZHKR15]. Two-Fold [FGGT^{+}12, JC09, KH15b, CJ11]. Two-Group [CHH24]. Two-Layer [CFR04]. Two-Layered [WWC^{+}18]. Two-Parameter [CLBdB09, GKS03]. Two-Patch [KPR15]. Two-Species [HL22, KLK10]. Two-Spike [SWR05]. Two-Timescale [PR22]. Twofold [NSTZ20]. Type [AEHV05, AVV23, CH10, GK22, GHH^{+}24, GK09, IS17, KRW13, Kri15, Mat18, MvB18, RCG12, SRS09, Ste23, Wil10, WZ16, YNT14, RSRT21].

U.S. [AJB^{+}16]. Ulam [BFGTM14, Yos17]. Unbinding [MV14]. Unbounded [CLN24, Kri20, OBK18]. Uncertain [KM10, KPW17]. Uncertainty [BU23, DKZ17, KL21, WL22]. Uncovering [Kri21]. Underlie [KVB22]. Underlying [LCDS12]. Understand [BW09]. Understanding [ABBC20, AJB^{+}16, DRH19, GYAU24, NWKR15, NCA^{+}21, RWZ21]. Underwater [Pat03]. Unfolded [ABPM22]. Unfolding [CKK^{+}09, GS16, KOP07, NVC18]. Unfoldings [GM12]. Unidirectionally [MV14]. Unification [Chi09, KPR11]. Unified [VCNSD24]. Uniform [CW17, HKM21, SCD07]. Uniform-in-Time [HKM21]. Uniformly [Wil10]. Unique [VCNSD24]. Uniqueness [Brö17, Dys20]. Unitary [Kim20]. Universal [GS16]. Universality [OP08]. Unpeeling [CS18]. Unstable [BJM20, CL13, CW22a, GS07, GJ17, Las18, PSSJ23, ZDG19]. Unsteady [AJ14]. Unstructured [KBGD^{+}23]. Unsupervised [Jud20]. Unveil [AJB^{+}16]. Urban [GW24, SAA^{+}18, SBB10, TW18b]. Use [KH15a, SDW15]. User [MRS14]. Using [AY23b, AST07, BW09, BBJ21, BK20a, Chu21, CZSV23, FG20, FJ18, GCD^{+}21, GCY22, Gie19, KPW17, Kri15, KH15b, Las18, LV14, MSM17, MH17, McC15, MK24, PSV22, Pos09, PN20, TV14, VSABM23, Wil21, WS24, YW10, FGHC16, Kri21, POR20]. Utkin [DEL14].

V [AEL08]. Vaccination

[ABM^{+}04, GHH^{+}21, KPR15]. Vacillating [Smi24]. Validated [BJ22, CJ17, SW16]. Validation [Bre23, SW14]. Validity [ETS21]. Value [AEHV05, Law16, LD18, SDW15]. Values [BFH22, BH23, GRS^{+}22]. Valve [BCH14]. Valves [EPCL05]. Variable [AK18, HKLN13]. Variables [FGH14, Guc08]. Variant [TKKCG16]. Variants [CL11]. Variation [APBB22, FH22, KRW13, MW10]. Variational [BM18, Brö17, DHP23, FMOW03, LP22]. Variations [HMP02, ML12]. Varying [AMNB06, BCBD20, CLW24, GMB16, HARB21, Kri21, LA18, SBR06, YCL08, ZRDC19]. Vector [AM23, BSKR16, BdCT12, Chi08, JC09, Jud20, KO03, NSS19]. Vectors [GMM08]. Vegetation [GLS23, TT20]. Vehicles [Pat03]. Vehicular [KM08]. Velocity [BN13, PK17]. Verge [FH12]. Verging [SPCT12]. Verification [FGLdlL17, GL17, GH15a, Ipp11, Mat11, SA13, vdBLQ21]. Version [CFST08, TAtN09]. Versus [BHLZ18, KKP15, WGCT21]. Vertically [DL10]. via [AEL08, ABBC20, AMBV22, BS18, BKS06, BJL^{+}17, BJ22, BD11, CLN24, CLJ15, CJAMV20, CGHM18, DF19, DsMS24, HMN09, LT17, LR19, LdST09, MPC^{+}22, Mat11, MW16, Ste23, TS23, YNN^{+}23, ZB20, ZL20, vdBL08, vdBKV11, vdBDLJ15, vdBLQ21]. Vibrated [DL10, WCM08]. Vibration [ESZ04]. Vibrations [GAL20]. Vibro [GG18]. Vibro-Impact [GG18]. Viewpoints [VBW13]. Viral [SLF24]. Virus [WWZ19, ZW24, GYAU24]. Viscosity [BC21, BW09]. Viscous [CM03, CZ15, JZ11, MW16]. Visual [CB16, GST03, TPC24]. Vivo [ZW24]. Vlasov [CH10]. Voltage [PR22]. Voltage-Dependent [PR22]. Volterra [BR23, CJN15, Sla20, TC21]. Volume [BF18, DM09, DM12, Fer20, FM16, Jam10, JL10, LRR08, MSM17, OY09]. Volume-Filling [OY09]. Volume-Preserving [DM09, DM12, FM16, LRR08, MSM17]. Vortex [BHL16, GKCG15, KTK20, LT21, MR18, MST03, NSUW09, Rob13, TKKCG16]. Vortices [DL21]. Vorticities [MR18]. Vulnerability [FH23].

Wake [APBB22, BXB17, GKSV23]. Walking [GO15, RK23a]. Wall [OZM11, RS21]. Wandering [BK15, CC06a, CEK22, KE13]. Water [CFR04, PRY23, WSWK12]. Wave [BKM22, CL09, DL22, GAL20, Guo10, HvHM^{+}14, HS03, JZ11, KE08, KFB08, MHB07, NP15, PYVG14, Rad06, ACJ24, SCD07, TDK18, Van06, VSABM23, WYM^{+}24, YHM^{+}02, ZYO05, ZME19]. Wave-Pinned [ACJ24]. Wavefronts [BHV11, OY09]. Wavenumber [ANR18, GMS11]. Waves [AAK12, BvdBV18, BJM20, BJSW08, CL14, CCD^{+}10, Com06, CJ18b, DG05, DE22, DS19, DL22, EHLW15, HSS22, HH19, HH20, HS22, HG10, HMS19, IM16, KKC06, KL17, Lai05, Lan16, LP22, LMNT09, MSB^{+}14, MRS14, PSW12, PTK09, SS07, She14, SS14, TS07, Tro08, TZKS12, UE15, VH08, WSWK12, Zha07]. Wavetrain [HP17]. Wavetrains [SRS09]. Weak [BM16, Bat17, BGT10, IW23, KL17, RL24, XCC07]. Weak-SINDy [RL24]. Weakly [BH20, DLP21, Des23, EW09, GST03, TKKCG16, WW20, ZL14]. Web [LHRK04]. Weighted [AD21, BJ24, MKM23, SAH21, SHA23]. Weingarten [BB24]. Well [AHW21, MRS14]. Well-Posedness [AHW21, MRS14]. Wells [KKC06]. West [WWZ19]. Wheel [HKLN13]. Which [BGZ16, JS17]. Whiskered [DGG16]. Whiskers [HdlL07]. White [Lu16, ZYO05]. Who [WFM^{+}14]. Wild [HKO13, Pat03]. Williams [Wil10]. Wilson [HE15]. Wind [BC21]. Winding [CP17]. Winfree [HKM21]. Within [aAA10]. Within-Burst [aAA10]. Without [FKS20, VC17, AY20, AGG^{+}19, EKO04, VC19, Wil22]. Witness [ABMS15]. Woodpile [DLP21]. Working [PK17]. Works [SGW09]. Wright [BCKN14, SP03]. Wrinkles [AGG^{+}19].

Yamada [HS05, TKB17]. Yield [BSKR16].

Yoccoz [NPV12]. Yorke [BFGTM14].

Zakharov [EMNT15]. Zebrafish [CZSV23]. Zeeman [DKB^{+}23]. Zero [BCKN14, BLL12, CF20, GS16, QCARL21, SSR10, Tak16, TW23, TD12]. Zero-One [TW23]. Zhabotinskii [GS13]. Zip [SSS06]. Zone [PK05a]. Zwanzig [LTLA21, LTPL23].

References

alAzad:2010:WBS

[aAA10] Abul Kalam al Azad and Peter Ashwin. Within-burst synchrony changes for coupled elliptic bursters. SIAM Journal on Applied Dynamical Systems, 9(1):261–281, ???? 2010. CODEN SJA-DAY. ISSN 1536-0040.

Antunes:2023:MCM

[AAC23] Felipe J. P. Antunes, M. Soledad Aronna, and Cláudia T. Codeço. Modeling and control of malaria dynamics in fish farming regions. SIAM Journal on Applied Dynamical Systems, 22(3):2123–2149, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/20M1376698.

Ambrosi:2012:HSE

[AAK12] D. Ambrosi, G. Arioli, and H. Koch. A homoclinic solution for excitation waves on a contractile substratum. SIAM Journal on Applied Dynamical Systems, 11(4):1533– 1542, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Abreu:2005:SCL

[AAM05] Stella Abreu, Philip Aston, and Ian Melbourne. Symmetric chaos in a local codimension two bifurcation with the symmetry group of a square. SIAM Journal on Applied Dynamical Systems, 4(1):32–52, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60331.

Alla:2020:UMT

[ABBC20] Alessandro Alla, Caterina Balzotti, Maya Briani, and Emiliano Cristiani. Understanding mass transfer directions via data-driven models with application to mobile phone data. SIAM Journal on Applied Dynamical Systems, 19(2):1372–1391, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Albeaik:2022:LII

[ABC^{+}22] Saleh Albeaik, Alexandre Bayen, Maria Teresa Chiri, Xiaoqian Gong, Amaury Hayat, Nicolas Kardous, Alexander Keimer, Sean T. McQuade, Benedetto Piccoli, and Yiling You. Limitations and improvements of the intelligent driver model (IDM). SIAM Journal on Applied Dynamical Systems, 21(3):1862–1892, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1406477.

Ascencio:2023:SEP

[ABCV23] Mauricio Ascencio, Esther Barrabés, Josep M. Cors, and Claudio Vidal. Stability of equilibrium points in the spatially restricted N +1body problem with Manev potential. SIAM Journal on Applied Dynamical Systems, 22(4):2732–2760, October 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1551912.

Arnold:2017:DDT

[ABG^{+}17] Maxim Arnold, Nikita Begun, Pavel Gurevich, Eyram Kwame, Harbir Lamba, and Dmitrii Rachinskii. Dynamics of discrete time systems with a hysteresis stop operator. SIAM Journal on Applied Dynamical Systems, 16 (1):91–119, ???? 2017. CO-DEN SJADAY. ISSN 1536-0040.

Abdalla:2024:GSS

[ABI24] Pedro Abdalla, Afonso S.

Bandeira, and Clara Invernizzi. Guarantees for spontaneous synchronization on random geometric graphs. SIAM Journal on Applied Dynamical Systems, 23(1):779– 790, March 2024. ISSN 1536-0040.

Alexander:2004:VMT

[ABM^{+}04] M. E. Alexander, C. Bowman, S. M. Moghadas, R. Summers, A. B. Gumel, and B. M. Sahai. A vaccination model for transmission dynamics of influenza. SIAM Journal on Applied Dynamical Systems, 3(4):503– 524, 2004. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60037.

Alexander:2015:SMM

[ABMS15] Zachary Alexander, Elizabeth Bradley, James D. Meiss, and Nicole F. Sanderson. Simplicial multivalued maps and the witness complex for dynamical analysis of time series. SIAM Journal on Applied Dynamical Systems, 14(3):1278– 1307, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Adimy:2022:NSM

[ABPM22] Mostafa Adimy, Louis Babin, and Laurent Pujo-Menjouet. Neuron scale modeling of prion production with the unfolded protein response. SIAM Journal on Applied Dynamical Systems, 21(4): 2487–2517, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1443157.

Abarbanel:2009:DSP

[ACFK09] Henry D. I. Abarbanel, Daniel R. Creveling, Reza Farsian, and Mark Kostuk. Dynamical state and parameter estimation. SIAM Journal on Applied Dynamical Systems, 8(4):1341–1381, ????

2009. CODEN SJADAY. ISSN 1536-0040.

Saadi:2024:WPP

[ACJ24] Fahad Al Saadi, Alan Champneys, and Mike R. Jeffrey. Wave-pinned patterns for cell polarity — a catastrophe theory explanation. SIAM Journal on Applied Dynamical Systems, 23(1):721–747, February 2024. ISSN 1536-0040.

Anderson:2017:FTD

[ACK17] David F. Anderson, Daniele Cappelletti, and Thomas G. Kurtz. Finite time distributions of stochastically modeled chemical systems with absolute concentration robustness. SIAM Journal on Applied Dynamical Systems, 16(3):1309–1339, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Abo:2023:CCT

[ACL23] Stéphanie M. C. Abo, José A.

Carrillo, and Anita T. Layton. Can the clocks tick together despite the noise? Stochastic simulations and analysis. SIAM Journal on Applied Dynamical Systems, 22(2):850–877, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M147788X.

Antonana:2020:GTR

[ACMM20] Mikel Antoñana, Philippe Chartier, Joseba Makazaga, and Ander Murua. Global time-renormalization of the gravitational N-body problem. SIAM Journal on Applied Dynamical Systems, 19 (4):2658–2681, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Austin:2015:MBC

[AD15] Daniel Austin and Ian H. Dinwoodie. Monomials and basin cylinders for network dynamics. SIAM Journal on Applied Dynamical Systems, 14 (1):25–42, ???? 2015. CO-DEN SJADAY. ISSN 1536-0040.

Aguiar:2021:SAW

[AD21] Manuela Aguiar and Ana Dias. Synchrony and antisynchrony in weighted networks. SIAM Journal on Applied Dynamical Systems, 20(3):1382– 1420, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Adams:2023:AFS

[Ada23] Zachary P. Adams. The asymptotic frequency of stochastic oscillators. SIAM Journal on Applied Dynamical Systems, 22(1):311–338, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1439584.

Alzate:2009:ENI

[AdBG^{+}09] R. Alzate, M. di Bernardo, G. Giordano, G. Rea, and S. Santini. Experimental and numerical investigation of coexistence, novel bifurcations, and chaos in a cam-follower system. SIAM Journal on Applied Dynamical Systems, 8(2):592–623, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Alfaro:2020:QET

[ADF20] Matthieu Alfaro, Arnaud Ducrot, and Grégory Faye. Quantitative estimates of the threshold phenomena for propagation in reaction– diffusion equations. SIAM Journal on Applied Dynamical Systems, 19(2):1291–1311, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Alfaro:2021:EQE

[ADF21] Matthieu Alfaro, Arnaud Ducrot, and Grégory Faye. Erratum: Quantitative estimates of the threshold phenomena for propagation in reaction-diffusion equations. SIAM Journal on Applied Dynamical Systems, 20(1):571– 572, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Antoneli:2008:HBC

[ADP08] Fernando Antoneli, Ana Paula S. Dias, and Rui C. Paiva. Hopf bifurcation in coupled cell networks with interior symmetries. SIAM Journal on Applied Dynamical Systems, 7(1):220–248, 2008. CODEN SJADAY. ISSN 1536-0040.

Aguiar:2016:SEO

[ADR16] M. A. D. Aguiar, A. P. S. Dias, and H. Ruan. Synchrony and elementary operations on coupled cell networks. SIAM Journal on Applied Dynamical Systems, 15 (1):322–337, ???? 2016. CO-DEN SJADAY. ISSN 1536-0040.

Andreuzzi:2023:DMD

[ADR23] Francesco Andreuzzi, Nicola Demo, and Gianluigi Rozza. A dynamic mode decomposition extension for the forecasting of parametric dynamical systems. SIAM Journal on Applied Dynamical Systems, 22(3):2432–2458, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1481658.

Ando:2024:NPD

[AEG24] Alessia Andò, Roderick Edwards, and Nicola Guglielmi. Nonuniqueness phenomena in discontinuous dynamical systems and their regularizations. SIAM Journal on Applied Dynamical Systems, 23(2):1345–1371, May 2024. ISSN 1536-0040.

Abell:2005:CMT

[AEHV05] Kate A. Abell, Christopher E. Elmer, A. R. Humphries, and Erik S. Van Vleck. Computation of mixed type functional differential boundary value problems. SIAM Journal on Applied Dynamical Systems, 4(3):755–781, 2005. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60342.

Abaid:2008:AEV

[AEL08] Nicole Abaid, Robert S.

Eisenberg, and Weishi Liu. Asymptotic expansions of I– V relations via a Poisson– Nernst–Planck system. SIAM Journal on Applied Dynamical Systems, 7(4):1507–1526, ???? 2008. CODEN SJA-DAY. ISSN 1536-0040.

Ashwin:2005:ICE

[AG05] Peter Ashwin and Arek Goetz. Invariant curves and explosion of periodic islands in systems of piecewise rotations. SIAM Journal on Applied Dynamical Systems, 4 (2):437–458, 2005. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60539.

Anwar:2023:STS

[AG23] Md Sayeed Anwar and Dibakar Ghosh. Synchronization in temporal simplicial complexes. SIAM Journal on Applied Dynamical Systems, 22(3):2054–2081, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1525909.

Agarwal:2018:SLD

[Aga18] Nikita Agarwal. A simple loop dwell time approach for stability of switched systems. SIAM Journal on Applied Dynamical Systems, 17(2):1377– 1394, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Avery:2019:GSW

[AGG^{+}19] M. Avery, R. Goh, O. Goodloe, A. Milewski, and A. Scheel. Growing stripes, with and without wrinkles. SIAM Journal on Applied Dynamical Systems, 18(2):1078–1117, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Akingbade:2023:ADM

[AGMS23] Samuel W. Akingbade, Marian Gidea, and Tere M-Seara. Arnold diffusion in a model of dissipative system. SIAM Journal on Applied Dynamical Systems, 22(3):1983–2023, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M1525508.

Aguirre:2015:BTD

[Agu15] Pablo Aguirre. Bifurcations of two-dimensional global invariant manifolds near a noncentral saddle-node homoclinic orbit. SIAM Journal on Applied Dynamical Systems, 14(3):1600–1643, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Atay:2006:NFD

[AH06] Fatihcan M. Atay and Axel Hutt. Neural fields with distributed transmission speeds and long-range feedback delays. SIAM Journal on Applied Dynamical Systems, 5 (4):670–698, 2006. CODEN SJADAY. ISSN 1536-0040.

Acosta-Humanez:2009:NHS

[AH09] Primitivo B. Acosta-Humánez.

Nonautonomous Hamiltonian systems and Morales–Ramis theory I. the case ẍ = f(x, t). SIAM Journal on Applied Dynamical Systems, 8(1):279– 297, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Aminzare:2019:HIC

[AH19] Zahra Aminzare and Philip Holmes. Heterogeneous inputs to central pattern generators can shape insect gaits. SIAM Journal on Applied Dynamical Systems, 18(2):1037– 1059, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Acosta-Humanez:2018:NAG

[AHARS18] P. Acosta-Humánez, M. Alvarez-Ramírez, and T. J. Stuchi. Nonintegrability of the Armbruster– Guckenheimer–Kim quartic Hamiltonian through Morales–Ramis theory. SIAM Journal on Applied Dynamical Systems, 17(1):78–96, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Al-Hdaibat:2016:IHS

[AHGKM16] B. Al-Hdaibat, W. Govaerts, Yu. A. Kuznetsov, and H. G. E. Meijer. Initialization of homoclinic solutions near Bogdanov–Takens points: Lindstedt–Poincaré compared with regular perturbation method. SIAM Journal on Applied Dynamical Systems, 15(2):952–980, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Avitabile:2014:NRC

[AHS14] Daniele Avitabile, Rebecca Hoyle, and Giovanni Samaey. Noise reduction in coarse bifurcation analysis of stochastic agent-based models: an example of consumer lock-in. SIAM Journal on Applied Dynamical Systems, 13(4):1583– 1619, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Ambrose:2021:WPA

[AHW21] David M. Ambrose, Fazel Hadadifard, and J. Douglas Wright. Well-posedness and asymptotics of a coordinatefree model of flame fronts. SIAM Journal on Applied Dynamical Systems, 20(4):2261– 2294, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Arai:2018:BHL

[AIT18] Zin Arai, Yutaka Ishii, and Hiroki Takahasi. Boundary of the horseshoe locus for the Hénon family. SIAM Journal on Applied Dynamical Systems, 17(3):2234–2248, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Abedi:2014:CDI

[AJ14] Maryam Abedi and Mir Abbas Jalali. Collective dynamics of interacting particles in unsteady flows. SIAM Journal on Applied Dynamical Systems, 13(1):194–209, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Anderson:2016:UPD

[AJB^{+}16] Ross P. Anderson, Geronimo Jimenez, Jin Yung Bae, Diana Silver, James Macinko, and Maurizio Porfiri. Understanding policy diffusion in the U.S.: an informationtheoretical approach to unveil connectivity structures in slowly evolving complex systems. SIAM Journal on Applied Dynamical Systems, 15 (3):1384–1409, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Atay:2006:SCM

[AK06] Fatihcan M. Atay and Özkan Karabacak. Stability of coupled map networks with delays. SIAM Journal on Applied Dynamical Systems, 5 (3):508–527, 2006. CODEN SJADAY. ISSN 1536-0040.

Arioli:2010:IDP

[AK10] Gianni Arioli and Hans Koch.

Integration of dissipative partial differential equations: a case study. SIAM Journal on Applied Dynamical Systems, 9(3):1119–1133, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Arioli:2017:FPS

[AK17] Gianni Arioli and Hans Koch. Families of periodic solutions for some Hamiltonian PDEs. SIAM Journal on Applied Dynamical Systems, 16(1):1–15, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Askham:2018:VPM

[AK18] Travis Askham and J. Nathan Kutz. Variable projection methods for an optimized dynamic mode decomposition. SIAM Journal on Applied Dynamical Systems, 17(1):380– 416, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Arai:2009:DSA

[AKK^{+}09] Zin Arai, William Kalies, Hiroshi Kokubu, Konstantin Mischaikow, Hiroe Oka, and Paweł Pilarczyk. A database schema for the analysis of global dynamics of multiparameter systems. SIAM Journal on Applied Dynamical Systems, 8(3):757–789, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Aguirre:2013:GIM

[AKO13] Pablo Aguirre, Bernd Krauskopf, and Hinke M. Osinga. Global invariant manifolds near homoclinic orbits to a real saddle: (non)orientability and flip bifurcation. SIAM Journal on Applied Dynamical Systems, 12(4):1803–1846, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Avitabile:2010:SSP

[ALB^{+}10] Daniele Avitabile, David J. B.

Lloyd, John Burke, Edgar Knobloch, and Björn Sandstede. To snake or not to snake in the planar Swift– Hohenberg equation. SIAM Journal on Applied Dynamical Systems, 9(3):704–733, ???? 2010. CODEN SJA-DAY. ISSN 1536-0040.

Arai:2006:RCH

[AM06] Zin Arai and Konstantin Mischaikow. Rigorous computations of homoclinic tangencies. SIAM Journal on Applied Dynamical Systems, 5 (2):280–292, 2006. CODEN SJADAY. ISSN 1536-0040.

Arbabi:2017:ETD

[AM17] Hassan Arbabi and Igor Mezić. Ergodic theory, dynamic mode decomposition, and computation of spectral properties of the Koopman operator. SIAM Journal on Applied Dynamical Systems, 16(4):2096–2126, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Avila:2023:SPP

[AM23] Allan M. Avila and Igor Mezi.

Spectral properties of pullback operators on vector bundles of a dynamical system.

SIAM Journal on Applied Dynamical Systems, 22(4):3059– 3092, November 2023. ISSN 1536-0040.

Alonso:2022:ASI

[AMBV22] Arantxa Alonso, Isabel Mercader, Oriol Batiste, and José M. Vega. Analyzing slightly inclined cylindrical binary fluid convection via higher order dynamic mode decomposition. SIAM Journal on Applied Dynamical Systems, 21(3):2148–2186, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1447416.

Allen:2024:EDH

[AMDH24] L. I. Allen, T. G. Molnár, Z. Dombóvári, and S. J. Hogan. The effects of delay on the HKB model of human motor coordination. SIAM Journal on Applied Dynamical Systems, 23(1):1–25, January 2024. ISSN 1536-0040.

Ambrosio-Mouser:2006:EVT

[AMNB06] Christina Ambrosio-Mouser, Farzan Nadim, and Amitabha Bose. The effects of varying the timing of inputs on a neural oscillator. SIAM Journal on Applied Dynamical Systems, 5(1):108–139, 2006. CODEN SJADAY. ISSN 1536-0040.

Alnahdi:2014:LPP

[ANR14] A. S. Alnahdi, J. Niesen, and A. M. Rucklidge. Localized patterns in periodically forced systems. SIAM Journal on Applied Dynamical Systems, 13(3):1311–1327, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Alnahdi:2018:LPP

[ANR18] A. S. Alnahdi, J. Niesen, and A. M. Rucklidge. Localized patterns in periodically forced systems: II. Patterns with nonzero wavenumber. SIAM Journal on Applied Dynamical Systems, 17(2):1478–1502, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Ashwin:2007:DNC

[AOWT07] Peter Ashwin, Gábor Orosz, John Wordsworth, and Stuart Townley. Dynamics on networks of cluster states for globally coupled phase oscillators. SIAM Journal on Applied Dynamical Systems, 6 (4):728–758, 2007. CODEN SJADAY. ISSN 1536-0040.

Ashwin:2016:QNA

[AP16] Peter Ashwin and Claire Postlethwaite. Quantifying noisy attractors: From heteroclinic to excitable networks. SIAM Journal on Applied Dynamical Systems, 15(4):1989– 2016, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Athanasouli:2022:BSP

[APBB22] Christina Athanasouli, Sofia H.

Piltz, Cecilia G. Diniz Behn, and Victoria Booth. Bifurcations of sleep patterns due to homeostatic and circadian variation in a sleep-wake flipflop model. SIAM Journal on Applied Dynamical Systems, 21(3):1893–1929, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1446149.

Aguiar:2012:ISM

[AR12] M. A. D. Aguiar and H. Ruan. Interior symmetries and multiple eigenvalues for homogeneous networks. SIAM Journal on Applied Dynamical Systems, 11(4):1231–1269, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Antali:2018:SCD

[AS18] Mate Antali and Gabor Stepan. Sliding and crossing dynamics in extended Filippov systems. SIAM Journal on Applied Dynamical Systems, 17(1):823–858, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Aminzare:2018:GTP

[ASH18] Zahra Aminzare, Vaibhav Srivastava, and Philip Holmes. Gait transitions in a phase oscillator model of an insect central pattern generator. SIAM Journal on Applied Dynamical Systems, 17 (1):626–671, ???? 2018. CO-DEN SJADAY. ISSN 1536-0040.

Aoi:2007:MMR

[AST07] Shinya Aoi, Hitoshi Sasaki, and Kazuo Tsuchiya. A multilegged modular robot that meanders: Investigation of turning maneuvers using its inherent dynamic characteristics. SIAM Journal on Applied Dynamical Systems, 6 (2):348–377, 2007. CODEN SJADAY. ISSN 1536-0040.

Alberti:2019:SDA

[AV19] Angelo Alberti and Claudio Vidal. Singularities and dynamics aspects of a particle in a gravitational field of a central punctual body surrounded by a solid circular ring. SIAM Journal on Applied Dynamical Systems, 18 (1):1–32, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Alberti:2023:PSK

[AVV23] Angelo Alberti, Claudio Vidal, and Jhon Vidarte. Periodic solutions, KAM tori, and bifurcations in the planar anisotropic Schwarzschildtype problem. SIAM Journal on Applied Dynamical Systems, 22(2):1053–1075, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1450409.

Akhavan:2020:PCE

[AY20] Naghmeh Akhavan and James A.

Yorke. Population collapse in elite-dominated societies: a differential equations model without differential equations. SIAM Journal on Applied Dynamical Systems, 19 (3):1736–1757, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Akhavan:2023:EMP

[AY23a] Naghmeh Akhavan and James A. Yorke. Extinction of multiple populations and a team of die-out Lyapunov functions. SIAM Journal on Applied Dynamical Systems, 22 (3):2382–2407, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1500253.

Anderson:2023:EAN

[AY23b] Brian D. O. Anderson and Mengbin Ye. Equilibria analysis of a networked bivirus epidemic model using Poincaré–Hopf and manifold theory. SIAM Journal on Applied Dynamical Systems, 22(4):2856–2889, October 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1529981.

Azencot:2019:CDM

[AYB19] Omri Azencot, Wotao Yin, and Andrea Bertozzi. Consistent dynamic mode decomposition. SIAM Journal on Applied Dynamical Systems, 18(3):1565–1585, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Askham:2022:RSM

[AZAK22] Travis Askham, Peng Zheng, Aleksandr Aravkin, and J. Nathan Kutz. Robust and scalable methods for the dynamic mode decomposition. SIAM Journal on Applied Dynamical Systems, 21(1):60– 79, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/21M1417405.

Bramburger:2019:LRR

[BAA^{+}19] Jason J. Bramburger, Dylan Altschuler, Chloe I. Avery, Tharathep Sangsawang, Margaret Beck, Paul Carter, and Björn Sandstede. Localized radial roll patterns in higher space dimensions. SIAM Journal on Applied Dynamical Systems, 18(3):1420– 1453, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Behn:2013:CER

[BAB13] Cecilia G. Diniz Behn, Aparna Ananthasubramaniam, and Victoria Booth. Contrasting existence and robustness of REM/Non-REM cycling in physiologically based models of REM sleep regulatory networks. SIAM Journal on Applied Dynamical Systems, 12(1):279–314, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Balasuriya:2005:DCF

[Bal05] Sanjeeva Balasuriya. Direct chaotic flux quantification in perturbed planar flows: General time-periodicity. SIAM Journal on Applied Dynamical Systems, 4(2):282–311, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60324.

Balasuriya:2011:TDT

[Bal11] Sanjeeva Balasuriya. A tangential displacement theory for locating perturbed saddles and their manifolds. SIAM Journal on Applied Dynamical Systems, 10(3): 1100–1126, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i3/p1100_s1.

Balasuriya:2017:TBT

[Bal17] Sanjeeva Balasuriya. Transport between two fluids across their mutual flow interface: The streakline approach. SIAM Journal on Applied Dynamical Systems, 16(2):1015– 1044, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Batko:2017:WIP

[Bat17] Bogdan Batko. Weak index pairs and the Conley index for discrete multivalued dynamical systems. Part II: Properties of the index. SIAM Journal on Applied Dynamical Systems, 16(3):1587–1617, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Behn:2012:FSA

[BB12] Cecilia G. Diniz Behn and Victoria Booth. A fastslow analysis of the dynamics of REM sleep. SIAM Journal on Applied Dynamical Systems, 11(1):212– 242, ???? 2012. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v11/i1/p212_s1.

Brooks:2016:MTP

[BB16] Heather A. Brooks and Paul C. Bressloff. A mechanism for Turing pattern formation with active and passive transport. SIAM Journal on Applied Dynamical Systems, 15(4):1823–1843, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Baginski:2024:CGW

[BB24] Frank E. Baginski and Valério Ramos Batista. Closed geodesics on Weingarten surfaces with κ_{1}/κ_{2} = c0. SIAM Journal on Applied Dynamical Systems, 23(3):1705–1719, ???? 2024. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/23M1608616.

Biswas:2021:DAN

[BBJ21] Animikh Biswas, Zachary Bradshaw, and Michael S. Jolly. Data assimilation for the Navier–Stokes equations using local observables. SIAM Journal on Applied Dynamical Systems, 20(4):2174–2203, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Bossolini:2017:CSS

[BBK17] Elena Bossolini, Morten Brøns, and Kristian Uldall Kristiansen. Canards in stiction: On solutions of a friction oscillator by regularization. SIAM Journal on Applied Dynamical Systems, 16 (4):2233–2258, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Bick:2023:PON

[BBK23] Christian Bick, Tobias Böhle, and Christian Kuehn. Phase oscillator networks with nonlocal higher-order interactions: Twisted states, stability, and bifurcations. SIAM Journal on Applied Dynamical Systems, 22(3):1590–1638, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M1500940.

Best:2005:DRB

[BBR^{+}05] Janet Best, Alla Borisyuk, Jonathan Rubin, David Terman, and Martin Wechselberger. The dynamic range of bursting in a model respiratory pacemaker network. SIAM Journal on Applied Dynamical Systems, 4(4): 1107–1139, 2005. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.

org/sam-bin/dbq/article/
62554.

Birtea:2009:ASD

[BC09] Petre Birtea and Dan Comănescu.

Asymptotic stability of dissipated Hamilton–Poisson systems. SIAM Journal on Applied Dynamical Systems, 8 (3):967–976, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Bressloff:2014:SDN

[BC14] Paul C. Bressloff and Samuel R.

Carroll. Spatiotemporal dynamics of neural fields on product spaces. SIAM Journal on Applied Dynamical Systems, 13(4):1620–1653, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Buono:2015:SBB

[BC15] Pietro-Luciano Buono and Juancho A. Collera. Symmetrybreaking bifurcations in rings of delay-coupled semiconductor lasers. SIAM Journal on Applied Dynamical Systems, 14(4):1868–1898, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Bauer:2021:ETS

[BC21] Adam Bauer and Paul Carter.

Existence of transonic solutions in the stellar wind problem with viscosity and heat conduction. SIAM Journal on Applied Dynamical Systems, 20(1):262–298, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Bustamante:2023:NCC

[BC23] Adrián P. Bustamante and Cristel Chandre. Numerical computation of critical surfaces for the breakup of invariant tori in Hamiltonian systems. SIAM Journal on Applied Dynamical Systems, 22(1):483–500, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1448501.

Bastiaansen:2020:PSE

[BCBD20] Robbin Bastiaansen, Martina Chirilus-Bruckner, and Arjen Doelman. Pulse solutions for an extended Klausmeier model with spatially varying coefficients. SIAM Journal on Applied Dynamical Systems, 19(1):1–57, ???? 2020. CO-DEN SJADAY. ISSN 1536-0040.

Bizzarri:2016:NSC

[BCDG16] Federico Bizzarri, Alessandro Colombo, Fabio Dercole, and Giancarlo Storti Gajani. Necessary and sufficient conditions for the noninvertibility of fundamental solution matrices of a discontinuous system. SIAM Journal on Applied Dynamical Systems, 15 (1):84–105, ???? 2016. CO-DEN SJADAY. ISSN 1536-0040.

Buono:2018:SBB

[BCF^{+}18] Pietro-Luciano Buono, Bernard Chan, Jocirei Ferreira, Antonio Palacios, Steven Reeves, Patrick Longhini, and Visarath In. Symmetry-breaking bifurcations and patterns of oscillations in rings of crystal oscillators. SIAM Journal on Applied Dynamical Systems, 17(2):1310–1352, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Berry:2013:TSS

[BCGFS13] T. Berry, J. R. Cressman, Z. Gregurić-Ferencek, and T. Sauer. Time-scale separation from diffusion-mapped delay coordinates. SIAM Journal on Applied Dynamical Systems, 12(2):618–649, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Banhelyi:2008:CAP

[BCGH08] Balázs Bánhelyi, Tibor Csendes, Barnabas M. Garay, and László Hatvani. A computerassisted proof of Σ_{3}-chaos in the forced damped pendulum equation. SIAM Journal on Applied Dynamical Systems, 7(3):843–867, 2008. CODEN SJADAY. ISSN 1536-0040.

Barrabes:2010:LCR

[BCH10] E. Barrabés, J. M. Cors, and G. R. Hall. A limit case of the “ring problem”: The planar circular restricted 1 + n body problem. SIAM Journal on Applied Dynamical Systems, 9(2):634–658, ???? 2010. CO-DEN SJADAY. ISSN 1536-

Bazso:2014:BAS

[BCH14] Csaba Bazsó, Alan R. Champneys, and Csaba J. Hös. Bifurcation analysis of a simplified model of a pressure relief valve attached to a pipe. SIAM Journal on Applied Dynamical Systems, 13 (2):704–721, ???? 2014. CO-DEN SJADAY. ISSN 1536-0040.

Bush:2016:CMD

[BCHM16] Justin Bush, Wes Cowan, Shaun Harker, and Konstantin Mischaikow. Conley– Morse databases for the angular dynamics of Newton’s method on the plane. SIAM Journal on Applied Dynamical Systems, 15(2):736–766, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Belykh:2019:SMN

[BCJ19] Igor Belykh, Douglas Carter, and Russell Jeter. Synchronization in multilayer networks: When good links go bad. SIAM Journal on Applied Dynamical Systems, 18 (4):2267–2302, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Banhelyi:2014:GAZ

[BCKN14] Balázs Bánhelyi, Tibor Csendes, Tibor Krisztin, and Arnold Neumaier. Global attractivity of the zero solution for Wright’s equation. SIAM Journal on Applied Dynamical Systems, 13(1):537–563, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Brooks:2024:EPS

[BCP24] Heather Z. Brooks, Philip S.

Chodrow, and Mason A. Porter. Emergence of polarization in a sigmoidal bounded-confidence model of opinion dynamics. SIAM Journal on Applied Dynamical Systems, 23(2):1442–1470, ???? 2024. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M1527258.

Barrabes:2008:EEI

[BCPS08] E. Barrabés, J. M. Cors, C. Pinyol, and J. Soler. J_{2} effect and elliptic inclined periodic orbits in the collision restricted three-body problem. SIAM Journal on Applied Dynamical Systems, 7(1):1–17, 2008. CODEN SJADAY. ISSN 1536-0040.

Bel:2021:PST

[BCRR21] Andrea Bel, Romina Cobiaga, Walter Reartes, and Horacio G. Rotstein. Periodic solutions in threshold-linear networks and their entrainment. SIAM Journal on Applied Dynamical Systems, 20(3):1177– 1208, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Blank:2011:FSG

[BD11] Elisabeth Blank and Tomás Dohnal. Families of surface gap solitons and their stability via the numerical Evans function method. SIAM Journal on Applied Dynamical Systems, 10(2): 667–706, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i2/p667_s1.

Burke:2012:LSE

[BD12] John Burke and Jonathan H. P. Dawes. Localized states in an extended Swift– Hohenberg equation. SIAM Journal on Applied Dynamical Systems, 11(1):261–284, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Berry:2023:LTD

[BD23] Tyrus Berry and Suddhasattwa Das. Learning theory for dynamical systems. SIAM Journal on Applied Dynamical Systems, 22(3):2082–2122, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M1516865.

Buzzi:2012:PFP

[BdCT12] Claudio A. Buzzi, Tiago de Carvalho, and Marco A. Teixeira. On 3-parameter families of piecewise smooth vector fields in the plane. SIAM Journal on Applied Dynamical Systems, 11(4):1402– 1424, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Breda:2016:PDN

[BDG^{+}16] D. Breda, O. Diekmann, M. Gyllenberg, F. Scarabel, and R. Vermiglio. Pseudospectral discretization of nonlinear delay equations: New prospects for numerical bifurcation analysis. SIAM Journal on Applied Dynamical Systems, 15(1):1–23, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Betts:2003:OLT

[BE03] John T. Betts and Sven O.

Erb. Optimal low thrust trajectories to the moon. SIAM Journal on Applied Dynamical Systems, 2(2):144–170, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40908.

Buono:2014:CTB

[BE14] Pietro-Luciano Buono and Raluca Eftimie. Codimensiontwo bifurcations in animal aggregation models with symmetry. SIAM Journal on Applied Dynamical Systems, 13 (4):1542–1582, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Bold:2003:FVP

[BEG^{+}03] Katherine Bold, Chantal Edwards, John Guckenheimer, Sabyasachi Guharay, Kathleen Hoffman, Judith Hubbard, Ricardo Oliva, and Warren Weckesser. The forced van der Pol equation II: Canards in the reduced system. SIAM Journal on Applied Dynamical Systems, 2(4):570–608, 2003. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
41913.

Blackbeard:2011:SIB

[BEW11] Nicholas Blackbeard, Hartmut Erzgräber, and Sebastian Wieczorek. Shear-induced bifurcations and chaos in models of three coupled lasers. SIAM Journal on Applied Dynamical Systems, 10(2): 469–509, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i2/p469_s1.

Bronski:2018:VBP

[BF18] Jared C. Bronski and Timothy Ferguson. Volume bounds for the phase-locking region in the Kuramoto model. SIAM Journal on Applied Dynamical Systems, 17(1):128–156, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Berchio:2021:SNN

[BFG21] Elvise Berchio, Alessio Falocchi, and Maurizio Garrione. On the stability of a nonlinear nonhomogeneous multiply hinged beam. SIAM Journal on Applied Dynamical Systems, 20(2):908–940, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Bose:2014:UML

[BFGTM14] Christopher Bose, Gary Froyland, Cecilia González-Tokman, and Rua Murray. Ulam’s method for Lasota– Yorke maps with holes. SIAM Journal on Applied Dynamical Systems, 13(2):1010–1032, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Burger:2014:SSA

[BFH14] Martin Burger, Razvan Fetecau, and Yanghong Huang. Stationary states and asymptotic behavior of aggregation models with nonlinear local repulsion. SIAM Journal on Applied Dynamical Systems, 13(1):397–424, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Beyn:2022:AVN

[BFH22] Wolf-Jürgen Beyn, Gary Froyland, and Thorsten Hüls. Angular values of nonautonomous and random linear dynamical systems: Part I — fundamentals. SIAM Journal on Applied Dynamical Systems, 21(2):1245–1286, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/20M1387730.

Bates:2018:GDM

[BFK18] Peter Bates, Giorgio Fusco, and Georgia Karali. Gradient dynamics: Motion near a manifold of quasi-equilibria. SIAM Journal on Applied Dynamical Systems, 17(3):2106– 2145, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Bayen:2022:MMT

[BFK^{+}22] Alexandre Bayen, Jan Friedrich, Alexander Keimer, Lukas Pflug, and Tanya Veeravalli. Modeling multilane traffic with moving obstacles by nonlocal balance laws. SIAM Journal on Applied Dynamical Systems, 21(2):1495–1538, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/20M1366654.

Bahsoun:2005:MSP

[BGB05] Wael Bahsoun, Pawel Góra, and Abraham Boyarsky. Markov switching for position dependent random maps with application to forecasting. SIAM Journal on Applied Dynamical Systems, 4 (2):391–406, 2005. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60404.

Batko:2024:IND

[BGH^{+}24] Bogdan Batko, Marcio Gameiro, Ying Hung, William Kalies, Konstantin Mischaikow, and Ewerton Vieira. Identifying nonlinear dynamics with high confidence from sparse data. SIAM Journal on Applied Dynamical Systems, 23(1):383– 409, January 2024. ISSN 1536-0040.

Borek:2011:CRP

[BGO11] Bartlomiej Borek, Leon Glass, and Bart E. Oldeman. Continuity of resetting a pacemaker in an excitable medium. SIAM Journal on Applied Dynamical Systems, 10(4):1502– 1524, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i4/p1502_s1.

Bollt:2008:BMP

[BGOŻ08] Erik Bollt, Pawe Góra, Andrzej Ostruszka, and Karol Życzkowski. Basis Markov partitions and transition matrices for stochastic systems. SIAM Journal on Applied Dynamical Systems, 7(2):341– 360, 2008. CODEN SJADAY. ISSN 1536-0040.

Belbruno:2010:WSB

[BGT10] Edward Belbruno, Marian Gidea, and Francesco Topputo. Weak stability boundary and invariant manifolds. SIAM Journal on Applied Dynamical Systems, 9(3):1061– 1089, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Berchio:2016:WRM

[BGZ16] Elvise Berchio, Filippo Gazzola, and Chiara Zanini. Which residual mode captures the energy of the dominating mode in second order Hamiltonian systems? SIAM Journal on Applied Dynamical Systems, 15(1):338–355, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Boros:2020:PWR

[BH20] Balázs Boros and Josef Hofbauer. Permanence of weakly reversible mass-action systems with a single linkage class. SIAM Journal on Applied Dynamical Systems, 19 (1):352–365, ???? 2020. CO-DEN SJADAY. ISSN 1536-0040.

Beyn:2023:AVN

[BH23] Wolf-Jürgen Beyn and Thorsten Hüls. Angular values of nonautonomous linear dynamical systems: Part II — reduction theory and algorithm. SIAM Journal on Applied Dynamical Systems, 22(1):162–198, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/20M1387766.

Barry:2016:ESS

[BHL16] Anna M. Barry and Alanna Hoyer-Leitzel. Existence, stability, and symmetry of relative equilibria with a dominant vortex. SIAM Journal on Applied Dynamical Systems, 15(4):1783–1805, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Bakhtin:2021:SID

[BHLM21] Yuri Bakhtin, Tobias Hurth, Sean D. Lawley, and Jonathan C.

Mattingly. Singularities of invariant densities for random switching between two linear ODEs in 2D. SIAM Journal on Applied Dynamical Systems, 20(4):1917–1958, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Bradshaw-Hajek:2024:GSP

[BHLMW24] Bronwyn H. Bradshaw-Hajek, Ian Lizarraga, Robert Marangell, and Martin Wechselberger. A geometric singular perturbation analysis of shock selection rules in composite regularized reaction–nonlinear diffusion models. SIAM Journal on Applied Dynamical Systems, 23(3):2099–2137, ???? 2024. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/23M1591803.

Barker:2018:EVL

[BHLZ18] Blake Barker, Jeffrey Humpherys, Gregory Lyng, and Kevin Zumbrun. Euler versus Lagrange: The role of coordinates in practical Evansfunction computations. SIAM Journal on Applied Dynamical Systems, 17(2):1766–1785, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Barcelona:2024:SCH

[BHM24] Miquel Barcelona, Alex Haro, and Josep-Maria Mondelo. Semianalytical computation of heteroclinic connections between center manifolds with the parameterization method. SIAM Journal on Applied Dynamical Systems, 23(1):98–126, January 2024. ISSN 1536-0040.

Barranca:2021:DDR

[BHP^{+}21] Victor J. Barranca, Yolanda Hu, Zoe Porterfield, Samuel Rothstein, and Alex Xuan. Data-driven reconstruction and encoding of sparse stimuli across convergent sensory layers from downstream neuronal network dynamics. SIAM Journal on Applied Dynamical Systems, 20(4):2602– 2629, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Brucal-Hallare:2011:TWA

[BHV11] Maila Brucal-Hallare and Erik Van Vleck. Traveling wavefronts in an antidiffusion lattice Nagumo model. SIAM Journal on Applied Dynamical Systems, 10(3): 921–959, ???? 2011. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.
                                        siam.org/siads/resource/
                                        1/sjaday/v10/i3/p921_s1.

Bychkov:2024:EOQ

[BIPK24] Andrey Bychkov, Opal Issan, Gleb Pogudin, and Boris Kramer. Exact and optimal quadratization of nonlinear finite-dimensional nonautonomous dynamical systems. SIAM Journal on Applied Dynamical Systems, 23(1):982– 1016, March 2024. ISSN 1536-0040.

Beck:2022:VSS

[BJ22] Margaret Beck and Jonathan Jaquette. Validated spectral stability via conjugate points. SIAM Journal on Applied Dynamical Systems, 21(1):366–404, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1420095.

Blessing:2024:WBA

[BJ24] David Blessing and J. D. Mireles James. Weighted Birkhoff averages and the parameterization method. SIAM Journal on Applied Dynamical Systems, 23(3):1766–1804, ???? 2024. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/23M1579546.

Bosschaert:2020:SNC

[BJK20] Maikel M. Bosschaert, Sebastiaan G. Janssens, and Yu. A. Kuznetsov. Switching to nonhyperbolic cycles from codimension two bifurcations of equilibria of delay differential equations. SIAM Journal on Applied Dynamical Systems, 19(1):252–303, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Bates:2017:IFS

[BJL^{+}17] Peter W. Bates, Yusheng Jia, Guojian Lin, Hong Lu, and Mingji Zhang. Individual flux study via steadystate Poisson–Nernst–Planck systems: Effects from boundary conditions. SIAM Journal on Applied Dynamical Systems, 16(1):410–430, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Barker:2020:PMU

[BJM20] Blake Barker, Jason Mireles James, and Jalen Morgan. Parameterization method for unstable manifolds of standing waves on the line. SIAM Journal on Applied Dynamical Systems, 19(3):1758–1797, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Beck:2008:EWO

[BJSW08] Margaret Beck, Christopher K. R. T. Jones, David Schaeffer, and Martin Wechselberger. Electrical waves in a one-dimensional model of cardiac tissue. SIAM Journal on Applied Dynamical Systems, 7(4):1558–1581, ???? 2008. CODEN SJADAY. ISSN 1536-0040.

Bressloff:2015:NLE

[BK15] Paul C. Bressloff and Zachary P. Kilpatrick. Nonlinear Langevin equations for wandering patterns in stochastic neural fields. SIAM Journal on Applied Dynamical Systems, 14 (1):305–334, ???? 2015. CO-DEN SJADAY. ISSN 1536-0040.

Breden:2020:CIS

[BK20a] Maxime Breden and Christian Kuehn. Computing invariant sets of random differential equations using polynomial chaos. SIAM Journal on Applied Dynamical Systems, 19(1):577–618, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Briat:2020:EAA

[BK20b] Corentin Briat and Mustafa Khammash. Ergodicity analysis and antithetic integral control of a class of stochastic reaction networks with delays. SIAM Journal on Applied Dynamical Systems, 19 (3):1575–1608, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Bosschaert:2024:BAB

[BK24a] M. M. Bosschaert and Yu. A.

Kuznetsov. Bifurcation analysis of Bogdanov–Takens bifurcations in delay differential equations. SIAM Journal on Applied Dynamical Systems, 23(1):553–591, January 2024. ISSN 1536-0040.

Bosschaert:2024:IBN

[BK24b] Maikel M. Bosschaert and Yuri A. Kuznetsov. Interplay between normal forms and center manifold reduction for homoclinic predictors near Bogdanov–Takens bifurcation. SIAM Journal on Applied Dynamical Systems, 23(1):410–439, January 2024. ISSN 1536-0040.

Bittracher:2015:PST

[BKJ15] Andreas Bittracher, Péter Koltai, and Oliver Junge. Pseudogenerators of spatial transfer operators. SIAM Journal on Applied Dynamical Systems, 14(3):1478–1517, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Barclay:2022:SPS

[BKM22] Bryce M. Barclay, Eric J. Kostelich, and Alex Mahalov. Sensor placement sensitivity and robust reconstruction of wave dynamics from multiple sensors. SIAM Journal on Applied Dynamical Systems, 21(4):2297–2313, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1479270.

Borovykh:2024:SMD

[BKPP24] A. Borovykh, N. Kantas, P. Parpas, and G. A. Pavliotis. Stochastic mirror descent for convex optimization with consensus constraints. SIAM Journal on Applied Dynamical Systems, 23(3):2208–2241, ???? 2024. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M1515197.

Bayen:2019:TCI

[BKPS19] Alexandre Bayen, Alexander Keimer, Emily Porter, and Michele Spinola. Timecontinuous instantaneous and past memory routing on traffic networks: a mathematical analysis on the basis of the link-delay model. SIAM Journal on Applied Dynamical Systems, 18(4):2143–2180, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Barkley:2006:MMN

[BKS06] D. Barkley, I. G. Kevrekidis, and A. M. Stuart. The moment map: Nonlinear dynamics of density evolution via a few moments. SIAM Journal on Applied Dynamical Systems, 5(3):403–434, 2006. CODEN SJADAY. ISSN 1536-0040.

Buono:2008:RCE

[BL08] P.-L. Buono and V. G.

LeBlanc. Realization of critical eigenvalues for scalar and symmetric linear delaydifferential equations. SIAM Journal on Applied Dynamical Systems, 7(4):1323–1354, 2008. CODEN SJADAY. ISSN 1536-0040.

Bollt:2018:MER

[BLDK18] Erik M. Bollt, Qianxiao Li, Felix Dietrich, and Ioannis Kevrekidis. On matching, and even rectifying, dynamical systems through Koopman operator eigenfunctions. SIAM Journal on Applied Dynamical Systems, 17(2):1925– 1960, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Baraviera:2012:SGS

[BLL12] A. T. Baraviera, R. Leplaideur, and A. O. Lopes. Selection of ground states in the zero temperature limit for a one-parameter family of potentials. SIAM Journal on Applied Dynamical Systems, 11(1):243– 260, ???? 2012. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.
                                        siam.org/siads/resource/
                                        1/sjaday/v11/i1/p243_s1.

Blomker:2003:AEL

[Blö03] Dirk Blömker. Amplitude equations for locally cubic nonautonomous nonlinearities. SIAM Journal on Applied Dynamical Systems, 2 (3):464–486, 2003. CODEN SJADAY. ISSN 1536-0040.

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        42135.

Baldoma:2012:IEG

[BM12] I. Baldomá and P. Martín. The inner equation for generalized standard maps. SIAM Journal on Applied Dynamical Systems, 11(3):1062–1097, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Briane:2015:IRC

[BM15] M. Briane and G. W. Milton. Isotropic realizability of current fields in R^{3}. SIAM Journal on Applied Dynamical Systems, 14(2):1165–1188, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Batko:2016:WIP

[BM16] Bogdan Batko and Marian Mrozek. Weak index pairs and the Conley index for discrete multivalued dynamical systems. SIAM Journal on Applied Dynamical Systems, 15(2):1143–1162, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Bressloff:2018:VMA

[BM18] Paul C. Bressloff and James N.

MacLaurin. A variational method for analyzing stochastic limit cycle oscillators. SIAM Journal on Applied Dynamical Systems, 17(3):2205– 2233, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Backer:2020:EBM

[BM20a] Arnd Bäcker and James D.

Meiss. Elliptic bubbles in Moser’s 4D quadratic map: The quadfurcation. SIAM Journal on Applied Dynamical Systems, 19(1):442–479, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Bressloff:2020:PRS

[BM20b] Paul C. Bressloff and James N.

MacLaurin. Phase reduction of stochastic biochemical oscillators. SIAM Journal on Applied Dynamical Systems, 19(1):151–180, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Bortner:2024:GBS

[BM24] Cashous Bortner and Nicolette Meshkat. Graphbased sufficient conditions for the indistinguishability of linear compartmental models. SIAM Journal on Applied Dynamical Systems, 23 (3):2179–2207, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1614663.

Brenna-Medina:2014:MMP

[BMCGW14] V. Brenña-Medina, A. R. Champneys, C. Grierson, and M. J. Ward. Mathematical modeling of plant root hair initiation: Dynamics of localized patches. SIAM Journal on Applied Dynamical Systems, 13(1):210–248, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Batko:2020:CIA

[BMMP20] Bogdan Batko, Konstantin Mischaikow, Marian Mrozek, and Mateusz Przybylski. Conley index approach to sampled dynamics. SIAM Journal on Applied Dynamical Systems, 19(1):665–704, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Baldoma:2024:SSR

[BMSM24] Inmaculada Baldomá, Tere M-Seara, and Román Moreno. Splitting of separatrices for rapid degenerate perturbations of the classical pendulum. SIAM Journal on Applied Dynamical Systems, 23(2):1159–1198, May 2024. ISSN 1536-0040.

Berner:2021:MDA

[BMSY21] Rico Berner, Volker Mehrmann, Eckehard Schöll, and Serhiy Yanchuk. The multiplex decomposition: an analytic framework for multilayer dynamical networks. SIAM Journal on Applied Dynamical Systems, 20(4):1752–1772, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Burylko:2018:CHL

[BMWY18] Oleksandr Burylko, Alexander Mielke, Matthias Wolfrum, and Serhiy Yanchuk. Coexistence of Hamiltonianlike and dissipative dynamics in rings of coupled phase oscillators with skew-symmetric coupling. SIAM Journal on Applied Dynamical Systems, 17(3):2076–2105, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Bressloff:2013:MSN

[BN13] Paul C. Bressloff and Jay M.

Newby. Metastability in a stochastic neural network modeled as a velocity jump Markov process. SIAM Journal on Applied Dynamical Systems, 12(3):1394–1435, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Bao:2019:SCR

[BN19] Jinrong Bao and Anatoly Neishtadt. Separatrix crossing in rotation of a body with changing geometry of masses. SIAM Journal on Applied Dynamical Systems, 18(1):150– 171, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Blomker:2023:BTS

[BN23] Dirk Blömker and Alexandra Neamtu. Bifurcation theory for SPDEs: Finite-time Lyapunov exponents and amplitude equations. SIAM Journal on Applied Dynamical Systems, 22(3):2150–2179, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/23M1549638.

Bokes:2022:STD

[Bok22]  Pavol Bokes.  Stationary

and time-dependent molecular distributions in slowfast feedback circuits. SIAM Journal on Applied Dynamical Systems, 21(2):903–931, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1404338.

Bretti:2008:TAC

[BP08] Gabriella Bretti and Benedetto Piccoli. A tracking algorithm for car paths on road networks. SIAM Journal on Applied Dynamical Systems, 7 (2):510–531, 2008. CODEN SJADAY. ISSN 1536-0040.

Banaji:2016:SRI

[BP16] Murad Banaji and Casian Pantea. Some results on injectivity and multistationarity in chemical reaction networks. SIAM Journal on Applied Dynamical Systems, 15 (2):807–869, ???? 2016. CO-DEN SJADAY. ISSN 1536-0040.

Bi:2013:BDD

[BR13] Ping Bi and Shigui Ruan. Bifurcations in delay differential equations and applications to tumor and immune system interaction models. SIAM Journal on Applied Dynamical Systems, 12(4):1847–1888, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Bonnard:2023:OCC

[BR23] Bernard Bonnard and Jérémy Rouot. Optimal control of the controlled Lotka–Volterra equations with applications. The permanent case. SIAM Journal on Applied Dynamical Systems, 22(4):2761–2791, October 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M151978X.

Breden:2023:PVG

[Bre23] Maxime Breden. A posteriori validation of generalized polynomial chaos expansions. SIAM Journal on Applied Dynamical Systems, 22(2):765–801, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1493197.

Briane:2019:IRF

[Bri19] Marc Briane. Isotropic realizability of fields and reconstruction of invariant measures under positivity properties. asymptotics of the flow by a non-ergodic approach. SIAM Journal on Applied Dynamical Systems, 18(4):1846– 1866, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Briat:2020:BIA

[Bri20] Corentin Briat. A biologyinspired approach to the positive integral control of positive systems: The antithetic, exponential, and logistic integral controllers. SIAM Journal on Applied Dynamical Systems, 19(1):619–664, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Bou-Rabee:2004:TTI

[BRMR04] Nawaf M. Bou-Rabee, Jerrold E. Marsden, and Louis A. Romero. Tippe top inversion as a dissipation-induced instability. SIAM Journal on Applied Dynamical Systems, 3(3):352–377, 2004. CODEN SJADAY. ISSN 1536-0040.

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        60135.

Brocker:2017:EUF

[Brö17] Jochen Bröcker. Existence and uniqueness for fourdimensional variational data assimilation in discrete time. SIAM Journal on Applied Dynamical Systems, 16(1):361– 374, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Bou-Rabee:2002:MNS

[BRRS02] Nawaf M. Bou-Rabee, Louis A.

Romero, and Andrew G. Salinger. A multiparameter, numerical stability analysis of a standing cantilever conveying fluid. SIAM Journal on Applied Dynamical Systems, 1(2):190–214, 2002. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
40075.

Budd:2005:MBS

[BRW05] C. J. Budd, V. Rottschäfer, and J. F. Williams. Multibump, blow-up, self-similar solutions of the complex Ginzburg–Landau equation. SIAM Journal on Applied Dynamical Systems, 4(3):649– 678, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/61086.

Bandegi:2018:AGM

[BS18] Mahdi Bandegi and David Shirokoff. Approximate global minimizers to pairwise interaction problems via convex relaxation. SIAM Journal on Applied Dynamical Systems, 17(1):417–456, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Blanchard:2019:ADO

[BS19] Antoine Blanchard and Themistoklis P. Sapsis. Analytical description of optimally timedependent modes for reducedorder modeling of transient instabilities. SIAM Journal on Applied Dynamical Systems, 18(2):1143–1162, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Burden:2016:ESV

[BSKR16] Samuel A. Burden, S. Shankar Sastry, Daniel E. Koditschek, and Shai Revzen. Eventselected vector field discontinuities yield piecewisedifferentiable flows. SIAM Journal on Applied Dynamical Systems, 15(2):1227–1267, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Bar-Shalom:2020:RFM

[BSOM20] Eyal Bar-Shalom, Alexander Ovseevich, and Michael Margaliot. Ribosome flow model with different site sizes. SIAM Journal on Applied Dynamical Systems, 19(1):541–576, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Beyn:2008:FMM

[BST08] Wolf-Jürgen Beyn, Sabrina Selle, and Vera Thümmler. Freezing multipulses and multifronts. SIAM Journal on Applied Dynamical Systems, 7(2):577–608, 2008. CODEN SJADAY. ISSN 1536-0040.

Blomker:2016:DNC

[BSW16] Dirk Blömker, Evelyn Sander, and Thomas Wanner. Degenerate nucleation in the Cahn– Hilliard–Cook model. SIAM Journal on Applied Dynamical Systems, 15(1):459–494, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Berner:2019:MNA

[BSY19] Rico Berner, Eckehard Schöll, and Serhiy Yanchuk. Multiclusters in networks of adaptively coupled phase oscillators. SIAM Journal on Applied Dynamical Systems, 18 (4):2227–2266, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Beyn:2004:FSE

[BT04] W.-J. Beyn and V. Thümmler.

Freezing solutions of equivariant evolution equations. SIAM Journal on Applied Dynamical Systems, 3(2):85– 116, 2004. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60051.

Beltrame:2010:TIS

[BT10] Philippe Beltrame and Uwe Thiele. Time integration and steady-state continuation for 2d lubrication equations. SIAM Journal on Applied Dynamical Systems, 9(2):484– 518, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Bernoff:2011:PSE

[BT11] Andrew J. Bernoff and Chad M. Topaz. A primer of swarm equilibria. SIAM Journal on Applied Dynamical Systems, 10(1):212– 250, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i1/p212_s1.

Bernoff:2016:BAD

[BT16] Andrew J. Bernoff and Chad M. Topaz. Biological aggregation driven by social and environmental factors: a nonlocal model and its degenerate Cahn–Hilliard approximation. SIAM Journal on Applied Dynamical Systems, 15(3):1528–1562, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Baccelli:2019:RMF

[BT19] François Baccelli and Thibaud Taillefumier. Replica-mean-field limits for intensity-based neural networks. SIAM Journal on Applied Dynamical Systems, 18(4):1756–1797, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Baccelli:2021:PRM

[BT21] François Baccelli and Thibaud Taillefumier. The pairreplica-mean-field limit for intensity-based neural networks. SIAM Journal on Applied Dynamical Systems, 20 (1):165–207, ???? 2021. CO-DEN SJADAY. ISSN 1536-0040.

Brunton:2014:CSL

[BTBK14] Steven L. Brunton, Jonathan H.

Tu, Ido Bright, and J. Nathan Kutz. Compressive sensing and low-rank libraries for classification of bifurcation regimes in nonlinear dynamical systems. SIAM Journal on Applied Dynamical Systems, 13(4):1716–1732, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Bick:2012:APF

[BTK12] Christian Bick, Marc Timme, and Christoph Kolodziejski. Adapting predictive feedback chaos control for optimal convergence speed. SIAM Journal on Applied Dynamical Systems, 11(4):1310–1324, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Ben-Tal:2016:CGS

[BTK16] A. Ben-Tal and I. G.

Kevrekidis. Coarse-graining and simplification of the dynamics seen in bursting neurons. SIAM Journal on Applied Dynamical Systems, 15 (2):1193–1226, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Branicki:2023:PBD

[BU23] Michał Branicki and Kenneth Uda. Path-based divergence rates and Lagrangian uncertainty in stochastic flows. SIAM Journal on Applied Dynamical Systems, 22(1):419–482, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1466530.

Butler:2020:HST

[But20] Leo T. Butler. Horseshoes for singly thermostated Hamiltonians. SIAM Journal on Applied Dynamical Systems, 19 (4):2268–2285, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Buza:2024:SSN

[Buz24] Gergely Buza. Spectral submanifolds of the Navier– Stokes equations. SIAM Journal on Applied Dynamical Systems, 23(2):1052–1089, April 2024. ISSN 1536-0040.

Bakker:2018:FHA

[BvdBV18] Bente Bakker, Jan Bouwe van den Berg, and Rob Vandervorst. A Floer homology approach to traveling waves in reaction–diffusion equations on cylinders. SIAM Journal on Applied Dynamical Systems, 17(4):2634–2706, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Beck:2009:UGI

[BW09] Margaret Beck and C. Eugene Wayne. Using global invariant manifolds to understand metastability in the Burgers equation with small viscosity.

SIAM Journal on Applied Dynamical Systems, 8(3):1043– 1065, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Bressloff:2012:FPS

[BW12] Paul C. Bressloff and Matthew A.

Webber. Front propagation in stochastic neural fields. SIAM Journal on Applied Dynamical Systems, 11(2):708–740, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Booth:2017:ODM

[BXB17] Victoria Booth, Ismael Xique, and Cecilia G. Diniz Behn. One-dimensional map for the circadian modulation of sleep in a sleep-wake regulatory network model for human sleep. SIAM Journal on Applied Dynamical Systems, 16 (2):1089–1112, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Burke:2008:CSL

[BYK08] J. Burke, A. Yochelis, and E. Knobloch. Classification of spatially localized oscillations in periodically forced dissipative systems. SIAM Journal on Applied Dynamical Systems, 7(3):651–711, 2008. CODEN SJADAY. ISSN 1536-0040.

Capinski:2012:CAE

[Cap12] Maciej J. Capiński. Computer assisted existence proofs of Lyapunov orbits at L_{2} and transversal intersections of invariant manifolds in the Jupiter–Sun PCR3BP. SIAM Journal on Applied Dynamical Systems, 11(4):1723–1753, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Creaser:2020:SES

[CATA20] Jennifer Creaser, Peter Ashwin, and Krasimira Tsaneva-Atanasova. Sequential escapes and synchrony breaking for networks of bistable oscillatory nodes. SIAM Journal on Applied Dynamical Systems, 19(4):2829–2846, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Carroll:2016:PEP

[CB16] Samuel R. Carroll and Paul C. Bressloff. Phase equation for patterns of orientation selectivity in a neural field model of visual cortex. SIAM Journal on Applied Dynamical Systems, 15(1):60– 83, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Carroll:2018:SBN

[CB18] Samuel R. Carroll and Paul C. Bressloff. Symmetric bifurcations in a neural field model for encoding the direction of spatial contrast gradients. SIAM Journal on Applied Dynamical Systems, 17 (1):1–51, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Champion:2019:DNM

[CBK19] Kathleen P. Champion, Steven L. Brunton, and J. Nathan Kutz. Discovery of nonlinear multiscale systems: Sampling strategies and embeddings. SIAM Journal on Applied Dynamical Systems, 18 (1):312–333, ???? 2019. CO-DEN SJADAY. ISSN 1536-0040.

Chossat:2005:MSP

[CBR05] Pascal Chossat and Nawaf M.

Bou-Rabee. The motion of the spherical pendulum subjected to a D_{n} symmetric perturbation. SIAM Journal on Applied Dynamical Systems, 4(4):1140–1158, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/61668.

Chen:2019:DCT

[CBR19] Ricky X. F. Chen, Andrei C.

Bura, and Christian M. Reidys. D-chain tomography of networks: a new structure spectrum and an application to the SIR process. SIAM Journal on Applied Dynamical Systems, 18(4):2181–2201, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Cannon:2024:SCB

[CBWS24] Olivia Cannon, Ty Bondurant, Malindi Whyte, and Arnd Scheel. Shifting consensus in a biased compromise model. SIAM Journal on Applied Dynamical Systems, 23(1):297–324, January 2024. ISSN 1536-0040.

Chow:2006:EWB

[CC06a] Carson C. Chow and S. Coombes. Existence and wandering of bumps in a spiking neural network model. SIAM Journal on Applied Dynamical Systems, 5(4):552–574, 2006. CODEN SJADAY. ISSN 1536-0040.

Chu:2006:RMQ

[CC06b] Delin Chu and Moody Chu. Reachable matrices by a QR step with shift. SIAM Journal on Applied Dynamical Systems, 5(1):91–107, 2006. CO-DEN SJADAY. ISSN 1536-0040.

Chen:2010:SSS

[CCD^{+}10] Min Chen, Christopher W. Curtis, Bernard Deconinck, Crystal W. Lee, and Nghiem Nguyen. Spectral stability of stationary solutions of a Boussinesq system describing long waves in dispersive media. SIAM Journal on Applied Dynamical Systems, 9 (3):999–1018, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Cen:2020:SSP

[CCHZ20] Xiuli Cen, Xuhua Cheng, Zaitang Huang, and Meirong Zhang. On the stability of symmetric periodic orbits of the elliptic Sitnikov problem. SIAM Journal on Applied Dynamical Systems, 19(2):1271– 1290, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Caraballo:2024:CAI

[CCL24] Tomas Caraballo, Zhang Chen, and Lingyu Li. Convergence and approximation of invariant measures for neural field lattice models under noise perturbation. SIAM Journal on Applied Dynamical Systems, 23(1):358–382, January 2024. ISSN 1536-0040.

Colombo:2010:DIB

[CD10] Alessandro Colombo and Fabio Dercole. Discontinuity induced bifurcations of nonhyperbolic cycles in nonsmooth systems. SIAM Journal on Applied Dynamical Systems, 9(1):62–83, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Caginalp:2017:DDG

[CD17] Reginald J. Caginalp and Brent Doiron. Decision dynamics in groups with interacting members. SIAM Journal on Applied Dynamical Systems, 16(3):1543–1562, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Craciun:2020:ENT

[CD20] Gheorghe Craciun and Abhishek Deshpande. Endotactic networks and toric differential inclusions. SIAM Journal on Applied Dynamical Systems, 19(3):1798–1822, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Cho:2023:CCA

[CDH23] Hangjun Cho, Jiu-Gang Dong, and Seung-Yeal Ha. Cardinality of collisions in asymptotic phase-locking for the Kuramoto model with inertia. SIAM Journal on Applied Dynamical Systems, 22 (2):1472–1501, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1500010.

Carrillo:2019:SAL

[CDKS19] José A. Carrillo, Bertram Düring, Lisa Maria Kreusser, and Carola-Bibiane Schönlieb. Stability analysis of line patterns of an anisotropic interaction model. SIAM Journal on Applied Dynamical Systems, 18(4):1798–1845, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Carcamo-Diaz:2021:NSS

[CDPVY21] Daniela Cárcamo-Díaz, Jesús F. Palacián, Claudio Vidal, and Patricia Yanguas. Nonlinear stability in the spatial attitude motion of a satellite in a circular orbit. SIAM Journal on Applied Dynamical Systems, 20(3):1421–1463, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Cvitanovic:2010:SSG

[CDS10] Predrag Cvitanović, Ruslan L. Davidchack, and Evangelos Siminos. On the state space geometry of the Kuramoto–Sivashinsky flow in a periodic domain. SIAM Journal on Applied Dynamical Systems, 9(1):1–33, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Creedon:2021:HFI

[CDT21] Ryan Creedon, Bernard Deconinck, and Olga Trichtchenko. High-frequency instabilities of the Kawahara equation: a perturbative approach. SIAM Journal on Applied Dynamical Systems, 20(3):1571–1595, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Chao:2024:SNI

[CDW24] Ying Chao, Jinqiao Duan, and Pingyuan Wei. Smallnoise-induced metastable transition of periodically perturbed systems. SIAM Journal on Applied Dynamical Systems, 23(1):961–981, March 2024. ISSN 1536-0040.

Curtu:2004:PFN

[CE04] Rodica Curtu and Bard Ermentrout. Pattern formation in a network of excitatory and inhibitory cells with adaptation. SIAM Journal on Applied Dynamical Systems, 3 (3):191–231, 2004. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60050.

Cihak:2022:DEI

[CEK22] Heather L. Cihak, Tahra L.

Eissa, and Zachary P. Kilpatrick. Distinct excitatory and inhibitory bump wandering in a stochastic neural field. SIAM Journal on Applied Dynamical Systems, 21(4):2579–2609, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1482329.

Crespo:2024:RRE

[CEvdM24] F. Crespo, D. E. Espejo, and J. C. van der Meer. R^{3} × SO(3) × T-reduction, relative equilibria, and bifurcations for the full averaged model of two interacting rigid bodies. SIAM Journal on Applied Dynamical Systems, 23 (1):668–695, February 2024. ISSN 1536-0040.

Clement:2007:MMG

[CF07] Frédérique Clément and Jean-Pierre Françoise. Mathematical modeling of the GnRH pulse and surge generator. SIAM Journal on Applied Dynamical Systems, 6 (2):441–456, 2007. CODEN SJADAY. ISSN 1536-0040.

Conradi:2012:SMA

[CF12] Carsten Conradi and Dietrich Flockerzi. Switching in mass action networks based on linear inequalities. SIAM Journal on Applied Dynamical Systems, 11(1): 110–134, ???? 2012. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.

siam.org/siads/resource/
1/sjaday/v11/i1/p110_s1.

Crespo:2020:ARS

[CF20] Francisco Crespo and Sebastián Ferrer. Alternative reduction by stages of Keplerian systems. positive, negative, and zero energy. SIAM Journal on Applied Dynamical Systems, 19(2):1525–1539, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Castro:2022:SCG

[CFdSL22] Sofia B. S. D. Castro, Ana Ferreira, Liliana Garrido da Silva, and Isabel S. Labouriau. Stability of cycles in a game of rock–scissors– paper–lizard–spock. SIAM Journal on Applied Dynamical Systems, 21(4):2393–2431, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/21M1435215.

Chumley:2021:KDR

[CFG21] Timothy Chumley, Renato Feres, and Luis Alberto Garcia German. Knudsen diffusivity in random billiards: Spectrum, geometry, and computation. SIAM Journal on Applied Dynamical Systems, 20(3):1655–1682, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Cotter:2004:HPM

[CFR04] Colin J. Cotter, Jason Frank, and Sebastian Reich. Hamiltonian particle-mesh method for two-layer shallow-water equations subject to the rigidlid approximation. SIAM Journal on Applied Dynamical Systems, 3(1):69–83, 2004. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60007.

Carmona:2008:ERP

[CFST08] Victoriano Carmona, Fernando Fernández-Sánchez, and Antonio E. Teruel. Existence of a reversible T-point heteroclinic cycle in a piecewise linear version of the Michelson system. SIAM Journal on Applied Dynamical Systems, 7(3):1032–1048, 2008. CODEN SJADAY. ISSN 1536-0040.

Celletti:2018:DRE

[CG18] Alessandra Celletti and Catalin Gales. Dynamics of resonances and equilibria of low Earth objects. SIAM Journal on Applied Dynamical Systems, 17(1):203–235, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Caby:2024:TLE

[CG24] Théophile Caby and Pierre Guiraud. Transverse Lyapunov exponent and chimeras in globally coupled maps. SIAM Journal on Applied Dynamical Systems, 23(3): 1946–1965, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1603339.

Cummins:2022:ECR

[CGG^{+}22] Bree Cummins, Marcio Gameiro, Tomas Gedeon, Shane Kepley, Konstantin Mischaikow, and Lun Zhang. Extending combinatorial regulatory network modeling to include activity control and decay modulation. SIAM Journal on Applied Dynamical Systems, 21(3):2096–2125, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1456832.

Cummins:2016:CRP

[CGH^{+}16] Bree Cummins, Tomas Gedeon, Shaun Harker, Konstantin Mischaikow, and Kafung Mok. Combinatorial representation of parameter space for switching networks. SIAM Journal on Applied Dynamical Systems, 15(4):2176–2212, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Cummins:2018:MRP

[CGHM18] Bree Cummins, Tomas Gedeon, Shaun Harker, and Konstantin Mischaikow. Model rejection and parameter reduction via time series. SIAM Journal on Applied Dynamical Systems, 17(2):1589–1616, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Colonius:2008:NIM

[CGK08] Fritz Colonius, Tobias Gayer, and Wolfgang Kliemann. Near invariance for Markov diffusion systems. SIAM Journal on Applied Dynamical Systems, 7(1):79–107, 2008. CODEN SJADAY. ISSN 1536-0040.

Chartrand:2019:SEC

[CGL19] Thomas Chartrand, Mark S. Goldman, and Timothy J. Lewis. Synchronization of electrically coupled resonateand-fire neurons. SIAM Journal on Applied Dynamical Systems, 18(3):1643–1693, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Celletti:2016:BLS

[CGP16] Alessandra Celletti, Catalin Gales, and Giuseppe Pucacco. Bifurcation of lunisolar secular resonances for space debris orbits. SIAM Journal on Applied Dynamical Systems, 15(3):1352–1383, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Cummins:2015:ESS

[CGS15] Bree Cummins, Tomás Gedeon, and Kelly Spendlove. On the efficacy of state space reconstruction methods in determining causality. SIAM Journal on Applied Dynamical Systems, 14(1):335–381, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Cors:2003:CPO

[CH03] Josep M. Cors and Glen R.

Hall. Coorbital periodic orbits in the three body problem. SIAM Journal on Applied Dynamical Systems, 2 (2):219–237, 2003. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
41130.

Calogero:2010:OTS

[CH10] Simone Calogero and J. Mark Heinzle. Oscillations toward the singularity of locally rotationally symmetric Bianchi type IX cosmological models with Vlasov matter. SIAM Journal on Applied Dynamical Systems, 9(4):1244–1262, ???? 2010. CODEN SJA-DAY. ISSN 1536-0040.

Chicone:2013:PLL

[CH13] Carmen Chicone and Michael T.

Heitzman. Phase-locked loops, demodulation, and averaging approximation timescale extensions. SIAM Journal on Applied Dynamical Systems, 12(2):674–721, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Choi:2014:CEL

[CH14] Sun-Ho Choi and Seung-Yeal Ha. Complete entrainment of Lohe oscillators under attractive and repulsive couplings. SIAM Journal on Applied Dynamical Systems, 13(4):1417– 1441, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Collins:2023:RCR

[CHH23] B. W. Collins, C. L. Hall, and S. J. Hogan. The rocking can: a reduced equation of motion and a matched asymptotic solution. SIAM Journal on Applied Dynamical Systems, 22(4):3358–3389, December 2023. ISSN 1536-0040.

Chen:2024:CPS

[CHH24] Shih-Hsin Chen, Chun-Hsiung Hsia, and Ting-Yang Hsiao. Complete and partial synchronization of two-group and three-group Kuramoto oscillators. SIAM Journal on Applied Dynamical Systems, 23 (3):1720–1765, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1586227.

Chiba:2008:AVF

[Chi08] Hayato Chiba. C^{1} approximation of vector fields based on the renormalization group method. SIAM Journal on Applied Dynamical Systems, 7(3):895–932, 2008. CODEN SJADAY. ISSN 1536-0040.

Chiba:2009:EUS

[Chi09] Hayato Chiba. Extension and unification of singular perturbation methods for ODEs based on the renormalization group method. SIAM Journal on Applied Dynamical Systems, 8(3):1066–1115, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Chiba:2017:CMR

[Chi17] Hayato Chiba. A center manifold reduction of the Kuramoto–Daido model with a phase-lag. SIAM Journal on Applied Dynamical Systems, 16(3):1235–1259, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Calleja:2017:RPS

[CHK17] R. C. Calleja, A. R. Humphries, and B. Krauskopf. Resonance phenomena in a scalar delay differential equation with two state-dependent delays. SIAM Journal on Applied Dynamical Systems, 16(3):1474– 1513, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Cheesman:2022:GPP

[CHK22] Noah Cheesman, S. J. Hogan, and Kristian Uldall Kristiansen. The geometry of the Painlevé paradox. SIAM Journal on Applied Dynamical Systems, 21(3):1798–1831, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/21M1455590.

Chen:2017:TCM

[CHP17] Kuo-Chang Chen, Ku-Jung Hsu, and Bo-Yu Pan. A theory of central measures for celestial mechanics. SIAM Journal on Applied Dynamical Systems, 16(1):204–225, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Corradi:2012:EFD

[CHS12] Olivier Corradi, Poul G. Hjorth, and Jens Starke. Equation-free detection and continuation of a Hopf bifurcation point in a particle model of pedestrian flow. SIAM Journal on Applied Dynamical Systems, 11(3):1007– 1032, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Church:2021:APC

[Chu21] Kevin E. M. Church. Analysis of pandemic closingreopening cycles using rigorous homotopy continuation: a case study with Montreal COVID-19 data. SIAM Journal on Applied Dynamical Systems, 20(2):745–783, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Chiba:2023:GEP

[CII23] Hayato Chiba, Masahiro Ikeda, and Isao Ishikawa. Generalized eigenvalues of the Perron–Frobenius operators of symbolic dynamical systems. SIAM Journal on Applied Dynamical Systems, 22(4):2825–2855, October 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1476204.

Chang:2008:SSG

[CJ08] Yu-Chuan Chang and Jonq Juang. Stable synchrony in globally coupled integrateand-fire oscillators. SIAM Journal on Applied Dynamical Systems, 7(4):1445–1476, ???? 2008. CODEN SJA-DAY. ISSN 1536-0040.

Colombo:2011:NCT

[CJ11] Alessandro Colombo and Mike R. Jeffrey. Nondeterministic chaos, and the two-fold singularity in piecewise smooth flows. SIAM Journal on Applied Dynamical Systems, 10(2):423– 451, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i2/p423_s1.

Capinski:2017:VCH

[CJ17] Maciej J. Capiński and J. D. Mireles James. Validated computation of heteroclinic sets. SIAM Journal on Applied Dynamical Systems, 16(1):375–409, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Cappelletti:2018:GBE

[CJ18a] Daniele Cappelletti and Badal Joshi. Graphically balanced equilibria and stationary measures of reaction networks. SIAM Journal on Applied Dynamical Systems, 17 (3):2146–2175, ???? 2018.

CODEN SJADAY. ISSN 1536-0040.

Cornwell:2018:ESF

[CJ18b] Paul Cornwell and Christopher K. R. T. Jones. On the existence and stability of fast traveling waves in a doubly diffusive FitzHugh– Nagumo system. SIAM Journal on Applied Dynamical Systems, 17(1):754–787, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Contreras-Julio:2020:FSR

[CJAMV20] Dana Contreras-Julio, Pablo Aguirre, José Mujica, and Olga Vasilieva. Finding strategies to regulate propagation and containment of dengue via invariant manifold analysis. SIAM Journal on Applied Dynamical Systems, 19(2):1392–1437, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Chen:2015:LVE

[CJN15] Xiaojing Chen, Jifa Jiang, and Lei Niu. On Lotka– Volterra equations with identical minimal intrinsic growth rate. SIAM Journal on Applied Dynamical Systems, 14 (3):1558–1599, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Cannon:2015:LOP

[CK15] Jonathan Cannon and Nancy Kopell. The leaky oscillator: Properties of inhibition-based rhythms revealed through the singular phase response curve. SIAM Journal on Applied Dynamical Systems, 14 (4):1930–1977, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Crawford-Kahrl:2019:CCS

[CKCG19] Peter Crawford-Kahrl, Bree Cummins, and Tomas Gedeon. Comparison of combinatorial signatures of global network dynamics generated by two classes of ODE models. SIAM Journal on Applied Dynamical Systems, 18(1):418–457, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Cheesman:2021:RIC

[CKH21] Noah Cheesman, Kristian Uldall Kristiansen, and S. J. Hogan. Regularization of isolated codimension-2 discontinuity sets. SIAM Journal on Applied Dynamical Systems, 20(4):2630–2670, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Champneys:2007:WSM

[CKK^{+}07] Alan R. Champneys, Vivien Kirk, Edgar Knobloch, Bart E. Oldeman, and James Sneyd. When Shil’nikov meets Hopf in excitable systems. SIAM Journal on Applied Dynamical Systems, 6(4):663–693, 2007. CODEN SJADAY. ISSN 1536-0040.

Champneys:2009:UTE

[CKK^{+}09] Alan R. Champneys, Vivien Kirk, Edgar Knobloch, Bart E.

Oldeman, and J. D. M. Rademacher. Unfolding a tangent equilibrium-to-periodic heteroclinic cycle. SIAM Journal on Applied Dynamical Systems, 8(3):1261– 1304, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Champneys:2012:HSB

[CKMW12] A. R. Champneys, E. Knobloch, Y.-P. Ma, and T. Wagenknecht. Homoclinic snakes bounded by a saddle-center periodic orbit. SIAM Journal on Applied Dynamical Systems, 11(4):1583–1613, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Creaser:2017:FFF

[CKO17] Jennifer L. Creaser, Bernd Krauskopf, and Hinke M. Osinga. Finding first foliation tangencies in the Lorenz system. SIAM Journal on Applied Dynamical Systems, 16(4):2127–2164, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Choi:2019:CSC

[CKPP19] Young-Pil Choi, Dante Kalise, Jan Peszek, and Andrés A. Peters. A collisionless singular Cucker–Smale model with decentralized formation control. SIAM Journal on Applied Dynamical Systems, 18 (4):1954–1981, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Castro:2008:NEN

[CL08] Mario Castro and Grant Lythe. Numerical experiments on noisy chains: From collective transitions to nucleation-diffusion. SIAM Journal on Applied Dynamical Systems, 7(1):207–219, 2008. CODEN SJADAY. ISSN 1536-0040.

Cytrynbaum:2009:GBA

[CL09] Eric N. Cytrynbaum and Timothy J. Lewis. A global bifurcation and the appearance of a one-dimensional spiral wave in excitable media. SIAM Journal on Applied Dynamical Systems, 8(1):348– 370, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Chu:2011:DSC

[CL11] Moody T. Chu and Matthew M.

Lin. Dynamical system characterization of the central path and its variants — a revisit. SIAM Journal on Applied Dynamical Systems, 10 (3):887–905, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i3/p887_s1.

Castelli:2013:RNF

[CL13] Roberto Castelli and Jean-Philippe Lessard. Rigorous numerics in Floquet theory: Computing stable and unstable bundles of periodic orbits. SIAM Journal on Applied Dynamical Systems, 12(1):204– 245, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Charette:2014:LSB

[CL14] Laurent Charette and Victor G. LeBlanc. Lattice symmetry-breaking perturbations for spiral waves. SIAM Journal on Applied Dynamical Systems, 13(4):1694–1715, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Castro:2016:SHN

[CL16] Sofia B. S. D. Castro and Alexander Lohse. Switching in heteroclinic networks. SIAM Journal on Applied Dynamical Systems, 15(2):1085– 1103, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Colombo:2009:TPB

[CLBdB09] Alessandro Colombo, Paola Lamiani, Luis Benadero, and Mario di Bernardo. Twoparameter bifurcation analysis of the Buck converter. SIAM Journal on Applied Dynamical Systems, 8(4):1507– 1522, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Castelli:2015:PIM

[CLJ15] Roberto Castelli, Jean-Philippe Lessard, and J. D. Mireles James. Parameterization of invariant manifolds for periodic orbits I: Efficient numerics via the Floquet normal form. SIAM Journal on Applied Dynamical Systems, 14 (1):132–167, ???? 2015. CO-DEN SJADAY. ISSN 1536-0040.

Caraballo:2012:GID

[CLL12] Tomás Caraballo, José A.

Langa, and Zhenxin Liu. Gradient infinite-dimensional random dynamical systems. SIAM Journal on Applied Dynamical Systems, 11(4):1817– 1847, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Cadiot:2024:RCS

[CLN24] Matthieu Cadiot, Jean-Philippe Lessard, and Jean-Christophe Nave. Rigorous computation of solutions of semilinear PDEs on unbounded domains via spectral methods. SIAM Journal on Applied Dynamical Systems, 23(3):1966–2017, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1607507.

Christov:2014:STD

[CLOS14] Ivan C. Christov, Richard M.

Lueptow, Julio M. Ottino, and Rob Sturman. A study in three-dimensional chaotic dynamics: Granular flow and transport in a bi-axial spherical tumbler. SIAM Journal on Applied Dynamical Systems, 13(2):901–943, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Chen:2024:EDA

[CLW24] Shanshan Chen, Jie Liu, and Yixiang Wu. Evolution of dispersal in advective patchy environments with varying drift rates. SIAM Journal on Applied Dynamical Systems, 23 (1):696–720, February 2024. ISSN 1536-0040.

Cen:2021:PSC

[CLZ21] Xiuli Cen, Changjian Liu, and Meirong Zhang. A proof for a stability conjecture on symmetric periodic solutions of the elliptic Sitnikov problem. SIAM Journal on Applied Dynamical Systems, 20 (2):941–952, ???? 2021. CO-DEN SJADAY. ISSN 1536-0040.

Canic:2003:EEM

[CM03] Suncica Canic and Andro Mikelic. Effective equations modeling the flow of a viscous incompressible fluid through a long elastic tube arising in the study of blood flow through small arteries. SIAM Journal on Applied Dynamical Systems, 2(3):431–463, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41128.

Colijn:2007:BBM

[CM07] Caroline Colijn and Michael C. Mackey. Bifurcation and bistability in a model of hematopoietic regulation. SIAM Journal on Applied Dynamical Systems, 6(2):378– 394, 2007. CODEN SJADAY. ISSN 1536-0040.

Cargill:2016:IRN

[CM16] Daniel S. Cargill and Richard O.

Moore. Incorporating radiation in noise-induced phase evolution of optical solitons. SIAM Journal on Applied Dynamical Systems, 15(2):1025– 1061, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Cohen:2023:DLN

[CMV23] Nadav Cohen, Govind Menon, and Zsolt Veraszto. Deep linear networks for matrix completion — an infinite depth limit. SIAM Journal on Applied Dynamical Systems, 22(4):3208–3232, November 2023. ISSN 1536-0040.

Choksi:2011:PDM

[CMW11] Rustum Choksi, Mirjana Maras, and J. F. Williams. 2D phase diagram for minimizers of a Cahn–Hilliard functional with long-range interactions. SIAM Journal on Applied Dynamical Systems, 10(4):1344– 1362, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i4/p1344_s1.

Coombes:2004:EFI

[CO04] S. Coombes and M. R. Owen.

Evans functions for integral neural field equations with Heaviside firing rate function. SIAM Journal on Applied Dynamical Systems, 3(4):574– 600, 2004. CODEN SJADAY.

ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/60595.

Comanici:2006:TRW

[Com06] Adela N. Comanici. Transition from rotating waves to modulated rotating waves on the sphere. SIAM Journal on Applied Dynamical Systems, 5(4):759–782, 2006. CODEN SJADAY. ISSN 1536-0040.

Coombes:2008:NNG

[Coo08] S. Coombes. Neuronal networks with gap junctions: a study of piecewise linear planar neuron models. SIAM Journal on Applied Dynamical Systems, 7(3):1101–1129, 2008. CODEN SJADAY. ISSN 1536-0040.

Celik:2019:SFI

[COT19] Emine Celik, Eric Olson, and Edriss S. Titi. Spectral filtering of interpolant observables for a discrete-in-time downscaling data assimilation algorithm. SIAM Journal on Applied Dynamical Systems, 18(2):1118–1142, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Chugunova:2006:BDS

[CP06] Marina Chugunova and Dmitry Pelinovsky. Block-diagonalization of the symmetric first-order coupled-mode system. SIAM Journal on Applied Dynamical Systems, 5(1):66–83, 2006. CODEN SJADAY. ISSN 1536-0040.

Colombo:2012:CSM

[CP12] R. M. Colombo and N. Pogodaev. Confinement strategies in a model for the interaction between individuals and a continuum. SIAM Journal on Applied Dynamical Systems, 11(2):741–770, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Cytrynbaum:2017:IWN

[CP17] Eric N. Cytrynbaum and Kelly M. Paton. An invariant winding number for the FitzHugh–Nagumo system with applications to cardiac dynamics. SIAM Journal on Applied Dynamical Systems, 16(4):1893–1922, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Chu:2023:NMM

[CP23] Weiqi Chu and Mason A.

Porter. Non-Markovian models of opinion dynamics on temporal networks. SIAM Journal on Applied Dynamical Systems, 22(3):2624–2647, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M151858X.

Cors:2019:COQ

[CPY19] Josep M. Cors, Jesús F.

Palacián, and Patricia Yanguas. On co-orbital quasiperiodic motion in the threebody problem. SIAM Journal on Applied Dynamical Systems, 18(1):334–353, ????

2019. CODEN SJADAY. ISSN 1536-0040.

Conway:2009:SPC

[CR09] Jessica M. Conway and Hermann Riecke. Superlattice patterns in the complex Ginzburg–Landau equation with multiresonant forcing. SIAM Journal on Applied Dynamical Systems, 8 (3):977–1004, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Curtu:2011:ICS

[CR11] R. Curtu and J. Rubin. Interaction of Canard and singular Hopf mechanisms in a neural model. SIAM Journal on Applied Dynamical Systems, 10(4):1443– 1479, ???? 2011. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.
                                        siam.org/siads/resource/
                                        1/sjaday/v10/i4/p1443_s1.

Capinski:2012:ECM

[CR12] Maciej J. Capiński and Pablo Roldán. Existence of a center manifold in a practical domain around L_{1} in the restricted three-body problem. SIAM Journal on Applied Dynamical Systems, 11(1):285– 318, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Cappelletti:2023:IMD

[CR23] Daniele Cappelletti and Grzegorz A. Rempala. Individual molecules dynamics in reaction network models. SIAM Journal on Applied Dynamical Systems, 22(2):1344–1382, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/21M1459563.

Casas:2011:FMB

[CRR11] Pablo S. Casas and Rafael Ramírez-Ros. The frequency map for billiards inside ellipsoids. SIAM Journal on Applied Dynamical Systems, 10 (1):278–324, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i1/p278_s1.

Chickarmane:2007:MPD

[CRSN07] Vijay Chickarmane, Animesh Ray, Herbert M. Sauro, and Ali Nadim. A model for p53 dynamics triggered by DNA damage. SIAM Journal on Applied Dynamical Systems, 6(1):61–78, 2007. CODEN SJADAY. ISSN 1536-0040.

Carter:2018:UHB

[CS18] Paul Carter and Björn Sandstede. Unpeeling a homoclinic banana in the FitzHugh– Nagumo system. SIAM Journal on Applied Dynamical Systems, 17(1):236–349, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Ciocanel:2018:MMB

[CSJM18]  Maria-Veronica  Ciocanel,

Bjorn Sandstede, Samantha P. Jeschonek, and Kimberly L. Mowry. Modeling microtubule-based transport and anchoring of mRNA. SIAM Journal on Applied Dynamical Systems, 17(4):2855– 2881, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Carr:2006:DMC

[CSKR06] T. W. Carr, I. B. Schwartz, Min-Young Kim, and Rajarshi Roy. Delayed-mutual coupling dynamics of lasers: Scaling laws and resonances. SIAM Journal on Applied Dynamical Systems, 5(4):699– 725, 2006. CODEN SJADAY. ISSN 1536-0040.

Curtu:2008:MFC

[CSRR08] Rodica Curtu, Asya Shpiro, Nava Rubin, and John Rinzel. Mechanisms for frequency control in neuronal competition models. SIAM Journal on Applied Dynamical Systems, 7(2):609–649, 2008. CODEN SJADAY. ISSN 1536-0040.

Cher:2017:LSS

[CSS17] Yuri Cher, Gideon Simpson, and Catherine Sulem. Local structure of singular profiles for a derivative nonlinear Schrödinger equation. SIAM Journal on Applied Dynamical Systems, 16(1):514–545, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Cardin:2017:FTM

[CT17] Pedro Toniol Cardin and Marco Antonio Teixeira. Fenichel theory for multiple time scale singular perturbation problems. SIAM Journal on Applied Dynamical Systems, 16(3):1425–1452, ???? 2017. CODEN SJADAY. ISSN 1536-0040. See corrigendum [CT19].

Cardin:2019:CFT

[CT19] Pedro Toniol Cardin and Marco Antonio Teixeira. Corrigendum: Fenichel theory for multiple time scale singular perturbation problems. SIAM Journal on Applied Dynamical Systems, 18(2):1223, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040. See [CT17].

Crespo:2022:PSR

[CT22] Francisco Crespo and Edward A. Turner. Poisson structure and reduction by stages of the full gravitational N-body problem. SIAM Journal on Applied Dynamical Systems, 21(3):1778–1797, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/21M1416783.

Creaser:2018:SNI

[CTAA18] Jennifer Creaser, Krasimira Tsaneva-Atanasova, and Peter Ashwin. Sequential noiseinduced escapes for oscillatory network dynamics. SIAM Journal on Applied Dynamical Systems, 17(1):500– 525, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Clement:2009:FBP

[CV09] Frédérique Clément and Alexandre Vidal. Foliationbased parameter tuning in a model of the GnRH pulse and surge generator. SIAM Journal on Applied Dynamical Systems, 8(4):1591–1631, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Cuesta:2014:ESD

[CV14] C. M. Cuesta and J. J. L. Velázquez. Existence of solutions describing accumulation in a thin-film flow. SIAM Journal on Applied Dynamical Systems, 13(1):47– 93, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Chen:2011:SDL

[CW11] W. Chen and M. J. Ward. The stability and dynamics of localized spot patterns in the two-dimensional Gray–Scott model. SIAM Journal on Applied Dynamical Systems, 10 (2):582–666, ???? 2011. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.
                                        siam.org/siads/resource/
                                        1/sjaday/v10/i2/p582_s1.

Cappelletti:2017:UAS

[CW17] Daniele Cappelletti and Carsten Wiuf. Uniform approximation of solutions by elimination of intermediate species in deterministic reaction networks. SIAM Journal on Applied Dynamical Systems, 16(4):2259–2286, ????

2017. CODEN SJADAY. ISSN 1536-0040.

Cyranka:2018:CAP

[CW18] Jacek Cyranka and Thomas Wanner. Computer-assisted proof of heteroclinic connections in the one-dimensional Ohta–Kawasaki model. SIAM Journal on Applied Dynamical Systems, 17(1):694–731, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Chandramoorthy:2022:ECL

[CW22a] Nisha Chandramoorthy and Qiqi Wang. Efficient computation of linear response of chaotic attractors with onedimensional unstable manifolds. SIAM Journal on Applied Dynamical Systems, 21(2):735–781, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1405599.

Chen:2022:GAA

[CW22b] Changbo Chen and Wenyuan Wu. A geometric approach for analyzing parametric biological systems by exploiting block triangular structure. SIAM Journal on Applied Dynamical Systems, 21 (2):1573–1596, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1436373.

Capinski:2017:CAP

[CWZ17] Maciej J. Capiński and Anna Wasieczko-Zajac. Computerassisted proof of Shil’nikov homoclinics: With application to the Lorenz-84 model. SIAM Journal on Applied Dynamical Systems, 16(3):1453– 1473, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Cyranka:2015:EGA

[CZ15] Jacek Cyranka and Piotr Zgliczyński. Existence of globally attracting solutions for one-dimensional viscous Burgers equation with nonautonomous forcing — a computer assisted proof. SIAM Journal on Applied Dynamical Systems, 14(2):787–821, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Czechowski:2016:EPS

[CZ16] Aleksander Czechowski and Piotr Zgliczyński. Existence of periodic solutions of the FitzHugh–Nagumo equations for an explicit range of the small parameter. SIAM Journal on Applied Dynamical Systems, 15(3):1615–1655, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Cleveland:2023:QDM

[CZSV23] Electa Cleveland, Angela Zhu, Björn Sandstede, and Alexandria Volkening. Quantifying different modeling frameworks using topological data analysis: a case study with zebrafish patterns. SIAM Journal on Applied Dynamical Systems, 22(4):3233– 3266, November 2023. ISSN 1536-0040.

DeSmet:2012:CCC

[DA12] F. De Smet and D. Aeyels.

Clustering conditions and the cluster formation process in a dynamical model of multidimensional attracting agents. SIAM Journal on Applied Dynamical Systems, 11(1):392– 415, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Duncan:2019:FSD

[DAFM19] Jacob P. Duncan, Teresa Aubele-Futch, and Monica McGrath. A fast-slow dynamical system model of addiction: Predicting relapse frequency. SIAM Journal on Applied Dynamical Systems, 18(2):881–903, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Dawes:2008:LPF

[Daw08] J. H. P. Dawes. Localized pattern formation with a largescale mode: Slanted snaking. SIAM Journal on Applied Dynamical Systems, 7(1):186– 206, 2008. CODEN SJADAY. ISSN 1536-0040.

Dawes:2009:MLS

[Daw09] J. H. P. Dawes. Modulated and localized states in a finite domain. SIAM Journal on Applied Dynamical Systems, 8(3):909–930, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Dorfler:2011:CCK

[DB11] Florian Dörfler and Francesco Bullo. On the critical coupling for Kuramoto oscillators. SIAM Journal on Applied Dynamical Systems, 10(3):1070– 1099, ???? 2011. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.
                                        siam.org/siads/resource/
                                        1/sjaday/v10/i3/p1070_s1.

Donnell:2013:LGS

[DB13] Pete Donnell and Murad Banaji. Local and global stability of equilibria for a class of chemical reaction networks. SIAM Journal on Applied Dynamical Systems, 12(2):899– 920, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Ding:2016:PEA

[DC16] Xiong Ding and Predrag Cvitanović. Periodic eigendecomposition and its application to Kuramoto–Sivashinsky system. SIAM Journal on Applied Dynamical Systems, 15 (3):1434–1454, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

DeMaesschalck:2013:NCT

[DD13] P. De Maesschalck and M. Desroches. Numerical continuation techniques for planar slow-fast systems. SIAM Journal on Applied Dynamical Systems, 12(3):1159–1180, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Dercole:2011:TDB

[DDCK11] Fabio Dercole, Fabio Della Rossa, Alessandro Colombo, and Yuri A. Kuznetsov. Two degenerate boundary equilibrium bifurcations in planar Filippov systems. SIAM Journal on Applied Dynamical Systems, 10(4):1525– 1553, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i4/p1525_s1.

DeVille:2016:MCF

[DDDGZ16] Lee DeVille, Sairaj Dhople, Alejandro D. Domínguez-García, and Jiangmeng Zhang. Moment closure and finitetime blowup for piecewise deterministic Markov processes. SIAM Journal on Applied Dynamical Systems, 15(1):526– 556, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

DeWitte:2013:NPN

[DDGK13] V. De Witte, F. Della Rossa, W. Govaerts, and Yu. A. Kuznetsov. Numerical periodic normalization for Codim 2 bifurcations of limit cycles: Computational formulas, numerical implementation, and examples. SIAM Journal on Applied Dynamical Systems, 12(2):722–788, ???? 2013.

CODEN SJADAY. ISSN 1536-0040.

Detrixhe:2016:FEA

[DDMG16] Miles Detrixhe, Marion Doubeck, Jeff Moehlis, and Frédéric Gibou. A fast Eulerian approach for computation of global isochrons in high dimensions. SIAM Journal on Applied Dynamical Systems, 15(3):1501– 1527, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Degond:2022:BTS

[DDN22] Pierre Degond, Antoine Diez, and Mingye Na. Bulk topological states in a new collective dynamics model. SIAM Journal on Applied Dynamical Systems, 21(2):1455–1494, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1393935.

Dunbar:2022:EIM

[DDSW22] Oliver R. A. Dunbar, Andrew B. Duncan, Andrew M. Stuart, and Marie-Therese Wolfram. Ensemble inference methods for models with noisy and expensive likelihoods. SIAM Journal on Applied Dynamical Systems, 21 (2):1539–1572, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1410853.

Derks:2007:SAK

[DDvGS07] G. Derks, A. Doelman, S. A. van Gils, and H. Susanto.

Stability analysis of π-kinks in a 0-π Josephson junction. SIAM Journal on Applied Dynamical Systems, 6 (1):99–141, 2007. CODEN SJADAY. ISSN 1536-0040.

DeLeo:2003:NAN

[De 03] Roberto De Leo. Numerical analysis of the Novikov problem of a normal metal in a strong magnetic field. SIAM Journal on Applied Dynamical Systems, 2(4):517– 545, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40664.

DeSterck:2007:CPA

[De 07] H. De Sterck. Critical point analysis of transonic flow profiles with heat conduction. SIAM Journal on Applied Dynamical Systems, 6(3):645– 662, 2007. CODEN SJADAY. ISSN 1536-0040.

Drover:2006:PBE

[DE06] Jonathan D. Drover and G. Bard Ermentrout. Phase boundaries as electrically induced phosphenes. SIAM Journal on Applied Dynamical Systems, 5(4):529–551, 2006. CODEN SJADAY. ISSN 1536-0040.

Dunworth:2016:HOS

[DE16] Jeffrey B. Dunworth and G. Bard Ermentrout. Heterogeneity and oscillations in small swarms. SIAM Journal on Applied Dynamical Systems, 15(3):1455–1484, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Ding:2022:RWN

[DE22] Yujie Ding and Bard Ermentrout. Rotating waves of nonlocally coupled oscillators on the annulus. SIAM Journal on Applied Dynamical Systems, 21(3):2047–2079, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1452202.

DelBuono:2014:EBS

[DEL14] Nicoletta Del Buono, Cinzia Elia, and Luciano Lopez. On the equivalence between the sigmoidal approach and Utkin’s approach for piecewise-linear models of gene regulatory networks. SIAM Journal on Applied Dynamical Systems, 13(3):1270– 1292, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Desi:2011:DNS

[DEP^{+}11] Jonathan P. Desi, Hanein H. Edrees, Joseph J. Price, Evelyn Sander, and Thomas Wanner. The dynamics of nucleation in stochastic Cahn–Morral systems. SIAM Journal on Applied Dynamical Systems, 10(2): 707–743, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i2/p707_s1.

Deshpande:2023:SOR

[Des23] Abhishek Deshpande. Sourceonly realizations, weakly reversible deficiency one networks, and dynamical equivalence. SIAM Journal on Applied Dynamical Systems, 22 (2):1502–1521, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1494932.

Dumas:2004:FOA

[DEV04] H. Scott Dumas, James A.

Ellison, and Mathias Vogt. First-order averaging principles for maps with applications to accelerator beam dynamics. SIAM Journal on Applied Dynamical Systems, 3(4):409–432, 2004. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60043.

Day:2019:SSC

[DF19] Sarah Day and Rafael Frongillo.

Sofic shifts via Conley index theory: Computing lower bounds on recurrent dynamics for maps. SIAM Journal on Applied Dynamical Systems, 18(3):1610–1642, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Day:2008:ARE

[DFT08] Sarah Day, Rafael Frongillo, and Rodrigo Treviño. Algorithms for rigorous entropy bounds and symbolic dynamics. SIAM Journal on Applied Dynamical Systems, 7 (4):1477–1506, ???? 2008. CODEN SJADAY. ISSN 1536-0040.

Derks:2005:RNM

[DG05] Gianne Derks and Georg A. Gottwald. A robust numerical method to study oscillatory instability of gap solitary waves. SIAM Journal on Applied Dynamical Systems, 4 (1):140–158, 2005. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60530.

Delshams:2016:ESS

[DGG16] Amadeu Delshams, Marina Gonchenko, and Pere Gutiérrez. Exponentially small splitting of separatrices and transversality associated to whiskered tori with quadratic frequency ratio. SIAM Journal on Applied Dynamical Systems, 15(2):981– 1024, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Duncan:2021:ETS

[DGK^{+}21] William Duncan, Tomas Gedeon, Hiroshi Kokubu, Konstantin Mischaikow, and Hiroe Oka. Equilibria and their stability in networks with steep sigmoidal nonlinearities. SIAM Journal on Applied Dynamical Systems, 20(4):2108–2141, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Diekman:2012:RDG

[DGMW12] Casey Diekman, Martin Golubitsky, Tyler McMillen, and Yunjiao Wang. Reduction and dynamics of a generalized rivalry network with two learned patterns. SIAM Journal on Applied Dynamical Systems, 11(4):1270–1309, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

DeSouza:2019:DMH

[DH19] Daniel C. De Souza and A. R.

Humphries. Dynamics of a mathematical hematopoietic stem-cell population model. SIAM Journal on Applied Dynamical Systems, 18(2):808– 852, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Dong:2020:EBK

[DHK20] Jiu-Gang Dong, Seung-Yeal Ha, and Doheon Kim. Emergent behaviors of the Kuramoto model with a time delay on a general digraph. SIAM Journal on Applied Dynamical Systems, 19(1):304– 328, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Day:2005:RNG

[DHMO05] Sarah Day, Yasuaki Hiraoka, Konstantin Mischaikow, and Toshiyuki Ogawa. Rigorous numerics for global dynamics: a study of the Swift– Hohenberg equation. SIAM Journal on Applied Dynamical Systems, 4(1):1–31, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60447.

Diamantakis:2023:VPG

[DHP23] Theo Diamantakis, Darryl D. Holm, and Grigorios A. Pavliotis. Variational principles on geometric rough paths and the Lévy area correction. SIAM Journal on Applied Dynamical Systems, 22 (2):1182–1218, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1522164.

Deng:2022:SAC

[DHW22] Yiyang Deng, Marshall Hampton, and Zhiqiang Wang. Symmetry and asymmetry in the 1 + N coorbital problem. SIAM Journal on Applied Dynamical Systems, 21 (3):2080–2095, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1475053.

Desoeuvres:2024:CAP

[DIL^{+}24a] Aurélien Desoeuvres, Alexandru Iosif, Christoph Lüders, Ovidiu Radulescu, Hamid Rahkooy, Matthias Seiß, and Thomas Sturm. A computational approach to polynomial conservation laws. SIAM Journal on Applied Dynamical Systems, 23(1):813–854, March 2024. ISSN 1536-0040.

Desoeuvres:2024:RCR

[DIL^{+}24b] Aurélien Desoeuvres, Alexandru Iosif, Christoph Lüders, Ovidiu Radulescu, Hamid Rahkooy, Matthias Seiß, and Thomas Sturm. Reduction of chemical reaction networks with approximate conservation laws. SIAM Journal on Applied Dynamical Systems, 23(1):256–296, January 2024. ISSN 1536-0040.

Dubois:2022:DOR

[DJCJC22] J. Dubois, M. Jorba-Cuscó, À. Jorba, and C. Chandre. Dynamical organization of recollisions by a family of invariant tori. SIAM Journal on Applied Dynamical Systems, 21(1):523–541, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1400912.

Delabays:2019:KMO

[DJD19] Robin Delabays, Philippe Jacquod, and Florian Dörfler. The Kuramoto model on oriented and signed graphs. SIAM Journal on Applied Dynamical Systems, 18(1):458– 480, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Dey:2019:PHM

[DJK^{+}19] Tamal K. Dey, Mateusz Juda, Tomasz Kapela, Jacek Kubica, Michał Lipiński, and Marian Mrozek. Persistent homology of Morse decompositions in combinatorial dynamics. SIAM Journal on Applied Dynamical Systems, 18(1):510–530, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Day:2004:RNM

[DJM04] S. Day, O. Junge, and K. Mischaikow. A rigorous numerical method for the global analysis of infinitedimensional discrete dynamical systems. SIAM Journal on Applied Dynamical Systems, 3(2):117–160, 2004. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60021.

Doelman:2003:SPI

[DK03] Arjen Doelman and Tasso J. Kaper. Semistrong pulse interactions in a class of coupled reaction–diffusion equations. SIAM Journal on Applied Dynamical Systems, 2 (1):53–96, 2003. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
40571.

Desroches:2018:SAC

[DK18] Mathieu Desroches and Vivien Kirk. Spike-adding in a canonical three-time-scale model: Superslow explosion and folded-saddle canards. SIAM Journal on Applied Dynamical Systems, 17(3):1989– 2017, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Deniz:2008:CDE

[DKaK^{+}08] Ali Deniz, Judy Kennedy, ahin Koçak, Andrei V. Ratiu, Cevat Üstün, and James A. Yorke. Chaotic n-dimensional Euclidean and hyperbolic open billiards and chaotic spinning planar billiards. SIAM Journal on Applied Dynamical Systems, 7(2):421– 436, 2008. CODEN SJADAY. ISSN 1536-0040.

Dietrich:2016:NMC

[DKB16] Felix Dietrich, Gerta Köster, and Hans-Joachim Bungartz. Numerical model construction with closed observables. SIAM Journal on Applied Dynamical Systems, 15(4):2078– 2108, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Dittus:2023:DSB

[DKB^{+}23] Anna Dittus, Niklas Kruse, Ingo Barke, Sylvia Speller, and Jens Starke. Detecting stability and bifurcation points in control-based continuation for a physical experiment of the Zeeman catastrophe machine. SIAM Journal on Applied Dynamical Systems, 22(2):1275–1299, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1503245.

Desroches:2008:GSM

[DKO08] M. Desroches, B. Krauskopf, and H. M. Osinga. The geometry of slow manifolds near a folded node. SIAM Journal on Applied Dynamical Systems, 7(4):1131–1162, 2008. CODEN SJADAY. ISSN 1536-0040.

Dipoppa:2012:SSF

[DKTG12] M. Dipoppa, M. Krupa, A. Torcini, and B. S. Gutkin. Splay states in finite pulsecoupled networks of excitable neurons. SIAM Journal on Applied Dynamical Systems, 11(3):864–894, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Dellnitz:2017:SON

[DKZ17] Michael Dellnitz, Stefan Klus, and Adrian Ziessler. A setoriented numerical approach for dynamical systems with parameter uncertainty. SIAM Journal on Applied Dynamical Systems, 16(1):120–138, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Dawes:2010:LSM

[DL10] J. H. P. Dawes and S. Lilley. Localized states in a model of pattern formation in a vertically vibrated layer. SIAM Journal on Applied Dynamical Systems, 9(1):238– 260, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Dai:2021:GLP

[DL21] Jia-Yuan Dai and Phillipo Lappicy. Ginzburg–Landau patterns in circular and spherical geometries: Vortices, spirals, and attractors.

SIAM Journal on Applied Dynamical Systems, 20(4):1959– 1984, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Dodson:2022:WRE

[DL22] Stephanie Dodson and Timothy J. Lewis. Wave reflections in excitable media linked to existence and stability of one-dimensional spiral waves. SIAM Journal on Applied Dynamical Systems, 21(2):1631–1659, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1425025.

deLeeuw:2022:SBD

[dLD22] Bart M. de Leeuw and Svetlana Dubinkina. Shadowingbased data assimilation method for partially observed models. SIAM Journal on Applied Dynamical Systems, 21 (2):879–902, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/18M1223897.

deLeeuw:2018:PSB

[dLDF^{+}18] Bart de Leeuw, Svetlana Dubinkina, Jason Frank, Andrew Steyer, Xuemin Tu, and Erik Van Vleck. Projected shadowing-based data assimilation. SIAM Journal on Applied Dynamical Systems, 17(4):2446–2477, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

delaLlave:2016:COC

[dlLJ16] R. de la Llave and J. D. Mireles James. Connecting orbits for compact infinite dimensional maps: Computer assisted proofs of existence. SIAM Journal on Applied Dynamical Systems, 15(2):1268– 1323, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

delaLlave:2019:GPL

[dlLK19] Rafael de la Llave and Florian Kogelbauer. Global persistence of Lyapunov subcenter manifolds as spectral submanifolds under dissipative perturbations. SIAM Journal on Applied Dynamical Systems, 18(4):2099–2142, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

delaLlave:2012:IMA

[dlLL12] Rafael de la Llave and Héctor E. Lomelí. Invariant manifolds for analytic difference equations. SIAM Journal on Applied Dynamical Systems, 11(4):1614–1651, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Deng:2021:NWN

[DLP21] Guo Deng, Christopher J. Lustri, and Mason A. Porter. Nanoptera in weakly nonlinear woodpile chains and diatomic granular chains. SIAM Journal on Applied Dynamical Systems, 20(4):2412– 2449, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Dudley:2019:SBA

[DLRB19] Harry J. Dudley, Lu Lu, Zhiyong Jason Ren, and David M. Bortz. Sensitivity and bifurcation analysis of a differential–algebraic equation model for a microbial electrolysis cell. SIAM Journal on Applied Dynamical Systems, 18(2):709–728, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Dullin:2009:QVP

[DM09] H. R. Dullin and J. D. Meiss.

Quadratic volume-preserving maps: Invariant circles and bifurcations. SIAM Journal on Applied Dynamical Systems, 8(1):76–128, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Dullin:2012:RTV

[DM12] H. R. Dullin and J. D.

Meiss. Resonances and twist in volume-preserving mappings. SIAM Journal on Applied Dynamical Systems, 11 (1):319–349, ???? 2012. CO-DEN SJADAY. ISSN 1536-0040.

Dilao:2020:PPO

[DM20] Rui Dilão and Manuel Murteira. Principal periodic orbits of the Keplerian dumbbell system. SIAM Journal on Applied Dynamical Systems, 19(1):181–207, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Dean:2015:ODP

[DMCK15] A. D. Dean, P. C. Matthews, S. M. Cox, and J. R. King. Orientation-dependent pinning and homoclinic snaking on a planar lattice. SIAM Journal on Applied Dynamical Systems, 14(1):481–521, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Dullin:2005:SCR

[DMS05] H. R. Dullin, J. D. Meiss, and D. G. Sterling. Symbolic codes for rotational orbits. SIAM Journal on Applied Dynamical Systems, 4(3):515– 562, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/61287.

Dey:2022:PCM

[DMS22] Tamal K. Dey, Marian Mrozek, and Ryan Slechta. Persistence of Conley–Morse graphs in combinatorial dynamical systems. SIAM Journal on Applied Dynamical Systems, 21(2):817–839, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M143162X.

Dieu:2016:CAB

[DNDY16] N. T. Dieu, D. H. Nguyen, N. H. Du, and G. Yin. Classification of asymptotic behavior in a stochastic SIR model. SIAM Journal on Applied Dynamical Systems, 15(2):1062– 1084, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Duenas:2023:CTD

[DNO23a] Jesús Dueñas, Carmen Núñez, and Rafael Obaya. Critical transitions in d-concave nonautonomous scalar ordinary differential equations appearing in population dynamics. SIAM Journal on Applied Dynamical Systems, 22 (4):2649–2692, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1542830.

Duenas:2023:CTC

[DNO23b] Jesús Dueñas, Carmen Núñez, and Rafael Obaya. Critical transitions in d-concave nonautonomous scalar ordinary differential equations appearing in population dynamics. SIAM Journal on Applied Dynamical Systems, 22 (4):2649–2692, October 2023. ISSN 1536-0040.

Dong:2024:FNI

[Don24] Jiu-Gang Dong. Flocks with nonlinear inherent dynamics under fixed and switching digraphs. SIAM Journal on Applied Dynamical Systems, 23(2):1242–1271, May 2024. ISSN 1536-0040.

Dohnal:2008:SGS

[DP08] Tomáš Dohnal and Dmitry Pelinovsky. Surface gap solitons at a nonlinearity interface. SIAM Journal on Applied Dynamical Systems, 7 (2):249–264, 2008. CODEN SJADAY. ISSN 1536-0040.

Dias:2009:SPP

[DP09] Ana Paula S. Dias and Eliana Manuel Pinho. Spatially periodic patterns of synchrony in lattice networks. SIAM Journal on Applied Dynamical Systems, 8(2):641– 675, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Dunmyre:2010:OID

[DR10] Justin R. Dunmyre and Jonathan E. Rubin. Optimal intrinsic dynamics for bursting in a three-cell network. SIAM Journal on Applied Dynamical Systems, 9(1):154– 187, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Djokam:2022:GMF

[DR22] Guy A. Djokam and Muruhan Rathinam. A generalized model of flocking with steering. SIAM Journal on Applied Dynamical Systems, 21(2):1352–1381, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1398793.

Day:2009:CBT

[DRC09] Judy Day, Jonathan E. Rubin, and Carson C. Chow. Competition between transients in the rate of approach to a fixed point. SIAM Journal on Applied Dynamical Systems, 8(4):1523–1563, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Doelman:2018:DMP

[DRdRV18] Arjen Doelman, Jens Rademacher, Björn de Rijk, and Frits Veerman. Destabilization mechanisms of periodic pulse patterns near a homoclinic limit. SIAM Journal on Applied Dynamical Systems, 17(2):1833– 1890, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Drubi:2019:UBC

[DRH19] Fatima Drubi and Alfonso Ruiz-Herrera. Understanding bacterial cheating: Biological and practical implications. SIAM Journal on Applied Dynamical Systems, 18(4):2303– 2324, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Dodson:2019:DSP

[DS19] Stephanie Dodson and Björn Sandstede. Determining the source of period-doubling instabilities in spiral waves. SIAM Journal on Applied Dynamical Systems, 18(4):2202– 2226, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Darabi:2023:CBT

[DS23] Atefe Darabi and Milad Siami. Centrality-based traffic restriction in delayed epidemic networks. SIAM Journal on Applied Dynamical Systems, 22(4):3165–3207, November 2023. ISSN 1536-

Das:2024:ETD

[DS24] Haridas K. Das and Lucas M. Stolerman. Epidemic thresholds and disease dynamics in metapopulations: The role of network structure and human mobility. SIAM Journal on Applied Dynamical Systems, 23(2):1579–1609, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1579522.

DeSantis:2012:NDS

[DSC12] M. DeSantis, D. Swigon, and G. Caginalp. Nonlinear dynamics and stability in a multigroup asset flow model. SIAM Journal on Applied Dynamical Systems, 11(3):1114– 1148, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Dey:2024:CCM

[DsMS24] Tamal K. Dey, Micha Lipi ski, Marian Mrozek, and Ryan Slechta. Computing connection matrices via persistencelike reductions. SIAM Journal on Applied Dynamical Systems, 23(1):81–97, January 2024. ISSN 1536-0040.

Dekker:2013:SPT

[DT13] Anthony H. Dekker and Richard Taylor. Synchronization properties of trees in the Kuramoto model. SIAM Journal on Applied Dynamical Systems, 12(2):596–617, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Dullin:2016:TS

[DT16] Holger R. Dullin and William Tong. Twisting somersault. SIAM Journal on Applied Dynamical Systems, 15(4):1806– 1822, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Dsilva:2016:DDR

[DTG^{+}16] Carmeline J. Dsilva, Ronen Talmon, C. William Gear, Ronald R. Coifman, and Ioannis G. Kevrekidis. Datadriven reduction for a class of multiscale fast–slow stochastic dynamical systems. SIAM Journal on Applied Dynamical Systems, 15(3):1327–1351, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Dietrich:2020:KOA

[DTK20] Felix Dietrich, Thomas N.

Thiem, and Ioannis G. Kevrekidis. On the Koopman operator of algorithms. SIAM Journal on Applied Dynamical Systems, 19(2):860–885, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Delgado:2019:DPR

[DV19] Joaquín Delgado and Claudio Vidal. Dynamics of the parabolic restricted collinear three-body problem. SIAM Journal on Applied Dynamical Systems, 18(1):172–204, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Doelman:2002:HSP

[DvdP02] Arjen Doelman and Harmen van der Ploeg. Homoclinic stripe patterns. SIAM Journal on Applied Dynamical Systems, 1(1):65–104, 2002. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39283.

Diekmann:2009:CRE

[DvG09] O. Diekmann and S. A. van Gils. On the cyclic replicator equation and the dynamics of semelparous populations. SIAM Journal on Applied Dynamical Systems, 8(3):1160– 1189, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Doelman:2016:GAS

[DvHX16] A. Doelman, P. van Heijster, and F. Xie. A geometric approach to stationary defect solutions in one space dimension. SIAM Journal on Applied Dynamical Systems, 15 (2):655–712, ???? 2016. CO-DEN SJADAY. ISSN 1536-0040.

deWiljes:2018:LTS

[dWRS18] Jana de Wiljes, Sebastian Reich, and Wilhelm Stannat. Long-time stability and accuracy of the ensemble Kalman– Bucy filter for fully observed processes and small measurement noise. SIAM Journal on Applied Dynamical Systems, 17(2):1152–1181, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

deWolff:2021:PAH

[dWSLD21] B. A. J. de Wolff, F. Scarabel, S. M. Verduyn Lunel, and O. Diekmann. Pseudospectral approximation of Hopf bifurcation for delay differential equations. SIAM Journal on Applied Dynamical Systems, 20(1):333–370, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Das:2017:MQ

[DY17] Suddhasattwa Das and James A.

Yorke. Multichaos from quasiperiodicity. SIAM Journal on Applied Dynamical Systems, 16(4):2196–2212, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Dyson:2020:EUT

[Dys20] Alan Dyson. Existence and uniqueness of traveling fronts in lateral inhibition neural fields with sigmoidal firing rates. SIAM Journal on Applied Dynamical Systems, 19 (3):2194–2231, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

DeVille:2014:SPE

[DZ14] Lee DeVille and Yi Zeng.

Synchrony and periodicity in excitable neural networks with multiple subpopulations. SIAM Journal on Applied Dynamical Systems, 13(3):1060– 1081, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Ersoz:2016:CMS

[EDKC16] Elif Köksal Ersöz, Mathieu Desroches, Martin Krupa, and Frédérique Clément. Canard-mediated (de)synchronization in coupled phantom bursters. SIAM Journal on Applied Dynamical Systems, 15(1):580– 608, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Evers:2017:EAM

[EFK17] Joep H. M. Evers, Razvan C. Fetecau, and Theodore Kolokolnikov. Equilibria for an aggregation model with two species. SIAM Journal on Applied Dynamical Systems, 16(4):2287–2338, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Etnyre:2005:GHI

[EG05] John Etnyre and Robert Ghrist. Generic hydrodynamic instability of curl eigenfields. SIAM Journal on Applied Dynamical Systems, 4(2):377–390, 2005. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60651.

Elmhirst:2006:NHB

[EG06] Toby Elmhirst and Martin Golubitsky. Nilpotent Hopf bifurcations in coupled cell systems. SIAM Journal on Applied Dynamical Systems, 5(2):205–251, 2006. CODEN SJADAY. ISSN 1536-0040.

Estrada:2018:LRI

[EGF18] Ernesto Estrada, Lucia Valentina Gambuzza, and Mattia Frasca. Long-range interactions and network synchronization. SIAM Journal on Applied Dynamical Systems, 17(1):672–693, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Eigel:2024:LIF

[EGS24] Martin Eigel, Robert Gruhlke, and David Sommer. Less interaction with forward models in Langevin dynamics: Enrichment and homotopy. SIAM Journal on Applied Dynamical Systems, 23 (3):1870–1908, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1546841.

Escher:2015:TWT

[EHLW15] Joachim Escher, Matthieu Hillairet, Philippe Laurençot, and Christoph Walker. Traveling waves for a thin film with gravity and insoluble surfactant. SIAM Journal on Applied Dynamical Systems, 14(4):1991–2012, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Eldering:2016:RSD

[EJ16] Jaap Eldering and Henry O.

Jacobs. The role of symmetry and dissipation in biolocomotion. SIAM Journal on Applied Dynamical Systems, 15 (1):24–59, ???? 2016. CO-DEN SJADAY. ISSN 1536-0040.

Engel:2020:ESL

[EJK20] Maximilian Engel and Hildeberto Jardón-Kojakhmetov. Extended and symmetric loss of stability for canards in planar fast-slow maps. SIAM Journal on Applied Dynamical Systems, 19(4):2530–2566, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Earnshaw:2010:GAS

[EK10a] Berton A. Earnshaw and James P. Keener. Global asymptotic stability of solutions of nonautonomous master equations. SIAM Journal on Applied Dynamical Systems, 9(1):220–237, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Earnshaw:2010:IMB

[EK10b] Berton A. Earnshaw and James P. Keener. Invariant manifolds of binomial-like nonautonomous master equations. SIAM Journal on Applied Dynamical Systems, 9 (2):568–588, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Evers:2016:MSA

[EK16] Joep H. M. Evers and Theodore Kolokolnikov. Metastable states for an aggregation model with noise. SIAM Journal on Applied Dynamical Systems, 15(4):2213–2226, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Erzgraber:2006:CLM

[EKL06] Hartmut Erzgräber, Bernd Krauskopf, and Daan Lenstra. Compound laser modes of mutually delay-coupled lasers. SIAM Journal on Applied Dynamical Systems, 5(1):30–65, 2006. CODEN SJADAY. ISSN 1536-0040.

Erzgraber:2007:BAS

[EKL07] Hartmut Erzgräber, Bernd Krauskopf, and Daan Lenstra. Bifurcation analysis of a semiconductor laser with filtered optical feedback. SIAM Journal on Applied Dynamical Systems, 6(1):1–28, 2007. CODEN SJADAY. ISSN 1536-0040.

England:2004:COD

[EKO04] J. P. England, B. Krauskopf, and H. M. Osinga. Computing one-dimensional stable manifolds and stable sets of planar maps without the inverse. SIAM Journal on Applied Dynamical Systems, 3 (2):161–190, 2004. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60013.

England:2005:COD

[EKO05] J. P. England, B. Krauskopf, and H. M. Osinga. Computing one-dimensional global manifolds of Poincaré maps by continuation. SIAM Journal on Applied Dynamical Systems, 4(4):1008–1041, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/62408.

Efstathiou:2015:BHF

[ELB15] Konstantinos Efstathiou, Xia Liu, and Henk W. Broer. The boundary-Hopf-fold bifurcation in Filippov systems. SIAM Journal on Applied Dynamical Systems, 14(2):914– 941, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Ercolani:2022:DNS

[ELT22] Nicholas Ercolani, Joceline Lega, and Brandon Tippings. Dynamics of nonpolar solutions to the discrete Painlevé I equation. SIAM Journal on Applied Dynamical Systems, 21(2):1322–1351, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1445156.

Erichson:2019:RDM

[EMKB19] N. Benjamin Erichson, Lionel Mathelin, J. Nathan Kutz, and Steven L. Brunton. Randomized dynamic mode decomposition. SIAM Journal on Applied Dynamical Systems, 18(4):1867–1891, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Erdogan:2015:SGA

[EMNT15] M. B. Erdogan, J. L. Marzuola, K. Newhall, and N. Tzirakis. The structure of global attractors for dissipative Zakharov systems with forcing on the torus. SIAM Journal on Applied Dynamical Systems, 14(4):1978–1990, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Eyres:2005:GBC

[EPCL05] R. D. Eyres, P. T. Piiroinen, A. R. Champneys, and N. A. J. Lieven. Grazing bifurcations and chaos in the dynamics of a hydraulic damper with relief valves. SIAM Journal on Applied Dynamical Systems, 4 (4):1076–1106, 2005. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
61969.

Escudero:2011:COC

[ERT11] Carlos Escudero, Andrés M.

Rivera, and Pedro J. Torres. Chemical oscillations out of chemical noise. SIAM Journal on Applied Dynamical Systems, 10(3):960– 986, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i3/p960_s1.

Erhardt:2022:BAM

[ES22] André H. Erhardt and Susanne Solem. Bifurcation analysis of a modified cardiac cell model. SIAM Journal on Applied Dynamical Systems, 21(1):231–247, ????

2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1425359.

Ebers:2024:DMF

[ESK24] Megan R. Ebers, Katherine M. Steele, and J. Nathan Kutz. Discrepancy modeling framework: Learning missing physics, modeling systematic residuals, and disambiguating between deterministic and random effects. SIAM Journal on Applied Dynamical Systems, 23(1):440– 469, January 2024. ISSN 1536-0040.

Efstathiou:2004:ARV

[ESZ04] K. Efstathiou, D. A. Sadovskii, and B. I. Zhilinskii. Analysis of rotation–vibration relative equilibria on the example of a tetrahedral four atom molecule. SIAM Journal on Applied Dynamical Systems, 3(3):261–351, 2004. CODEN SJADAY. ISSN 1536-0040.

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        60001.

Eilertsen:2021:HOV

[ETS21] Justin Eilertsen, Małgorzata Anna Tyczynska, and Santiago Schnell. Hunting ε: The origin and validity of quasisteady-state reductions in enzyme kinetics. SIAM Journal on Applied Dynamical Systems, 20(4):2450–2481, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Ersoz:2018:CMT

[EVC18] Elif Köksal Ersöz, Alexandre Vidal, and Frédérique Clément. Coupled multiple timescale dynamics in populations of endocrine neurons: Pulsatile and surge patterns of GnRH secretion. SIAM Journal on Applied Dynamical Systems, 17(1):1052–1090, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Ermentrout:2009:CCS

[EW09] Bard Ermentrout and Martin Wechselberger. Canards, clusters, and synchronization in a weakly coupled interneuron model. SIAM Journal on Applied Dynamical Systems, 8(1):253–278, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Eckhoff:2011:DRD

[EWLH11] Philip Eckhoff, KongFatt Wong-Lin, and Philip Holmes. Dimension reduction and dynamics of a spiking neural network model for decision making under neuromodulation. SIAM Journal on Applied Dynamical Systems, 10 (1):148–188, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i1/p148_s1.

Ferreira:2013:GPA

[FA13] Ana S. Rufino Ferreira and Murat Arcak. A graph partitioning approach to predicting patterns in lateral inhibition systems. SIAM Journal on Applied Dynamical Systems, 12(4):2012–2031, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Faye:2013:EST

[Fay13] Grégory Faye. Existence and stability of traveling pulses in a neural field equation with synaptic depression. SIAM Journal on Applied Dynamical Systems, 12(4):2032–2067, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Folias:2004:BPE

[FB04] Stefanos E. Folias and Paul C. Bressloff. Breathing pulses in an excitatory neural network. SIAM Journal on Applied Dynamical Systems, 3(3):378– 407, 2004. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/60262.

Figueras:2017:NCC

[FdlL17] Jordi-Lluís Figueras and Rafael de la Llave. Numerical computations and computer assisted proofs of periodic orbits of the Kuramoto– Sivashinsky equation. SIAM Journal on Applied Dynamical Systems, 16(2):834–852, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Franci:2012:OCP

[FDS12] Alessio Franci, Guillaume Drion, and Rodolphe Sepulchre. An organizing center in a planar model of neuronal excitability. SIAM Journal on Applied Dynamical Systems, 11(4):1698–1722, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Franci:2014:MMN

[FDS14] Alessio Franci, Guillaume Drion, and Rodolphe Sepulchre. Modeling the modulation of neuronal bursting: a singularity theory approach. SIAM Journal on Applied Dynamical Systems, 13(2):798– 829, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Folias:2010:SLS

[FE10] Stefanos E. Folias and G. Bard Ermentrout. Spatially localized synchronous oscillations in synaptically coupled neuronal networks: Conductance-based models and discrete maps. SIAM Journal on Applied Dynamical Systems, 9(3):1019–1060, ???? 2010. CODEN SJA-DAY. ISSN 1536-0040.

Folias:2012:BSS

[FE12] Stefanos E. Folias and G. Bard Ermentrout. Bifurcations of stationary solutions in an interacting pair of E-I neural fields. SIAM Journal on Applied Dynamical Systems, 11(3):895–938, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Foxall:2012:CAT

[FEIvdD12] Eric Foxall, Roderick Edwards, Slim Ibrahim, and P. van den Driessche. A contraction argument for twodimensional spiking neuron models. SIAM Journal on Applied Dynamical Systems, 11(1):540–566, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Ferguson:2018:TSK

[Fer18] Timothy Ferguson. Topological states in the Kuramoto model. SIAM Journal on Applied Dynamical Systems, 17 (1):484–499, ???? 2018. CO-DEN SJADAY. ISSN 1536-0040.

Ferguson:2020:VBP

[Fer20]  Timothy Ferguson.  Vol-

ume bounds for the phaselocking region in the Kuramoto model with asymmetric coupling. SIAM Journal on Applied Dynamical Systems, 19(4):2322–2342, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Filipski:2010:AHT

[FG10] Natasha Filipski and Martin Golubitsky. The Abelian Hopf H mod K theorem. SIAM Journal on Applied Dynamical Systems, 9(2):283– 291, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Fang:2019:LFP

[FG19] Zhou Fang and Chuanhou Gao. Lyapunov function partial differential equations for chemical reaction networks: Some special cases. SIAM Journal on Applied Dynamical Systems, 18(2):1163–1199, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Fantuzzi:2020:BEE

[FG20] Giovanni Fantuzzi and David Goluskin. Bounding extreme events in nonlinear dynamics using convex optimization. SIAM Journal on Applied Dynamical Systems, 19(3):1823– 1864, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Fernandez-Garcia:2015:MTS

[FGDKC15] S. Fernández-García, M. Desroches, M. Krupa, and F. Clément. A multiple time scale coupling of piecewise linear oscillators. application to a neuroendocrine system. SIAM Journal on Applied Dynamical Systems, 14(2):643–673, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Fernandez-Garcia:2012:SST

[FGGT^{+}12] S. Fernández-García, D. Angulo García, G. Olivar Tost, M. di Bernardo, and M. R. Jeffrey. Structural stability of the two-fold singularity. SIAM Journal on Applied Dynamical Systems, 11(4):1215– 1230, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Froyland:2014:CME

[FGH14] Gary Froyland, Georg A.

Gottwald, and Andy Hammerlindl. A computational method to extract macroscopic variables and their dynamics in multiscale systems. SIAM Journal on Applied Dynamical Systems, 13(4):1816– 1846, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Fantuzzi:2016:BDS

[FGHC16] G. Fantuzzi, D. Goluskin, D. Huang, and S. I. Chernyshenko. Bounds for deterministic and stochastic dynamical systems using sum-of-squares optimization. SIAM Journal on Applied Dynamical Systems, 15(4):1962–1988, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Farhat:2020:DAL

[FGHM^{+}20] A. Farhat, N. E. Glatt-Holtz, V. R. Martinez, S. A. Mc-Quarrie, and J. P. Whitehead. Data assimilation in large Prandtl Rayleigh– Bénard convection from thermal measurements. SIAM Journal on Applied Dynamical Systems, 19(1):510–540, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Figueras:2017:FNC

[FGLdlL17] Jordi-Lluís Figueras, Marcio Gameiro, Jean-Philippe Lessard, and Rafael de la Llave. A framework for the numerical computation and a posteriori verification of invariant objects of evolution equations. SIAM Journal on Applied Dynamical Systems, 16(2):1070–1088, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Falconer:2007:ATC

[FGMW07] Ian Falconer, Georg A.

Gottwald, Ian Melbourne, and Kjetil Wormnes. Application of the 0-1 test for chaos to experimental data. SIAM Journal on Applied Dynamical Systems, 6(2):395– 402, 2007. CODEN SJADAY. ISSN 1536-0040.

Franci:2023:BIM

[FGS^{+}23] Alessio Franci, Martin Golubitsky, Ian Stewart, Anastasia Bizyaeva, and Naomi E. Leonard. Breaking indecision in multiagent, multioption dynamics. SIAM Journal on Applied Dynamical Systems, 22(3):1780–1817, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1507826.

Figueras:2012:RCR

[FH12] Jordi-Lluís Figueras and Àlex Haro. Reliable computation of robust response tori on the verge of breakdown. SIAM Journal on Applied Dynamical Systems, 11(2):597–628, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Franco:2014:SPA

[FH14] Daniel Franco and Frank M.

Hilker. Stabilizing populations with adaptive limiters: Prospects and fallacies. SIAM Journal on Applied Dynamical Systems, 13(1):447–465, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Fisher:2022:HCR

[FH22] Michael W. Fisher and Ian A. Hiskens. Hausdorff continuity of region of attraction boundary under parameter variation with application to disturbance recovery. SIAM Journal on Applied Dynamical Systems, 21(1):327–365, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/20M1371944.

Fisher:2023:SNS

[FH23] Michael W. Fisher and Ian A. Hiskens. Stability of the nonwandering set in the region of attraction boundary under perturbations with application to vulnerability assessment. SIAM Journal on Applied Dynamical Systems, 22(4):3390–3430, December 2023. ISSN 1536-0040.

Fiedler:2021:GHR

[FHKK21] Bernold Fiedler, Sindre W. Haugland, Felix P. Kemeth, and Katharina Krischer. Global heteroclinic rebel dynamics among large 2-clusters in permutation equivariant systems. SIAM Journal on Applied Dynamical Systems, 20(3):1277–1319, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Fetecau:2022:EBR

[FHP22] Razvan C. Fetecau, Seung-Yeal Ha, and Hansol Park. Emergent behaviors of rotation matrix flocks. SIAM Journal on Applied Dynamical Systems, 21(2):1382–1425, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/21M1404569.

Froyland:2018:RFB

[FJ18] Gary Froyland and Oliver Junge. Robust FEM-Based extraction of finite-time coherent sets using scattered, sparse, and incomplete trajectories. SIAM Journal on Applied Dynamical Systems, 17(2):1891–1924, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Fetecau:2017:SED

[FK17] R. C. Fetecau and M. Kovacic. Swarm equilibria in domains with boundaries. SIAM Journal on Applied Dynamical Systems, 16(3):1260–1308, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Feliu:2023:PRM

[FKdWY23] Elisenda Feliu, Nidhi Kaihnsa, Timo de Wolff, and Oguzhan Yürük. Parameter region for multistationarity in n- site phosphorylation networks. SIAM Journal on Applied Dynamical Systems, 22(3):2024–2053, ???? 2023. CODEN SJADAY.

ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1504548.

Frauenfelder:2023:SMN

[FKM23] Urs Frauenfelder, Dayung Koh, and Agustin Moreno. Symplectic methods in the numerical search of orbits in real-life planetary systems. SIAM Journal on Applied Dynamical Systems, 22(4):3284– 3319, December 2023. ISSN 1536-0040.

Farjami:2018:CSM

[FKO18] Saeed Farjami, Vivien Kirk, and Hinke M. Osinga. Computing the stable manifold of a saddle slow manifold. SIAM Journal on Applied Dynamical Systems, 17(1):350–379, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Froyland:2020:COP

[FKS20] Gary Froyland, Péter Koltai, and Martin Stahn. Computation and optimal perturbation of finite-time coherent sets for aperiodic flows without trajectory integration. SIAM Journal on Applied Dynamical Systems, 19 (3):1659–1700, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Feliu:2022:QSS

[FLWW22] Elisenda Feliu, Christian Lax, Sebastian Walcher, and Carsten Wiuf. Quasi-steady-state and singular perturbation reduction for reaction networks with noninteracting species. SIAM Journal on Applied Dynamical Systems, 21(2):782–816, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/20M1364503.

Fox:2016:CCI

[FM16] Adam M. Fox and James D.

Meiss. Computing the conjugacy of invariant tori for volume-preserving maps. SIAM Journal on Applied Dynamical Systems, 15(1):557– 579, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Fernandez-Mora:2024:FMP

[FMHM24] Álvaro Fernández-Mora, Alex Haro, and J. M. Mondelo. Flow map parameterization methods for invariant tori in quasi-periodic Hamiltonian systems. SIAM Journal on Applied Dynamical Systems, 23(1):127–166, January 2024. ISSN 1536-0040.

Fetecau:2003:NLM

[FMOW03] R. C. Fetecau, J. E. Marsden, M. Ortiz, and M. West. Nonsmooth Lagrangian mechanics and variational collision integrators. SIAM Journal on Applied Dynamical Systems, 2(3):381–416, 2003. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
40603.

Foias:2016:DDA

[FMT16] Ciprian Foias, Cecilia F. Mondaini, and Edriss S. Titi. A discrete data assimilation scheme for the solutions of the two-dimensional Navier– Stokes equations and their statistics. SIAM Journal on Applied Dynamical Systems, 15(4):2109–2142, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Folias:2011:NAB

[Fol11] Stefanos E. Folias. Nonlinear analysis of breathing pulses in a synaptically coupled neural network. SIAM Journal on Applied Dynamical Systems, 10(2): 744–787, ???? 2011. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.
                                        siam.org/siads/resource/
                                        1/sjaday/v10/i2/p744_s1.

Foster:2016:SSS

[FP16] Jamie M. Foster and Dmitry E. Pelinovsky. Self-similar solutions for reversing interfaces in the slow diffusion equation with strong absorption. SIAM Journal on Applied Dynamical Systems, 15(4):2017– 2050, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Freire:2012:CDP

[FPT12] Emilio Freire, Enrique Ponce, and Francisco Torres. Canonical discontinuous planar piecewise linear systems. SIAM Journal on Applied Dynamical Systems, 11(1): 181–211, ???? 2012. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v11/i1/p181_s1.

Fan:2019:NNN

[FRB19] Gaoyang Fan, Giovanni Russo, and Paul C. Bressloff. Node-to-node and node-to-medium synchronization in quorum sensing networks affected by state-dependent noise. SIAM Journal on Applied Dynamical Systems, 18 (4):1934–1953, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Forgoston:2009:ERS

[FS09] Eric Forgoston and Ira B.

Schwartz. Escape rates in a stochastic environment with multiple scales. SIAM Journal on Applied Dynamical Systems, 8(3):1190–1217, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Franci:2016:TSM

[FS16] Alessio Franci and Rodolphe Sepulchre. A three-scale model of spatio-temporal bursting. SIAM Journal on Applied Dynamical Systems, 15(4):2143–2175, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Fekih-Salem:2021:MMA

[FSDAS21] Radhouane Fekih-Salem, Yessmine Daoud, Nahla Abdellatif, and Tewfik Sari. A mathematical model of anaerobic digestion with syntrophic relationship, substrate inhibition, and distinct removal rates. SIAM Journal on Applied Dynamical Systems, 20 (3):1621–1654, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Fekih-Salem:2019:PCM

[FSS19] Radhouane Fekih-Salem and Tewfik Sari. Properties of the chemostat model with aggregated biomass and distinct removal rates. SIAM Journal on Applied Dynamical Systems, 18(1):481–509, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Finn:2007:TEB

[FT07] Matthew D. Finn and Jean-Luc Thiffeault. Topological entropy of braids on the torus. SIAM Journal on Applied Dynamical Systems, 6(1):79–98, 2007. CODEN SJADAY. ISSN 1536-0040.

Frongillo:2012:EAI

[FT12] Rafael Frongillo and Rodrigo Treviño. Efficient automation of index pairs in computational Conley index theory. SIAM Journal on Applied Dynamical Systems, 11 (1):82–109, ???? 2012. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v11/i1/p82_s1.

Fang:2020:GFN

[FvdSG20] Zhou Fang, Arjan van der Schaft, and Chuanhou Gao. A graphic formulation of nonisothermal chemical reaction systems and the analysis of detailed balanced networks. SIAM Journal on Applied Dynamical Systems, 19(4):2594– 2627, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Farcot:2013:HAS

[FY13] Etienne Farcot and Yuan Yuan. Homogeneous Auxin steady states and spontaneous oscillations in fluxbased Auxin transport models. SIAM Journal on Applied Dynamical Systems, 12 (3):1330–1353, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Gao:2021:RPT

[GAG^{+}21] H. Gao, M. J. Garrido Atienza, A. Gu, K. Lu, and B. Schmalfuß. Rough path theory to approximate random dynamical systems. SIAM Journal on Applied Dynamical Systems, 20(2):997– 1021, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Ghigliazza:2003:SSR

[GAHK03] R. M. Ghigliazza, R. Altendorfer, P. Holmes, and D. Koditschek. A simply stabilized running model. SIAM Journal on Applied Dynamical Systems, 2(2):187–218, 2003. CODEN SJADAY.

ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40831.

Garcia-Azpeitia:2021:GNN

[GAKTWW21] Carlos García-Azpeitia, Wieslaw Krawcewicz, Manuel Tejada-Wriedt, and Haopin Wu. Global nonlinear normal modes in the fullerene molecule C_{60}. SIAM Journal on Applied Dynamical Systems, 20(1):94–129, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Garcia-Azpeitia:2020:FVW

[GAL20] Carlos García-Azpeitia and Jean-Philippe Lessard. Free vibrations in a wave equation modeling MEMS. SIAM Journal on Applied Dynamical Systems, 19(4):2749–2782, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Garrido-Atienza:2016:RDS

[GALS16] María J. Garrido-Atienza, Kening Lu, and Björn Schmalfuss. Random dynamical systems for stochastic evolution equations driven by multiplicative fractional Brownian noise with Hurst parameters H ∈ (1/3, 1/2]. SIAM Journal on Applied Dynamical Systems, 15(1):625– 654, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Garrido-Atienza:2018:LSD

[GAS18] María J. Garrido-Atienza and Björn Schmalfuss. Local stability of differential equations driven by Hölder–continuous paths with Hölder index in (1/3, 1/2). SIAM Journal on Applied Dynamical Systems, 17(3):2352–2380, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Gorman:2009:DSH

[GB09] Michael Gorman and Robert Brockman. Dynamic states of heavy hydrocarbon-oxygen premixed flames on an annular burner. SIAM Journal on Applied Dynamical Systems, 8(2):676–688, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Ginsberg:2023:MFF

[GB23] Alexander G. Ginsberg and Victoria Booth. A meanfield firing-rate model for the suprachiasmatic nucleus. SIAM Journal on Applied Dynamical Systems, 22(1): 90–128, ???? 2023. CO-DEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1496256.

Guardia:2020:ASS

[GBCV20] Daniel Balagué Guardia, Alethea Barbaro, Jose A. Carrillo, and Robert Volkin. Analysis of spherical shell solutions for the radially symmetric aggregation equation. SIAM Journal on Applied Dynamical Systems, 19(4):2628– 2657, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Green:2011:BPM

[GBH11] S. C. Green, C. J. Budd, and G. W. Hunt. Breathers in a pinned mechanical lattice. SIAM Journal on Applied Dynamical Systems, 10 (1):66–91, ???? 2011. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.
                                        siam.org/siads/resource/
                                        1/sjaday/v10/i1/p66_s1.

Gora:2006:ACI

[GBIB06] Pawel Góra, Abraham Boyarsky, MD Shafiqul Islam, and Wael Bahsoun. Absolutely continuous invariant measures that cannot be observed experimentally. SIAM Journal on Applied Dynamical Systems, 5(1):84–90, 2006. CODEN SJADAY. ISSN 1536-0040.

Gandhi:2015:NRM

[GBK15] Punit Gandhi, Cédric Beaume, and Edgar Knobloch. A new resonance mechanism in the Swift–Hohenberg equation with time-periodic forcing. SIAM Journal on Applied Dynamical Systems, 14 (2):860–892, ???? 2015. CO-DEN SJADAY. ISSN 1536-0040.

Guo:2005:ESSa

[GC05a] Yixin Guo and Carson C. Chow. Existence and stability of standing pulses in neural networks: I. existence. SIAM Journal on Applied Dynamical Systems, 4(2):217– 248, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60947.

Guo:2005:ESSb

[GC05b] Yixin Guo and Carson C.

Chow. Existence and stability of standing pulses in neural networks: II. stability. SIAM Journal on Applied Dynamical Systems, 4(2):249– 281, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60948.

Gromov:2020:SAL

[GC20] Dmitry Gromov and Fernando Castaños. Sensitivity analysis of limit cycles in an alpha Stirling engine: a bifurcation-theory approach. SIAM Journal on Applied Dynamical Systems, 19(3):1865– 1883, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

George:2021:DSD

[GCD^{+}21] Erin George, Colleen E.

Chan, Gal Dimand, Ryan M. Chakmak, Claudia Falcon, Daniel Eckhardt, and Robert Martin. Decomposing signals from dynamical systems using shadow manifold interpolation. SIAM Journal on Applied Dynamical Systems, 20(4):2236–2260, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Geddes:2007:OOM

[GCKW07] John B. Geddes, Russell T. Carr, Nathaniel J. Karst, and Fan Wu. The onset of oscillations in microvascular blood flow. SIAM Journal on Applied Dynamical Systems, 6 (4):694–727, 2007. CODEN SJADAY. ISSN 1536-0040.

Gibson:2022:KLQ

[GCY22] Andrew J. Gibson, Michael L. Calvisi, and Xin C. Yee. Koopman linear quadratic regulator using complex eigenfunctions for nonlinear dynamical systems. SIAM Journal on Applied Dynamical Systems, 21(4):2463–2486, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1456078.

Grunberg:2023:ICR

[GD23] Theodore W. Grunberg and Domitilla Del Vecchio. Identifiability of chemical reaction networks with intrinsic and extrinsic noise from stationary distributions. SIAM Journal on Applied Dynamical Systems, 22(3):2206–2241, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M1517202.

Gedeon:2010:OMS

[Ged10]  Tomáš Gedeon.  Oscilla-

tions in monotone systems with a negative feedback. SIAM Journal on Applied Dynamical Systems, 9(1):84– 112, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Geddes:2003:PDA

[GFB03] John B. Geddes, Willie J.

Firth, and Kelly Black. Pulse dynamics in an actively mode-locked laser. SIAM Journal on Applied Dynamical Systems, 2(4):647–671, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41659.

Gambuzza:2020:HAL

[GFE20] Lucia Valentina Gambuzza, Mattia Frasca, and Ernesto Estrada. Hubs-attracting Laplacian and related synchronization on networks. SIAM Journal on Applied Dynamical Systems, 19(2):1057– 1079, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Grinberg:2018:LFA

[GG18] Itay Grinberg and Oleg V.

Gendelman. Localization in finite asymmetric vibroimpact chains. SIAM Journal on Applied Dynamical Systems, 17(3):1961–1988, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Gandhi:2020:BFI

[GGP^{+}20] Punit Gandhi, Martin Golubitsky, Claire Postlethwaite, Ian Stewart, and Yangyang Wang. Bifurcations on fully inhomogeneous networks. SIAM Journal on Applied Dynamical Systems, 19 (1):366–411, ???? 2020. CO-DEN SJADAY. ISSN 1536-0040.

Ghigliazza:2004:MMC

[GH04a] R. M. Ghigliazza and P. Holmes. A minimal model of a central pattern generator and motoneurons for insect locomotion. SIAM Journal on Applied Dynamical Systems, 3 (4):671–700, 2004. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60756.

Ghigliazza:2004:MMB

[GH04b] R. M. Ghigliazza and P. Holmes. Minimal models of bursting neurons: How multiple currents, conductances, and timescales affect bifurcation diagrams. SIAM Journal on Applied Dynamical Systems, 3(4):636–670, 2004. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60230.

Goodman:2005:KAC

[GH05] Roy H. Goodman and Richard Haberman. Kink-antikink collisions in the ϕ^{4} equation: The n-bounce resonance and the separatrix map. SIAM Journal on Applied Dynamical Systems, 4(4):1195–1228, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/63298.

Guillamon:2009:CGA

[GH09] Antoni Guillamon and Gemma Huguet. A computational and geometric approach to phase resetting curves and surfaces. SIAM Journal on Applied Dynamical Systems, 8(3):1005– 1042, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Giesl:2015:CVL

[GH15a] Peter Giesl and Sigurdur Hafstein. Computation and verification of Lyapunov functions. SIAM Journal on Applied Dynamical Systems, 14 (4):1663–1698, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Guglielmi:2015:CHD

[GH15b] Nicola Guglielmi and Ernst Hairer. Classification of hidden dynamics in discontinuous dynamical systems. SIAM Journal on Applied Dynamical Systems, 14(3):1454– 1477, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Geiger:2021:IAS

[GH21] Brett Geiger and Kathryn Hedrick. Integration of additive self-motion inputs in a general bump attractor model. SIAM Journal on Applied Dynamical Systems, 20 (4):1715–1751, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Goldsztejn:2011:TC

[GHC11] Alexandre Goldsztejn, Wayne Hayes, and Pieter Collins.

Tinkerbell is chaotic. SIAM Journal on Applied Dynamical Systems, 10(4):1480– 1501, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i4/p1480_s1.

Grundel:2021:HCV

[GHH^{+}21] Sara M. Grundel, Stefan Heyder, Thomas Hotz, Tobias K. S. Ritschel, Philipp Sauerteig, and Karl Worthmann. How to coordinate vaccination and social distancing to mitigate SARS-CoV-2 outbreaks. SIAM Journal on Applied Dynamical Systems, 20(2):1135–1157, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Glendinning:2024:CFT

[GHH^{+}24] P. Glendinning, S. J. Hogan, M. E. Homer, M. R. Jeffrey, and R. Szalai. Classification of Filippov type 3 singular points in planar bimodal piecewise smooth systems. SIAM Journal on Applied Dynamical Systems, 23 (3):2242–2261, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1622842.

Giunta:2022:LGE

[GHLP22] Valeria Giunta, Thomas Hillen, Mark Lewis, and Jonathan R. Potts. Local and global existence for nonlocal multispecies advection– diffusion models. SIAM Journal on Applied Dynamical Systems, 21(3):1686–1708, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/21M1425992.

Guardia:2010:AAC

[GHS10] Marcel Guardia, S. John Hogan, and Tere M. Seara. An analytical approach to codimension-2 sliding bifurcations in the dry-friction oscillator. SIAM Journal on Applied Dynamical Systems, 9 (3):769–798, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Granados:2012:MMS

[GHS12] A. Granados, S. J. Hogan, and T. M. Seara. The Melnikov method and subharmonic orbits in a piecewisesmooth system. SIAM Journal on Applied Dynamical Systems, 11(3):801–830, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Giannakis:2023:LFD

[GHTW23] Dimitrios Giannakis, Amelia Henriksen, Joel A. Tropp, and Rachel Ward. Learning to forecast dynamical systems from streaming data. SIAM Journal on Applied Dynamical Systems, 22(2):527–558, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/21M144983X.

Guckenheimer:2003:FVP

[GHW03] John Guckenheimer, Kathleen Hoffman, and Warren Weckesser. The forced van der Pol equation I: The Slow flow and its bifurcations. SIAM Journal on Applied Dynamical Systems, 2(1):1–35, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40473.

Giannakis:2015:DAC

[Gia15] Dimitrios Giannakis. Dynamicsadapted cone kernels. SIAM Journal on Applied Dynamical Systems, 14(2):556–608, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Giesl:2019:CCM

[Gie19] Peter Giesl. Computation of a contraction metric for a periodic orbit using meshfree collocation. SIAM Journal on Applied Dynamical Systems, 18(3):1536–1564, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Garbuno-Inigo:2020:ILD

[GIHLS20] Alfredo Garbuno-Inigo, Franca Hoffmann, Wuchen Li, and Andrew M. Stuart. Interacting Langevin diffusions: Gradient structure and ensemble Kalman sampler. SIAM Journal on Applied Dynamical Systems, 19(1):412–441, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Gai:2020:SDP

[GIKR20] Chunyi Gai, David Iron, Theodore Kolokolnikov, and John Rumsey. Spike dynamics in the presence of noise. SIAM Journal on Applied Dynamical Systems, 19(4):2783– 2802, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Garbuno-Inigo:2020:AII

[GINR20] Alfredo Garbuno-Inigo, Nikolas Nüsken, and Sebastian Reich. Affine invariant interacting Langevin dynamics for Bayesian inference. SIAM Journal on Applied Dynamical Systems, 19(3):1633–1658, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Gonzalez:2017:HOP

[GJ17] J. L. Gonzalez and J. D. Mireles James. High-order parameterization of stable/unstable manifolds for long periodic orbits of maps. SIAM Journal on Applied Dynamical Systems, 16(3):1748–1795, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Guckenheimer:2012:RES

[GJM12] John Guckenheimer, Tomas Johnson, and Philipp Meerkamp. Rigorous enclosures of a slow manifold. SIAM Journal on Applied Dynamical Systems, 11(3):831–863, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Gimeno:2022:NCH

[GJNO22] Joan Gimeno, Àngel Jorba, Begoña Nicolás, and Estrella Olmedo. Numerical computation of high-order expansions of invariant manifolds of high-dimensional tori. SIAM Journal on Applied Dynamical Systems, 21(3):1832–1861, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/21M1458363.

Guckenheimer:2009:CSM

[GK09] John Guckenheimer and Christian Kuehn. Computing slow manifolds of saddle type. SIAM Journal on Applied Dynamical Systems, 8 (3):854–879, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Guckenheimer:2010:HOF

[GK10] John Guckenheimer and Christian Kuehn. Homoclinic orbits of the FitzHugh– Nagumo equation: Bifurcations in the full system. SIAM Journal on Applied Dynamical Systems, 9(1):138–153, ???? 2010. CODEN SJA-DAY. ISSN 1536-0040.

Gupta:2018:CII

[GK18] Ankit Gupta and Mustafa Khammash. Computational identification of irreducible state-spaces for stochastic reaction networks. SIAM Journal on Applied Dynamical Systems, 17(2):1213–1266, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Gkogkas:2022:GMF

[GK22] Marios Antonios Gkogkas and Christian Kuehn. Graphop mean-field limits for Kuramototype models. SIAM Journal on Applied Dynamical Systems, 21(1):248–283, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1391523.

Granados:2014:BCB

[GKC14] A. Granados, M. Krupa, and F. Clément. Border collision bifurcations of stroboscopic maps in periodically driven spiking models. SIAM Journal on Applied Dynamical Systems, 13(4):1387–1416, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Goodman:2015:DVD

[GKCG15] Roy H. Goodman, P. G.

Kevrekidis, and R. Carretero-González. Dynamics of vortex dipoles in anisotropic Bose– Einstein condensates. SIAM Journal on Applied Dynamical Systems, 14(2):699–729, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Gear:2005:PSM

[GKKZ05] C. W. Gear, T. J. Kaper, I. G.

Kevrekidis, and A. Zagaris. Projecting to a slow manifold: Singularly perturbed systems and legacy codes.

SIAM Journal on Applied Dynamical Systems, 4(3):711– 732, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60829.

Gonchenko:2005:GHM

[GKM05] V. S. Gonchenko, Yu. A. Kuznetsov, and H. G. E. Meijer. Generalized Hénon map and bifurcations of homoclinic tangencies. SIAM Journal on Applied Dynamical Systems, 4(2):407–436, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60487.

Gruene:2021:RAT

[GKM21] Lars Gruene, Thomas Kriecherbauer, and Michael Margaliot. Random attraction in the TASEP model. SIAM Journal on Applied Dynamical Systems, 20 (1):65–93, ???? 2021. CO-DEN SJADAY. ISSN 1536-0040.

Green:2009:BAS

[GKML09] Kirk Green, Bernd Krauskopf, Frank Marten, and Daan Lenstra. Bifurcation analysis of a spatially extended laser with optical feedback. SIAM Journal on Applied Dynamical Systems, 8(1):222–252, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Gabern:2006:BAO

[GKMS06] F. Gabern, W. S. Koon, J. E. Marsden, and D. J. Scheeres.

Binary asteroid observation orbits from a global dynamical perspective. SIAM Journal on Applied Dynamical Systems, 5(2):252–279, 2006. CODEN SJADAY. ISSN 1536-0040.

Giraldo:2017:SIO

[GKO17] Andrus Giraldo, Bernd Krauskopf, and Hinke M. Osinga. Saddle invariant objects and their global manifolds in a neighborhood of a homoclinic flip bifurcation of Case B. SIAM Journal on Applied Dynamical Systems, 16(1):640–686, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Giraldo:2018:CGB

[GKO18] Andrus Giraldo, Bernd Krauskopf, and Hinke M. Osinga. Cascades of global bifurcations and chaos near a homoclinic flip bifurcation: a case study. SIAM Journal on Applied Dynamical Systems, 17(4):2784– 2829, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Guan:2024:GSS

[GKR24] Ruofei Guan, Jana Kopfová, and Dmitrii Rachinskii. Global stability of SIR model with heterogeneous transmission rate modeled by the Preisach operator. SIAM Journal on Applied Dynamical Systems, 23(2):1199–1241, May 2024. ISSN 1536-0040.

Green:2003:TPS

[GKS03] Kirk Green, Bernd Krauskopf, and Giovanni Samaey. A twoparameter study of the locking region of a semiconductor laser subject to phaseconjugate feedback. SIAM Journal on Applied Dynamical Systems, 2(2):254–276, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41657.

Goh:2023:FWP

[GKSV23] Ryan Goh, Tasso J. Kaper, Arnd Scheel, and Theodore Vo. Fronts in the wake of a parameter ramp: Slow passage through pitchfork and fold bifurcations. SIAM Journal on Applied Dynamical Systems, 22(3):2312–2356, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M1541812.

Golubitsky:2009:BSH

[GL09] Martin Golubitsky and Reiner Lauterbach. Bifurcations from synchrony in homogeneous networks: Linear theory. SIAM Journal on Applied Dynamical Systems, 8 (1):40–75, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Glasner:2013:SEC

[GL13] Karl B. Glasner and Alan E. Lindsay. The stability and evolution of curved domains arising from one-dimensional localized patterns. SIAM Journal on Applied Dynamical Systems, 12(2):650–673, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Guckenheimer:2015:SHB

[GL15] John Guckenheimer and Ian Lizarraga. Shilnikov homoclinic bifurcation of mixedmode oscillations. SIAM Journal on Applied Dynamical Systems, 14(2):764–786, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Gameiro:2017:PVI

[GL17] Marcio Gameiro and Jean-Philippe Lessard. A posteriori verification of invariant objects of evolution equations: Periodic orbits in the Kuramoto–Sivashinsky PDE. SIAM Journal on Applied Dynamical Systems, 16(1):687– 728, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Gilbert:2024:AOC

[GL24] Rhys E. Gilbert and Davide Lasagna. On the application of optimal control techniques to the shadowing approach for time averaged sensitivity analysis of chaotic systems. SIAM Journal on Applied Dynamical Systems, 23(1):505– 552, January 2024. ISSN 1536-0040.

Glendinning:2014:BCN

[Gle14] Paul Glendinning. The border collision normal form with stochastic switching surface. SIAM Journal on Applied Dynamical Systems, 13(1):181– 193, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Gimeno:2023:PPO

[GLJY23] Joan Gimeno, Jean-Philippe Lessard, J. D. Mireles James, and Jiaqi Yang. Persistence of periodic orbits under statedependent delayed perturbations: Computer-assisted proofs. SIAM Journal on Applied Dynamical Systems, 22 (3):1743–1779, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1499418.

Gou:2015:SOD

[GLNW15] J. Gou, Y. X. Li, W. Nagata, and M. J. Ward. Synchronized oscillatory dynamics for a 1-D model of membrane kinetics coupled by linear bulk diffusion. SIAM Journal on Applied Dynamical Systems, 14(4):2096–2137, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Giunta:2021:PFT

[GLS21] Valeria Giunta, Maria Carmela Lombardo, and Marco Sammartino. Pattern formation and transition to chaos in a chemotaxis model of acute inflammation. SIAM Journal on Applied Dynamical Systems, 20(4):1844–1881, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Gandhi:2023:PPM

[GLS23] Punit Gandhi, Lily Liu, and Mary Silber. A pulsedprecipitation model of dryland vegetation pattern formation. SIAM Journal on Applied Dynamical Systems, 22(2):657–693, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1469572.

Gourley:2010:SDM

[GLW10] Stephen A. Gourley, Rongsong Liu, and Jianhong Wu. Spatiotemporal distributions of migratory birds: Patchy models with delay. SIAM Journal on Applied Dynamical Systems, 9(2):589–610, ???? 2010. CODEN SJA-DAY. ISSN 1536-0040.

Gottwald:2009:ITC

[GM09] Georg A. Gottwald and Ian Melbourne. On the implementation of the 0–1 test for chaos. SIAM Journal on Applied Dynamical Systems, 8 (1):129–145, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Guckenheimer:2012:USH

[GM12] John Guckenheimer and Philipp Meerkamp. Unfoldings of singular Hopf bifurcation. SIAM Journal on Applied Dynamical Systems, 11 (4):1325–1359, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Ghazaryan:2015:CSP

[GM15] Anna Ghazaryan and Vahagn Manukian. Coherent structures in a population model for mussel-algae interaction. SIAM Journal on Applied Dynamical Systems, 14(2):893– 913, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Gomez:2016:ETV

[GMB16] Marcella M. Gomez, Richard M. Murray, and Matthew R. Bennett. The effects of time-varying temperature on delays in genetic networks. SIAM Journal on Applied Dynamical Systems, 15(3):1734– 1752, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Govindarajan:2019:AKS

[GMCM19] Nithin Govindarajan, Ryan Mohr, Shivkumar Chandrasekaran, and Igor Mezic. On the approximation of Koopman spectra for measure preserving transformations. SIAM Journal on Applied Dynamical Systems, 18(3):1454– 1497, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Govindarajan:2021:AKS

[GMCM21] Nithin Govindarajan, Ryan Mohr, Shivkumar Chandrasekaran, and Igor Mezic. On the approximation of Koopman spectra of measurepreserving flows. SIAM Journal on Applied Dynamical Systems, 20(1):232–261, ????

2021. CODEN SJADAY. ISSN 1536-0040.

Gerhard:2008:NVC

[GMM08] Johannes Gerhard, Wolfgang Marquardt, and Martin Mönnigmann. Normal vectors on critical manifolds for robust design of transient processes in the presence of fast disturbances. SIAM Journal on Applied Dynamical Systems, 7(2):461–490, 2008. CODEN SJADAY. ISSN 1536-0040.

Goh:2011:SWS

[GMS11] Ryan N. Goh, Samantha Mesuro, and Arnd Scheel. Spatial wavenumber selection in recurrent precipitation. SIAM Journal on Applied Dynamical Systems, 10 (1):360–402, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i1/p360_s1.

Gopalkrishnan:2014:GAG

[GMS14] Manoj Gopalkrishnan, Ezra Miller, and Anne Shiu. A geometric approach to the global attractor conjecture. SIAM Journal on Applied Dynamical Systems, 13(2):758–797, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Gong:2018:OLF

[GMY18] Xue Gong, Martin J. Mohlenkamp, and Todd R. Young. The optimization landscape for fitting a rank-2 tensor with a rank-1 tensor. SIAM Journal on Applied Dynamical Systems, 17(2):1432–1477, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Galatolo:2014:EAR

[GN14] Stefano Galatolo and Isaia Nisoli. An elementary approach to rigorous approximation of invariant measures. SIAM Journal on Applied Dynamical Systems, 13(2):958– 985, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Galan:2018:QSC

[GNR18] Jorge Galán, Daniel Núñez, and Andrés Rivera. Quantitative stability of certain families of periodic solutions in the Sitnikov problem. SIAM Journal on Applied Dynamical Systems, 17(1):52–77, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Guckenheimer:2002:CHH

[GO02] John Guckenheimer and Ricardo A. Oliva. Chaos in the Hodgkin–Huxley model. SIAM Journal on Applied Dynamical Systems, 1(1):105– 114, 2002. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/39404.

Gamus:2015:DBW

[GO15] Benny Gamus and Yizhar Or. Dynamic bipedal walking under stick-slip transitions. SIAM Journal on Applied Dynamical Systems, 14 (2):609–642, ???? 2015. CO-DEN SJADAY. ISSN 1536-0040.

Giesl:2020:ADC

[GOH20] Peter Giesl, Conor Osborne, and Sigurdur Hafstein. Automatic determination of connected sublevel sets of CPA Lyapunov functions. SIAM Journal on Applied Dynamical Systems, 19(2):1029–1056, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Gorban:2013:TTS

[Gor13] Alexander N. Gorban. Thermodynamic tree: The space of admissible paths. SIAM Journal on Applied Dynamical Systems, 12(1):246–278, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Gilson:2020:PSM

[GP20] Matthieu Gilson and Jean-Pascal Pfister. Propagation of spiking moments in linear Hawkes networks. SIAM Journal on Applied Dynamical Systems, 19(2):828–859, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Guillamon:2017:ESC

[GPTV17] A. Guillamon, R. Prohens, A. E. Teruel, and C. Vich. Estimation of synaptic conductance in the spiking regime for the McKean neuron model. SIAM Journal on Applied Dynamical Systems, 16(3):1397– 1424, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Gardini:2022:PFV

[GRS^{+}22] Laura Gardini, Davide Radi, Noemi Schmitt, Iryna Sushko, and Frank Westerhoff. Perception of fundamental values and financial market dynamics: Mathematical insights from a 2D piecewise linear map. SIAM Journal on Applied Dynamical Systems, 21(4):2314–2337, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1456339.

Griffin:2019:CIC

[GRSB19] Christopher Griffin, Sarah Rajtmajer, Anna Squicciarini, and Andrew Belmonte. Consensus and information cascades in game-theoretic imitation dynamics with static and dynamic network topologies. SIAM Journal on Applied Dynamical Systems, 18(2):597–628, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Ghazaryan:2007:NCI

[GS07] Anna Ghazaryan and Björn Sandstede. Nonlinear convective instability of Turingunstable fronts near onset: a case study. SIAM Journal on Applied Dynamical Systems, 6(2):319–347, 2007. CODEN SJADAY. ISSN 1536-0040.

Giraldo:2009:SCA

[GS09a] A. Giraldo and J. M. R. Sanjurjo. Singular continuations of attractors. SIAM Journal on Applied Dynamical Systems, 8(2):554–575, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Goel:2009:GBD

[GS09b] Pranay Goel and Arthur Sherman. The geometry of bursting in the dual oscillator model of pancreatic β-cells. SIAM Journal on Applied Dynamical Systems, 8(4):1664– 1693, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Guckenheimer:2011:GMM

[GS11] John Guckenheimer and Chris Scheper. A geometric model for mixed-mode oscillations in a chemical system. SIAM Journal on Applied Dynamical Systems, 10 (1):92–128, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i1/p92_s1.

Guckenheimer:2013:MTS

[GS13] John Guckenheimer and Chris Scheper. Multiple time scale analysis of a model Belousov–Zhabotinskii reaction. SIAM Journal on Applied Dynamical Systems, 12 (4):1968–1996, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Gazor:2016:BCU

[GS16] Majid Gazor and Nasrin Sadri. Bifurcation control and universal unfolding for Hopfzero singularities with leading solenoidal terms. SIAM Journal on Applied Dynamical Systems, 15(2):870–903, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Golubitsky:2018:HMI

[GS18] Martin Golubitsky and Ian Stewart. Homeostasis with multiple inputs. SIAM Journal on Applied Dynamical Systems, 17(2):1816–1832, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Gurevich:2020:DSA

[GS20] Pavel Gurevich and Hannes Stuke. Dynamical systems approach to outlier robust deep neural networks for regression. SIAM Journal on Applied Dynamical Systems, 19(4):2567–2593, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Gholami:2022:IDS

[GS22] Amin Gholami and Xu A. Sun. The impact of damping in second-order dynamical systems with applications to power grid stability. SIAM Journal on Applied Dynamical Systems, 21(1):405–437, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/20M1370392.

Gomez:2016:SSS

[GSB^{+}16] Marcella M. Gomez, Mehdi Sadeghpour, Matthew R.

Bennett, Gábor Orosz, and Richard M. Murray. Stability of systems with stochastic delays and applications to genetic regulatory networks. SIAM Journal on Applied Dynamical Systems, 15(4):1844– 1873, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Garcia:2015:OCR

[GSDN15] F. Garcia, J. Sánchez, E. Dormy, and M. Net. Oscillatory convection in rotating spherical shells: Low Prandtl number and non-slip boundary conditions. SIAM Journal on Applied Dynamical Systems, 14(4):1787–1807, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Gou:2023:SSB

[GSJ23] Wei Gou, Yongli Song, and Zhen Jin. The steady state bifurcation for general networkorganized reaction–diffusion systems and its application in a metapopulation epidemic model. SIAM Journal on Applied Dynamical Systems, 22(2):559–602, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1439092.

Golubitsky:2003:BVC

[GST03] Martin Golubitsky, Lie June Shiau, and Andrei Török. Bifurcation on the visual cortex with weakly anisotropic lateral coupling. SIAM Journal on Applied Dynamical Systems, 2(2):97–143, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40988.

Golubitsky:2005:PSC

[GST05] Martin Golubitsky, Ian Stewart, and Andrei Török. Patterns of synchrony in coupled cell networks with multiple arrows. SIAM Journal on Applied Dynamical Systems, 4(1):78–100, 2005. CODEN SJADAY. ISSN 1536-0040.

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        61263.

Goodman:2018:SFS

[GT18] Benjamin Goodman and Paul F. Tupper. Stability and fluctuations in a simple model of phonetic category change. SIAM Journal on Applied Dynamical Systems, 17(3):2332– 2351, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Guckenheimer:2008:SHB

[Guc08] John Guckenheimer. Singular Hopf bifurcation in systems with two slow variables. SIAM Journal on Applied Dynamical Systems, 7(4):1355– 1377, 2008. CODEN SJA-DAY. ISSN 1536-0040.

Guo:2010:DSS

[Guo10] Yujin Guo. Dynamical solutions of singular wave equations modeling electrostatic MEMS. SIAM Journal on Applied Dynamical Systems, 9(4):1135–1163, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Guo:2012:EST

[Guo12] Yixin Guo. Existence and stability of traveling fronts in a lateral inhibition neural network. SIAM Journal on Applied Dynamical Systems, 11 (4):1543–1582, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Guckenheimer:2004:FMA

[GV04] John Guckenheimer and Alexander Vladimirsky. A fast method for approximating invariant manifolds. SIAM Journal on Applied Dynamical Systems, 3(3):232– 260, 2004. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60017.

Germay:2009:NDD

[GVNS09] Christophe Germay, Nathan Van de Wouw, Henk Nijmeijer, and Rodolphe Sepulchre. Nonlinear drillstring dynamics analysis. SIAM Journal on Applied Dynamical Systems, 8(2):527–553, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Gavish:2017:SLS

[GVY17] Nir Gavish, Idan Versano, and Arik Yochelis. Spatially localized self-assembly driven by electrically charged phase separation. SIAM Journal on Applied Dynamical Systems, 16(4):1946–1968, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Gai:2024:NAD

[GW24] Chunyi Gai and Michael J. Ward. The nucleation– annihilation dynamics of hotspot patterns for a reaction– diffusion system of urban crime with police deployment. SIAM Journal on Applied Dynamical Systems, 23 (3):2018–2060, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1562330.

Gomez:2019:LSS

[GWW19] Daniel Gomez, Michael J. Ward, and Juncheng Wei. The linear stability of symmetric spike patterns for a bulk-membrane coupled Gierer–Meinhardt model. SIAM Journal on Applied Dynamical Systems, 18(2):729– 768, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Gracy:2024:TUE

[GYAU24] Sebin Gracy, Mengbin Ye, Brian D. O. Anderson, and Cesar A. Uribe. Towards understanding the endemic behavior of a competitive trivirus SIS networked model. SIAM Journal on Applied Dynamical Systems, 23(2): 1372–1410, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1563074.

Gimeno:2021:NCP

[GYdlL21] Joan Gimeno, Jiaqi Yang, and Rafael de la Llave. Numerical computation of periodic orbits and isochrons for state-dependent delay perturbation of an ODE in the plane. SIAM Journal on Applied Dynamical Systems, 20 (3):1514–1543, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Gerlach:2020:SOP

[GZED20] Raphael Gerlach, Adrian Ziessler, Bruno Eckhardt, and Michael Dellnitz. A set-oriented path following method for the approximation of parameter dependent attractors. SIAM Journal on Applied Dynamical Systems, 19(1):705–723, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Hsieh:2022:MML

[HACY22] Dai-Ni Hsieh, Sylvain Arguillère, Nicolas Charon, and Laurent Younes. Mechanistic modeling of longitudinal shape changes: Equations of motion and inverse problems. SIAM Journal on Applied Dynamical Systems, 21(1):80–101, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1423099.

Harris:2021:TVA

[HARB21] Kameron Decker Harris, Aleksandr Aravkin, Rajesh Rao, and Bingni Wen Brunton. Time-varying autoregression with low-rank tensors. SIAM Journal on Applied Dynamical Systems, 20 (4):2335–2358, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Hill:2016:AAS

[HAS16] Kaitlin Hill, Dorian S. Abbot, and Mary Silber. Analysis of an Arctic sea ice loss model in the limit of a discontinuous albedo. SIAM Journal on Applied Dynamical Systems, 15(2):1163–1192, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Haskovec:2021:SPA

[Has21] Jan Haskovec. A simple proof of asymptotic consensus in the Hegselmann–Krause and Cucker–Smale models with normalization and delay. SIAM Journal on Applied Dynamical Systems, 20 (1):130–148, ???? 2021. CO-DEN SJADAY. ISSN 1536-0040.

Hasler:2013:DSBa

[HBB13a] Martin Hasler, Vladimir Belykh, and Igor Belykh. Dynamics of stochastically blinking systems. Part I: Finite time properties. SIAM Journal on Applied Dynamical Systems, 12(2):1007–1030, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Hasler:2013:DSBb

[HBB13b] Martin Hasler, Vladimir Belykh, and Igor Belykh. Dynamics of stochastically blinking systems. Part II: Asymptotic properties. SIAM Journal on Applied Dynamical Systems, 12(2):1031–1084, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Hasan:2023:RIT

[HCW23] Cris R. Hasan, Ruaidhrí Mac Cárthaigh, and Sebastian Wieczorek. Rate-induced tipping in heterogeneous reaction-diffusion systems: an invariant manifold framework and geographically shifting ecosystems. SIAM Journal on Applied Dynamical Systems, 22(4):2991–3024, October 2023. ISSN 1536-0040.

Homs-Dones:2021:NCN

[HDDL21] Marc Homs-Dones, Karel Devriendt, and Renaud Lambiotte. Nonlinear consensus on networks: Equilibria, effective resistance, and trees of motifs. SIAM Journal on Applied Dynamical Systems, 20(3):1544–1570, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Horenko:2008:MDC

[HDL^{+}08] Illia Horenko, Evelyn Dittmer, Filip Lankas, John Maddocks, Philipp Metzner, and Christof Schütte. Macroscopic dynamics of complex metastable systems: Theory, algorithms, and application to B-DNA. SIAM Journal on Applied Dynamical Systems, 7(2):532–560, 2008. CODEN SJADAY. ISSN 1536-0040.

Haro:2007:PMC

[HdlL07] A. Haro and R. de la Llave. A parameterization method for the computation of invariant tori and their whiskers in quasi-periodic maps: Explorations and mechanisms for the breakdown of hyperbolicity. SIAM Journal on Applied Dynamical Systems, 6 (1):142–207, 2007. CODEN SJADAY. ISSN 1536-0040.

Huguet:2013:CLC

[HdlL13] Gemma Huguet and Rafael de la Llave. Computation of limit cycles and their isochrons: Fast algorithms and their convergence. SIAM Journal on Applied Dynamical Systems, 12(4):1763–1802, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

He:2023:RPL

[HdlL23] Xiaolong He and Rafael de la Llave. Resonances and phase locking phenomena for foliation preserving torus maps. SIAM Journal on Applied Dynamical Systems, 22(1):382–418, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1485103.

Hilder:2023:MMP

[HdRS23] Bastian Hilder, Björn de Rijk, and Guido Schneider. Moving modulating pulse and front solutions of permanent form in a FPU model with nearest and next-to-nearest neighbor interaction. SIAM Journal on Applied Dynamical Systems, 22(2):1076–1113, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1502902.

Harris:2015:BWC

[HE15] Jeremy Harris and Bard Ermentrout. Bifurcations in the Wilson–Cowan equations with nonsmooth firing rate. SIAM Journal on Applied Dynamical Systems, 14(1):43– 72, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Henderson:2005:CIM

[Hen05] Michael E. Henderson. Computing invariant manifolds by integrating fat trajectories. SIAM Journal on Applied Dynamical Systems, 4(4):832– 882, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60289.

Henderson:2011:FBT

[Hen11] Michael E. Henderson. Flow box tiling methods for compact invariant manifolds.

SIAM Journal on Applied Dynamical Systems, 10(3): 1154–1176, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i3/p1154_s1.

Hermann:2010:LCL

[HG10] Sebastian Hermann and Georg A. Gottwald. The large core limit of spiral waves in excitable media: a numerical approach. SIAM Journal on Applied Dynamical Systems, 9(2):536–567, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Hernandez-Gardunno:2015:LRE

[HGS15] Antonio Hernández-Gardunño and Cristina Stoica. Lagrangian relative equilibria in a modified three-body problem with a rotationally symmetric ellipsoid. SIAM Journal on Applied Dynamical Systems, 14(1):221–252, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Hamed:2015:IMC

[HGT15] Mohammad Abu Hamed, Yanqiu Guo, and Edriss S. Titi. Inertial manifolds for certain subgrid-scale αmodels of turbulence. SIAM Journal on Applied Dynamical Systems, 14(3):1308–1325, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Hamster:2019:STW

[HH19] C. H. S. Hamster and H. J.

Hupkes. Stability of traveling waves for reaction–diffusion equations with multiplicative noise. SIAM Journal on Applied Dynamical Systems, 18 (1):205–278, ???? 2019. CO-DEN SJADAY. ISSN 1536-0040.

Hamster:2020:STW

[HH20] C. H. S. Hamster and H. J.

Hupkes. Stability of traveling waves on exponentially long timescales in stochastic reaction–diffusion equations. SIAM Journal on Applied Dynamical Systems, 19(4):2469– 2499, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Ha:2022:DSP

[HHBS22] Kyung Ha, Brendan Harding, Andrea L. Bertozzi, and Yvonne M. Stokes. Dynamics of small particle inertial migration in curved square ducts. SIAM Journal on Applied Dynamical Systems, 21(1):714–734, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1430935.

Halper:2009:STD

[HHHY09] Russell D. Halper, Eric J.

Harder, Brian R. Hunt, and James A. Yorke. Stability of TCP dynamics in large data networks. SIAM Journal on Applied Dynamical Systems, 8(1):146–159, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Hirsh:2020:CDI

[HHKB20] Seth M. Hirsh, Kameron Decker Harris, J. Nathan Kutz, and Bingni W. Brunton. Centering data improves the dynamic mode decomposition. SIAM Journal on Applied Dynamical Systems, 19(3):1920– 1955, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Han:2021:DTB

[HHW21] Daozhi Han, Marco Hernandez, and Quan Wang. Dynamic transitions and bifurcations for a class of axisymmetric geophysical fluid flow. SIAM Journal on Applied Dynamical Systems, 20 (1):38–64, ???? 2021. CO-DEN SJADAY. ISSN 1536-0040.

Herty:2023:FMC

[HI23] Michael Herty and Elisa Iacomini. Filtering methods for coupled inverse problems. SIAM Journal on Applied Dynamical Systems, 22 (2):1234–1252, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1483839.

Huang:2016:ABD

[HJL16] Qihua Huang, Yu Jin, and Mark A. Lewis. R_{0} analysis of a benthic-drift model for a stream population. SIAM Journal on Applied Dynamical Systems, 15(1):287–321, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040. See erratum [HJL17].

Huang:2017:EAB

[HJL17] Qihua Huang, Yu Jin, and Mark A. Lewis. Erratum: R_{0} analysis of a benthic-drift model for a stream population. SIAM Journal on Applied Dynamical Systems, 16 (1):770, ???? 2017. CODEN SJADAY. ISSN 1536-0040. See [HJL16].

Hupkes:2023:PRB

[HJSS23] Hermen Jan Hupkes, Mia Jukić, Petr Stehlík, and Vladimír Svígler. Propagation reversal for bistable differential equations on trees. SIAM Journal on Applied Dynamical Systems, 22(3): 1906–1944, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1502203.

Hewitt:2005:DOP

[HK05] Sarah E. Hewitt and J. Nathan Kutz. Dynamics of the optical parametric oscillator near resonance detuning. SIAM Journal on Applied Dynamical Systems, 4(4):808–831, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/61029.

Holzel:2015:SLT

[HK15] Robert W. Hölzel and Katharina Krischer. Stability and long term behavior of a Hebbian network of Kuramoto oscillators. SIAM Journal on Applied Dynamical Systems, 14(1):188–201, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Hofherr:2018:LTT

[HK18] Florian Hofherr and Daniel Karrasch. Lagrangian transport through surfaces in compressible flows. SIAM Journal on Applied Dynamical Systems, 17(1):526–546, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Ha:2019:SOP

[HK19] Seung-Yeal Ha and Dohyun Kim. A second-order particle swarm model on a sphere and emergent dynamics. SIAM Journal on Applied Dynamical Systems, 18(1):80–116, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Hickok:2022:BCM

[HKB^{+}22a] Abigail Hickok, Yacoub Kureh, Heather Z. Brooks, Michelle Feng, and Mason A. Porter. A boundedconfidence model of opinion dynamics on hypergraphs. SIAM Journal on Applied Dynamical Systems, 21(1):1–32, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/

doi/10.1137/21M1399427.

See erratum [HKB^{+}22b].

Hickok:2022:EBC

[HKB^{+}22b] Abigail Hickok, Yacoub Kureh, Heather Zinn Brooks, Michelle Feng, and Mason A. Porter. Erratum: A bounded-confidence model of opinion dynamics on hypergraphs. SIAM Journal on Applied Dynamical Systems, 21(2):1660–1661, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M147267X.  See

[HKB^{+}22a].

Haller:2020:BTD

[HKK20] George Haller, Daniel Karrasch, and Florian Kogelbauer. Barriers to the transport of diffusive scalars in compressible flows. SIAM Journal on Applied Dynamical Systems, 19(1):85–123, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Ha:2014:LTD

[HKL14] Seung-Yeal Ha, Yongduck Kim, and Zhuchun Li. Largetime dynamics of Kuramoto oscillators under the effects of inertia and frustration. SIAM Journal on Applied Dynamical Systems, 13(1):466–492, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Howcroft:2013:IVS

[HKLN13] Chris Howcroft, Bernd Krauskopf, Mark H. Lowenberg, and Simon A. Neild. Influence of variable side-stay geometry on the shimmy dynamics of an aircraft dual-wheel main landing gear. SIAM Journal on Applied Dynamical Systems, 12(3):1181–1209, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Ha:2020:EBA

[HKLN20] Seung-Yeal Ha, Dohyun Kim, Jaeseung Lee, and Se Eun Noh. Emergence of bicluster aggregation for the swarm sphere model with attractive– repulsive couplings. SIAM Journal on Applied Dynamical Systems, 19(2):1225–1270, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Ha:2021:UTC

[HKM21] Seung-Yeal Ha, Myeongju Kang, and Bora Moon. Uniform-in-time continuum limit of the winfree model on an infinite cylinder and emergent dynamics. SIAM Journal on Applied Dynamical Systems, 20(2):1104–1134, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Hittmeyer:2013:IGI

[HKO13] Stefanie Hittmeyer, Bernd Krauskopf, and Hinke M. Osinga. Interacting global invariant sets in a planar map model of wild chaos. SIAM Journal on Applied Dynamical Systems, 12(3):1280–1329, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Hasan:2017:MMO

[HKO17] Cris R. Hasan, Bernd Krauskopf, and Hinke M. Osinga. Mixedmode oscillations and twin Canard orbits in an autocatalytic chemical reaction. SIAM Journal on Applied Dynamical Systems, 16(4):2165– 2195, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Ha:2022:EED

[HKP22] Seung-Yeal Ha, Dohyun Kim, and Hansol Park. Existence and emergent dynamics of quadratically separable states to the Lohe tensor model. SIAM Journal on Applied Dynamical Systems, 21(2):1166–1208, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1409664.

Ha:2018:EPL

[HKZ18] Seung-Yeal Ha, Dongnam Ko, and Yinglong Zhang. Emergence of phase-locking in the Kuramoto model for identical oscillators with frustration. SIAM Journal on Applied Dynamical Systems, 17 (1):581–625, ???? 2018. CO-DEN SJADAY. ISSN 1536-0040.

Holm:2002:SPR

[HL02] Darryl D. Holm and Peter Lynch. Stepwise precession of the resonant swinging spring. SIAM Journal on Applied Dynamical Systems, 1(1):44–64, 2002. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38857.

Heiter:2018:TDG

[HL18] Pascal Heiter and Dirk Lebiedz. Towards differential geometric characterization of slow invariant manifolds in extended phase space: Sectional curvature and flow invariance. SIAM Journal on Applied Dynamical Systems, 17(1):732–753, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

He:2022:CRA

[HL22] Xiaoqing He and Liu Liu. On the conjecture of the role of advection in a twospecies competition–diffusion model. SIAM Journal on Applied Dynamical Systems, 21 (3):1663–1685, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1451713.

Henot:2023:NCT

[HLJ23] Olivier Hénot, Jean-Philippe Lessard, and Jason D. Mireles James. Numerical computation of transverse homoclinic orbits for periodic solutions of delay differential equations. SIAM Journal on Applied Dynamical Systems, 22(4):3093– 3129, November 2023. ISSN 1536-0040.

Ha:2018:EDK

[HLLP18] Seung-Yeal Ha, Jaeseung Lee, Zhuchun Li, and Jinyeong Park. Emergent dynamics of Kuramoto oscillators with adaptive couplings: Conservation law and fast learning. SIAM Journal on Applied Dynamical Systems, 17(2):1560– 1588, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Harrington:2022:RAT

[HLvdD22] Peter D. Harrington, Mark A.

Lewis, and P. van den Driessche. Reactivity, attenuation, and transients in metapopulations. SIAM Journal on Applied Dynamical Systems, 21(2):1287–1321, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M140451X.

Hampton:2022:AIT

[HM22] Amanda E. Hampton and James D. Meiss. Antiintegrability for three-dimensional quadratic maps. SIAM Journal on Applied Dynamical Systems, 21(1):650–675, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1433289.

Hampton:2024:CAI

[HM24] Amanda E. Hampton and James D. Meiss. Connecting anti-integrability to attractors for three-dimensional quadratic diffeomorphisms.

SIAM Journal on Applied Dynamical Systems, 23(1):616– 640, February 2024. ISSN 1536-0040.

Hamdan:2023:TDP

[HMD^{+}23] Taleb Bou Hamdan, Guillaume Mercère, Thibault Dairay, Raphael Meunier, Pascal Tremblay, and Patrick Coirault. Tracking distributed parameters system dynamics with recursive dynamic mode decomposition with control. SIAM Journal on Applied Dynamical Systems, 22(1):37–64, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1478665.

Huijberts:2009:STD

[HMN09] Henri Huijberts, Wim Michiels, and Henk Nijmeijer. Stabilizability via time-delayed feedback: An eigenvalue optimization approach. SIAM Journal on Applied Dynamical Systems, 8(1):1–20, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Hoffman:2002:CSI

[HMP02] Kathleen A. Hoffman, Robert S. Manning, and Randy C. Paffenroth. Calculation of the stability index in parameterdependent calculus of variations problems: Buckling of a twisted elastic strut. SIAM Journal on Applied Dynamical Systems, 1(1):115–145, 2002. CODEN SJADAY.

ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39662.

Hupkes:2019:BTW

[HMS19] Hermen Jan Hupkes, Leonardo Morelli, and Petr Stehlík. Bichromatic travelling waves for lattice Nagumo equations. SIAM Journal on Applied Dynamical Systems, 18(2):973– 1014, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Han:2014:DSS

[HN14] Xiaoying Han and Habib N.

Najm. Dynamical structures in stochastic chemical reaction systems. SIAM Journal on Applied Dynamical Systems, 13(3):1328–1351, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Ha:2016:SKO

[HNP16] Seung-Yeal Ha, Se Eun Noh, and Jinyeong Park. Synchronization of Kuramoto oscillators with adaptive couplings. SIAM Journal on Applied Dynamical Systems, 15(1):162– 194, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Honda:2021:RCI

[Hon21] Hirotada Honda. Reservoir computing with an inertial form. SIAM Journal on Applied Dynamical Systems, 20 (3):1320–1347, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Huls:2014:QAN

[HP14] Thorsten Hüls and Christian Pötzsche. Qualitative analysis of a nonautonomous Beverton–Holt Ricker model. SIAM Journal on Applied Dynamical Systems, 13(4):1442– 1488, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Holzer:2017:WSR

[HP17] Matt Holzer and Nikola Popović. Wavetrain solutions of a reaction–diffusion– advection model of mussel– algae interaction. SIAM Journal on Applied Dynamical Systems, 16(1):431–478, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Ha:2020:LTM

[HP20] Seung-Yeal Ha and Hansol Park. From the Lohe tensor model to the Lohe Hermitian sphere model and emergent dynamics. SIAM Journal on Applied Dynamical Systems, 19(2):1312–1342, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Huang:2020:SDT

[HQW^{+}20] Wen Huang, Hong Qian, Shirou Wang, Felix X.-F. Ye, and Yingfei Yi. Synchronization in discrete-time, discretestate random dynamical systems. SIAM Journal on Applied Dynamical Systems, 19 (1):233–251, ???? 2020. CO-DEN SJADAY. ISSN 1536-0040.

Henson:2003:PIP

[HRR^{+}03] Shandelle M. Henson, James R.

Reilly, Suzanne L. Robertson, Matthew C. Schu, Eric W. D. Rozier, and J. M. Cushing. Predicting irregularities in population cycles. SIAM Journal on Applied Dynamical Systems, 2(2):238– 253, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41126.

Higuera:2004:NRS

[HRS04] María Higuera, Hermann Riecke, and Mary Silber. Near-resonant steady mode interaction: Periodic, quasiperiodic, and localized patterns. SIAM Journal on Applied Dynamical Systems, 3 (4):463–502, 2004. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60055.

Huang:2019:BAM

[HRYZ19] Jicai Huang, Shigui Ruan, Pei Yu, and Yuyue Zhang. Bifurcation analysis of a mosquito population model with a saturated release rate of sterile mosquitoes. SIAM Journal on Applied Dynamical Systems, 18(2):939–972, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Holm:2003:WSN

[HS03] Darryl D. Holm and Martin F. Staley. Wave structure and nonlinear balances in a family of evolutionary PDEs. SIAM Journal on Applied Dynamical Systems, 2(3):323– 380, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41094.

Huber:2005:GSP

[HS05] A. Huber and P. Szmolyan. Geometric singular perturbation analysis of the Yamada model. SIAM Journal on Applied Dynamical Systems, 4 (3):607–648, 2005. CODEN SJADAY. ISSN 1536-0040.

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        60482.

Huang:2009:SSP

[HS09] Jianhua Huang and Wenxian Shen. Speeds of spread and propagation for KPP models in time almost and space periodic media. SIAM Journal on Applied Dynamical Systems, 8(3):790–821, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Haller:2010:LIA

[HS10a] George Haller and Themistoklis Sapsis. Localized instability and attraction along invariant manifolds. SIAM Journal on Applied Dynamical Systems, 9(2):611–633, ???? 2010. CODEN SJA-DAY. ISSN 1536-0040.

Hupkes:2010:TPS

[HS10b] Hermen Jan Hupkes and Björn Sandstede. Traveling pulse solutions for the discrete FitzHugh–Nagumo system. SIAM Journal on Applied Dynamical Systems, 9 (3):827–882, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Hsu:2015:TTM

[HS15] Sheng-Yi Hsu and Mau-Hsiang Shih. The tendency toward a moving equilibrium. SIAM Journal on Applied Dynamical Systems, 14(4):1699– 1730, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Hastings:2022:ISW

[HS22] S. P. Hastings and M. M.

Sussman. On the initiation of spiral waves in excitable media. SIAM Journal on Applied Dynamical Systems, 21(3):2241–2267, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M144791X.

Hakkaev:2013:SSS

[HSS13] Sevdzhan Hakkaev, Milena Stanislavova, and Atanas Stefanov. Spectral stability for subsonic traveling pulses of the Boussinesq “abc” system. SIAM Journal on Applied Dynamical Systems, 12(2):878– 898, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Hakkaev:2022:SPW

[HSS22] Sevdzhan Hakkaev, Milena Stanislavova, and Atanas Stefanov. On the stability of the periodic waves for the Benney system. SIAM Journal on Applied Dynamical Systems, 21(3):1726–1747, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1461630.

Haque:2023:DTL

[HSSY23] Sabina J. Haque, Matthew Satriano, Miruna-Stefana Sorea, and Polly Y. Yu. The disguised toric locus and affine equivalence of reaction networks. SIAM Journal on Applied Dynamical Systems, 22(2):1423–1444, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M149853X.

Hsu:2019:NSR

[Hsu19] Ting-Hao Hsu. Number and stability of relaxation oscillations for predator–prey systems with small death rates. SIAM Journal on Applied Dynamical Systems, 18(1):33– 67, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Harding:2024:IFD

[HSV24] Brendan Harding, Yvonne M. Stokes, and Rahil N. Valani. Inertial focusing dynamics of spherical particles in curved microfluidic ducts with a trapezoidal cross section. SIAM Journal on Applied Dynamical Systems, 23(3): 1805–1835, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1613220.

Huang:2020:NLC

[HTV20] Jianfeng Huang, Joan Torregrosa, and Jordi Villadelprat. On the number of limit cycles in generalized Abel equations. SIAM Journal on Applied Dynamical Systems, 19(4):2343– 2370, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Hu:2023:QPS

[Hu23]  Shengqing Hu.  Quasi-

periodic solutions for Schrödinger equation with finite smooth quasi-periodic forcing. SIAM Journal on Applied Dynamical Systems, 22(3):1945–1982, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M1523649.

Huls:2005:BCO

[Hül05] Thorsten Hüls. Bifurcation of connecting orbits with one nonhyperbolic fixed point for maps. SIAM Journal on Applied Dynamical Systems, 4 (4):985–1007, 2005. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
61467.

Huls:2016:CAC

[Hül16]  Thorsten Hüls.  A con-

tour algorithm for computing stable fiber bundles of nonautonomous, noninvertible maps. SIAM Journal on Applied Dynamical Systems, 15(2):923–951, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Harley:2014:ETW

[HvHM^{+}14] K. Harley, P. van Heijster, R. Marangell, G. J. Pettet, and M. Wechselberger. Existence of traveling wave solutions for a model of tumor invasion. SIAM Journal on Applied Dynamical Systems, 13 (1):366–396, ???? 2014. CO-DEN SJADAY. ISSN 1536-0040.

Ilak:2010:MRN

[IBB^{+}10] Miloš Ilak, Shervin Bagheri, Luca Brandt, Clarence W. Rowley, and Dan S. Henningson. Model reduction of the nonlinear complex Ginzburg– Landau equation. SIAM Journal on Applied Dynamical Systems, 9(4):1284–1302, ???? 2010. CODEN SJA-DAY. ISSN 1536-0040.

Ikeda:2023:ESS

[Ike23] Hideo Ikeda. Existence and stability of standing spot solutions in a three-component FitzHugh–Nagumo system in R^{2}. SIAM Journal on Applied Dynamical Systems, 22 (2):951–995, ???? 2023.

CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M144027X.

Iannelli:2020:CBM

[ILM20] Andrea Iannelli, Mark Lowenberg, and Andrés Marcos. Computation of bifurcation margins based on robust control concepts. SIAM Journal on Applied Dynamical Systems, 19(3):1956–1992, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Inglis:2016:GFS

[IM16] J. Inglis and J. MacLaurin. A general framework for stochastic traveling waves and patterns, with application to neural field equations. SIAM Journal on Applied Dynamical Systems, 15(1):195– 234, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Insperger:2015:STD

[IMS15] Tamas Insperger, John Milton, and Gabor Stepan. Semidiscretization for timedelayed neural balance control. SIAM Journal on Applied Dynamical Systems, 14 (3):1258–1277, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Ippei:2011:CAV

[Ipp11] Obayashi Ippei. Computerassisted verification method for invariant densities and rates of decay of correlations. SIAM Journal on Applied Dynamical Systems, 10 (2):788–816, ???? 2011. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.
                                        siam.org/siads/resource/
                                        1/sjaday/v10/i2/p788_s1.

Iuorio:2019:SPA

[IPS19] Annalisa Iuorio, Nikola Popović, and Peter Szmolyan. Singular perturbation analysis of a regularized MEMS model. SIAM Journal on Applied Dynamical Systems, 18(2):661–708, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Iooss:2022:PQS

[IR22] Gerard Iooss and Alastair M. Rucklidge. Patterns and quasipatterns from the superposition of two hexagonal lattices. SIAM Journal on Applied Dynamical Systems, 21(2):1119–1165, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1372780.

Issa:2017:DCM

[IS17] Tahir Bachar Issa and Wenxian Shen. Dynamics in chemotaxis models of parabolicelliptic type on bounded domain with time and space dependent logistic sources. SIAM Journal on Applied Dynamical Systems, 16(2):926– 973, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Ivanov:2024:TCD

[IS24] Milen Ivanov and Björn Sandstede. Truncation of contact defects in reaction-diffusion systems. SIAM Journal on Applied Dynamical Systems, 23(1):26–49, January 2024. ISSN 1536-0040.

Iyaniwura:2021:SOD

[IW21] Sarafa A. Iyaniwura and Michael J. Ward. Synchrony and oscillatory dynamics for a 2-D PDE-ODE model of diffusion-mediated communication between small signaling compartments. SIAM Journal on Applied Dynamical Systems, 20(1):438–499, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Iturriaga:2023:DWK

[IW23] Renato Iturriaga and Kaizhi Wang. A discrete weak KAM method for first-order stationary mean field games. SIAM Journal on Applied Dynamical Systems, 22(2): 1253–1274, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1508212.

Jackson:2006:PNC

[Jac06] Russell K. Jackson. Pulses in nonlinearly coupled Schrödinger equations I. A homoclinic flip bifurcation. SIAM Journal on Applied Dynamical Systems, 5(4):575–597, 2006. CODEN SJADAY. ISSN 1536-0040.

James:2010:ASO

[Jam10] J. D. Mireles James. Adaptive set-oriented computation of topological horseshoe factors in area and volume preserving maps. SIAM Journal on Applied Dynamical Systems, 9(4):1164–1200, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Jeffrey:2009:TFS

[JC09] M. R. Jeffrey and A. Colombo. The two-fold singularity of discontinuous vector fields. SIAM Journal on Applied Dynamical Systems, 8(2):624– 640, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Joshi:2023:RNM

[JC23] Badal Joshi and Gheorghe Craciun. Reaction network motifs for static and dynamic absolute concentration robustness. SIAM Journal on Applied Dynamical Systems, 22(2):501–526, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1476162.

Jeffrey:2014:DSI

[Jef14] Mike R. Jeffrey. Dynamics at a switching intersection: Hierarchy, isonomy, and multiple sliding. SIAM Journal on Applied Dynamical Systems, 13(3):1082–1105, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Jelbart:2024:RBI

[Jel24] Samuel Jelbart. Rate and bifurcation induced transitions in asymptotically slowfast systems. SIAM Journal on Applied Dynamical Systems, 23(3):1836–1869, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/24M1632000.

Jung:2020:ECS

[JJ20] Jinwook Jung and Shi Jin.

Emergence of the consensus and separation in an agentbased model with attractive and singular repulsive forces. SIAM Journal on Applied Dynamical Systems, 19(3):2103– 2134, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Jardon-Kojakhmetov:2024:PSH

[JKKL24] Hildeberto Jardón-Kojakhmetov, Christian Kuehn, and Iacopo P. Longo. Persistent synchronization of heterogeneous networks with timedependent linear diffusive coupling. SIAM Journal on Applied Dynamical Systems, 23(2):1540–1578, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1602024.

Juang:2008:SCC

[JL08] Jonq Juang and Yu-Hao Liang. Synchronous chaos in coupled map lattices with general connectivity topology.

SIAM Journal on Applied Dynamical Systems, 7(3):755– 765, 2008. CODEN SJADAY. ISSN 1536-0040.

James:2010:CHA

[JL10] J. D. Mireles James and Hector Lomelí. Computation of heteroclinic arcs with application to the volume preserving Hénon family. SIAM Journal on Applied Dynamical Systems, 9(3):919–953, ???? 2010. CODEN SJA-DAY. ISSN 1536-0040.

Johnson:2021:NMG

[JLG21] Carter L. Johnson, Timothy J. Lewis, and Robert Guy. Neuromechanical mechanisms of gait adaptation in C. elegans: Relative roles of neural and mechanical coupling. SIAM Journal on Applied Dynamical Systems, 20 (2):1022–1052, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

James:2013:RPC

[JM13] J. D. Mireles James and Konstantin Mischaikow. Rigorous a posteriori computation of (un)stable manifolds and connecting orbits for analytic maps. SIAM Journal on Applied Dynamical Systems, 12 (2):957–1006, ???? 2013. CO-DEN SJADAY. ISSN 1536-0040.

Jimenez:2013:LCD

[JMB^{+}13] Nicolas D. Jimenez, Stefan Mihalas, Richard Brown, Ernst Niebur, and Jonathan Rubin. Locally contractive dynamics in generalized integrate-and-fire neurons. SIAM Journal on Applied Dynamical Systems, 12 (3):1474–1514, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Jorba:2009:CRI

[JO09] Àngel Jorba and Estrella Olmedo. On the computation of reducible invariant tori on a parallel computer. SIAM Journal on Applied Dynamical Systems, 8(4):1382–1404, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Jelbart:2022:POG

[JPK^{+}22] Samuel Jelbart, Nathan Pages, Vivien Kirk, James Sneyd, and Martin Wechselberger. Process-oriented geometric singular perturbation theory and calcium dynamics. SIAM Journal on Applied Dynamical Systems, 21(2):982–1029, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1412402.

Josic:2005:DIA

[JR05] Kresimir Josic and Jonathan Rubin. Deriving information about architecture from activity patterns in coupled cell systems. SIAM Journal on Applied Dynamical Systems, 4(1):53–77, 2005. CODEN SJADAY. ISSN 1536-0040.

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        60758.

Jacobs:2006:RPD

[JS06] Frans Jacobs and Sebastian J. Schreiber. Random perturbations of dynamical systems with absorbing states. SIAM Journal on Applied Dynamical Systems, 5(2):293–312, 2006. CODEN SJADAY. ISSN 1536-0040.

Joshi:2017:WSR

[JS17] Badal Joshi and Anne Shiu. Which small reaction networks are multistationary? SIAM Journal on Applied Dynamical Systems, 16(2):802– 833, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Juda:2020:UFL

[Jud20] Mateusz Juda. Unsupervised features learning for sampled vector fields. SIAM Journal on Applied Dynamical Systems, 19(4):2720–2736, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Johnson:2011:NSP

[JZ11] Mathew A. Johnson and Kevin Zumbrun. Nonlinear stability of periodic traveling-wave solutions of viscous conservation laws in dimensions one and two. SIAM Journal on Applied Dynamical Systems, 10(1): 189–211, ???? 2011. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.

siam.org/siads/resource/
1/sjaday/v10/i1/p189_s1.

Kilpatrick:2010:BRC

[KB10] Zachary P. Kilpatrick and Paul C. Bressloff. Binocular rivalry in a competitive neural network with synaptic depression. SIAM Journal on Applied Dynamical Systems, 9(4):1303–1347, ???? 2010. CODEN SJADAY. ISSN 1536-0040. URL http:

//epubs.siam.org/siads/
resource/1/sjaday/v9/i4/
p1303_s1.

Kirtland:2023:UMA

[KBGD^{+}23] Aaron Kirtland, Jonah Botvinick-Greenhouse, Marianne DeBrito, Megan Osborne, Casey Johnson, Robert S. Martin, Samuel J. Araki, and Daniel Q. Eckhardt. An unstructured mesh approach to nonlinear noise reduction for coupled systems. SIAM Journal on Applied Dynamical Systems, 22(4):2927–2944, October 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M152092X.

Kristiansen:2014:IMA

[KBS14] K. Uldall Kristiansen, M. Brøns, and J. Starke. An iterative method for the approximation of fibers in slow-fast systems. SIAM Journal on Applied Dynamical Systems, 13 (2):861–900, ???? 2014. CO-DEN SJADAY. ISSN 1536-0040.

Kamei:2013:CBE

[KC13] Hiroko Kamei and Peter J. A. Cock. Computation of balanced equivalence relations and their lattice for a coupled Cell network. SIAM Journal on Applied Dynamical Systems, 12(1):352–382, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Kan:2013:SSI

[KDKR13] Xingye Kan, Jinqiao Duan, Ioannis G. Kevrekidis, and Anthony J. Roberts. Simulating stochastic inertial manifolds by a backward-forward approach. SIAM Journal on Applied Dynamical Systems, 12(1):487–514, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Kazanci:2008:WFT

[KE08] Fatma Gürel Kazanci and Bard Ermentrout. Wave formation through the interactions between clustered states and local coupling in arrays of neural oscillators. SIAM Journal on Applied Dynamical Systems, 7(2):491–509, 2008. CODEN SJADAY. ISSN 1536-0040.

Kilpatrick:2013:WBS

[KE13] Zachary P. Kilpatrick and Bard Ermentrout. Wandering bumps in stochastic neural fields. SIAM Journal on Applied Dynamical Systems, 12(1):61–94, ???? 2013. CO-DEN SJADAY. ISSN 1536-0040.

Kilpatrick:2014:PBS

[KF14] Zachary P. Kilpatrick and Grégory Faye. Pulse bifurcations in stochastic neural fields. SIAM Journal on Applied Dynamical Systems, 13 (2):830–860, ???? 2014. CO-DEN SJADAY. ISSN 1536-0040.

Kilpatrick:2008:TPW

[KFB08] Zachary P. Kilpatrick, Stefanos E. Folias, and Paul C. Bressloff. Traveling pulses and wave propagation failure in inhomogeneous neural media. SIAM Journal on Applied Dynamical Systems, 7 (1):161–185, 2008. CODEN SJADAY. ISSN 1536-0040.

Kutz:2016:MDM

[KFB16] J. Nathan Kutz, Xing Fu, and Steven L. Brunton. Multiresolution dynamic mode decomposition. SIAM Journal on Applied Dynamical Systems, 15(2):713–735, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Kramer:2017:SSD

[KGB^{+}17] Boris Kramer, Piyush Grover, Petros Boufounos, Saleh Nabi, and Mouhacine Benosman. Sparse sensing and DMD-based identification of flow regimes and bifurcations in complex flows. SIAM Journal on Applied Dynamical Systems, 16(2):1164–1196, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Kvalheim:2021:CFT

[KGK21] Matthew D. Kvalheim, Paul Gustafson, and Daniel E. Koditschek. Conley’s fundamental theorem for a class of hybrid systems. SIAM Journal on Applied Dynamical Systems, 20(2):784–825, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Kristiansen:2015:UBS

[KH15a] K. Uldall Kristiansen and S. J. Hogan. On the use of blowup to study regularizations of singularities of piecewise smooth dynamical systems in R^{3}. SIAM Journal on Applied Dynamical Systems, 14(1):382–422, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Kristiansen:2015:RTF

[KH15b] K. Uldall Kristiansen and S. J. Hogan. Regularizations of two-fold bifurcations in planar piecewise smooth systems using blowup. SIAM Journal on Applied Dynamical Systems, 14(4):1731–1786, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Kristiansen:2018:ECP

[KH18a] K. Uldall Kristiansen and S. J. Hogan. Erratum: Le Canard de Painlevé. SIAM Journal on Applied Dynamical Systems, 17(4):2915–2916, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040. See [KH18b].

Kristiansen:2018:CP

[KH18b] K. Uldall Kristiansen and S. J. Hogan. Le canard de Painlevé. SIAM Journal on Applied Dynamical Systems, 17(1):859–908, ???? 2018. CODEN SJADAY. ISSN 1536-0040. See erratum [KH18a].

Kawan:2021:SAD

[KHG21] Christoph Kawan, Sigurdur Hafstein, and Peter Giesl. A subgradient algorithm for data-rate optimization in the remote state estimation problem. SIAM Journal on Applied Dynamical Systems, 20 (4):2142–2173, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Kiers:2020:RIT

[Kie20] Claire Kiers. Rate-induced tipping in discrete-time dynamical systems. SIAM Journal on Applied Dynamical Systems, 19(2):1200–1224, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Kim:2020:SDD

[Kim20] Dohyun Kim. State-dependent dynamics of the Lohe matrix ensemble on the unitary group under the gradient flow. SIAM Journal on Applied Dynamical Systems, 19(2):1080–1123, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Kaklamanos:2019:RGP

[KK19] Panagiotis Kaklamanos and Kristian U. Kristiansen. Regularization and geometry of piecewise smooth systems with intersecting discontinuity sets. SIAM Journal on Applied Dynamical Systems, 18(3):1225–1264, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Kamb:2020:TDO

[KKBK20] Mason Kamb, Eurika Kaiser, Steven L. Brunton, and J. Nathan Kutz. Timedelay observables for Koopman: Theory and applications. SIAM Journal on Applied Dynamical Systems, 19 (2):886–917, ???? 2020. CO-DEN SJADAY. ISSN 1536-0040.

Kapitula:2006:TCS

[KKC06] Todd Kapitula, P. G. Kevrekidis, and Zhigang Chen. Three is a crowd: Solitary waves in photorefractive media with three potential wells. SIAM Journal on Applied Dynamical Systems, 5(4):598–633, 2006. CODEN SJADAY. ISSN 1536-0040.

Kalies:2018:ACL

[KKJ18] William D. Kalies, Shane Kepley, and J. D. Mireles James.

Analytic continuation of local (un)stable manifolds with rigorous computer assisted error bounds. SIAM Journal on Applied Dynamical Systems, 17(1):157–202, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Kalise:2020:RFC

[KKK20] Dante Kalise, Sudeep Kundu, and Karl Kunisch. Robust feedback control of nonlinear PDEs by numerical approximation of high-dimensional Hamilton– Jacobi–Isaacs equations. SIAM Journal on Applied Dynamical Systems, 19(2):1496–1524, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Keane:2015:DFV

[KKP15] Andrew Keane, Bernd Krauskopf, and Claire Postlethwaite. Delayed feedback versus seasonal forcing: Resonance phenomena in an El Ninño southern oscillation model. SIAM Journal on Applied Dynamical Systems, 14(3):1229–1257, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Keane:2016:IIB

[KKP16] Andrew Keane, Bernd Krauskopf, and Claire Postlethwaite. Investigating irregular behavior in a model for the El Niño Southern Oscillation with positive and negative delayed feedback. SIAM Journal on Applied Dynamical Systems, 15(3):1656–1689, ???? 2016.

CODEN SJADAY. ISSN 1536-0040.

Kalies:2018:AAL

[KKV18] William D. Kalies, Dinesh Kasti, and Robert Vandervorst. An algorithmic approach to lattices and order in dynamics. SIAM Journal on Applied Dynamical Systems, 17(2):1617–1649, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Kitanov:2017:DMS

[KL17] Petko M. Kitanov and Victor G. LeBlanc. Dynamics of meandering spiral waves with weak lattice perturbations. SIAM Journal on Applied Dynamical Systems, 16 (1):16–53, ???? 2017. CO-DEN SJADAY. ISSN 1536-0040.

Kuehn:2021:UQB

[KL21] Christian Kuehn and Kerstin Lux. Uncertainty quantification of bifurcations in random ordinary differential equations. SIAM Journal on Applied Dynamical Systems, 20(4):2295–2334, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Kapitula:2010:IES

[KLK10] Todd Kapitula, Kody J. H. Law, and Panayotis G. Kevrekidis. Interaction of excited states in two-species Bose–Einstein condensates: a case study. SIAM Journal on Applied Dynamical Systems, 9(1):34–61, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Kitanov:2013:DHB

[KLW13] P. M. Kitanov, W. Langford, and A. R. Willms. Double Hopf bifurcation with Huygens symmetry. SIAM Journal on Applied Dynamical Systems, 12(1):126–174, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Kimura:2008:NVT

[KM08] Margot Kimura and Jeff Moehlis. Novel vehicular trajectories for collective motion from coupled oscillator steering control. SIAM Journal on Applied Dynamical Systems, 7(4):1191–1212, 2008. CO-DEN SJADAY. ISSN 1536-0040.

Kastsian:2010:ROF

[KM10] Darya Kastsian and Martin Mönnigmann. Robust optimization of fixed points of nonlinear discrete time systems with uncertain parameters. SIAM Journal on Applied Dynamical Systems, 9 (2):357–390, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Kuehn:2017:GPH

[KM17] Christian Kuehn and Christian Münch. Generalized play hysteresis operators in limits of fast-slow systems. SIAM Journal on Applied Dynamical Systems, 16(3):1650–1685, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Kuehn:2024:PBT

[KM24] Christian Kuehn and Jan Mölter. Preserving bifurcations through moment closures. SIAM Journal on Applied Dynamical Systems, 23(1):791–812, March 2024. ISSN 1536-0040.

Kintos:2014:MEH

[KN14] Nickolas Kintos and Farzan Nadim. A modeling exploration of how synaptic feedback to descending projection neurons shapes the activity of an oscillatory network. SIAM Journal on Applied Dynamical Systems, 13(3):1239–1269, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Kalmar-Nagy:2011:DHR

[KNWH11] Tamás Kalmár-Nagy, Pankaj Wahi, and Abhishek Halder. Dynamics of a hysteretic relay oscillator with periodic forcing. SIAM Journal on Applied Dynamical Systems, 10 (2):403–422, ???? 2011. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.
                                        siam.org/siads/resource/
                                        1/sjaday/v10/i2/p403_s1.

Krauskopf:2003:CGL

[KO03] Bernd Krauskopf and Hinke M. Osinga. Computing geodesic level sets on global (un)stable manifolds of vector fields. SIAM Journal on Applied Dynamical Systems, 2(4):546– 569, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60018.

Kristiansen:2024:CDK

[KO24] K. U. Kristiansen and R. Ortega. Circularization in the damped Kepler problem. SIAM Journal on Applied Dynamical Systems, 23 (2):1372–1410, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1623720.

Krauskopf:2007:UCC

[KOP07] Bernd Krauskopf, Hinke M.

Osinga, and Bruce B. Peckham. Unfolding the cusp-cusp bifurcation of planar endomorphisms. SIAM Journal on Applied Dynamical Systems, 6(2):403–440, 2007. CODEN SJADAY. ISSN 1536-0040.

Kristiansen:2023:MMO

[KP23] Kristian Uldall Kristiansen and Morten Gram Pedersen. Mixed-mode oscillations in coupled FitzHugh– Nagumo oscillators: Blowup analysis of cusped singularities. SIAM Journal on Applied Dynamical Systems, 22(2):1383–1422, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1480495.

Kairzhan:2019:DSS

[KPG19] Adilbek Kairzhan, Dmitry E. Pelinovsky, and Roy H. Goodman. Drift of spectrally stable shifted states on star graphs. SIAM Journal on Applied Dynamical Systems, 18(4):1723–1755, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Krupa:2008:MMO

[KPK08] Martin Krupa, Nikola Popović, and Nancy Kopell. Mixedmode oscillations in three time-scale systems: a prototypical example. SIAM Journal on Applied Dynamical Systems, 7(2):361–420, 2008. CODEN SJADAY. ISSN 1536-0040.

Kaklamanos:2023:GSP

[KPK23] Panagiotis Kaklamanos, Nikola Popović, and Kristian U. Kristiansen. Geometric singular perturbation analysis of the multiple-timescale Hodgkin–Huxley equations. SIAM Journal on Applied Dynamical Systems, 22(3): 1552–1589, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1477477.

Kristiansen:2011:UMT

[KPR11] K. Uldall Kristiansen, P. Palmer, and M. Roberts. A unification of models of tethered satellites. SIAM Journal on Applied Dynamical Systems, 10(3):1042– 1069, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i3/p1042_s1.

Kirk:2012:RBR

[KPR12a] Vivien Kirk, Claire Postlethwaite, and Alastair M. Rucklidge. Resonance bifurcations of robust heteroclinic networks. SIAM Journal on Applied Dynamical Systems, 11(4):1360–1401, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Kristiansen:2012:PSM

[KPR12b] K. Uldall Kristiansen, P. Palmer, and R. M. Roberts. The persistence of a slow manifold with bifurcation. SIAM Journal on Applied Dynamical Systems, 11(2):661–683, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Knipl:2015:RBS

[KPR15] Diána H. Knipl, Pawel Pilarczyk, and Gergely Röst. Rich bifurcation structure in a two-patch vaccination model. SIAM Journal on Applied Dynamical Systems, 14(2):980– 1017, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Kevrekidis:2013:NSS

[KPT13] Panayotis G. Kevrekidis, Dmitry E. Pelinovsky, and Dmitry Tyugin. Nonlinear stationary states in PTsymmetric lattices. SIAM Journal on Applied Dynamical Systems, 12(3):1210–1236, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Kramer:2017:FCS

[KPW17] Boris Kramer, Benjamin Peherstorfer, and Karen Willcox. Feedback control for systems with uncertain parameters using online-adaptive reduced models. SIAM Journal on Applied Dynamical Systems, 16(3):1563–1586, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Kolokolnikov:2011:SRS

[KR11] Theodore Kolokolnikov and Xiaofeng Ren. Smokering solutions of Gierer– Meinhardt system in R^{3}. SIAM Journal on Applied Dynamical Systems, 10(1): 251–277, ???? 2011. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.
                                        siam.org/siads/resource/
                                        1/sjaday/v10/i1/p251_s1.

Kamei:2021:RLS

[KR21] Hiroko Kamei and Haibo Ruan. Reduced lattices of synchrony subspaces and their indices. SIAM Journal on Applied Dynamical Systems, 20(2):636–670, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Kramer:2021:SDQ

[Kra21] Boris Kramer. Stability domains for quadratic-bilinear reduced-order models. SIAM Journal on Applied Dynamical Systems, 20(2):981–996, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Kristiansen:2015:CST

[Kri15] K. Uldall Kristiansen. Computation of saddle-type slow manifolds using iterative methods. SIAM Journal on Applied Dynamical Systems, 14(2):1189–1227, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Kristiansen:2020:PBF

[Kri20] Kristian U. Kristiansen. On the pitchfork bifurcation of the folded node and other unbounded time-reversible connection problems in R^{3}. SIAM Journal on Applied Dynamical Systems, 19(3):2059– 2102, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Kristiansen:2021:SOU

[Kri21] Kristian U. Kristiansen.

A stiction oscillator under slowly varying forcing: Uncovering small scale phenomena using blowup. SIAM Journal on Applied Dynamical Systems, 20(4):2359–2390, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Kloc:2014:SHS

[KRK14] M. Kloc and V. Rom-Kedar.

Smooth Hamiltonian systems with soft impacts. SIAM Journal on Applied Dynamical Systems, 13(3):1033–1059, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Kitavtsev:2012:AST

[KRW12] Georgy Kitavtsev, Lutz Recke, and Barbara Wagner. Asymptotics for the spectrum of a thin film equation in a singular limit. SIAM Journal on Applied Dynamical Systems, 11(4):1425–1457, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Knipl:2013:ESV

[KRW13] Diána H. Knipl, Gergely Röst, and Jianhong Wu. Epidemic spread and variation of peak times in connected regions due to travel-related infections — dynamics of an antigravity-type delay differential model. SIAM Journal on Applied Dynamical Systems, 12(4):1722–1762, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Kollar:2007:CSG

[KS07] Richard Kollár and Arnd Scheel. Coherent structures generated by inhomogeneities in oscillatory media. SIAM Journal on Applied Dynamical Systems, 6(1):236–262, 2007. CODEN SJADAY. ISSN 1536-0040.

Kosiuk:2011:SSP

[KS11] Ilona Kosiuk and Peter Szmolyan. Scaling in singular perturbation problems: Blowing up a relaxation oscillator. SIAM Journal on Applied Dynamical Systems, 10(4):1307– 1343, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i4/p1307_s1.

Kruger:2014:FPS

[KS14] J. Krüger and W. Stannat.

Front propagation in stochastic neural fields: a rigorous mathematical framework. SIAM Journal on Applied Dynamical Systems, 13(3):1293– 1310, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Karst:2014:SOS

[KSG14] Nathaniel J. Karst, Brian D.

Storey, and John B. Geddes. Spontaneous oscillations in simple fluid networks. SIAM Journal on Applied Dynamical Systems, 13(1):157– 180, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Karamched:2020:BEA

[KSKJ20] Bhargav Karamched, Simon Stolarczyk, Zachary P. Kilpatrick, and Kresimir Josić. Bayesian evidence accumulation on social networks. SIAM Journal on Applied Dynamical Systems, 19(3):1884– 1919, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Kolokolnikov:2006:SSG

[KSWW06] Theodore Kolokolnikov, Wentao Sun, Michael Ward, and Juncheng Wei. The stability of a stripe for the Gierer– Meinhardt model and the effect of saturation. SIAM Journal on Applied Dynamical Systems, 5(2):313–363, 2006. CODEN SJADAY. ISSN 1536-0040.

Kreider:2024:RBI

[KT24] Max Kreider and Peter J. Thomas. On the relation between infinitesimal shape response curves and phaseamplitude reduction for single and coupled limit-cycle oscillators. SIAM Journal on Applied Dynamical Systems, 23(2):1636–1676, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1575159.

Kolokolnikov:2020:TVF

[KTK20] Theodore Kolokolnikov, Chris Ticknor, and Panayotis Kevrekidis. Toroidal vortex filament knots and links: Existence, stability and dynamics. SIAM Journal on Applied Dynamical Systems, 19(4):2403–2427, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Kulkarni:2016:TES

[Kul16] Varsha S. Kulkarni. Temporal evolution of social innovation: What matters? SIAM Journal on Applied Dynamical Systems, 15(3):1485–1500, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Kurihara:2017:DEO

[Kur17] Eru Kurihara. Dynamical equations for oscillating nonspherical bubbles with nonlinear interactions. SIAM Journal on Applied Dynamical Systems, 16(1):139–158, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Kimrey:2020:BDH

[KVB20] Joshua Kimrey, Theodore Vo, and Richard Bertram. Big ducks in the heart: Canard analysis can explain large early afterdepolarizations in cardiomyocytes. SIAM Journal on Applied Dynamical Systems, 19(3):1701–1735, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Kimrey:2022:CUB

[KVB22] Joshua Kimrey, Theodore Vo, and Richard Bertram. Canards underlie both electrical and Ca^{2+}-induced early afterdepolarizations in a model for cardiac myocytes. SIAM Journal on Applied Dynamical Systems, 21(2):1059–1091, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M147757X.

Krupa:2012:MMO

[KVDC12] Martin Krupa, Alexandre Vidal, Mathieu Desroches, and Frédérique Clément. Mixedmode oscillations in a multiple time scale phantom bursting system. SIAM Journal on Applied Dynamical Systems, 11(4):1458–1498, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Kunisch:2004:HPB

[KVX04] K. Kunisch, S. Volkwein, and L. Xie. HJB-POD-based feedback design for the optimal control of evolution problems. SIAM Journal on Applied Dynamical Systems, 3(4):701– 722, 2004. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/60048.

Knobloch:2008:SMH

[KW08] J. Knobloch and T. Wagenknecht. Snaking of multiple homoclinic orbits in reversible systems. SIAM Journal on Applied Dynamical Systems, 7(4):1397–1420, 2008. CODEN SJADAY. ISSN 1536-0040.

Kalita:2021:RFO

[KZ21] Piotr Kalita and Piotr Zgliczyński. Rigorous FEM for one-dimensional Burgers equation. SIAM Journal on Applied Dynamical Systems, 20(2):853–907, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Louca:2013:SSI

[LA13] Stilianos Louca and Fatihcan M. Atay. Stationary states in infinite networks of spiking oscillators with noise. SIAM Journal on Applied Dynamical Systems, 12(1):415– 449, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Lu:2018:SPD

[LA18] Wenlian Lu and Fatihcan M.

Atay. Stability of phase difference trajectories of networks of Kuramoto oscillators with time-varying couplings and intrinsic frequencies. SIAM Journal on Applied Dynamical Systems, 17 (1):457–483, ???? 2018. CO-DEN SJADAY. ISSN 1536-0040.

Laing:2005:SWN

[Lai05] Carlo R. Laing. Spiral waves in nonlocal equations. SIAM Journal on Applied Dynamical Systems, 4(3):588–606, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/61289.

Laing:2015:ENF

[Lai15] Carlo R. Laing. Exact neural fields incorporating gap junctions. SIAM Journal on Applied Dynamical Systems, 14(4):1899–1929, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Laing:2017:CTD

[Lai17] Carlo R. Laing. Chimeras in two-dimensional domains: Heterogeneity and the continuum limit. SIAM Journal on Applied Dynamical Systems, 16(2):974–1014, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Lang:2016:MAT

[Lan16] Eva Lang. A multiscale analysis of traveling waves in stochastic neural fields. SIAM Journal on Applied Dynamical Systems, 15(3):1581–1614, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Lasagna:2018:SAC

[Las18] Davide Lasagna. Sensitivity analysis of chaotic systems using unstable periodic orbits. SIAM Journal on Applied Dynamical Systems, 17 (1):547–580, ???? 2018. CO-DEN SJADAY. ISSN 1536-0040.

Lawley:2016:BVP

[Law16] Sean D. Lawley. Boundary value problems for statistics of diffusion in a randomly switching environment: PDE and SDE perspectives. SIAM Journal on Applied Dynamical Systems, 15(3):1410–1433, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Lopez:2005:RPS

[LBHM05] Vanessa López, Philip Boyland, Michael T. Heath, and Robert D. Moser. Relative periodic solutions of the complex Ginzburg–Landau equation. SIAM Journal on Applied Dynamical Systems, 4 (4):1042–1075, 2005. CODEN SJADAY. ISSN 1536-0040.

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        61897.

Lu:2018:ECB

[LBR18] Jiajun Lu, Frank Baginski, and Xiaofeng Ren. Equilibrium configurations of boundary droplets in a selforganizing inhibitory system. SIAM Journal on Applied Dynamical Systems, 17(2):1353– 1376, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Linaro:2012:CTH

[LCDS12] Daniele Linaro, Alan Champneys, Mathieu Desroches, and Marco Storace. Codimensiontwo homoclinic bifurcations underlying spike adding in the Hindmarsh–Rose burster. SIAM Journal on Applied Dynamical Systems, 11(3):939– 962, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Lee:2008:TBG

[LCKO08] Clare M. Lee, Pieter J.

Collins, Bernd Krauskopf, and Hinke M. Osinga. Tangency bifurcations of global Poincaré maps. SIAM Journal on Applied Dynamical Systems, 7(3):712–754, 2008. CODEN SJADAY. ISSN 1536-0040.

Landry:2005:DIP

[LCMA05] Maria Landry, Sue Ann Campbell, Kirsten Morris, and Cesar O. Aguilar. Dynamics of an inverted pendulum with delayed feedback control. SIAM Journal on Applied Dynamical Systems, 4(2):333–351, 2005. CODEN SJADAY. ISSN 1536-0040.

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        60046.

Li:2018:SCA

[LD18] Mingwu Li and Harry Dankowicz. Staged construction of adjoints for constrained optimization of integro-differential boundaryvalue problems. SIAM Journal on Applied Dynamical Systems, 17(2):1117–1151, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Liao:2020:EDF

[LDB20] Guangyuan Liao, Casey Diekman, and Amitabha Bose. Entrainment dynamics of forced hierarchical circadian systems revealed by 2dimensional maps. SIAM Journal on Applied Dynamical Systems, 19(3):2135–2161, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Llibre:2009:SSN

[LdST09] Jaume Llibre, Paulo R. da Silva, and Marco A. Teixeira. Study of singularities in nonsmooth dynamical systems via singular perturbation. SIAM Journal on Applied Dynamical Systems, 8 (1):508–526, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Ly:2010:ARN

[LE10] Cheng Ly and G. Bard Ermentrout. Analysis of recurrent networks of pulsecoupled noisy neural oscillators. SIAM Journal on Applied Dynamical Systems, 9 (1):113–137, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Lee:2022:SSP

[Lee22] Euiwoo Lee. Stable synchronized periodic solutions in a pair of mutually inhibitory model neurons with conduction delays. SIAM Journal on Applied Dynamical Systems, 21(1):166–203, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M140064X.

Lega:2011:CBT

[Leg11] Joceline Lega. Collective behaviors in two-dimensional systems of interacting particles. SIAM Journal on Applied Dynamical Systems, 10 (4):1213–1231, ???? 2011. CODEN SJADAY. ISSN 1536-0040. URL http:

//epubs.siam.org/siads/
resource/1/sjaday/v10/i4/

p1213_s1. See erratum [Leg13].

Lega:2013:ECB

[Leg13]  Joceline Lega.  Erratum:

Collective behaviors in twodimensional systems of interacting particles. SIAM Journal on Applied Dynamical Systems, 12(4):2093, ???? 2013. CODEN SJADAY. ISSN 1536-0040. See [Leg11].

Lakshmi:2020:FEP

[LFFC^{+}20] Mayur V. Lakshmi, Giovanni Fantuzzi, Jesús D. Fernández-Caballero, Yongyun Hwang, and Sergei I. Chernyshenko. Finding extremal periodic orbits with polynomial optimization, with application to a nine-mode model of shear flow. SIAM Journal on Applied Dynamical Systems, 19 (2):763–787, ???? 2020. CO-DEN SJADAY. ISSN 1536-0040.

Langfield:2017:BPS

[LFOG17] Peter Langfield, Wilson L. C. Façanha, Bart Oldeman, and Leon Glass. Bifurcations in a periodically stimulated limit cycle oscillator with finite relaxation times. SIAM Journal on Applied Dynamical Systems, 16(2):1045–1069, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Ling:2015:SBA

[LGLC15] Guang Ling, Zhi-Hong Guan, Rui-Quan Liao, and Xin-Ming Cheng. Stability and bifurcation analysis of cyclic genetic regulatory networks with mixed time delays. SIAM Journal on Applied Dynamical Systems, 14(1):202– 220, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Litvak-Hinenzon:2004:ESR

[LHRK04] Anna Litvak-Hinenzon and Vered Rom-Kedar. On energy surfaces and the resonance web. SIAM Journal on Applied Dynamical Systems, 3(4):525–573, 2004. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60010.

Lu:2023:OCC

[LHW23] Min Lu, Jicai Huang, and Hao Wang. An organizing center of codimension four in a predator-prey model with generalist predator: From tristability and quadristability to transients in a nonlinear environmental change. SIAM Journal on Applied Dynamical Systems, 22(2):694–729, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M1488466.

Lin:2006:ECP

[Lin06]  Kevin K. Lin.  Entrain-

ment and chaos in a pulsedriven Hodgkin–Huxley oscillator. SIAM Journal on Applied Dynamical Systems, 5 (2):179–204, 2006. CODEN SJADAY. ISSN 1536-0040.

Lawley:2015:NDR

[LK15] Sean D. Lawley and James P.

Keener. A new derivation of Robin boundary conditions through homogenization of a stochastically switching boundary. SIAM Journal on Applied Dynamical Systems, 14(4):1845–1867, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Louzeiro:2022:PSM

[LKH^{+}22] Mauricio Louzeiro, Christoph Kawan, Sigurdur Hafstein, Peter Giesl, and Jinyun Yuan. A projected subgradient method for the computation of adapted metrics for dynamical systems. SIAM Journal on Applied Dynamical Systems, 21(4):2610–2641, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M1475776.

Langfield:2015:FTB

[LKO15] Peter Langfield, Bernd Krauskopf, and Hinke M. Osinga. Forwardtime and backward-time isochrons and their interactions. SIAM Journal on Applied Dynamical Systems, 14 (3):1418–1453, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Lewis:2008:HRD

[LL08] Gregory M. Lewis and William F. Langford. Hysteresis in a rotating differentially heated spherical shell of Boussinesq fluid. SIAM Journal on Applied Dynamical Systems, 7(4):1421–1444, 2008. CODEN SJADAY. ISSN 1536-0040.

Lloyd:2019:IFO

[Llo19] David J. B. Lloyd. Invasion fronts outside the homoclinic snaking region in the planar Swift–Hohenberg equation. SIAM Journal on Applied Dynamical Systems, 18(4):1892–1933, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Lloyd:2021:HIF

[Llo21] David J. Lloyd. Hexagon invasion fronts outside the homoclinic snaking region in the planar Swift–Hohenberg equation. SIAM Journal on Applied Dynamical Systems, 20(2):671–700, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Lin:2013:PNP

[LLYZ13] Guojian Lin, Weishi Liu, Yingfei Yi, and Mingji Zhang. Poisson–Nernst–Planck systems for ion flow with a local hard-sphere potential for ion size effects. SIAM Journal on Applied Dynamical Systems, 12(3):1613–1648, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Li:2017:DBI

[LLZ17] Chunqiu Li, Desheng Li, and Zhijun Zhang. Dynamic bifurcation from infinity of nonlinear evolution equations. SIAM Journal on Applied Dynamical Systems, 16(4):1831–1868, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Labouriau:2016:HBT

[LM16] Isabel Salgado Labouriau and Adrian Calin Murza. Hopf bifurcation with tetrahedral and octahedral symmetry.

SIAM Journal on Applied Dynamical Systems, 15(1):125– 141, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Lauriere:2019:PND

[LM19] Mathieu Laurière and Laurent Mertz. Penalization of nonsmooth dynamical systems with noise: Ergodicity and asymptotic formulae for threshold crossings probabilities. SIAM Journal on Applied Dynamical Systems, 18 (2):853–880, ???? 2019. CO-DEN SJADAY. ISSN 1536-0040.

Li:2023:ASK

[LMKY23] Bian Li, Yian Ma, J. Nathan Kutz, and Xiu Yang. The adaptive spectral Koopman method for dynamical systems. SIAM Journal on Applied Dynamical Systems, 22 (3):1523–1551, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1487941.

Ledoux:2009:CSM

[LMNT09] Veerle Ledoux, Simon J. A. Malham, Jitse Niesen, and Vera Thümmler. Computing stability of multidimensional traveling waves. SIAM Journal on Applied Dynamical Systems, 8(1):480–507, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Longo:2021:RIT

[LNOR21] Iacopo P. Longo, Carmen Núñez, Rafael Obaya, and Martin Rasmussen. Rateinduced tipping and saddlenode bifurcation for quadratic differential equations with nonautonomous asymptotic dynamics. SIAM Journal on Applied Dynamical Systems, 20(1):500–540, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Llibre:2008:FPO

[LO08] Jaume Llibre and Rafael Ortega. On the families of periodic orbits of the Sitnikov problem. SIAM Journal on Applied Dynamical Systems, 7(2):561–576, 2008. CODEN SJADAY. ISSN 1536-0040.

Lin:2010:OID

[LO10] Jessica Lin and William Ott. Observing infinitedimensional dynamical systems. SIAM Journal on Applied Dynamical Systems, 9 (4):1229–1243, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Lustri:2018:NPT

[LP18] Christopher J. Lustri and Mason A. Porter. Nanoptera in a period-2 Toda chain. SIAM Journal on Applied Dynamical Systems, 17(2):1182–1212, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Leventides:2021:KOD

[LP21] John Leventides and Costas Poulios. Koopman operators and the 3x+1-dynamical system. SIAM Journal on Applied Dynamical Systems, 20 (4):1773–1813, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Le:2022:PWM

[LP22] Uyen Le and Dmitry E. Pelinovsky. Periodic waves of the modified KdV equation as minimizers of a new variational problem. SIAM Journal on Applied Dynamical Systems, 21(4):2518–2534, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/21M1465329.

Lee:2023:ODM

[LP23] Minwoo Lee and Jongho Park. An optimized dynamic mode decomposition model robust to multiplicative noise. SIAM Journal on Applied Dynamical Systems, 22(1):235–268, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1443832.

Lois-Prados:2022:BSD

[LPH22] Cristina Lois-Prados and Frank M. Hilker. Bifurcation sequences in a discontinuous piecewise-smooth map combining constant-catch and threshold-based harvesting strategies. SIAM Journal on Applied Dynamical Systems, 21(1):470–499, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1416515.

Lucken:2015:CCD

[LPK15] L. Lücken, J. P. Pade, and K. Knauer. Classification of coupled dynamical systems with multiple delays: Finding the minimal number of delays. SIAM Journal on Applied Dynamical Systems, 14 (1):286–304, ???? 2015. CO-DEN SJADAY. ISSN 1536-0040.

Lakrad:2022:QDI

[LPR22] Faouzi Lakrad, Jamol Pender, and Richard Rand. Queues with delayed information: A dynamical systems perspective. SIAM Journal on Applied Dynamical Systems, 21(1):676–713, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/20M1334358.

Luque:2013:MCP

[LPS13] Alejandro Luque and Daniel Peralta-Salas. Motion of charged particles in ABC magnetic fields. SIAM Journal on Applied Dynamical Systems, 12(4):1889–1947, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Letson:2018:NFO

[LR18] Benjamin Letson and Jonathan E. Rubin. A new frame for an old (phase) portrait: Finding rivers and other flow features in the plane. SIAM Journal on Applied Dynamical Systems, 17(4):2414–2445, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Leiser:2019:NRI

[LR19] Randolph J. Leiser and Horacio G. Rotstein. Network resonance: Impedance interactions via a frequency response alternating map (FRAM). SIAM Journal on Applied Dynamical Systems, 18(2):769– 807, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Letson:2020:LAP

[LR20] Benjamin Letson and Jonathan E. Rubin. LOR for analysis of periodic dynamics: a onestop shop approach. SIAM Journal on Applied Dynamical Systems, 19(1):58–84, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Letson:2022:NTC

[LR22] Benjamin Letson and Jonathan E. Rubin. On the nonexistence of terrestrial canards: Linking canards and rivers. SIAM Journal on Applied Dynamical Systems, 21(4):2432–2462, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1421957.

Liz:2012:CDS

[LRH12] Eduardo Liz and Alfonso Ruiz-Herrera. Chaos in discrete structured population models. SIAM Journal on Applied Dynamical Systems, 11(4):1200–1214, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Lerman:2012:SCM

[LRK12] L. Lerman and V. Rom-Kedar. A saddle in a corner — a model of collinear triatomic chemical reactions. SIAM Journal on Applied Dynamical Systems, 11(1):416– 446, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Lomeli:2008:SSV

[LRR08] Héctor E. Lomelí and Rafael Ramírez-Ros. Separatrix splitting in 3D volumepreserving maps. SIAM Journal on Applied Dynamical Systems, 7(4):1527–1557, ???? 2008. CODEN SJA-DAY. ISSN 1536-0040.

Leimkuhler:2005:HFR

[LS05] Benedict J. Leimkuhler and Christopher R. Sweet. A Hamiltonian formulation for recursive multiple thermostats in a common timescale. SIAM Journal on Applied Dynamical Systems, 4(1):187–216, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60609.

Levine:2015:DDS

[LS15] Howard A. Levine and Yeon-Jung Seo. Discrete dynamical systems in multiple target and alternate SELEX. SIAM Journal on Applied Dynamical Systems, 14(2):1048–1101, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040. See erratum [LS16].

Levine:2016:EDD

[LS16] Howard A. Levine and Yeon-Jung Seo. Erratum: Discrete Dynamical Systems in Multiple Target and Alternate SE-LEX. SIAM Journal on Applied Dynamical Systems, 15 (2):1324–1325, ???? 2016. CODEN SJADAY. ISSN 1536-0040. See [LS15].

Lloyd:2017:CBG

[LS17] David J. B. Lloyd and Arnd Scheel. Continuation and bifurcation of grain boundaries in the Swift–Hohenberg equation. SIAM Journal on Applied Dynamical Systems, 16 (1):252–293, ???? 2017. CO-DEN SJADAY. ISSN 1536-0040.

Lloyd:2008:LHP

[LSAC08] David J. B. Lloyd, Björn Sandstede, Daniele Avitabile, and Alan R. Champneys. Localized hexagon patterns of the planar Swift–Hohenberg equation. SIAM Journal on Applied Dynamical Systems, 7(3):1049–1100, 2008. CO-DEN SJADAY. ISSN 1536-0040.

Lajoie:2011:SIE

[LSB11] Guillaume Lajoie and Eric Shea-Brown. Shared inputs, entrainment, and desynchrony in elliptic bursters: From slow passage to discontinuous circle maps. SIAM Journal on Applied Dynamical Systems, 10(4):1232– 1271, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i4/p1232_s1.

Laing:2003:PMN

[LT03] Carlo R. Laing and William C.

Troy. PDE methods for nonlocal models. SIAM Journal on Applied Dynamical Systems, 2(3):487–516, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60004.

Lee:2013:SAO

[LT13] E. Lee and D. Terman. Stable antiphase oscillations in a network of electrically coupled model neurons. SIAM Journal on Applied Dynamical Systems, 12(1):1–27, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Lee:2015:SAO

[LT15] E. Lee and D. Terman. Stability of antiphase oscillations in a network of inhibitory neurons. SIAM Journal on Applied Dynamical Systems, 14 (1):448–480, ???? 2015. CO-DEN SJADAY. ISSN 1536-0040.

Lee:2017:ELS

[LT17] Min-Gi Lee and Athanasios Tzavaras. Existence of localizing solutions in plasticity via geometric singular perturbation theory. SIAM Journal on Applied Dynamical Systems, 16(1):337–360, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Lee:2019:ORM

[LT19] Euiwoo Lee and David Terman. Oscillatory rhythms in a model network of excitatory and inhibitory neurons. SIAM Journal on Applied Dynamical Systems, 18(1):354– 392, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Liu:2021:SMI

[LT21] Qihuai Liu and Pedro J. Torres. Stability of motion induced by a point vortex under arbitrary polynomial perturbations. SIAM Journal on Applied Dynamical Systems, 20(1):149–164, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Leverentz:2009:ADA

[LTB09] Andrew J. Leverentz, Chad M. Topaz, and Andrew J. Bernoff. Asymptotic dynamics of attractive-repulsive swarms. SIAM Journal on Applied Dynamical Systems, 8(3):880–908, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Lin:2021:DDL

[LTLA21] Yen Ting Lin, Yifeng Tian, Daniel Livescu, and Marian Anghel. Data-driven learning for the Mori–Zwanzig formalism: a generalization of the Koopman learning framework. SIAM Journal on Applied Dynamical Systems, 20 (4):2558–2601, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Lin:2023:RBP

[LTPL23] Yen Ting Lin, Yifeng Tian, Danny Perez, and Daniel Livescu. Regression-based projection for learning Mori– Zwanzig operators. SIAM Journal on Applied Dynamical Systems, 22(4):2890–2926, October 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M1506146.

Lu:2016:PHO

[Lu16] Nan Lu. Persistence of homoclinic orbits under white noise perturbation. SIAM Journal on Applied Dynamical Systems, 15(2):1124–1142, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Luque:2014:QPF

[LV14] Alejandro Luque and Jordi Villanueva. Quasi-periodic frequency analysis using averaging-extrapolation methods. SIAM Journal on Applied Dynamical Systems, 13 (1):1–46, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

LeClainche:2017:HOD

[LV17] Soledad Le Clainche and José M. Vega. Higher order dynamic mode decomposition. SIAM Journal on Applied Dynamical Systems, 16 (2):882–925, ???? 2017. CO-DEN SJADAY. ISSN 1536-0040.

Lin:2002:CSC

[LW02] Wen-Wei Lin and Yi-Qian Wang. Chaotic synchronization in coupled map lattices with periodic boundary conditions. SIAM Journal on Applied Dynamical Systems, 1 (2):175–189, 2002. CODEN SJADAY. ISSN 1536-0040.

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        39541.

Lizarraga:2020:CSP

[LW20] Ian Lizarraga and Martin Wechselberger. Computational singular perturbation method for nonstandard slowfast systems. SIAM Journal on Applied Dynamical Systems, 19(2):994–1028, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Li:2023:ENA

[LWK23] Yangrong Li, Fengling Wang, and Peter Kloeden. Enlarged numerical attractors for lattice systems with porous media degeneracies. SIAM Journal on Applied Dynamical Systems, 22(3):2282–2311, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M1531725.

Ly:2014:DCN

[Ly14] Cheng Ly. Dynamics of coupled noisy neural oscillators with heterogeneous phase resetting curves. SIAM Journal on Applied Dynamical Systems, 13(4):1733–1755, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Lyu:2018:GSP

[Lyu18] Hanbaek Lyu. Global synchronization of pulse-coupled oscillators on trees. SIAM Journal on Applied Dynamical Systems, 17(2):1521–1559, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Liu:2023:SCC

[LZ23] Yangshanshan Liu and Shiqing Zhang. Stacked central configurations with a homogeneous potential in R^{3}. SIAM Journal on Applied Dynamical Systems, 22(2):635–656, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M1495032.

Lopez-Zazueta:2023:QBM

[LZF23] Claudia Lopez-Zazueta and Vincent Fromion. Qualitative behavior of a metabolic pathway with hybrid feedback. SIAM Journal on Applied Dynamical Systems, 22(1):339–381, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1451282.

Liu:2017:GSS

[LZH^{+}17] Shuai Liu, Wei Zou, MiaoZhuang He, Jürgen Kurths, and Meng Zhan. Global stability of the sync with amplitude effects. SIAM Journal on Applied Dynamical Systems, 16(4):1923– 1945, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Li:2022:HNK

[LZX22] Zhuchun Li, Xiaoxue Zhao, and Xiaoping Xue. Hebbian network of Kuramoto oscillators with second-order couplings for binary pattern retrieve: II. Nonorthogonal standard patterns and structural stability. SIAM Journal on Applied Dynamical Systems, 21(1):102–136, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1393224.

Makarenkov:2017:BLC

[Mak17] Oleg Makarenkov. Bifurcation of limit cycles from a fold-fold singularity in planar switched systems. SIAM Journal on Applied Dynamical Systems, 16(3):1340–1371, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Matsue:2011:RVB

[Mat11]  Kaname Matsue.  Rigor-

ous verification of bifurcations of differential equations via the Conley Index Theory. SIAM Journal on Applied Dynamical Systems, 10 (1):325–359, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i1/p325_s1.

Matsue:2018:BSD

[Mat18] Kaname Matsue. On blowup solutions of differential equations with Poincaré-type compactifications. SIAM Journal on Applied Dynamical Systems, 17(3):2249–2288, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Ma:2014:DGP

[MB14a] Tian Ma and Erik M. Bollt.

Differential geometry perspective of shape coherence and curvature evolution by finite-time nonhyperbolic splitting. SIAM Journal on Applied Dynamical Systems, 13(3):1106–1136, ???? 2014. CODEN SJADAY. ISSN 1536-0040. See erratum [MB14b].

Ma:2014:EDG

[MB14b] Tian Ma and Erik M. Bollt.

Erratum: Differential geometry perspective of shape coherence and curvature evolution by finite-time nonhyperbolic splitting. SIAM Journal on Applied Dynamical Systems, 13(4):1916, ???? 2014. CODEN SJADAY. ISSN 1536-0040. See [MB14a].

McCluskey:2015:ULF

[McC15] C. Connell McCluskey. Using Lyapunov functions to construct Lyapunov functionals for delay differential equations. SIAM Journal on Applied Dynamical Systems, 14 (1):1–24, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Ma:2020:DBR

[MCL^{+}20] Chuang Ma, Han-Shuang Chen, Xiang Li, Ying-Cheng Lai, and Hai-Feng Zhang. Data based reconstruction of duplex networks. SIAM Journal on Applied Dynamical Systems, 19(1):124–150, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Melvin:2009:DTS

[MCP09] Thomas R. O. Melvin, Alan R. Champneys, and Dmitry E. Pelinovsky. Discrete traveling solitons in the Salerno model. SIAM Journal on Applied Dynamical Systems, 8(2):689–709, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Macesic:2018:KOF

[MCZM18] Senka Maćesić, Nelida Crnjarić-Zic, and Igor Mezić. Koopman operator family spectrum for nonautonomous systems. SIAM Journal on Applied Dynamical Systems, 17(4):2478–2515, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Millan:2018:SMB

[MD18] Mercedes Pérez Millán and Alicia Dickenstein. The structure of MESSI biological systems. SIAM Journal on Applied Dynamical Systems, 17 (2):1650–1682, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Morrison:2024:DED

[MDIC24] Katherine Morrison, Anda Degeratu, Vladimir Itskov, and Carina Curto. Diversity of emergent dynamics in competitive threshold-linear networks. SIAM Journal on Applied Dynamical Systems, 23(1):855–884, March 2024. ISSN 1536-0040.

Moore:2023:MPD

[MDW23] Joshua W. Moore, Trevor C.

Dale, and Thomas E. Woolley. Modeling polarity-driven laminar patterns in bilayer tissues with mixed signaling mechanisms. SIAM Journal on Applied Dynamical Systems, 22(4):2945–2990, October 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1522565.

Machina:2013:SDG

[MEvdD13] A. Machina, R. Edwards, and P. van den Driessche. Singular dynamics in gene network models. SIAM Journal on Applied Dynamical Systems, 12(1):95–125, ???? 2013. CO-DEN SJADAY. ISSN 1536-0040.

Moehlis:2005:POC

[MFE05] Jeff Moehlis, Holger Faisst, and Bruno Eckhardt. Periodic orbits and chaotic sets in a low-dimensional model for shear flows. SIAM Journal on Applied Dynamical Systems, 4(2):352–376, 2005. CODEN SJADAY. ISSN 1536-0040.

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        60614.

Miller:2023:GMI

[MFN^{+}23] Pearson W. Miller, Daniel Fortunato, Matteo Novaga, Stanislav Y. Shvartsman, and Cyrill B. Muratov. Generation and motion of interfaces in a mass-conserving reaction–diffusion system. SIAM Journal on Applied Dynamical Systems, 22(3): 2408–2431, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M152548X.

Mahdi:2017:CLB

[MFVW17] Adam Mahdi, Antoni Ferragut, Claudia Valls, and Carsten Wiuf. Conservation laws in biochemical reaction networks. SIAM Journal on Applied Dynamical Systems, 16(4):2213–2232, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Montanelli:2016:CPS

[MG16] Hadrien Montanelli and Nikola I. Gushterov. Computing planar and spherical choreographies. SIAM Journal on Applied Dynamical Systems, 15(1):235–256, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Maro:2018:LTD

[MG18] Stefano Marò and Giovanni F. Gronchi. Long term dynamics for the restricted N-body problem with mean motion resonances and crossing singularities. SIAM Journal on Applied Dynamical Systems, 17(2):1786–1815, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Mauroy:2017:SIN

[MH17] A. Mauroy and J. Hendrickx.

Spectral identification of networks using sparse measurements. SIAM Journal on Applied Dynamical Systems, 16 (1):479–513, ???? 2017. CO-DEN SJADAY. ISSN 1536-0040.

Mehta:2007:SSB

[MHB07] Prashant G. Mehta, Gregory Hagen, and Andrzej Banaszuk. Symmetry and symmetry-breaking for a wave equation with feedback. SIAM Journal on Applied Dynamical Systems, 6(3):549– 575, 2007. CODEN SJADAY. ISSN 1536-0040.

Martel:2009:LDS

[MHC09] C. Martel, M. Higuera, and J. D. Carrasco. Localized dispersive states in nonlinear coupled mode equations for light propagation in fiber Bragg gratings. SIAM Journal on Applied Dynamical Systems, 8(2):576–591, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Morita:2019:TCA

[MIK19] Hidetoshi Morita, Masaru Inatsu, and Hiroshi Kokubu. Topological computation analysis of meteorological timeseries data. SIAM Journal on Applied Dynamical Systems, 18(2):1200–1222, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Mirtabatabaei:2014:EOD

[MJB14] Anahita Mirtabatabaei, Peng Jia, and Francesco Bullo. Eulerian opinion dynamics with bounded confidence and exogenous inputs. SIAM Journal on Applied Dynamical Systems, 13(1):425–446, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Mckenzie:2012:ASM

[MJJL12] H. W. Mckenzie, Y. Jin, J. Jacobsen, and M. A. Lewis. R_{0} analysis of a spatiotemporal model for a stream population. SIAM Journal on Applied Dynamical Systems, 11 (2):567–596, ???? 2012. CO-DEN SJADAY. ISSN 1536-0040.

Mullowney:2005:BRC

[MJM05] P. Mullowney, K. Julien, and J. D. Meiss. Blinking rolls: Chaotic advection in a threedimensional flow with an invariant. SIAM Journal on Applied Dynamical Systems, 4(1):159–186, 2005. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60672.

McCarthy:2012:EPA

[MK12] Michelle M. McCarthy and Nancy Kopell. The effect of propofol anesthesia on rebound spiking. SIAM Journal on Applied Dynamical Systems, 11(4):1674–1697, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Morrison:2024:SNO

[MK24] Megan Morrison and J. Nathan Kutz. Solving nonlinear ordinary differential equations using the invariant manifolds and Koopman eigenfunctions. SIAM Journal on Applied Dynamical Systems, 23(1):924– 960, March 2024. ISSN 1536-0040.

Myers:2023:PWO

[MKM23] Audun Myers, Firas A.

Khasawneh, and Elizabeth Munch. Persistence of weighted ordinal partition networks for dynamic state detection. SIAM Journal on Applied Dynamical Systems, 22(1):65–89, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1476848.

Mujica:2018:TBG

[MKO18] José Mujica, Bernd Krauskopf, and Hinke M. Osinga. Tangencies between global invariant manifolds and slow manifolds near a singular Hopf bifurcation. SIAM Journal on Applied Dynamical Systems, 17(2):1395–1431, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

McGehee:2012:PMI

[ML12] Richard McGehee and Clarence Lehman. A paleoclimate model of ice-albedo feedback forced by variations in Earth’s orbit. SIAM Journal on Applied Dynamical Systems, 11(2):684–707, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Modhara:2021:NFR

[MLTC21] Sunil Modhara, Yi Ming Lai, Rüdiger Thul, and Stephen Coombes. Neural fields with rebound currents: Novel routes to patterning. SIAM Journal on Applied Dynamical Systems, 20(3):1596–1620, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

McLachlan:2006:KHI

[MM06] Robert I. McLachlan and Stephen R. Marsland. The Kelvin–Helmholtz instability of momentum sheets in the Euler equations for planar diffeomorphisms. SIAM Journal on Applied Dynamical Systems, 5(4):726–758, 2006. CODEN SJADAY. ISSN 1536-0040.

Mosovsky:2011:TTD

[MM11] B. A. Mosovsky and J. D.

Meiss. Transport in transitory dynamical systems. SIAM Journal on Applied Dynamical Systems, 10(1): 35–65, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i1/p35_s1.

Mosovsky:2012:TTT

[MM12] Brock A. Mosovsky and James D. Meiss. Transport in transitory, three-dimensional, Liouville flows. SIAM Journal on Applied Dynamical Systems, 11(4):1785–1816, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Mitchell:2017:DFT

[MM17] R. A. Mitchell and J. D.

Meiss. Designing a finite-time mixer: Optimizing stirring for two-dimensional maps. SIAM Journal on Applied Dynamical Systems, 16(3):1514–1542, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Meyer:2022:IMA

[MM22] Katherine J. Meyer and Richard P. McGehee. Intensity — a metric approach to quantifying attractor robustness in ODEs. SIAM Journal on Applied Dynamical Systems, 21(2):960–981, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M138689X.

Matano:2022:ABF

[MMNS22] Hiroshi Matano, Yoichiro Mori, Mitsunori Nara, and Koya Sakakibara. Asymptotic behavior of fronts and pulses of the bidomain model. SIAM Journal on Applied Dynamical Systems, 21(1):616–649, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1416904.

Maia:2016:PNS

[MMP16] Daniel M. N. Maia, Elbert E. N. Macau, and Tiago Pereira. Persistence of network synchronization under nonidentical coupling functions. SIAM Journal on Applied Dynamical Systems, 15 (3):1563–1580, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Morarescu:2007:SCC

[MNG07] Constantin-Irinel Morărescu, Silviu-Iulian Niculescu, and Keqin Gu. Stability crossing curves of shifted gammadistributed delay systems. SIAM Journal on Applied Dynamical Systems, 6(2):475– 493, 2007. CODEN SJADAY. ISSN 1536-0040.

Mohapatra:2015:HLH

[MO15] Anushaya Mohapatra and William Ott. Homoclinic loops, heteroclinic cycles, and rank one dynamics. SIAM Journal on Applied Dynamical Systems, 14(1):107–131, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Misquero:2020:SRR

[MO20] Mauricio Misquero and Rafael Ortega. Some rigorous results on the 1: 1 resonance of the spin–orbit problem. SIAM Journal on Applied Dynamical Systems, 19(4):2233–2267, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Moeckel:2015:FPT

[Moe15] Richard Moeckel. The Foucault pendulum (with a twist). SIAM Journal on Applied Dynamical Systems, 14 (3):1644–1662, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Mohlenkamp:2019:DSC

[Moh19] Martin J. Mohlenkamp. The dynamics of swamps in the canonical tensor approximation problem. SIAM Journal on Applied Dynamical Systems, 18(3):1293–1333, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Moreira:2015:SJS

[Mor15] Célia Sofia Moreira. Special Jordan subspaces and synchrony subspaces in coupled cell networks. SIAM Journal on Applied Dynamical Systems, 14(1):253–285, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Mason:2009:ECT

[MP09] Joanna F. Mason and Petri T. Piiroinen. The effect of codimension-two bifurcations on the global dynamics of a gear model. SIAM Journal on Applied Dynamical Systems, 8(4):1694–1711, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Mengers:2013:ODS

[MP13] J. D. Mengers and J. M. Powers. One-dimensional slow invariant manifolds for fully coupled reaction and microscale diffusion. SIAM Journal on Applied Dynamical Systems, 12(2):560–595, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Massaroli:2022:OES

[MPC^{+}22] Stefano Massaroli, Michael Poli, Federico Califano, Jinkyoo Park, Atsushi Yamashita, and Hajime Asama. Optimal energy shaping via neural approximators. SIAM Journal on Applied Dynamical Systems, 21(3):2126–2147, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1414279.

Ma:2004:DBG

[MPW04] Tian Ma, Jungho Park, and Shouhong Wang. Dynamic bifurcation of the Ginzburg– Landau equation. SIAM Journal on Applied Dynamical Systems, 3(4):620–635, 2004. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60374.

Meyer:2011:GAH

[MPY11] Kenneth R. Meyer, Jesús F.

Palacián, and Patricia Yanguas. Geometric averaging of Hamiltonian systems: Periodic solutions, stability, and KAM tori. SIAM Journal on Applied Dynamical Systems, 10(3):817– 856, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i3/p817_s1.

MacKenzie:2006:AMK

[MR06] T. MacKenzie and A. J.

Roberts. Accurately model the Kuramoto–Sivashinsky dynamics with holistic discretization. SIAM Journal on Applied Dynamical Systems, 5(3):365–402, 2006. CODEN SJADAY. ISSN 1536-0040.

Menezes:2018:ESF

[MR18] Brian Menezes and Gareth E.

Roberts. Existence and stability of four-vortex collinear relative equilibria with three equal vorticities. SIAM Journal on Applied Dynamical Systems, 17(1):1023–1051, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Muller:2024:SCL

[MR24] Stefan Müller and Georg Regensburger. Sufficient conditions for linear stability of complex-balanced equilibria in generalized mass-action systems. SIAM Journal on Applied Dynamical Systems, 23(1):325–357, January 2024. ISSN 1536-0040.

McCullen:2013:MMI

[MRB^{+}13] N. J. McCullen, A. M. Rucklidge, C. S. E. Bale, T. J. Foxon, and W. F. Gale. Multiparameter models of innovation diffusion on complex networks. SIAM Journal on Applied Dynamical Systems, 12 (1):515–532, ???? 2013. CO-DEN SJADAY. ISSN 1536-0040.

Mauroy:2014:GIP

[MRMM14] Alexandre Mauroy, Blane Rhoads, Jeff Moehlis, and Igor Mezić. Global isochrons and phase sensitivity of bursting neurons. SIAM Journal on Applied Dynamical Systems, 13(1):306–338, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Melli:2006:MPA

[MRR06] Juan B. Melli, Clarence W. Rowley, and Dzhelil S. Rufat.

Motion planning for an articulated body in a perfect planar fluid. SIAM Journal on Applied Dynamical Systems, 5(4):650–669, 2006. CODEN SJADAY. ISSN 1536-0040.

Meyries:2014:QLP

[MRS14] M. Meyries, J. D. M.

Rademacher, and E. Siero. Quasi-linear parabolic reaction– diffusion systems: a user’s guide to well-posedness, spectra, and stability of travelling waves. SIAM Journal on Applied Dynamical Systems, 13 (1):249–275, ???? 2014. CO-DEN SJADAY. ISSN 1536-0040.

Marzuola:2017:NBS

[MRS17] J. L. Marzuola, S. G.

Raynor, and G. Simpson. Nonlinear bound states in a Schrödinger–Poisson system with external potential. SIAM Journal on Applied Dynamical Systems, 16(1):226– 251, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

McCalla:2013:SSH

[MS13] S. G. McCalla and B. Sandstede. Spots in the Swift– Hohenberg equation. SIAM Journal on Applied Dynamical Systems, 12(2):831–877, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Morrissey:2015:CEB

[MS15] David Morrissey and Arnd Scheel. Characterizing the effect of boundary conditions on striped phases. SIAM Journal on Applied Dynamical Systems, 14(3):1387–1417, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Maruskin:2009:DSS

[MSB09] Jared M. Maruskin, Daniel J. Scheeres, and Anthony M. Bloch. Dynamics of symplectic subvolumes. SIAM Journal on Applied Dynamical Systems, 8(1):180–201, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Marschler:2014:IME

[MSB^{+}14] Christian Marschler, Jan Sieber, Rainer Berkemer, Atsushi Kawamoto, and Jens Starke. Implicit methods for equation-free analysis: Convergence results and analysis of emergent waves in microscopic traffic models. SIAM Journal on Applied Dynamical Systems, 13(3):1202–1238, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Maelfeyt:2017:UIM

[MSM17] Bryan Maelfeyt, Spencer A. Smith, and Kevin A. Mitchell. Using invariant manifolds to construct symbolic dynamics for three-dimensional volumepreserving maps. SIAM Journal on Applied Dynamical Systems, 16(1):729–769, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Montaldi:2003:VDC

[MST03] James Montaldi, Anik Soulière, and Tadashi Tokieda. Vortex dynamics on a cylinder. SIAM Journal on Applied Dynamical Systems, 2(3):417– 430, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41556.

Mrozek:2015:TAA

[MSW15] Marian Mrozek, Roman Srzednicki, and Frank Weilandt. A topological approach to the algorithmic computation of the Conley index for Poincaré maps. SIAM Journal on Applied Dynamical Systems, 14(3):1348–1386, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Munnier:2011:PSP

[Mun11] Alexandre Munnier. Passive and self-propelled locomotion of an elastic swimmer in a perfect fluid. SIAM Journal on Applied Dynamical Systems, 10(4):1363– 1403, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i4/p1363_s1.

Marschler:2014:UCM

[MV14] Christian Marschler and Jürgen Vollmer. Unidirectionally coupled map lattices with nonlinear coupling: Unbinding transitions and superlong transients. SIAM Journal on Applied Dynamical Systems, 13(3):1137–1151, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

MacLaurin:2021:SSB

[MV21] James N. MacLaurin and Pedro A. Vilanova. Synchronization in stochastic biochemical oscillators subject to common multiplicative extrinsic noise. SIAM Journal on Applied Dynamical Systems, 20(3):1253–1276, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

McCalla:2018:CDS

[MvB18] Scott G. McCalla and James H. von Brecht. Consistent dynamics of stripes formed by cell-type interfaces. SIAM Journal on Applied Dynamical Systems, 17(4):2615–2633, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Mileyko:2010:BAG

[MW10] Yuriy Mileyko and Joshua S. Weitz. Bifurcation analysis of gene regulatory circuits subject to copy number variation. SIAM Journal on Applied Dynamical Systems, 9 (3):799–826, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

McGehee:2014:QAB

[MW14] Richard McGehee and Esther Widiasih. A quadratic approximation to Budyko’s ice– albedo feedback model with ice line dynamics. SIAM Journal on Applied Dynamical Systems, 13(1):518–536, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

McQuighan:2016:EMV

[MW16] Kelly McQuighan and C. Eugene Wayne. An explanation of metastability in the viscous Burgers equation with periodic boundary conditions via a spectral analysis. SIAM Journal on Applied Dynamical Systems, 15(4):1916–1961, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Medvedev:2017:STS

[MW17a] Georgi S. Medvedev and J. Douglas Wright. Stability of twisted states in the continuum Kuramoto model. SIAM Journal on Applied Dynamical Systems, 16(1):188–203, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Mitry:2017:FSF

[MW17b] John Mitry and Martin Wechselberger. Folded saddles and faux canards. SIAM Journal on Applied Dynamical Systems, 16(1):546–596, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Moyles:2017:ESD

[MW17c] Iain R. Moyles and Michael J.

Ward. Existence, stability, and dynamics of ring and near-ring solutions to the saturated Gierer–Meinhardt model in the semistrong regime. SIAM Journal on Applied Dynamical Systems, 16 (1):597–639, ???? 2017. CO-DEN SJADAY. ISSN 1536-0040.

McMillen:2016:BOF

[MXYZ16] Tyler McMillen, Jack Xin, Yifeng Yu, and Andrej Zlatos. Ballistic orbits and front speed enhancement for ABC flows. SIAM Journal on Applied Dynamical Systems, 15 (3):1753–1782, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

McLachlan:2011:ABP

[MZ11] Robert I. McLachlan and Xingyou Zhang. Asymptotic blowup profiles for modified Camassa–Holm equations. SIAM Journal on Applied Dynamical Systems, 10 (2):452–468, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i2/p452_s1.

Nguyen:2023:EKO

[NB^{+}23] Tung T. Nguyen, Roberto C. Budzinski, Jacqueline oàn, Federico W. Pasini, Ján Minác, and Lyle E. Muller. Equilibria in Kuramoto oscillator networks: an algebraic approach. SIAM Journal on Applied Dynamical Systems, 22(2):802–824, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1457321.

Nicola:2016:NBM

[NC16] Wilten Nicola and Sue Ann Campbell. Nonsmooth bifurcations of mean field systems of two-dimensional integrate and fire neurons. SIAM Journal on Applied Dynamical Systems, 15(1):391–439, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Nicola:2021:NCS

[NC21] Wilten Nicola and Sue Ann Campbell. Normalized connectomes show increased synchronizability with age through their second largest eigenvalue. SIAM Journal on Applied Dynamical Systems, 20(2):1158–1176, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Nicks:2021:USI

[NCA^{+}21] Rachel Nicks, Abigail Cocks, Daniele Avitabile, Alan Johnston, and Stephen Coombes. Understanding sensory induced hallucinations: From neural fields to amplitude equations. SIAM Journal on Applied Dynamical Systems, 20(4):1683–1714, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Newby:2014:SEM

[New14] Jay M. Newby. Spontaneous excitability in the Morris– Lecar model with ion channel noise. SIAM Journal on Applied Dynamical Systems, 13(4):1756–1791, ???? 2014.

CODEN SJADAY. ISSN 1536-0040.

Ni:2023:FAA

[Ni23] Angxiu Ni. Fast adjoint algorithm for linear responses of hyperbolic chaos. SIAM Journal on Applied Dynamical Systems, 22(4):2792–2824, October 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M1522383.

Noakes:2008:ANL

[Noa08] Lyle Noakes. Asymptotics of null Lie quadratics in E^{3}. SIAM Journal on Applied Dynamical Systems, 7(2):437– 460, 2008. CODEN SJADAY. ISSN 1536-0040.

Natali:2015:FOD

[NP15] Fábio Natali and Ademir Pastor. The fourth-order dispersive nonlinear Schrödinger equation: Orbital stability of a standing wave. SIAM Journal on Applied Dynamical Systems, 14(3):1326–1347, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Novitzky:2019:NDQ

[NPRW19] Sophia Novitzky, Jamol Pender, Richard H. Rand, and Elizabeth Wesson. Nonlinear dynamics in queueing theory: Determining the size of oscillations in queues with delay. SIAM Journal on Applied Dynamical Systems, 18(1):279– 311, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Nieto:2012:LTS

[NPV12] J. J. Nieto, M. J. Pacifico, and J. L. Vieitez. Longterm and short-term dynamics of Microtus epiroticus: a Yoccoz–Birkeland model. SIAM Journal on Applied Dynamical Systems, 11(4):1499– 1532, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Nijholt:2020:QRD

[NRS20] Eddie Nijholt, Bob W. Rink, and Sören Schwenker. Quiver representations and dimension reduction in dynamical systems. SIAM Journal on Applied Dynamical Systems, 19(4):2428–2468, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Neishtadt:2013:ADP

[NS13] Anatoly Neishtadt and Tan Su. On asymptotic description of passage through a resonance in quasilinear Hamiltonian systems. SIAM Journal on Applied Dynamical Systems, 12(3):1436–1473, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Net:2015:CBP

[NS15] M. Net and J. Sánchez. Continuation of bifurcations of periodic orbits for large-scale systems. SIAM Journal on Applied Dynamical Systems, 14(2):674–698, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Neuberger:2006:SAB

[NSS06] John M. Neuberger, Nándor Sieben, and James W. Swift. Symmetry and automated branch following for a semilinear elliptic PDE on a fractal region. SIAM Journal on Applied Dynamical Systems, 5(3):476–507, 2006. CODEN SJADAY. ISSN 1536-0040.

Neuberger:2013:NMS

[NSS13] John M. Neuberger, Nándor Sieben, and James W. Swift. Newton’s method and symmetry for semilinear elliptic PDE on the cube. SIAM Journal on Applied Dynamical Systems, 12(3):1237–1279, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Neuberger:2019:SAD

[NSS19] John M. Neuberger, Nándor Sieben, and James W. Swift. Synchrony and antisynchrony for difference-coupled vector fields on graph network systems. SIAM Journal on Applied Dynamical Systems, 18 (2):904–938, ???? 2019. CO-DEN SJADAY. ISSN 1536-0040.

Neuberger:2020:ISS

[NSS20] John M. Neuberger, Nándor Sieben, and James W. Swift. Invariant synchrony subspaces of sets of matrices. SIAM Journal on Applied Dynamical Systems, 19(2):964– 993, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Neuberger:2024:BLI

[NSS24] John M. Neuberger, Nándor Sieben, and James W. Swift. A bifurcation lemma for invariant subspaces. SIAM Journal on Applied Dynamical Systems, 23(2):1610–1635, ???? 2024. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/23M1595540.

Novaes:2020:SPO

[NSTZ20] Douglas D. Novaes, Tere M.

Seara, Marco A. Teixeira, and Iris O. Zeli. Study of periodic orbits in periodic perturbations of planar reversible Filippov systems having a twofold cycle. SIAM Journal on Applied Dynamical Systems, 19(2):1343–1371, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Nagem:2009:GHK

[NSUW09] Raymond Nagem, Guido Sandri, David Uminsky, and C. Eugene Wayne. Generalized Helmholtz–Kirchhoff model for two-dimensional distributed vortex motion. SIAM Journal on Applied Dynamical Systems, 8(1):160– 179, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Nishiura:2005:CPD

[NUY05] Yasumasa Nishiura, Daishin Ueyama, and Tatsuo Yanagita. Chaotic pulses for discrete reaction diffusion systems. SIAM Journal on Applied Dynamical Systems, 4(3):733– 754, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60871.

Nordmark:2018:DBD

[NVC18] Arne Nordmark, Peter L. Várkonyi, and Alan R. Champneys. Dynamics beyond dynamic jam; unfolding the Painlevé paradox singularity. SIAM Journal on Applied Dynamical Systems, 17(2):1267–1309, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Nan:2015:UDT

[NWKR15] Pingyu Nan, Yangyang Wang, Vivien Kirk, and Jonathan E. Rubin. Understanding and distinguishing three-time-scale oscillations: Case study in a coupled Morris–Lecar system. SIAM Journal on Applied Dynamical Systems, 14(3):1518–1557, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Nadeau:2021:SGC

[NWW21] Alice Nadeau, James Walsh, and Esther Widaish. Synchronous glacial cycles in a nonsmooth conceptual climate model with asymmetric hemispheres. SIAM Journal on Applied Dynamical Systems, 20(4):2482–2515, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Nishiura:2022:DSR

[NX22] Yasumasa Nishiura and Shuangquan Xie. Dynamics of N-spot rings with oscillatory tails in a three-component reaction– diffusion system. SIAM Journal on Applied Dynamical Systems, 21(3):2268–2296, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M1492143.

Oljaca:2018:ASE

[OBK18] Lea Oljaca, Jochen Bröcker, and Tobias Kuna. Almost sure error bounds for data assimilation in dissipative systems with unbounded observation noise. SIAM Journal on Applied Dynamical Systems, 17(4):2882–2914, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Osorio:2008:CIB

[OdBS08] Gustavo Osorio, Mario di Bernardo, and Stefania Santini. Cornerimpact bifurcations: a novel class of discontinuity-induced bifurcations in cam-follower systems. SIAM Journal on Applied Dynamical Systems, 7(1):18–38, 2008. CODEN SJADAY. ISSN 1536-0040.

Orr:2021:NDO

[OE21] Derek Orr and G. Bard Ermentrout. Noise-driven oscillations in coupled excitable systems. SIAM Journal on Applied Dynamical Systems, 20(2):826–852, ???? 2021.

CODEN SJADAY. ISSN 1536-0040.

Owens:2023:DDD

[OK23] Katherine Owens and J. Nathan Kutz. Data-driven discovery of governing equations for coarse-grained heterogeneous network dynamics. SIAM Journal on Applied Dynamical Systems, 22(3):2601–2623, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M1497882.

Oliver:2024:DSS

[OK24] Marcel Oliver and Marc A. Tiofack Kenfack. Deterministic and stochastic surrogate models for a slowly driven fast oscillator. SIAM Journal on Applied Dynamical Systems, 23(2):1090–1107, April 2024. ISSN 1536-0040.

Osinga:2010:CBC

[OM10] Hinke M. Osinga and Jeff Moehlis. Continuationbased computation of global isochrons. SIAM Journal on Applied Dynamical Systems, 9(4):1201–1228, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Olvera:2008:RPC

[OP08] Arturo Olvera and Nikola P. Petrov. Regularity properties of critical invariant circles of twist maps, and their universality. SIAM Journal on Applied Dynamical Systems, 7 (3):962–987, 2008. CODEN SJADAY. ISSN 1536-0040.

Otkin:2021:NCM

[OPL21] Jason A. Otkin, Roland W. E.

Potthast, and Amos S. Lawless. Nonlinear conditional model bias estimation for data assimilation. SIAM Journal on Applied Dynamical Systems, 20(1):299–332, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Otto:2024:LBM

[OPR24] Samuel Otto, Sebastian Peitz, and Clarence Rowley. Learning bilinear models of actuated Koopman generators from partially observed trajectories. SIAM Journal on Applied Dynamical Systems, 23(1):885–923, March 2024. ISSN 1536-0040.

Otto:2019:LRA

[OR19] Samuel E. Otto and Clarence W.

Rowley. Linearly recurrent autoencoder networks for learning dynamics. SIAM Journal on Applied Dynamical Systems, 18(1):558–593, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Orosz:2014:DND

[Oro14] Gábor Orosz. Decomposition of nonlinear delayed networks around cluster states with applications to neurodynamics. SIAM Journal on Applied Dynamical Systems, 13(4):1353– 1386, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

OKeeffe:2020:TPP

[OW20] Paul E. O’Keeffe and Sebastian Wieczorek. Tipping phenomena and points of no return in ecosystems: Beyond classical bifurcations. SIAM Journal on Applied Dynamical Systems, 19(4):2371–2402, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Omelchenko:2018:SSC

[OWK18] Oleh E. Omel’chenko, Matthias Wolfrum, and Edgar Knobloch. Stability of spiral chimera states on a torus. SIAM Journal on Applied Dynamical Systems, 17(1):97–127, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Ott:2003:LAR

[OY03] William Ott and James A. Yorke. Learning about reality from observation. SIAM Journal on Applied Dynamical Systems, 2(3):297–322, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40742.

Ou:2009:TWV

[OY09] Chunhua Ou and Wei Yuan. Traveling wavefronts in a volume-filling chemotaxis model. SIAM Journal on Applied Dynamical Systems, 8 (1):390–416, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Or:2011:DSL

[OZM11] Yizhar Or, Sebastian Zhang, and Richard M. Murray. Dynamics and stability of low-Reynolds-number swimming near a wall. SIAM Journal on Applied Dynamical Systems, 10(3):1013– 1041, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i3/p1013_s1.

Perkins:2019:SFA

[PA19] Melinda Liu Perkins and Murat Arcak. A spatial filtering approach to biological patterning. SIAM Journal on Applied Dynamical Systems, 18(3):1694–1721, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Parmelee:2022:SAC

[PACM22] Caitlyn Parmelee, Juliana Londono Alvarez, Carina Curto, and Katherine Morrison. Sequential attractors in combinatorial threshold-linear networks. SIAM Journal on Applied Dynamical Systems, 21 (2):1597–1630, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1445120.

Patrick:2003:SKC

[Pat03] George W. Patrick. Stability by KAM confinement of certain wild, nongeneric relative equilibria of underwater vehicles with coincident centers of mass and buoyancy. SIAM Journal on Applied Dynamical Systems, 2(1):36–52, 2003. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40475.

Pring:2010:DRD

[PB10] S. R. Pring and C. J. Budd. The dynamics of regularized discontinuous maps with applications to impacting systems. SIAM Journal on Applied Dynamical Systems, 9 (1):188–219, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Porfiri:2017:MMS

[PB17] Maurizio Porfiri and Igor Belykh. Memory matters in synchronization of stochastically coupled maps. SIAM Journal on Applied Dynamical Systems, 16(3):1372–1396, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Parker:2022:BNN

[PB22] Ross Parker and Andrea K. Barreiro. Bifurcations of a neural network model with symmetry. SIAM Journal on Applied Dynamical Systems, 21(4):2535–2578, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1470451.

Priyankara:2022:MTT

[PBB22] K. G. D. Sulalitha Priyankara, Sanjeeva Balasuriya, and Erik M. Bollt. Melnikov theory for two-dimensional manifolds in three-dimensional flows. SIAM Journal on Applied Dynamical Systems, 21 (4):2642–2696, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1464300.

Potthast:2009:IPN

[PbG09] Roland Potthast and Peter beim Graben. Inverse problems in neural field theory. SIAM Journal on Applied Dynamical Systems, 8(4):1405– 1433, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Proctor:2016:DMD

[PBK16] Joshua L. Proctor, Steven L.

Brunton, and J. Nathan Kutz. Dynamic mode decomposition with control. SIAM Journal on Applied Dynamical Systems, 15(1):142–161, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Proctor:2018:GKT

[PBK18] Joshua L. Proctor, Steven L.

Brunton, and J. Nathan Kutz. Generalizing Koopman theory to allow for inputs and control. SIAM Journal on Applied Dynamical Systems, 17(1):909–930, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Poignard:2016:PON

[PCG16] Camille Poignard, Madalena Chaves, and Jean-Luc Gouzé. Periodic oscillations for nonmonotonic smooth negative feedback circuits. SIAM Journal on Applied Dynamical Systems, 15(1):257–286, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Poignard:2018:SRP

[PCG18] Camille Poignard, Madalena Chaves, and Jean-Luc Gouzé. A stability result for periodic solutions of nonmonotonic smooth negative feedback systems. SIAM Journal on Applied Dynamical Systems, 17(2):1091–1116, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Pais:2012:HBL

[PCNL12] Darren Pais, Carlos H. Caicedo-Núnñez, and Naomi E. Leonard. Hopf bifurcations and limit cycles in evolutionary network dynamics. SIAM Journal on Applied Dynamical Systems, 11(4):1754–1784, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Pan:2018:DDD

[PD18] Shaowu Pan and Karthik Duraisamy. Data-driven discovery of closure models. SIAM Journal on Applied Dynamical Systems, 17(4):2381–2413, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Pan:2020:PIP

[PD20] Shaowu Pan and Karthik Duraisamy. Physics-informed probabilistic learning of linear embeddings of nonlinear dynamics with guaranteed stability. SIAM Journal on Applied Dynamical Systems, 19 (1):480–509, ???? 2020. CO-DEN SJADAY. ISSN 1536-0040.

Pasour:2010:CAP

[PE10] V. B. Pasour and S. P. Ellner. Computational and analytic perspectives on the drift paradox. SIAM Journal on Applied Dynamical Systems, 9(2):333–356, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Park:2018:SRN

[PE18] Youngmin Park and G. Bard Ermentrout. Scalar reduction of a neural field model with spike frequency adaptation. SIAM Journal on Applied Dynamical Systems, 17 (1):931–981, ???? 2018. CO-DEN SJADAY. ISSN 1536-0040.

Pascual-Escudero:2022:AAP

[PEH22] Beatriz Pascual-Escudero and Linard Hoessly. An algebraic approach to productform stationary distributions for some reaction networks. SIAM Journal on Applied Dynamical Systems, 21(1):588–615, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1401498.

Penner:2012:PFM

[PES12] Kevin Penner, Bard Ermentrout, and David Swigon. Pattern formation in a model of acute inflammation. SIAM Journal on Applied Dynamical Systems, 11(2):629–660, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Paccosi:2014:BAS

[PFGV14] Rubén Gustavo Paccosi, Alejandra Figliola, and Jorge Galán-Vioque. A bifurcation approach to the synchronization of coupled van der Pol oscillators. SIAM Journal on Applied Dynamical Systems, 13(3):1152–1167, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Peter:2021:LNS

[PID21] Stephan Peter, Bashar Ibrahim, and Peter Dittrich. Linking network structure and dynamics to describe the set of persistent species in reaction diffusion systems. SIAM Journal on Applied Dynamical Systems, 20(4):2037–2076, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Pinto:2005:EST

[PJW05] David J. Pinto, Russell K. Jackson, and C. Eugene Wayne. Existence and stability of traveling pulses in a continuous neuronal network.

SIAM Journal on Applied Dynamical Systems, 4(4):954– 984, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/61302.

Peercy:2005:CCM

[PK05a] Bradford E. Peercy and James P. Keener. Coupled cell model of border zone arrhythmias. SIAM Journal on Applied Dynamical Systems, 4(3):679–710, 2005. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
61597.

Porter:2005:BEC

[PK05b] Mason A. Porter and P. G.

Kevrekidis. Bose–Einstein condensates in superlattices. SIAM Journal on Applied Dynamical Systems, 4(4):783– 807, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/61061.

Poll:2017:VIM

[PK17] Daniel B. Poll and Zachary P.

Kilpatrick. Velocity integration in a multilayer neural field model of spatial working memory. SIAM Journal on Applied Dynamical Systems, 16(3):1197–1234, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Paquin-Lefebvre:2019:PFO

[PLNW19] Frédéric Paquin-Lefebvre, Wayne Nagata, and Michael J.

Ward. Pattern formation and oscillatory dynamics in a twodimensional coupled bulksurface reaction–diffusion system. SIAM Journal on Applied Dynamical Systems, 18 (3):1334–1390, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Patterson:2020:PFS

[PLST20] Denis D. Patterson, Simon A. Levin, Carla Staver, and Jonathan D. Touboul. Probabilistic foundations of spatial mean-field models in ecology and applications. SIAM Journal on Applied Dynamical Systems, 19(4):2682–2719, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Pujo-Menjouet:2005:LPO

[PMBM05] Laurent Pujo-Menjouet, Samuel Bernard, and Michael C. Mackey. Long period oscillations in a G_{0} model of hematopoietic stem cells. SIAM Journal on Applied Dynamical Systems, 4(2):312– 332, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60047.

Prasadan:2020:TSS

[PN20] Arvind Prasadan and Raj Rao Nadakuditi. Time series source separation using dynamic mode decomposition. SIAM Journal on Applied Dynamical Systems, 19(2):1160– 1199, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Peitz:2020:DDM

[POR20] Sebastian Peitz, Samuel E.

Otto, and Clarence W. Rowley. Data-driven model predictive control using interpolated Koopman generators. SIAM Journal on Applied Dynamical Systems, 19(3):2162– 2193, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Postlethwaite:2009:SLP

[Pos09] Claire M. Postlethwaite. Stabilization of long-period periodic orbits using time-delayed feedback control. SIAM Journal on Applied Dynamical Systems, 8(1):21–39, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Porfiri:2008:MSG

[PP08] Maurizio Porfiri and Roberta Pigliacampo. Master-slave global stochastic synchronization of chaotic oscillators. SIAM Journal on Applied Dynamical Systems, 7(3):825– 842, 2008. CODEN SJADAY. ISSN 1536-0040.

Penati:2012:BBT

[PP12] T. Penati and S. Paleari.

Breathers and Q-breathers: Two sides of the same coin. SIAM Journal on Applied Dynamical Systems, 11(1): 1–30, ???? 2012. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v11/i1/p1_s1.

Parker:2020:KAI

[PP20] Jeremy P. Parker and Jacob Page. Koopman analysis of isolated fronts and solitons. SIAM Journal on Applied Dynamical Systems, 19(4):2803– 2828, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Purewal:2014:GBA

[PPK14] A. S. Purewal, C. M. Postlethwaite, and B. Krauskopf. A global bifurcation analysis of the subcritical Hopf normal form subject to pyragas time-delayed feedback control. SIAM Journal on Applied Dynamical Systems, 13 (4):1879–1915, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Piltz:2014:PSL

[PPM14] Sofia H. Piltz, Mason A. Porter, and Philip K. Maini. Prey switching with a linear preference trade-off. SIAM Journal on Applied Dynamical Systems, 13(2):658–682, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Park:2022:APT

[PR22] Choongseok Park and Jonathan E. Rubin. Activity patterns of a two-timescale neuronal ring model with voltagedependent, piecewise smooth inhibitory coupling. SIAM Journal on Applied Dynamical Systems, 21(3):1952–1999, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/21M1431679.

Pavlyukevich:2024:SEB

[PR24] Ilya Pavlyukevich and Marian Ritsch. Stochastic energybalance model with a moving ice line. SIAM Journal on Applied Dynamical Systems, 23(3):2061–2098, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1619873.

Perdomo:2023:POB

[PRAC23] Oscar Perdomo, Andrés Rivera, John A. Arredondo, and Nelson Castañeda. Periodic oscillations in a 2Nbody problem. SIAM Journal on Applied Dynamical Systems, 22(2):744–764, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1484407.

Parra-Rivas:2023:OSL

[PRCA^{+}23] Pedro Parra-Rivas, Alan R.

Champneys, Fahad Al Saadi, Damia Gomila, and Edgar Knobloch. Organization of spatially localized structures near a codimensionthree cusp-Turing bifurcation. SIAM Journal on Applied Dynamical Systems, 22 (4):2693–2731, October 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1514234.

Pnueli:2018:NIS

[PRK18] Michal Pnueli and Vered Rom-Kedar. On near integrability of some impact systems. SIAM Journal on Applied Dynamical Systems, 17(4):2707– 2732, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Pnueli:2022:NTD

[PRK22] Michal Pnueli and Vered Rom-Kedar. Near tangent dynamics in a class of Hamiltonian impact systems. SIAM Journal on Applied Dynamical Systems, 21(3):2000–2046, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1454596.

Propst:2020:PMF

[Pro20] Georg Propst. Pumping in models of flow in a loop of rigid pipes. SIAM Journal on Applied Dynamical Systems, 19(4):2737–2748, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Prugger:2023:RSW

[PRY23] Artur Prugger, Jens D. M. Rademacher, and Jichen Yang. Rotating shallow water equations with bottom drag: Bifurcations and growth due to kinetic energy backscatter. SIAM Journal on Applied Dynamical Systems, 22 (3):2490–2526, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M152222X.

Panagiotopoulos:2023:CNC

[PSSJ23] Ilias Panagiotopoulos, Jens Starke, Jan Sieber, and Wolfram Just. Continuation with noninvasive control schemes: Revealing unstable states in a pedestrian evacuation scenario. SIAM Journal on Applied Dynamical Systems, 22(1):1–36, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1482032.

Pavliotis:2022:DFB

[PSV22] G. A. Pavliotis, A. M. Stuart, and U. Vaes. Derivative-free Bayesian inversion using multiscale dynamics. SIAM Journal on Applied Dynamical Systems, 21(1):284–326, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1397416.

Pelinovsky:2012:PSW

[PSW12] Dmitry E. Pelinovsky, Gideon Simpson, and Michael I. Weinstein. Polychromatic solitary waves in a periodic and nonlinear Maxwell system. SIAM Journal on Applied Dynamical Systems, 11 (1):478–506, ???? 2012. CO-DEN SJADAY. ISSN 1536-

Pinto:2009:AAW

[PTK09] David J. Pinto, William Troy, and Timothy Kneezel. Asymmetric activity waves in synaptic cortical systems. SIAM Journal on Applied Dynamical Systems, 8(3):1218– 1233, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Piltz:2017:PPF

[PVMP17] Sofia H. Piltz, Frits Veerman, Philip K. Maini, and Mason A. Porter. A predator–2 prey fast–slow dynamical system for rapid predator evolution. SIAM Journal on Applied Dynamical Systems, 16 (1):54–90, ???? 2017. CO-DEN SJADAY. ISSN 1536-0040.

Poussot-Vassal:2021:IMG

[PVVCS21] Charles Poussot-Vassal, Pierre Vuillemin, Olivier Cantinaud, and Florian Sève. Interpolatory methods for generic Biz-Jet gust load alleviation function. SIAM Journal on Applied Dynamical Systems, 20 (4):2391–2411, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Palacian:2017:PSK

[PVVY17] J. F. Palacián, C. Vidal, J. Vidarte, and P. Yanguas. Periodic solutions and KAM tori in a triaxial potential. SIAM Journal on Applied Dynamical Systems, 16(1):159– 187, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Peletier:2007:SCB

[PW07] L. A. Peletier and J. F.

Williams. Some canonical bifurcations in the Swift– Hohenberg equation. SIAM Journal on Applied Dynamical Systems, 6(1):208–235, 2007. CODEN SJADAY. ISSN 1536-0040.

Park:2021:HOA

[PW21] Youngmin Park and Dan D.

Wilson. High-order accuracy computation of coupling functions for strongly coupled oscillators. SIAM Journal on Applied Dynamical Systems, 20(3):1464–1484, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Park:2024:BBO

[PW24] Youngmin Park and Dan Wilson. N-body oscillator interactions of higher-order coupling functions. SIAM Journal on Applied Dynamical Systems, 23(2):1471–1503, ???? 2024. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/23M1594182.

Promislow:2014:ECB

[PY14] K. Promislow and L. Yang.

Existence of compressible bilayers in the functionalized Cahn–Hilliard equation. SIAM Journal on Applied Dynamical Systems, 13(2):629– 657, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Palacian:2006:AIS

[PYGR06] Jesús F. Palacián, Patricia Yanguas, and Susana Gutiérrez-Romero. Approximating the invariant sets of a finite straight segment near its collinear equilibria. SIAM Journal on Applied Dynamical Systems, 5(1):12–29, 2006. CODEN SJADAY. ISSN 1536-0040.

Puzyrev:2014:SPW

[PYVG14] D. Puzyrev, S. Yanchuk, A. G. Vladimirov, and S. V. Gurevich. Stability of plane wave solutions in complex Ginzburg–Landau equation with delayed feedback. SIAM Journal on Applied Dynamical Systems, 13(2):986–1009, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Pavliotis:2022:EME

[PZ22] Grigorios A. Pavliotis and Andrea Zanoni. Eigenfunction martingale estimators for interacting particle systems and their mean field limit. SIAM Journal on Applied Dynamical Systems, 21 (4):2338–2370, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1464348.

Qin:2021:HOA

[QCARL21] B. W. Qin, K. W. Chung, A. Algaba, and A. J. Rodríguez-Luis. High-order approximation of heteroclinic bifurcations in truncated 2Dnormal forms for the generic cases of Hopf-zero and nonresonant double Hopf singularities. SIAM Journal on Applied Dynamical Systems, 20(1):403–437, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Quinn:2019:EPF

[QSvdH19] Courtney Quinn, Jan Sieber, and Anna S. von der Heydt. Effects of periodic forcing on a paleoclimate delay model. SIAM Journal on Applied Dynamical Systems, 18(2):1060– 1077, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Quininao:2020:CSS

[QT20] Cristobal Quiñinao and Jonathan D.

Touboul. Clamping and synchronization in the strongly coupled FitzHugh–Nagumo model. SIAM Journal on Applied Dynamical Systems, 19 (2):788–827, ???? 2020. CO-DEN SJADAY. ISSN 1536-0040.

Rudy:2019:DDI

[RABK19] Samuel Rudy, Alessandro Alla, Steven L. Brunton, and J. Nathan Kutz. Data-driven identification of parametric partial differential equations. SIAM Journal on Applied Dynamical Systems, 18(2):643– 660, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Rademacher:2006:GRA

[Rad06] Jens D. M. Rademacher. Geometric relations of absolute and essential spectra of wave trains. SIAM Journal on Applied Dynamical Systems, 5 (4):634–649, 2006. CODEN SJADAY. ISSN 1536-0040.

Rademacher:2013:FSO

[Rad13] Jens D. M. Rademacher. First and second order semistrong interaction in reaction–diffusion systems. SIAM Journal on Applied Dynamical Systems, 12(1):175– 203, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Revel:2015:DRH

[RAM15] G. Revel, D. M. Alonso, and J. L. Moiola. A degenerate 2: 3 resonant Hopf–Hopf bifurcation as organizing center of the dynamics: Numerical semiglobal results. SIAM Journal on Applied Dynamical Systems, 14(2):1130–1164, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Rabinovich:2021:EEC

[RB21] Dmitry Rabinovich and Alfred M. Bruckstein. Erratic extremism causes dynamic consensus: a new model for opinion dynamics. SIAM Journal on Applied Dynamical Systems, 20(4):2077–2107, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Rakshit:2020:ISE

[RBBG20] Sarbendu Rakshit, Bidesh K.

Bera, Erik M. Bollt, and Dibakar Ghosh. Intralayer synchronization in evolving multiplex hypernetworks: Analytical approach. SIAM Journal on Applied Dynamical Systems, 19(2):918–963, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Righetti:2021:SFD

[RBI21] Ludovic Righetti, Jonas Buchli, and Auke Jan Ijspeert. Slow-fast dynamics of strongly coupled adaptive frequency oscillators. SIAM Journal on Applied Dynamical Systems, 20(4):1985–2012, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Rahman:2015:DNS

[RBK15] B. Rahman, K. B. Blyuss, and Y. N. Kyrychko. Dynamics of neural systems with discrete and distributed time delays. SIAM Journal on Applied Dynamical Systems, 14 (4):2069–2095, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Rodrigues:2023:AIS

[RC23] Alexandre A. P. Rodrigues and Luisa Castro. Abundance of infinite switching. SIAM Journal on Applied Dynamical Systems, 22(3):2570–2600, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M151371X.

Rotstein:2012:CLE

[RCG12] Horacio G. Rotstein, Stephen Coombes, and Ana Maria Gheorghe. Canard-like explosion of limit cycles in twodimensional piecewise-linear models of FitzHugh–Nagumo type. SIAM Journal on Applied Dynamical Systems, 11 (1):135–180, ???? 2012. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v11/i1/p135_s1.

Ravazzi:2023:ASD

[RCG^{+}23] Chiara Ravazzi, Giacomo Como, Michele Garetto, Emilio Leonardi, and Alberto Tarable. Asynchronous semianonymous dynamics over large-scale networks. SIAM Journal on Applied Dynamical Systems, 22(2):1300–1343, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M1492155.

Rada:2021:RMR

[RCLR21] Alejandra Rada, Cristian Coletti, Elcio Lebensztayn, and Pablo M. Rodriguez. The role of multiple repetitions on the size of a rumor. SIAM Journal on Applied Dynamical Systems, 20(3):1209–1231, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Rendall:2012:MNS

[Ren12] Alan D. Rendall. Mathematics of the NFAT signaling pathway. SIAM Journal on Applied Dynamical Systems, 11(3):988–1006, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Razavi:2016:SCC

[RGAB16] Hamed Razavi, Rohit Gupta, Fred C. Adams, and Anthony M. Bloch. Stability of a class of coupled Hill’s equations and the Lorentz oscillator model. SIAM Journal on Applied Dynamical Systems, 15(2):1104–1123, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Ruiz-Herrera:2013:PSC

[RHT13] Alfonso Ruiz-Herrera and Pedro J. Torres. Periodic solutions and chaotic dynamics in forced impact oscillators. SIAM Journal on Applied Dynamical Systems, 12(1):383– 414, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Rivera:2013:PSG

[Riv13] Andrés Rivera. Periodic solutions in the generalized Sitnikov (N + 1)-body problem. SIAM Journal on Applied Dynamical Systems, 12(3):1515– 1540, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Reverdy:2018:DSP

[RK18] Paul Reverdy and Daniel E.

Koditschek. A dynamical system for prioritizing and coordinating motivations. SIAM Journal on Applied Dynamical Systems, 17(2):1683–1715, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Rahman:2023:WDD

[RK23a] Aminur Rahman and J. Nathan Kutz. Walking droplets as a damped-driven system. SIAM Journal on Applied Dynamical Systems, 22 (2):1219–1233, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1486042.

Rosenfeld:2023:SDM

[RK23b] Joel A. Rosenfeld and Rushikesh Kamalapurkar. Singular dynamic mode decomposition. SIAM Journal on Applied Dynamical Systems, 22 (3):2357–2381, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1475892.

Russo:2024:CWS

[RL24] Benjamin P. Russo and M. Paul Laiu. Convergence of weak-SINDy surrogate models. SIAM Journal on Applied Dynamical Systems, 23(2): 1017–1051, April 2024. ISSN 1536-0040.

Rossides:2023:DIM

[RLZT23] T. Rossides, D. J. B. Lloyd, S. Zelik, and M. R. Turner.

The dynamics of interacting multi-pulses in the onedimensional quintic complex Ginzburg–Landau equation. SIAM Journal on Applied Dynamical Systems, 22 (3):2242–2281, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1519195.

Roberts:2004:SDA

[Rob04] A. J. Roberts. Shear dispersion along circular pipes is affected by bends, but the torsion of the pipe is negligible. SIAM Journal on Applied Dynamical Systems, 3(4):433– 462, 2004. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60088.

Roberts:2013:SRE

[Rob13] Gareth E. Roberts. Stability of relative equilibria in the planar n-vortex problem. SIAM Journal on Applied Dynamical Systems, 12(2):1114– 1134, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Roberts:2016:CER

[Rob16] Andrew Roberts. Canard explosion and relaxation oscillation in planar, piecewisesmooth, continuous systems. SIAM Journal on Applied Dynamical Systems, 15(1):609– 624, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Rodrigues:2021:ASA

[Rod21] Alexandre A. P. Rodrigues. Abundance of strange attractors near an attracting periodically perturbed network. SIAM Journal on Applied Dynamical Systems, 20(1):541– 570, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Rotstein:2022:NSR

[Rot22] Horacio G. Rotstein. Nonlinearities shape the response patterns to oscillatory inputs in a caricature electrochemical cell model. SIAM Journal on Applied Dynamical Systems, 21(1):500–522, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1402121.

Rombouts:2020:CTE

[RPY20] Jan Rombouts, Kiattisak Prathom, and Todd R. Young. Clusters tend to be of equal size in a negative feedback population model of cell cycle dynamics. SIAM Journal on Applied Dynamical Systems, 19(2):1540–1573, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Rozada:2014:SLS

[RRW14] Ignacio Rozada, Steven J. Ruuth, and M. J. Ward. The stability of localized spot patterns for the Brusselator on the sphere. SIAM Journal on Applied Dynamical Systems, 13(1):564–627, ???? 2014.

CODEN SJADAY. ISSN 1536-0040.

Roberts:2015:AFS

[RRW15] Kerry-Lyn Roberts, Jonathan E.

Rubin, and Martin Wechselberger. Averaging, folded singularities, and torus canards: Explaining transitions between bursting and spiking in a coupled neuron model. SIAM Journal on Applied Dynamical Systems, 14(4):1808– 1844, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Ross:2007:MGA

[RS07] Shane D. Ross and Daniel J.

Scheeres. Multiple gravity assists, capture, and escape in the restricted three-body problem. SIAM Journal on Applied Dynamical Systems, 6(3):576–596, 2007. CODEN SJADAY. ISSN 1536-0040.

Rucklidge:2009:DPF

[RS09] A. M. Rucklidge and M. Silber. Design of parametrically forced patterns and quasipatterns. SIAM Journal on Applied Dynamical Systems, 8 (1):298–347, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Reis:2011:LBP

[RS11] Timo Reis and Tatjana Stykel. Lyapunov balancing for passivity-preserving model reduction of RC circuits. SIAM Journal on Applied Dynamical Systems, 10 (1):1–34, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i1/p1_s1.

Rink:2013:AHB

[RS13] Bob W. Rink and Jan A. Sanders. Amplified Hopf bifurcations in feed-forward networks. SIAM Journal on Applied Dynamical Systems, 12(2):1135–1157, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Rusu:2016:BCC

[RS16] David Rusu and Manuele Santoprete. Bifurcations of central configurations in the four-body problem with some equal masses. SIAM Journal on Applied Dynamical Systems, 15(1):440–458, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Rademacher:2021:DWM

[RS21] Jens D. M. Rademacher and Lars Siemer. Domain wall motion in axially symmetric spintronic nanowires. SIAM Journal on Applied Dynamical Systems, 20(4):2204–2235, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Rubin:2021:TIR

[RSRT21] Jonathan E. Rubin, Justyna Signerska-Rynkowska, and Jonathan D. Touboul. Type III responses to transient inputs in hybrid nonlinear neuron models. SIAM Journal on Applied Dynamical Systems, 20(2):953–980, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Recke:2012:FLE

[RSTY12] Lutz Recke, Anatoly Samoilenko, Viktor Tkachenko, and Serhiy Yanchuk. Frequency locking by external forcing in systems with rotational symmetry. SIAM Journal on Applied Dynamical Systems, 11 (3):771–800, ???? 2012. CO-DEN SJADAY. ISSN 1536-0040.

Rubin:2002:SAL

[RT02] Jonathan Rubin and David Terman. Synchronized activity and loss of synchrony among heterogeneous conditional oscillators. SIAM Journal on Applied Dynamical Systems, 1(1):146–174, 2002. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40323.

Rotstein:2008:CIM

[RWK08] Horacio G. Rotstein, Martin Wechselberger, and Nancy Kopell. Canard induced mixed-mode oscillations in a medial entorhinal cortex layer II stellate cell model. SIAM Journal on Applied Dynamical Systems, 7(4):1582–1611, ???? 2008. CODEN SJA-DAY. ISSN 1536-0040.

Rodriguez:2021:UEH

[RWZ21] Nancy Rodríguez, Qi Wang, and Lu Zhang. Understanding the effects of on- and offhotspot policing: Evidence of hotspot, oscillating, and chaotic activities. SIAM Journal on Applied Dynamical Systems, 20(4):1882–1916, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Soudjani:2013:ASG

[SA13] Sadegh Esmaeil Zadeh Soudjani and Alessandro Abate. Adaptive and sequential gridding procedures for the abstraction and verification of stochastic processes. SIAM Journal on Applied Dynamical Systems, 12(2):921–956, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Saldana:2018:ASP

[SAA^{+}18] Joan Saldaña, Maria Aguareles, Albert Avinyó, Marta Pellicer, and Jordi Ripoll. An age-structured population approach for the mathematical modeling of urban burglaries. SIAM Journal on Applied Dynamical Systems, 17(4):2733– 2760, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Sequeira:2021:CMF

[SAH21] Pedro M. Sequeira, António P. Aguiar, and João Hespanha. Commutative monoid formalism for weighted coupled cell networks and invariant synchrony patterns. SIAM Journal on Applied Dynamical Systems, 20(3):1485–1513, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Saadatpour:2013:RMB

[SAR13] Assieh Saadatpour, Réka Albert, and Timothy C. Reluga. A reduction method for Boolean network models proven to conserve attractors. SIAM Journal on Applied Dynamical Systems, 12(4):1997– 2011, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Slowinski:2020:NMC

[SARTA20] Piotr Słowiński, Sohaib Al-Ramadhani, and Krasimira Tsaneva-Atanasova. Neurologically motivated coupling functions in models of motor coordination. SIAM Journal on Applied Dynamical Systems, 19(1):208–232, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Shoval:2011:SIA

[SAS11] Oren Shoval, Uri Alon, and Eduardo Sontag. Symmetry invariance for adapting biological systems. SIAM Journal on Applied Dynamical Systems, 10(3):857– 886, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i3/p857_s1.

Schlomer:2012:NBS

[SAV12] Nico Schlömer, Daniele Avitabile, and Wim Vanroose. Numerical bifurcation study of superconducting patterns on a square. SIAM Journal on Applied Dynamical Systems, 11 (1):447–477, ???? 2012. CO-DEN SJADAY. ISSN 1536-0040.

Spiller:2010:ISD

[SB10] Elaine T. Spiller and Gino Biondini. Importance sampling for dispersion-managed solitons. SIAM Journal on Applied Dynamical Systems, 9(2):432–461, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Short:2010:NPU

[SBB10] M. B. Short, A. L. Bertozzi, and P. J. Brantingham. Nonlinear patterns in urban crime: Hotspots, bifurcations, and suppression. SIAM Journal on Applied Dynamical Systems, 9(2):462–483, ???? 2010. CODEN SJA-DAY. ISSN 1536-0040.

Selley:2015:DCM

[SBKS15] Fanni Sélley, Ádám Besenyei, Istvan Z. Kiss, and Péter L. Simon. Dynamic control of modern, network-based epidemic models. SIAM Journal on Applied Dynamical Systems, 14(1):168–187, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Sun:2009:CGS

[SBN09] Jie Sun, Erik M. Bollt, and Takashi Nishikawa. Constructing generalized synchronization manifolds by manifold equation. SIAM Journal on Applied Dynamical Systems, 8(1):202–221, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Stilwell:2006:SCF

[SBR06] Daniel J. Stilwell, Erik M.

Bollt, and D. Gray Roberson. Sufficient conditions for fast switching synchronization in time-varying network topologies. SIAM Journal on Applied Dynamical Systems, 5 (1):140–156, 2006. CODEN SJADAY. ISSN 1536-0040.

Stiadle:2023:ODC

[SBV23] Thomas I. Stiadle, Alvin Bayliss, and Vladimir A. Volpert. Order and disorder in a cyclically competitive ecological community. SIAM Journal on Applied Dynamical Systems, 22(1):129–161, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/21M145923X.

Sun:2007:UAN

[SCD07] Chunyou Sun, Daomin Cao, and Jinqiao Duan. Uniform attractors for nonautonomous wave equations with nonlinear damping. SIAM Journal on Applied Dynamical Systems, 6(2):293–318, 2007. CODEN SJADAY. ISSN 1536-0040.

Sclosa:2023:KNI

[Scl23] Davide Sclosa. Kuramoto networks with infinitely many stable equilibria. SIAM Journal on Applied Dynamical Systems, 22(4):3267–3283, November 2023. ISSN 1536-0040.

Scott:2013:DPF

[Sco13] Sherry E. Scott. Different perspectives and formulas for capturing deviation from ergodicity. SIAM Journal on Applied Dynamical Systems, 12(4):1948–1967, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Sewalt:2017:SPM

[SD17] Lotte Sewalt and Arjen Doelman. Spatially periodic multipulse patterns in a generalized Klausmeier–Gray–Scott model. SIAM Journal on Applied Dynamical Systems, 16 (2):1113–1163, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Smith:2022:TES

[SD22] Spencer A. Smith and Sierra Dunn. Topological entropy of surface braids and maximally efficient mixing. SIAM Journal on Applied Dynamical Systems, 21(2):1209–1244, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M142647X.

Sanford:2020:SRS

[SDK20] Nathan L. Sanford, Graham M. Donovan, and William L. Kath. Slip rates and slip modes in an actively mode-locked laser. SIAM Journal on Applied Dynamical Systems, 19(2):1472–1495, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Schmidt:2009:PSG

[SDR09] Sven Schmidt, Holger R.

Dullin, and Peter H. Richter. A Poincaré section for the general heavy rigid body. SIAM Journal on Applied Dynamical Systems, 8(1):371– 389, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Saghafi:2017:ETD

[SDT17] Mehdi Saghafi, Harry Dankowicz, and Whitney Tabor. Emergent task differentiation on network filters. SIAM Journal on Applied Dynamical Systems, 16(3):1686–1709, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Saghafi:2015:UNB

[SDW15] Mehdi Saghafi, Harry Dankowicz, and Matthew West. On the use of nonlinear boundary-value problems to estimate the cloud-formation potential of aerosol particles. SIAM Journal on Applied Dynamical Systems, 14(2):822– 859, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Sherwood:2010:DPR

[SG10] William Erik Sherwood and John Guckenheimer. Dissecting the phase response of a model bursting neuron. SIAM Journal on Applied Dynamical Systems, 9(3):659–703, ???? 2010. CODEN SJA-DAY. ISSN 1536-0040.

Stewart:2011:SBB

[SG11] Ian Stewart and Martin Golubitsky. Synchrony-breaking bifurcation at a simple real eigenvalue for regular networks 1: 1-dimensional cells. SIAM Journal on Applied Dynamical Systems, 10(4): 1404–1442, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i4/p1404_s1.

Stewart:2003:SGP

[SGP03] Ian Stewart, Martin Golubitsky, and Marcus Pivato. Symmetry groupoids and patterns of synchrony in coupled cell networks. SIAM Journal on Applied Dynamical Systems, 2(4):609–646, 2003. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
41989.

Seidel:2009:MCF

[SGW09] T. Seidel, I. Gasser, and B. Werner. Microscopic carfollowing models revisited: From road works to fundamental diagrams. SIAM Journal on Applied Dynamical Systems, 8(3):1305–1323, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Sequeira:2023:DAF

[SHA23] Pedro M. Sequeira, João P.

Hespanha, and A. Pedro Aguiar. Decomposition of admissible functions in weighted coupled cell networks. SIAM Journal on Applied Dynamical Systems, 22(2):1114–1152, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M1471328.

Shirani:2017:GDE

[SHdlL17] Farshad Shirani, Wassim M.

Haddad, and Rafael de la Llave. On the global dynamics of an electroencephalographic mean field model of the neocortex. SIAM Journal on Applied Dynamical Systems, 16(4):1969–2029, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Sherratt:2014:PTW

[She14] Jonathan A. Sherratt. Periodic traveling waves in integrodifferential equations for nonlocal dispersal. SIAM Journal on Applied Dynamical Systems, 13(4):1517–1541, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Simpson:2013:SRG

[SHK13] D. J. W. Simpson, S. J.

Hogan, and R. Kuske.

Stochastic regular grazing bifurcations. SIAM Journal on Applied Dynamical Systems, 12(2):533–559, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Sieber:2002:NBA

[Sie02] Jan Sieber. Numerical bifurcation analysis for multisection semiconductor lasers. SIAM Journal on Applied Dynamical Systems, 1(2):248– 270, 2002. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40174.

Song:2017:SDD

[SJLY17] Yongli Song, Heping Jiang, Quan-Xing Liu, and Yuan Yuan. Spatiotemporal dynamics of the diffusive mussel–algae model near Turing–Hopf bifurcation. SIAM Journal on Applied Dynamical Systems, 16(4):2030–2062, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Spiller:2008:MDM

[SK08] Elaine T. Spiller and William L. Kath. A method for determining most probable errors in nonlinear lightwave systems. SIAM Journal on Applied Dynamical Systems, 7 (3):868–894, 2008. CODEN SJADAY. ISSN 1536-0040.

Serenevy:2013:EHP

[SK13] A. K. Serenevy and N. J. Kopell. Effects of heterogeneous periodic forcing on inhibitory networks. SIAM Journal on Applied Dynamical Systems, 12(3):1649–1684, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Schlosser:2022:SSK

[SK22] Corbinian Schlosser and Milan Korda. Sparsity structures for Koopman and Perron–Frobenius operators. SIAM Journal on Applied Dynamical Systems, 21(3): 2187–2214, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1466608.

Singhal:2023:OPS

[SKL23] Bharat Singhal, István Z.

Kiss, and Jr-Shin Li. Optimal phase-selective entrainment of heterogeneous oscillator ensembles. SIAM Journal on Applied Dynamical Systems, 22(3):2180–2205, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1521201.

Schwemmer:2012:BLI

[SL12] Michael A. Schwemmer and Timothy J. Lewis. Bistability in a leaky integrate-and-fire neuron with a passive dendrite. SIAM Journal on Applied Dynamical Systems, 11 (1):507–539, ???? 2012. CO-DEN SJADAY. ISSN 1536-

Sun:2019:ATD

[SL19] Hao Sun and Junfeng Li. Application of topological degree theory to the equilibrium point in irregular gravitational field. SIAM Journal on Applied Dynamical Systems, 18(2):629–642, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Slavik:2020:LVC

[Sla20]  Antonín Slavík.  Lotka–

Volterra competition model on graphs. SIAM Journal on Applied Dynamical Systems, 19(2):725–762, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Sun:2024:AEC

[SLF24] Mengfeng Sun, Yijun Lou, and Xinchu Fu. Analysis of equilibria and connecting orbits in a nonlinear viral infection model. SIAM Journal on Applied Dynamical Systems, 23(2):1272–1312, May 2024. ISSN 1536-0040.

Schmutz:2023:FSN

[SLS23] Valentin Schmutz, Eva Löcherbach, and Tilo Schwalger. On a finite-size neuronal population equation. SIAM Journal on Applied Dynamical Systems, 22(2):996–1029, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1445041.

Simpson:2008:NSB

[SM08] D. J. W. Simpson and J. D.

Meiss. Neimark–Sacker bifurcations in planar, piecewisesmooth, continuous maps. SIAM Journal on Applied Dynamical Systems, 7(3):795– 824, 2008. CODEN SJADAY. ISSN 1536-0040.

Sigrist:2011:SSP

[SM11] Rachel Sigrist and Paul Matthews. Symmetric spiral patterns on spheres. SIAM Journal on Applied Dynamical Systems, 10(3):1177– 1211, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i3/p1177_s1.

Susuki:2020:ISQ

[SM20] Yoshihiko Susuki and Igor Mezić. Invariant sets in quasiperiodically forced dynamical systems. SIAM Journal on Applied Dynamical Systems, 19(1):329–351, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Smith:2024:SHH

[Smi24] Lauren D. Smith. Swarmalators with higher harmonic coupling: Clustering and vacillating. SIAM Journal on Applied Dynamical Systems, 23(2):1133–1158, May 2024. ISSN 1536-0040.

Susuki:2021:KRL

[SMM21] Yoshihiko Susuki, Alexandre Mauroy, and Igor Mezić. Koopman resolvent: a Laplacedomain analysis of nonlinear autonomous dynamical systems. SIAM Journal on Applied Dynamical Systems, 20 (4):2013–2036, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Singh:2011:ENR

[SMRB11] Thounaojam Umeshkanta Singh, Kaustubh Manchanda, Ramakrishna Ramaswamy, and Amitabha Bose. Excitable nodes on random graphs: Relating dynamics to network structure. SIAM Journal on Applied Dynamical Systems, 10(3): 987–1012, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i3/p987_s1.

Sieber:2018:CEF

[SMS18] Jan Sieber, Christian Marschler, and Jens Starke. Convergence of equation-free methods in the case of finite time scale separation with application to deterministic and stochastic systems. SIAM Journal on Applied Dynamical Systems, 17(4):2574–2614, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Szalai:2009:ATA

[SO09] Róbert Szalai and Hinke M.

Osinga. Arnol’d tongues arising from a grazing-sliding bifurcation. SIAM Journal on Applied Dynamical Systems, 8(4):1434–1461, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Soares:2017:SBL

[Soa17] Pedro Soares. Synchrony branching lemma for regular networks. SIAM Journal on Applied Dynamical Systems, 16(4):1869–1892, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Staerk-Ostergaard:2024:HDC

[SØRD24] Jacob Stærk-Østergaard, Anders Rahbek, and Susanne Ditlevsen. High-dimensional cointegration and Kuramoto inspired systems. SIAM Journal on Applied Dynamical Systems, 23(1):236–255, January 2024. ISSN 1536-0040.

Schilder:2005:CQP

[SOV05] Frank Schilder, Hinke M.

Osinga, and Werner Vogt. Continuation of quasi-periodic invariant tori. SIAM Journal on Applied Dynamical Systems, 4(3):459–488, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/61124.

Schanz:2003:SSA

[SP03] Michael Schanz and Axel Pelster. Synergetic system analysis for the delay-induced Hopf bifurcation in the Wright equation. SIAM Journal on Applied Dynamical Systems, 2(3):277–296, 2003. CODEN SJADAY. ISSN 1536-0040.

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        41280.

Schwarze:2021:MPN

[SP21] Alice C. Schwarze and Mason A. Porter. Motifs for processes on networks. SIAM Journal on Applied Dynamical Systems, 20(4):2516–2557, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Shaw:2012:PRA

[SPCT12] Kendrick M. Shaw, Young-Min Park, Hillel J. Chiel, and Peter J. Thomas. Phase resetting in an asymptotically phaseless system: On the phase response of limit cycles verging on a heteroclinic orbit. SIAM Journal on Applied Dynamical Systems, 11 (1):350–391, ???? 2012. CO-DEN SJADAY. ISSN 1536-0040.

Speetjens:2008:SAT

[SRMPM08] M. Speetjens, A. Reusken, S. Maier-Paape, and W. Marquardt. Stability analysis of two-dimensional pool-boiling systems. SIAM Journal on Applied Dynamical Systems, 7(3):933–961, 2008. CODEN SJADAY. ISSN 1536-0040.

Smith:2009:ASW

[SRS09] Matthew J. Smith, Jens D. M. Rademacher, and Jonathan A. Sherratt. Absolute stability of wavetrains can explain spatiotemporal dynamics in reaction– diffusion systems of lambdaomega type. SIAM Journal on Applied Dynamical Systems, 8(3):1136–1159, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Stanhope:2014:ILL

[SRS14] S. Stanhope, J. E. Rubin, and D. Swigon. Identifiability of linear and linearin-parameters dynamical systems from a single trajectory. SIAM Journal on Applied Dynamical Systems, 13(4):1792– 1815, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Sandstede:2004:DOM

[SS04] Björn Sandstede and Arnd Scheel. Defects in oscillatory media: Toward a classification. SIAM Journal on Applied Dynamical Systems, 3(1):1–68, 2004. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
60019.

Sandstede:2007:PDS

[SS07] Björn Sandstede and Arnd Scheel. Period-doubling of spiral waves and defects. SIAM Journal on Applied Dynamical Systems, 6(2):494– 547, 2007. CODEN SJADAY. ISSN 1536-0040.

Schecter:2009:PRU

[SS09] Stephen Schecter and Peter Szmolyan. Persistence of rarefactions under Dafermos regularization: Blowup and an exchange lemma for gain-of-stability turning points. SIAM Journal on Applied Dynamical Systems, 8 (3):822–853, ???? 2009. CO-DEN SJADAY. ISSN 1536-0040.

Sieber:2011:CML

[SS11] Jan Sieber and Robert Szalai. Characteristic matrices for linear periodic delay differential equations. SIAM Journal on Applied Dynamical Systems, 10(1): 129–147, ???? 2011. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.
                                        siam.org/siads/resource/
                                        1/sjaday/v10/i1/p129_s1.

Slater:2012:DSC

[SS12] D. M. Slater and P. H. Steen. Detecting symmetry in the chaotic and quasiperiodic motions of three coupled droplet oscillators. SIAM Journal on Applied Dynamical Systems, 11(3):1098– 1113, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Stepien:2014:TWO

[SS14] Tracy L. Stepien and David Swigon. Traveling waves in a one-dimensional elastic continuum model of cell layer migration with stretch-dependent proliferation. SIAM Journal on Applied Dynamical Systems, 13 (4):1489–1516, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Shoshani:2016:GPR

[SS16] O. Shoshani and S. W. Shaw.

Generalized parametric resonance. SIAM Journal on Applied Dynamical Systems, 15 (2):767–788, ???? 2016. CO-DEN SJADAY. ISSN 1536-0040.

Sherratt:2010:PSS

[SSR10] Jonathan A. Sherratt, Matthew J.

Smith, and Jens D. M. Rademacher. Patterns of sources and sinks in the complex Ginzburg–Landau equation with zero linear dispersion. SIAM Journal on Applied Dynamical Systems, 9 (3):883–918, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Schwieger:2018:CTS

[SSR18] Robert Schwieger, Heike Siebert, and Susanna Röblitz. Correspondence of trap spaces in different models of bioregulatory networks. SIAM Journal on Applied Dynamical Systems, 17(2):1742–1765, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Saez:2006:SZB

[SSS06] Eduardo Sáez, Eduardo Stange, and Iván Szántó. Simultaneous zip bifurcation and limit cycles in three dimensional competition models. SIAM Journal on Applied Dynamical Systems, 5(1):1– 11, 2006. CODEN SJADAY. ISSN 1536-0040.

Shih:2013:GAS

[ST13] Chih-Wen Shih and Jui-Pin Tseng. A general approach to synchronization of coupled cells. SIAM Journal on Applied Dynamical Systems, 12 (3):1354–1393, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Sun:2015:CNI

[STB15] Jie Sun, Dane Taylor, and Erik M. Bollt. Causal network inference by optimal causation entropy. SIAM Journal on Applied Dynamical Systems, 14(1):73–106, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Stewart:2014:SBB

[Ste14]  Ian Stewart.  Synchrony-

breaking bifurcation at a simple real eigenvalue for regular networks 2: Higherdimensional cells. SIAM Journal on Applied Dynamical Systems, 13(1):129–156, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Steinerberger:2023:MCK

[Ste23] Stefan Steinerberger. Maxcut via Kuramoto-type oscillators. SIAM Journal on Applied Dynamical Systems, 22(2):730–743, ???? 2023.

CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1432211.

Shih:2023:ASS

[STW23] Chih-Wen Shih, Jui-Pin Tseng, and Chang-Hong Wu. Absolute stability and synchronization in second-order neural fields with conduction delays. SIAM Journal on Applied Dynamical Systems, 22(2):878–917, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1441109.

Saiki:2023:HDC

[STY23] Yoshitaka Saiki, Hiroki Takahasi, and James A. Yorke. Hausdorff dimension of Cantor intersections and robust heterodimensional cycles for heterochaos horseshoe maps. SIAM Journal on Applied Dynamical Systems, 22(3):1852–1876, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1504986.

Sultanov:2023:LTB

[Sul23] Oskar A. Sultanov. Longterm behaviour of asymptotically autonomous Hamiltonian systems with multiplicative noise. SIAM Journal on Applied Dynamical Systems, 22(3):1818–1851, ???? 2023. CODEN SJADAY.

ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1504731.

Smith:2018:OMC

[SUOL18] Lachlan D. Smith, Paul B. Umbanhowar, Julio M. Ottino, and Richard M. Lueptow. Optimized mixing by cutting-and-shuffling. SIAM Journal on Applied Dynamical Systems, 17(4):2544–2573, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Steur:2016:CCP

[SÜvLM16] E. Steur, H. U. Ünal, C. van Leeuwen, and W. Michiels. Characterization and computation of partial synchronization manifolds for diffusive delay-coupled systems. SIAM Journal on Applied Dynamical Systems, 15(4):1874–1915, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Sahai:2009:NMA

[SV09] Tuhin Sahai and Alexander Vladimirsky. Numerical methods for approximating invariant manifolds of delayed systems. SIAM Journal on Applied Dynamical Systems, 8(3):1116–1135, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Spek:2024:HBT

[SvGKP24] Len Spek, Stephan A. van Gils, Yuri A. Kuznetsov, and Mónika Polner. Hopf bifurcations of two population neural fields on the sphere with diffusion and distributed delays. SIAM Journal on Applied Dynamical Systems, 23 (3):1909–1945, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1554011.

Stephens:2014:RVI

[SW14] Thomas Stephens and Thomas Wanner. Rigorous validation of isolating blocks for flows and their Conley indices. SIAM Journal on Applied Dynamical Systems, 13 (4):1847–1878, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Sander:2016:VSN

[SW16] Evelyn Sander and Thomas Wanner. Validated saddlenode bifurcations and applications to lattice dynamical systems. SIAM Journal on Applied Dynamical Systems, 15(3):1690–1733, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Sakajo:2021:SDR

[SW21] Takashi Sakajo and Penghao Wang. Spot dynamics of a reaction–diffusion system on the surface of a torus. SIAM Journal on Applied Dynamical Systems, 20(2):1053–1089, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Stewart:2024:SSP

[SW24] Ian Stewart and David Wood.

Stable synchronous propagation of signals by feedforward networks. SIAM Journal on Applied Dynamical Systems, 23(1):167–204, January 2024. ISSN 1536-0040.

Sun:2005:SDT

[SWR05] Wentao Sun, Michael J. Ward, and Robert Russell. The slow dynamics of twospike solutions for the Gray– Scott and Gierer–Meinhardt systems: Competition and oscillatory instabilities. SIAM Journal on Applied Dynamical Systems, 4(4):904–953, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/62099.

Szczelina:2013:HOP

[SZ13] Robert Szczelina and Piotr Zgliczyński. A homoclinic orbit in a planar singular ODE — a computer assisted proof. SIAM Journal on Applied Dynamical Systems, 12(3):1541– 1565, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Takahasi:2016:EMT

[Tak16] Hiroki Takahasi. Equilibrium measures at temperature zero for Hénon-like maps at the first bifurcation. SIAM Journal on Applied Dynamical Systems, 15(1):106–124, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Tudoran:2009:HVR

[TAtN09] Răzvan M. Tudoran, Anania Aron, and tefan Nicoară.

On a Hamiltonian version of the Rikitake system. SIAM Journal on Applied Dynamical Systems, 8(1):454–479, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Touboul:2009:SDB

[TB09] Jonathan Touboul and Romain Brette. Spiking dynamics of bidimensional integrateand-fire neurons. SIAM Journal on Applied Dynamical Systems, 8(4):1462–1506, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Tumelty:2023:TCS

[TBR23] J. Tumelty, Cedric Beaume, and Alastair M. Rucklidge. Toward convectons in the supercritical regime: Homoclinic snaking in natural doubly diffusive convection. SIAM Journal on Applied Dynamical Systems, 22 (3):1710–1742, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1510303.

Tang:2021:GDL

[TC21] De Tang and Yuming Chen.

Global dynamics of a Lotka– Volterra competition–diffusion system in advective heterogeneous environments. SIAM Journal on Applied Dynamical Systems, 20(3):1232–1252, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Tang:2023:BLC

[TC23] Hong Tang and Alan Champneys. Bifurcation of limit cycles from boundary equilibria in impacting hybrid systems. SIAM Journal on Applied Dynamical Systems, 22(4):3320– 3357, December 2023. ISSN 1536-0040.

Thota:2008:TNT

[TD08] Phanikrishna Thota and Harry Dankowicz. TC-HAT (T̂ C): a novel toolbox for the continuation of periodic trajectories in hybrid dynamical systems. SIAM Journal on Applied Dynamical Systems, 7(4):1283–1322, 2008. CO-DEN SJADAY. ISSN 1536-0040.

Tong:2012:EPZ

[TD12] William Tong and Holger R. Dullin. The equilateral pentagon at zero angular momentum: Maximal rotation through optimal deformation. SIAM Journal on Applied Dynamical Systems, 11(3):963– 987, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Trichtchenko:2018:SPT

[TDK18] Olga Trichtchenko, Bernard Deconinck, and Richard Kollár. Stability of periodic traveling wave solutions to the Kawahara equation. SIAM Journal on Applied Dynamical Systems, 17(4):2761– 2783, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Thorin:2017:NMA

[TDL17] Anders Thorin, Pierre Delezoide, and Mathias Legrand. Nonsmooth modal analysis of piecewise-linear impact oscillators. SIAM Journal on Applied Dynamical Systems, 16 (3):1710–1747, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Torres:2021:SPB

[TF21] Angélica Torres and Elisenda Feliu. Symbolic proof of bistability in reaction networks. SIAM Journal on Applied Dynamical Systems, 20(1):1–37, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Thorne:2021:MST

[TFBN21] Michael A. S. Thorne, Eric Forgoston, Lora Billings, and Anje-Margriet Neutel. Matrix scaling and tipping points. SIAM Journal on Applied Dynamical Systems, 20(2):1090– 1103, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Tonello:2019:LNC

[TFC19] Elisa Tonello, Etienne Farcot, and Claudine Chaouiya. Local negative circuits and cyclic attractors in Boolean networks with at most five components. SIAM Journal on Applied Dynamical Systems, 18(1):68–79, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Takeishi:2022:DDM

[TFTK22] Naoya Takeishi, Keisuke Fujii, Koh Takeuchi, and Yoshinobu Kawahara. Discriminant dynamic mode decomposition for labeled spatiotemporal data collections. SIAM Journal on Applied Dynamical Systems, 21(2): 1030–1058, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1399907.

Tomlin:2019:OCT

[TGPP19] Ruben J. Tomlin, Susana N. Gomes, Grigorios A. Pavliotis, and Demetrios T. Papageorgiou. Optimal control of thin liquid films and transverse mode effects. SIAM Journal on Applied Dynamical Systems, 18(1):117–149, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Touboul:2012:NIB

[THF12] Jonathan Touboul, Geoffroy Hermann, and Olivier Faugeras. Noise-induced behaviors in neural mean field dynamics. SIAM Journal on Applied Dynamical Systems, 11(1):49–81, ???? 2012. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v11/i1/p49_s1.

Terrien:2017:BAY

[TKB17] Soizic Terrien, Bernd Krauskopf, and Neil G. R. Broderick.

Bifurcation analysis of the Yamada model for a pulsing semiconductor laser with saturable absorber and delayed optical feedback. SIAM Journal on Applied Dynamical Systems, 16(2):771–801, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Tzou:2016:WNA

[TKKCG16] J. C. Tzou, P. G. Kevrekidis, T. Kolokolnikov, and R. Carretero-González. Weakly nonlinear analysis of vortex formation in a dissipative variant of the Gross–Pitaevskii equation. SIAM Journal on Applied Dynamical Systems, 15 (2):904–922, ???? 2016. CO-DEN SJADAY. ISSN 1536-0040.

Tian:2024:SSB

[TL24] Yu Tian and Renaud Lambiotte. Spreading and structural balance on signed networks. SIAM Journal on Applied Dynamical Systems, 23(1):50–80, January 2024. ISSN 1536-0040.

Terman:2011:SAP

[TLRB11] D. Terman, E. Lee, J. Rinzel, and T. Bem. Stability of anti-phase and in-phase locking by electrical coupling but not fast inhibition alone. SIAM Journal on Applied Dynamical Systems, 10(3): 1127–1153, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.

                                        siam.org/siads/resource/
                                        1/sjaday/v10/i3/p1127_s1.

Tonnelier:2010:PSS

[Ton10] Arnaud Tonnelier. Propagation of spike sequences in neural networks. SIAM Journal on Applied Dynamical Systems, 9(3):1090–1118, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Tonnelier:2019:SPA

[Ton19] Arnaud Tonnelier. Signal propagations along excitable chains. SIAM Journal on Applied Dynamical Systems, 18 (3):1391–1419, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Tonnelier:2023:SSC

[Ton23] Arnaud Tonnelier. Sustainability or societal collapse: Dynamics and bifurcations of the HANDY model. SIAM Journal on Applied Dynamical Systems, 22(3):1877–1905, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M1494336.

Tamekue:2024:MMV

[TPC24] Cyprien Tamekue, Dario Prandi, and Yacine Chitour. A mathematical model of the visual MacKay effect. SIAM Journal on Applied Dynamical Systems, 23(3):2138–2178, ???? 2024. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/23M1616686.

Tian:2019:PFS

[TR19] Canrong Tian and Shigui Ruan. Pattern formation and synchronism in an allelopathic plankton model with delay in a network. SIAM Journal on Applied Dynamical Systems, 18(1):531–557, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Troy:2008:TWS

[Tro08] William C. Troy. Traveling waves and synchrony in an excitable large-scale neuronal network with asymmetric connections. SIAM Journal on Applied Dynamical Systems, 7(4):1247–1282, 2008. CODEN SJADAY. ISSN 1536-0040.

Troy:2007:PFF

[TS07] William C. Troy and Vladimir Shusterman. Patterns and features of families of traveling waves in large-scale neuronal networks. SIAM Journal on Applied Dynamical Systems, 6(1):263–292, 2007. CODEN SJADAY. ISSN 1536-0040.

Torchinsky:2023:MME

[TS23] Jason L. Torchinsky and Samuel N. Stechmann. Mitigating model error via a multimodel method and application to tropical intraseasonal oscillations. SIAM Journal on Applied Dynamical Systems, 22(4):3025–3058, November 2023. ISSN 1536-0040.

Tzou:2020:ASP

[TT20] J. C. Tzou and L. Tzou. Analysis of spot patterns on a coordinate-invariant model for vegetation on a curved terrain. SIAM Journal on Applied Dynamical Systems, 19 (4):2500–2529, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Tupper:2005:ENS

[Tup05] P. F. Tupper. Ergodicity and the numerical simulation of Hamiltonian systems. SIAM Journal on Applied Dynamical Systems, 4(3):563–587, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/60380.

Tupper:2009:RBA

[Tup09] Paul Tupper. The relation between approximation in distribution and shadowing in molecular dynamics. SIAM Journal on Applied Dynamical Systems, 8(2):734–755, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Terragni:2014:CBD

[TV14] Filippo Terragni and José M. Vega. Construction of bifurcation diagrams using POD on the fly. SIAM Journal on Applied Dynamical Systems, 13(1):339–365, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Teramoto:2021:TPS

[TvH21] Takashi Teramoto and Peter van Heijster. Traveling pulse solutions in a three-component FitzHugh– Nagumo model. SIAM Journal on Applied Dynamical Systems, 20(1):371–402, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Thavanayagam:2018:MSD

[TW18a] Ealasukanthan Thavanayagam and David J. N. Wall. Modeling of spatial dynamical silence in the macro arterial domain. SIAM Journal on Applied Dynamical Systems, 17(3):2176–2204, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Tse:2018:AIC

[TW18b] Wang Hung Tse and Michael J.

Ward. Asynchronous instabilities of crime hotspots for a 1-D reaction–diffusion model of urban crime with focused police patrol. SIAM Journal on Applied Dynamical Systems, 17(3):2018–2075, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Tang:2023:HBR

[TW23] Xiaoxian Tang and Kaizhang Wang. Hopf bifurcations of reaction networks with zero-one stoichiometric coefficients. SIAM Journal on Applied Dynamical Systems, 22(3):2459–2489, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1519754.

Tzou:2018:ASH

[TWW18] J. C. Tzou, M. J. Ward, and J. C. Wei. Anomalous scaling of Hopf bifurcation thresholds for the stability of localized spot patterns for reaction–diffusion systems in two dimensions. SIAM Journal on Applied Dynamical Systems, 17(1):982–1022, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Tang:2021:MSR

[TX21] Xiaoxian Tang and Hao Xu. Multistability of small reaction networks. SIAM Journal on Applied Dynamical Systems, 20(2):608–635, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Tzou:2017:SSD

[TXKW17] J. C. Tzou, S. Xie, T. Kolokolnikov, and M. J. Ward. The stability and slow dynamics of localized spot patterns for the 3-D Schnakenberg reaction– diffusion model. SIAM Journal on Applied Dynamical Systems, 16(1):294–336, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Tempkin:2007:SLE

[TY07] Joshua A. Tempkin and James A. Yorke. Spurious Lyapunov exponents computed from data. SIAM Journal on Applied Dynamical Systems, 6(2):457–474, 2007. CODEN SJADAY. ISSN 1536-0040.

Tang:2022:MRN

[TZ22] Xiaoxian Tang and Zhishuo Zhang. Multistability of reaction networks with onedimensional stoichiometric subspaces. SIAM Journal on Applied Dynamical Systems, 21(2):1426–1454, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1424676.

Tsai:2012:TWS

[TZKS12] Je-Chiang Tsai, Wenjun Zhang, Vivien Kirk, and James Sneyd. Traveling waves in a simplified model of calcium dynamics. SIAM Journal on Applied Dynamical Systems, 11(4):1149–1199, ???? 2012. CODEN SJA-DAY. ISSN 1536-0040.

Udeigwe:2015:WPR

[UE15] Lawrence C. Udeigwe and G. Bard Ermentrout. Waves and patterns on regular graphs. SIAM Journal on Applied Dynamical Systems, 14(2):1102–1129, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Umbria:2021:CDH

[UN21] J. Sánchez Umbría and M. Net. Continuation of double Hopf points in thermal convection of rotating fluid spheres. SIAM Journal on Applied Dynamical Systems, 20(1):208–231, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

Ueda:2005:MTT

[USW05] Takahiro Ueda, Ken Sasaki, and Shinsuke Watanabe. Motion of the Tippe top: Gyroscopic balance condition and stability. SIAM Journal on Applied Dynamical Systems, 4(4):1159–1194, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/61598.

Uecker:2014:NRS

[UW14] Hannes Uecker and Daniel Wetzel. Numerical results for snaking of patterns over patterns in some 2D Selkov– Schnakenberg reaction–diffusion systems. SIAM Journal on Applied Dynamical Systems, 13(1):94–128, ???? 2014. CO-DEN SJADAY. ISSN 1536-0040.

Varkonyi:2021:DEC

[VA21] Peter L. Varkonyi and Mate Antali. On differential equations with codimension-n discontinuity sets. SIAM Journal on Applied Dynamical Systems, 20(3):1348–1381, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Vanneste:2006:WRB

[Van06] J. Vanneste. Wave radiation by balanced motion in a simple model. SIAM Journal on Applied Dynamical Systems, 5(4):783–807, 2006. CODEN SJADAY. ISSN 1536-0040.

Vanneste:2008:ASM

[Van08] J. Vanneste. Asymptotics of a slow manifold. SIAM Journal on Applied Dynamical Systems, 7(4):1163–1190, 2008. CODEN SJADAY. ISSN 1536-0040.

Vassena:2023:SCS

[Vas23] Nicola Vassena. Structural conditions for saddle-node bifurcations in chemical reaction networks. SIAM Journal on Applied Dynamical Systems, 22(3):1639–1672, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1527933.

Votel:2009:ECT

[VBG^{+}09] Ronald Votel, David A. W.

Barton, Takahide Gotou, Takeshi Hatanaka, Masayuki Fujita, and Jeff Moehlis. Equilibrium configurations for a territorial model. SIAM Journal on Applied Dynamical Systems, 8(3):1234–1260, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Vo:2013:MGV

[VBW13] Theodore Vo, Richard Bertram, and Martin Wechselberger. Multiple geometric viewpoints of mixed mode dynamics associated with pseudoplateau bursting. SIAM Journal on Applied Dynamical Systems, 12(2):789–830, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Veliz-Cuba:2012:AAR

[VC12] Alan Veliz-Cuba. An algebraic approach to reverse engineering finite dynamical systems arising from biology. SIAM Journal on Applied Dynamical Systems, 11 (1):31–48, ???? 2012. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.
                                        siam.org/siads/resource/
                                        1/sjaday/v11/i1/p31_s1.

Verschueren:2017:MCP

[VC17] Nicolas Verschueren and Alan Champneys. A model for cell polarization without mass conservation. SIAM Journal on Applied Dynamical Systems, 16(4):1797–1830, ???? 2017. CODEN SJADAY. ISSN 1536-0040. See erratum [VC19].

Verschueren:2019:EMC

[VC19] Nicolas Verschueren and Alan Champneys. Erratum: A Model for Cell Polarization without Mass Conservation. SIAM Journal on Applied Dynamical Systems, 18(1):594– 595, ???? 2019. CODEN SJADAY. ISSN 1536-0040. See [VC17].

Vierling-Claassen:2009:DPF

[VCK09] Dorea Vierling-Claassen and Nancy Kopell. The dynamics of a periodically forced cortical microcircuit, with an application to schizophrenia. SIAM Journal on Applied Dynamical Systems, 8(2):710– 733, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Veliz-Cuba:2024:UAR

[VCNSD24] Alan Veliz-Cuba, Vanessa Newsome-Slade, and Elena S. Dimitrova. A unified approach to reverse engineering and data selection for unique network identification. SIAM Journal on Applied Dynamical Systems, 23(1):592–615, February 2024. ISSN 1536-0040.

Veerman:2013:PGM

[VD13] Frits Veerman and Arjen Doelman. Pulses in a Gierer– Meinhardt equation with a slow nonlinearity. SIAM Journal on Applied Dynamical Systems, 12(1):28–60, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

vandenBerg:2015:SCH

[vdBDLJ15] Jan Bouwe van den Berg, Andréa Deschênes, Jean-Philippe Lessard, and Jason D. Mireles James. Stationary coexistence of hexagons and rolls via rigorous computations. SIAM Journal on Applied Dynamical Systems, 14 (2):942–979, ???? 2015. CO-DEN SJADAY. ISSN 1536-

vandenBerg:2015:RCR

[vdBGW15] J. B. van den Berg, C. M. Groothedde, and J. F. Williams. Rigorous computation of a radially symmetric localized solution in a Ginzburg–Landau problem. SIAM Journal on Applied Dynamical Systems, 14(1):423– 447, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

vandenBerg:2011:OBP

[vdBKV11] Jan Bouwe van den Berg, Miroslav Kramár, and Robert C. Vandervorst. The order of bifurcation points in fourth order conservative systems via braids. SIAM Journal on Applied Dynamical Systems, 10 (2):510–550, ???? 2011. CO-DEN SJADAY. ISSN 1536-

                                        0040. URL http://epubs.
                                        siam.org/siads/resource/
                                        1/sjaday/v10/i2/p510_s1.

vandenBerg:2008:CBS

[vdBL08] Jan Bouwe van den Berg and Jean-Philippe Lessard. Chaotic braided solutions via rigorous numerics: Chaos in the Swift–Hohenberg equation. SIAM Journal on Applied Dynamical Systems, 7 (3):988–1031, 2008. CODEN SJADAY. ISSN 1536-0040.

vandenBerg:2021:RVH

[vdBLQ21] Jan Bouwe van den Berg, Jean-Philippe Lessard, and Elena Queirolo. Rigorous verification of Hopf bifurcations via desingularization and continuation. SIAM Journal on Applied Dynamical Systems, 20(2):573–607, ???? 2021. CODEN SJADAY. ISSN 1536-0040.

vandenDriessche:2004:DIO

[vdDZ04] P. van den Driessche and M. L. Zeeman. Disease induced oscillations between two competing species. SIAM Journal on Applied Dynamical Systems, 3(4):601–619, 2004. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60039.

Veltz:2013:IBS

[Vel13] Romain Veltz. Interplay between synaptic delays and propagation delays in neural field equations. SIAM Journal on Applied Dynamical Systems, 12(3):1566–1612, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Vera:2008:EEG

[Ver08] J. A. Vera. Eulerian equilibria of a gyrostat in Newtonian interaction with two rigid bodies. SIAM Journal on Applied Dynamical Systems, 7 (4):1378–1396, 2008. CODEN SJADAY. ISSN 1536-0040.

Veltz:2010:LGA

[VF10] Romain Veltz and Olivier Faugeras. Local/global analysis of the stationary solutions of some neural field equations. SIAM Journal on Applied Dynamical Systems, 9 (3):954–998, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Varkonyi:2008:STW

[VH08] Péter L. Várkonyi and Philip Holmes. On synchronization and traveling waves in chains of relaxation oscillators with an application to lamprey CPG. SIAM Journal on Applied Dynamical Systems, 7 (3):766–794, 2008. CODEN SJADAY. ISSN 1536-0040.

vanHeijster:2010:FIT

[vHDKP10] P. van Heijster, A. Doelman, T. J. Kaper, and K. Promislow. Front interactions in a three-component system. SIAM Journal on Applied Dynamical Systems, 9(2):292– 332, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Valani:2022:BDI

[VHS22] Rahil N. Valani, Brendan Harding, and Yvonne M. Stokes. Bifurcations and dynamics in inertial focusing of particles in curved rectangular ducts. SIAM Journal on Applied Dynamical Systems, 21(4):2371–2392, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1451919.

Villanueva:2018:PML

[Vil18] Jordi Villanueva. A parameterization method for Lagrangian tori of exact symplectic maps of R^{2}r. SIAM Journal on Applied Dynamical Systems, 17(3):2289–2331, ???? 2018. CODEN SJA-DAY. ISSN 1536-0040.

Viswanath:2013:MEO

[VLS13] Divakar Viswanath, Xuan Liang, and Kirill Serkh. Metric entropy and the optimal prediction of chaotic signals. SIAM Journal on Applied Dynamical Systems, 12(2):1085– 1113, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Verwoerd:2008:GPL

[VM08] Mark Verwoerd and Oliver Mason. Global phase-locking in finite populations of phasecoupled oscillators. SIAM Journal on Applied Dynamical Systems, 7(1):134–160, 2008. CODEN SJADAY. ISSN 1536-0040.

Verwoerd:2009:CCC

[VM09] Mark Verwoerd and Oliver Mason. On computing the critical coupling coefficient for the Kuramoto model on a complete bipartite graph. SIAM Journal on Applied Dynamical Systems, 8(1):417– 453, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Verwoerd:2011:CRK

[VM11] Mark Verwoerd and Oliver Mason. A convergence result for the Kuramoto model with all-to-all coupling. SIAM Journal on Applied Dynamical Systems, 10(3):906–

920, ????  2011.  CO-

DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i3/p906_s1.

Volpert:2008:BSN

[VNSG08] V. A. Volpert, A. A. Nepomnyashchy, L. G. Stanton, and A. A. Golovin. Bounded solutions of nonlocal complex Ginzburg–Landau equations for a subcritical bifurcation. SIAM Journal on Applied Dynamical Systems, 7 (2):265–283, 2008. CODEN SJADAY. ISSN 1536-0040.

Vojkovic:2023:LOP

[VQPVV23] Tea Vojkovic, David Quero, Charles Poussot-Vassal, and Pierre Vuillemin. Low-order parametric state-space modeling of MIMO systems in the Loewner framework. SIAM Journal on Applied Dynamical Systems, 22(4):3130–3164, November 2023. ISSN 1536-0040.

Varzaneh:2022:DTS

[VRS22] M. Ghani Varzaneh, S. Riedel, and M. Scheutzow. A dynamical theory for singular stochastic delay differential equations i: Linear equations and a multiplicative ergodic theorem on fields of Banach spaces. SIAM Journal on Applied Dynamical Systems, 21(1):542–587, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1433435.

Villar-Sepulveda:2023:CSM

[VSABM23] Edgardo Villar-Sepúlveda, Pablo Aguirre, and Víctor F. Breña-Medina. A case study of multiple wave solutions in a reaction–diffusion system using invariant manifolds and global bifurcations. SIAM Journal on Applied Dynamical Systems, 22(2):918–950, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/22M1474709.

Villar-Sepulveda:2023:DTB

[VSC23] Edgardo Villar-Sepúlveda and Alan Champneys. Degenerate Turing bifurcation and the birth of localized patterns in activator-inhibitor systems. SIAM Journal on Applied Dynamical Systems, 22(3):1673–1709, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1509734.

Vainchtein:2015:PPP

[VVZ15] Anna Vainchtein, Erik S.

Van Vleck, and Aijun Zhang. Propagation of periodic patterns in a discrete system with competing interactions. SIAM Journal on Applied Dynamical Systems, 14(2):523– 555, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Wheeler:2006:CSS

[WB06] Paul Wheeler and Dwight Barkley. Computation of spiral spectra. SIAM Journal on Applied Dynamical Systems, 5(1):157–177, 2006. CODEN SJADAY. ISSN 1536-0040.

Welshman:2014:DRD

[WB14] Chris Welshman and John Martin Brooke. Dimensionality reduction of dynamical systems with parameters: a geometric approach. SIAM Journal on Applied Dynamical Systems, 13(1):493–517, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Wilczak:2017:SCA

[WB17] Daniel Wilczak and Roberto Barrio. Systematic computerassisted proof of branches of stable elliptic periodic orbits and surrounding invariant tori. SIAM Journal on Applied Dynamical Systems, 16(3):1618–1649, ???? 2017. CODEN SJADAY. ISSN 1536-0040. Winterbottom:2008:PFM

[WCM08] D. M. Winterbottom, S. M. Cox, and P. C. Matthews. Pattern formation in a model of a vibrated granular layer. SIAM Journal on Applied Dynamical Systems, 7(1):63–78, 2008. CODEN SJADAY. ISSN 1536-0040.

Wilson:2018:ODP

[WE18] Dan Wilson and Bard Ermentrout. An operational definition of phase characterizes the transient response of perturbed limit cycle oscillators. SIAM Journal on Applied Dynamical Systems, 17(4):2516– 2543, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Wechselberger:2005:EBC

[Wec05] Martin Wechselberger. Existence and bifurcation of canards in R^{3} in the case of a folded node. SIAM Journal on Applied Dynamical Systems, 4(1):101–139, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60199.

Weideman:2003:CDC

[Wei03] J. A. C. Weideman. Computing the dynamics of complex singularities of nonlinear PDEs. SIAM Journal on Applied Dynamical Systems, 2 (2):171–186, 2003. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
39830.

Wiuf:2013:PLK

[WF13] Carsten Wiuf and Elisenda Feliu. Power-law kinetics and determinant criteria for the preclusion of multistationarity in networks of interacting species. SIAM Journal on Applied Dynamical Systems, 12(4):1685–1721, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Watts:2014:CMO

[WFM^{+}14] Margaret Watts, Bernard Fendler, Matthew J. Merrins, Leslie S. Satin, Richard Bertram, and Arthur Sherman. Calcium and metabolic oscillations in pancreatic islets: Who’s driving the bus? SIAM Journal on Applied Dynamical Systems, 13(2):683– 703, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Wiser:2015:BFR

[WG15] Justin Wiser and Martin Golubitsky. Bifurcations in forced response curves. SIAM Journal on Applied Dynamical Systems, 14(4):2013–2029, ???? 2015. CODEN SJA-DAY. ISSN 1536-0040.

Wang:2021:SVT

[WGCT21] Yangyang Wang, Jeffrey P. Gill, Hillel J. Chiel, and Peter J. Thomas. Shape versus timing: Linear responses of a limit cycle with hard boundaries under instantaneous and static perturbation. SIAM Journal on Applied Dynamical Systems, 20(2):701–744, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Wittmann:2013:TCP

[WHT13] Dominik M. Wittmann, Sabrina Hock, and Fabian J. Theis. Truth-content and phase transitions of random Boolean networks with generic logics. SIAM Journal on Applied Dynamical Systems, 12(1):315–351, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Widiasih:2013:DBE

[Wid13] Esther R. Widiasih. Dynamics of the Budyko energy balance model. SIAM Journal on Applied Dynamical Systems, 12(4):2068–2092, ???? 2013. CODEN SJA-DAY. ISSN 1536-0040.

Wilczak:2005:SHC

[Wil05] Daniel Wilczak. Symmetric heteroclinic connections in the Michelson system: a computer assisted proof. SIAM Journal on Applied Dynamical Systems, 4(3):489–514, 2005. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/61111.

Wilczak:2010:UHA

[Wil10] Daniel Wilczak. Uniformly hyperbolic attractor of the Smale–Williams type for a Poincaré map in the Kuznetsov system. SIAM Journal on Applied Dynamical Systems, 9(4):1263–1283, ???? 2010. CODEN SJA-DAY. ISSN 1536-0040.

Wilson:2019:OFN

[Wil19] Dan D. Wilson. An optimal framework for nonfeedback stability control of chaos. SIAM Journal on Applied Dynamical Systems, 18(4):1982– 1999, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Wilson:2021:OCO

[Wil21] Dan Wilson. Optimal control of oscillation timing and entrainment using large magnitude inputs: an adaptive phase-amplitude-coordinate-based approach. SIAM Journal on Applied Dynamical Systems, 20(4):1814–1843, ???? 2021. CODEN SJA-DAY. ISSN 1536-0040.

Wilson:2022:APA

[Wil22]  Dan Wilson.  An adap-

tive phase-amplitude reduction framework without O(ϵ) constraints on inputs. SIAM Journal on Applied Dynamical Systems, 21(1):204–230, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1391791.

Wilson:2023:KOI

[Wil23a] Dan Wilson. Koopman operator inspired nonlinear system identification. SIAM Journal on Applied Dynamical Systems, 22(2):1445–1471, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1512272.

Wilson:2023:ROM

[Wil23b] Dan Wilson. A reduced order modeling framework for strongly perturbed nonlinear dynamical systems near arbitrary trajectory sets. SIAM Journal on Applied Dynamical Systems, 22(2):603–634, ???? 2023. CODEN SJA-DAY. ISSN 1536-0040. URL

https://epubs.siam.org/
doi/10.1137/21M1451154.

Watanabe:2016:SCD

[WIN16] Takeshi Watanabe, Makoto Iima, and Yasumasa Nishiura. A skeleton of collision dynamics: Hierarchical network structure among evensymmetric steady pulses in binary fluid convection. SIAM Journal on Applied Dynamical Systems, 15(2):789– 806, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Wang:2022:UPG

[WL22] Weixin Wang and Taeyoung Lee. Uncertainty propagation for general stochastic hybrid systems on compact Lie groups. SIAM Journal on Applied Dynamical Systems, 21(3):2215–2240, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M144147X.

Wang:2015:RAD

[WLW15] Xiaohu Wang, Kening Lu, and Bixiang Wang. Random attractors for delay parabolic equations with additive noise and deterministic nonautonomous forcing. SIAM Journal on Applied Dynamical Systems, 14(2):1018– 1047, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Wilson:2014:EOM

[WM14a] Dan Wilson and Jeff Moehlis. An energy-optimal methodology for synchronization of excitable media. SIAM Journal on Applied Dynamical Systems, 13(2):944–957, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Wilson:2014:OCD

[WM14b] Dan Wilson and Jeff Moehlis. Optimal chaotic desynchronization for neural populations. SIAM Journal on Applied Dynamical Systems, 13 (1):276–305, ???? 2014. CO-DEN SJADAY. ISSN 1536-0040.

Wanner:2022:RAS

[WM22] Mathias Wanner and Igor Mezić. Robust approximation of the stochastic Koopman operator. SIAM Journal on Applied Dynamical Systems, 21(3):1930–1951, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1414425.

Wanner:2024:HOD

[WM24] Mathias Wanner and Igor Mezić. On higher order drift and diffusion estimates for stochastic SINDy. SIAM Journal on Applied Dynamical Systems, 23(2):1504–1539, ???? 2024. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/23M1567011.

Wulff:2002:HSN

[WR02] Claudia Wulff and Mark Roberts. Hamiltonian systems near relative periodic orbits. SIAM Journal on Applied Dynamical Systems, 1(1):1–43, 2002. CODEN SJADAY. ISSN 1536-0040.

URL http://epubs.siam.
org/sam-bin/dbq/article/
38776.

Wang:2013:SSA

[WR13] Wei Wang and A. J. Roberts.

Self-similarity and attraction in stochastic nonlinear reaction–diffusion systems. SIAM Journal on Applied Dynamical Systems, 12(1):450– 486, ???? 2013. CODEN SJADAY. ISSN 1536-0040.

Wright:2010:IMR

[Wri10] J. Douglas Wright. Interaction manifolds for reaction diffusion equations in two dimensions. SIAM Journal on Applied Dynamical Systems, 9(3):734–768, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Wulff:2006:NCS

[WS06] Claudia Wulff and Andreas Schebesch. Numerical continuation of symmetric periodic orbits. SIAM Journal on Applied Dynamical Systems, 5 (3):435–475, 2006. CODEN SJADAY. ISSN 1536-0040.

Wulff:2009:NBH

[WS09] Claudia Wulff and Frank Schilder. Numerical bifurcation of Hamiltonian relative periodic orbits. SIAM Journal on Applied Dynamical Systems, 8(3):931–966, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Wulff:2014:RLC

[WS14] Claudia Wulff and Frank Schilder. Relative Lyapunov center bifurcations. SIAM Journal on Applied Dynamical Systems, 13(2):722–757, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Wilson:2024:ROC

[WS24] Dan Wilson and Kai Sun. Reduced order characterization of nonlinear oscillations using an adaptive phase-amplitude coordinate framework. SIAM Journal on Applied Dynamical Systems, 23(1):470–504, January 2024. ISSN 1536-0040.

Wilczak:2016:CDC

[WSB16] Daniel Wilczak, Sergio Serrano, and Roberto Barrio. Coexistence and dynamical connections between hyperchaos and chaos in the 4D Rössler system: a computerassisted proof. SIAM Journal on Applied Dynamical Systems, 15(1):356–390, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Williams:2012:LDT

[WSWK12] Matthew O. Williams, Eli Shlizerman, Jon Wilkening, and J. Nathan Kutz. The low dimensionality of timeperiodic standing waves in water of finite and infinite depth. SIAM Journal on Applied Dynamical Systems, 11 (3):1033–1061, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Wang:2023:TSA

[WSYTA23] Ai Wang, Jan Sieber, William R.

Young, and Krasimira Tsaneva-Atanasova. Time series analysis and modeling of the freezing of gait phenomenon. SIAM Journal on Applied Dynamical Systems, 22(2):825–849, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1484341.

Wulff:2008:HAM

[Wul08] Claudia Wulff. A Hamiltonian analogue of the meandering transition. SIAM Journal on Applied Dynamical Systems, 7(4):1213–1246, 2008. CODEN SJADAY. ISSN 1536-0040.

Wouters:2019:AET

[WV19] Michiel Wouters and Wim Vanroose. Automatic exploration techniques of numerical bifurcation diagrams illustrated by the Ginzburg– Landau equation. SIAM Journal on Applied Dynamical Systems, 18(4):2047–2098, ???? 2019. CODEN SJA-DAY. ISSN 1536-0040.

Wayne:2002:HOM

[WW02] C. Eugene Wayne and J. Douglas Wright. Higher order modulation equations for a Boussinesq equation. SIAM Journal on Applied Dynamical Systems, 1(2):271–302, 2002. CODEN SJADAY. ISSN 1536-0040. URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41129.

Wong:2020:WNA

[WW20] Tony Wong and Michael J. Ward. Weakly nonlinear analysis of peanut-shaped deformations for localized spots of singularly perturbed reaction–diffusion systems. SIAM Journal on Applied Dynamical Systems, 19(3):2030– 2058, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Wei:2018:CES

[WWC^{+}18] Xiang Wei, Xiaoqun Wu, Shihua Chen, Jun an Lu, and Guanrong Chen. Cooperative epidemic spreading on a two-layered interconnected network. SIAM Journal on Applied Dynamical Systems, 17(2):1503–1520, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Wang:2024:BHM

[WWXR24] Shaoli Wang, Tengfei Wang, Fei Xu, and Libin Rong. Bistability of an HIV model with immune impairment. SIAM Journal on Applied Dynamical Systems, 23(2):1108– 1132, May 2024. ISSN 1536-0040.

Wang:2019:WNV

[WWZ19] Feng-Bin Wang, Ruiwen Wu, and Xiao-Qiang Zhao. A West Nile virus transmission model with periodic incubation periods. SIAM Journal on Applied Dynamical Systems, 18(3):1498–1535, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Wolfrum:2022:MSL

[WYD22] Matthias Wolfrum, Serhiy Yanchuk, and Otti D’Huys. Multiple self-locking in the Kuramoto–Sakaguchi system with delay. SIAM Journal on Applied Dynamical Systems, 21(3):1709–1725, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1458971.

Wang:2024:DCM

[WYM^{+}24] Haotian Wang, Hujiang Yang, Xiankui Meng, Ye Tian, and Wenjun Liu. Dynamics of controllable matter–wave solitons and soliton molecules for a Rabi-coupled Gross– Pitaevskii equation with temporally and spatially modulated coefficients. SIAM Journal on Applied Dynamical Systems, 23(1):748–778, February 2024. ISSN 1536-

Wilczak:2009:CAP

[WZ09] Daniel Wilczak and Piotr Zgliczyński. Computer assisted proof of the existence of homoclinic tangency for the Hénon map and for the forced damped pendulum. SIAM Journal on Applied Dynamical Systems, 8(4):1632–1663, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Wang:2012:BRN

[WZ12] Wendi Wang and Xiao-Qiang Zhao. Basic reproduction numbers for reaction– diffusion epidemic models. SIAM Journal on Applied Dynamical Systems, 11(4):1652– 1673, ???? 2012. CODEN SJADAY. ISSN 1536-0040.

Wilczak:2016:COS

[WZ16] Daniel Wilczak and Piotr Zgliczyński. Connecting orbits for a singular nonautonomous real Ginzburg– Landau type equation. SIAM Journal on Applied Dynamical Systems, 15(1):495–525, ???? 2016. CODEN SJA-DAY. ISSN 1536-0040.

Wang:2017:DTD

[WZ17] Xiunan Wang and Xiao-Qiang Zhao. Dynamics of a time-delayed Lyme disease model with seasonality. SIAM Journal on Applied Dynamical Systems, 16(2):853–881, ???? 2017. CODEN SJA-DAY. ISSN 1536-0040.

Wang:2018:SLD

[WZ18] Cheng Wang and Xiang Zhang. Stability loss delay and smoothness of the return map in slow-fast systems. SIAM Journal on Applied Dynamical Systems, 17 (1):788–822, ???? 2018. CO-DEN SJADAY. ISSN 1536-0040.

Wang:2023:HNS

[WZM23] Rong-Nian Wang, Jia-Cheng Zhao, and Alain Miranville. Hyperdissipative Navier–Stokes equations driven by timedependent forces: Invariant manifolds. SIAM Journal on Applied Dynamical Systems, 22(1):199–234, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1470323.

Xu:2007:EMS

[XCC07] Jian Xu, Kwok-Wai Chung, and Chuen-Lit Chan. An efficient method for studying weak resonant double Hopf bifurcation in nonlinear systems with delayed feedbacks. SIAM Journal on Applied Dynamical Systems, 6(1):29–60, 2007. CODEN SJADAY. ISSN 1536-0040.

Xiao:2008:QCR

[Xia08] MingQing Xiao. Quantitative characteristic of rotating stall and surge for Moore–Greitzer PDE model of an axial flow compressor. SIAM Journal on Applied Dynamical Systems, 7(1):39–62, 2008. CO-DEN SJADAY. ISSN 1536-0040.

Xu:2020:SPQ

[XSS20] Xiaodan Xu, Wen Si, and Jianguo Si. Stoker’s problem for quasi-periodically forced reversible systems with multidimensional Liouvillean frequency. SIAM Journal on Applied Dynamical Systems, 19(4):2286–2321, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Xue:2024:STD

[XSW24] Shuyang Xue, Yongli Song, and Hao Wang. Spatiotemporal dynamics in a reaction-diffusion equation with nonlocal spatial memory. SIAM Journal on Applied Dynamical Systems, 23 (1):641–667, February 2024. ISSN 1536-0040.

Xi:2023:SCP

[XWC^{+}23] Kaihua Xi, Zhen Wang, Aijie Cheng, Hai Xiang Lin, Jan H. van Schuppen, and Chenghui Zhang. Synchronization of coupled phase oscillators with stochastic disturbances and the cycle space of the graph. SIAM Journal on Applied Dynamical Systems, 22(2):1030–1052, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1489502.

Yakubu:2008:ASD

[Yak08] Abdul-Aziz Yakubu. Asynchronous and synchronous dispersals in spatially discrete population models. SIAM Journal on Applied Dynamical Systems, 7(2):284–310, 2008. CODEN SJADAY. ISSN 1536-0040.

Yuan:2011:SHB

[YB11] Yuan Yuan and Jacques Bélair. Stability and Hopf bifurcation analysis for functional differential equation with distributed delay. SIAM Journal on Applied Dynamical Systems, 10(2):551– 581, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i2/p551_s1.

Yuan:2022:MAE

[YB22] Shenglan Yuan and Dirk Bloemker. Modulation and amplitude equations on bounded domains for nonlinear SPDEs driven by cylindrical α-stable Lévy processes. SIAM Journal on Applied Dynamical Systems, 21 (3):1748–1777, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1431333.

Yoon:2022:NED

[YBO22] Ryeongkyung Yoon, Harish S.

Bhat, and Braxton Osting. A nonautonomous equation discovery method for time signal classification. SIAM Journal on Applied Dynamical Systems, 21(1):33–59, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1405216.

Yulin:2010:DSM

[YC10] A. V. Yulin and A. R. Champneys. Discrete snaking: Multiple cavity solitons in saturable media. SIAM Journal on Applied Dynamical Systems, 9(2):391–431, ???? 2010. CODEN SJADAY. ISSN 1536-0040.

Yabo:2022:DAO

[YCG^{+}22] Agustín G. Yabo, Jean-Baptiste Caillau, Jean-Luc Gouzé, Hidde de Jong, and Francis Mairet. Dynamical analysis and optimization of a generalized resource allocation model of microbial growth. SIAM Journal on Applied Dynamical Systems, 21(1):137–165, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M141097X.

Yu:2008:GSL

[YCL08] Wenwu Yu, Jinde Cao, and Jinhu Lü. Global synchronization of linearly hybrid coupled networks with timevarying delay. SIAM Journal on Applied Dynamical Systems, 7(1):108–133, 2008.

CODEN SJADAY. ISSN 1536-0040.

Yu:2022:GTC

[YCMP22] Polly Y. Yu, Gheorghe Craciun, Maya Mincheva, and Casian Pantea. A graphtheoretic condition for delay stability of reaction systems. SIAM Journal on Applied Dynamical Systems, 21 (2):1092–1118, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1420307.

Yochelis:2002:DSW

[YHM^{+}02] Arik Yochelis, Aric Hagberg, Ehud Meron, Anna L. Lin, and Harry L. Swinney. Development of standing-wave labyrinthine patterns. SIAM Journal on Applied Dynamical Systems, 1(2):236–247, 2002. CODEN SJADAY. ISSN 1536-0040. URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39711.

Yang:2016:SBS

[YLWK16] Xianshan Yang, Xiong Li, Hao Wang, and Yang Kuang. Stability and bifurcation in a stoichiometric producergrazer model with knife edge. SIAM Journal on Applied Dynamical Systems, 15(4):2051– 2077, ???? 2016. CODEN SJADAY. ISSN 1536-0040.

Yela:2023:PPM

[YM23] Alberto López Yela and Joaquín Míguez. Polynomial propagation of moments in stochastic differential equations. SIAM Journal on Applied Dynamical Systems, 22 (2):1153–1181, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1431497.

Yang:2023:OTP

[YNN^{+}23] Yunan Yang, Levon Nurbekyan, Elisa Negrini, Robert Martin, and Mirjeta Pasha. Optimal transport for parameter identification of chaotic dynamics via invariant measures. SIAM Journal on Applied Dynamical Systems, 22(1):269–310, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/21M1421337.

Yadome:2014:RPG

[YNT14] Masaaki Yadome, Yasumasa Nishiura, and Takashi Teramoto. Robust pulse generators in an excitable medium with jump-type heterogeneity. SIAM Journal on Applied Dynamical Systems, 13 (3):1168–1201, ???? 2014. CODEN SJADAY. ISSN 1536-0040.

Yan:2022:CDA

[YNZ22] Xiao Yan, Hua Nie, and Peng Zhou. On a competitiondiffusion-advection system from river ecology: Mathematical analysis and numerical study. SIAM Journal on Applied Dynamical Systems, 21(1):438–469, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/20M1387924.

Yoshimura:2017:ESO

[Yos17] Kazuyuki Yoshimura. Existence and stability of odd and even parity discrete breathers in Fermi–Pasta–Ulam lattices. SIAM Journal on Applied Dynamical Systems, 16 (4):2063–2095, ???? 2017. CODEN SJADAY. ISSN 1536-0040.

Yanguas:2008:PSH

[YPMD08] Patricia Yanguas, Jesús F.

Palacián, Kenneth R. Meyer, and H. Scott Dumas. Periodic solutions in Hamiltonian systems, averaging, and the lunar problem. SIAM Journal on Applied Dynamical Systems, 7(2):311–340, 2008. CODEN SJADAY. ISSN 1536-0040.

Yu:2024:ECH

[YPTW24] Pei Yu, Pantea Pooladvand, Mark M. Tanaka, and Lindi M. Wahl. Extinctions caused by host-range expansion. SIAM Journal on Applied Dynamical Systems, 23(2):1677–1703, ???? 2024. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/23M1605582.

Ylinen:2022:ADS

[YvLKL22] Lauri Ylinen, Tuomo von Lerber, Franko Küppers, and Matti Lassas. Analysis of a dynamical system modeling lasers and applications for optical neural networks. SIAM Journal on Applied Dynamical Systems, 21(2):840–878, ???? 2022. CODEN SJA-DAY. ISSN 1536-0040. URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1405976.

Yanchuk:2010:MTS

[YW10] Serhiy Yanchuk and Matthias Wolfrum. A multiple time scale approach to the stability of external cavity modes in the Lang–Kobayashi system using the limit of large delay. SIAM Journal on Applied Dynamical Systems, 9 (2):519–535, ???? 2010. CO-DEN SJADAY. ISSN 1536-0040.

Yagasaki:2019:BRE

[YY19] Kazuyuki Yagasaki and Shotaro Yamazoe. Bifurcation of relative equilibria in infinitedimensional Hamiltonian systems. SIAM Journal on Applied Dynamical Systems, 18 (1):393–417, ???? 2019. CO-DEN SJADAY. ISSN 1536-0040.

Zhang:2020:CTG

[ZB20] Li Zhang and Sanjeeva Balasuriya. Controlling trajectories globally via spatiotemporal finite-time optimal control. SIAM Journal on Applied Dynamical Systems, 19 (3):1609–1632, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Zhang:2009:ICD

[ZBN09] Yu Zhang, Amitabha Bose, and Farzan Nadim. The influence of the A-current on the dynamics of an oscillatorfollower inhibitory network. SIAM Journal on Applied Dynamical Systems, 8(4):1564– 1590, ???? 2009. CODEN SJADAY. ISSN 1536-0040.

Ziessler:2019:NCU

[ZDG19] Adrian Ziessler, Michael Dellnitz, and Raphael Gerlach. The numerical computation of unstable manifolds for infinite dimensional dynamical systems by embedding techniques. SIAM Journal on Applied Dynamical Systems, 18(3):1265–1292, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Zhang:2011:PFH

[ZG11] Yanyan Zhang and Martin Golubitsky. Periodically forced Hopf bifurcation. SIAM Journal on Applied Dynamical Systems, 10(4): 1272–1306, ???? 2011. CO-DEN SJADAY. ISSN 1536-

0040. URL http://epubs.
siam.org/siads/resource/
1/sjaday/v10/i4/p1272_s1.

Zgliczynski:2002:AFP

[Zgl02] Piotr Zgliczynski. Attracting fixed points for the Kuramoto–Sivashinsky equation: a computer assisted proof. SIAM Journal on Applied Dynamical Systems, 1 (2):215–235, 2002. CODEN SJADAY. ISSN 1536-0040.

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        40176.

Zhang:2007:HDS

[Zha07] Linghai Zhang. How do synaptic coupling and spatial temporal delay influence traveling waves in nonlinear nonlocal neuronal networks? SIAM Journal on Applied Dynamical Systems, 6(3):597– 644, 2007. CODEN SJADAY. ISSN 1536-0040.

Zhelyazov:2015:GSL

[ZHKR15] D. Zhelyazov, D. Han-Kwan, and J. D. M. Rademacher. Global stability and local bifurcations in a two-fluid model for tokamak plasma. SIAM Journal on Applied Dynamical Systems, 14(2):730– 763, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Zhang:2019:LCR

[ZKCS19] Wei Zhang, Stefan Klus, Tim Conrad, and Christof Schütte. Learning chemical reaction networks from trajectory data. SIAM Journal on Applied Dynamical Systems, 18(4):2000–2046, ????

2019. CODEN SJADAY. ISSN 1536-0040.

Zhu:2015:TPN

[ZKE15] Jielin Zhu, Rachel Kuske, and Thomas Erneux. Tipping points near a delayed saddle node bifurcation with periodic forcing. SIAM Journal on Applied Dynamical Systems, 14(4):2030–2068, ???? 2015. CODEN SJADAY. ISSN 1536-0040.

Zlotnik:2014:OSE

[ZL14] Anatoly Zlotnik and Jr-Shin Li. Optimal subharmonic entrainment of weakly forced nonlinear oscillators. SIAM Journal on Applied Dynamical Systems, 13(4):1654–1693, ???? 2014. CODEN SJA-DAY. ISSN 1536-0040.

Zhang:2020:ELP

[ZL20] Liwei Zhang and Weishi Liu.

Effects of large permanent charges on ionic flows via Poisson–Nernst–Planck models. SIAM Journal on Applied Dynamical Systems, 19 (3):1993–2029, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Zhou:2022:SAC

[ZLL22] Shijie Zhou, Ying-Cheng Lai, and Wei Lin. Stochastically adaptive control and synchronization: From globally one-sided Lipschitzian to only locally Lipschitzian systems. SIAM Journal on Applied Dynamical Systems, 21(2):932–959, ???? 2022. CODEN SJADAY. ISSN 1536-0040. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1402042.

Zhao:2020:SHN

[ZLX20] Xiaoxue Zhao, Zhuchun Li, and Xiaoping Xue. Stability in a Hebbian network of Kuramoto oscillators with second-order couplings for binary pattern retrieve. SIAM Journal on Applied Dynamical Systems, 19(2):1124–1159, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Ziepke:2019:CNW

[ZME19] Alexander Ziepke, Steffen Martens, and Harald Engel. Control of nonlinear wave solutions to neural field equations. SIAM Journal on Applied Dynamical Systems, 18 (2):1015–1036, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Zazo:2020:LCH

[ZR20] Miriam Steinherr Zazo and Jens D. M. Rademacher. Lyapunov coefficients for Hopf bifurcations in systems with piecewise smooth nonlinearity. SIAM Journal on Applied Dynamical Systems, 19 (4):2847–2886, ???? 2020. CODEN SJADAY. ISSN 1536-0040.

Zhang:2019:ODM

[ZRDC19] Hao Zhang, Clarence W. Rowley, Eric A. Deem, and Louis N. Cattafesta. Online dynamic mode decomposition for time-varying systems. SIAM Journal on Applied Dynamical Systems, 18 (3):1586–1609, ???? 2019. CODEN SJADAY. ISSN 1536-0040.

Zino:2018:MME

[ZRP18] Lorenzo Zino, Alessandro Rizzo, and Maurizio Porfiri. Modeling memory effects in activity-driven networks. SIAM Journal on Applied Dynamical Systems, 17(4):2830– 2854, ???? 2018. CODEN SJADAY. ISSN 1536-0040.

Zheng:2023:SAI

[ZSL23] Pan Zheng, Wenhai Shan, and Guangyuan Liao. Stability analysis of the immune system induced by chemotaxis. SIAM Journal on Applied Dynamical Systems, 22 (3):2527–2569, ???? 2023. CODEN SJADAY. ISSN 1536-0040. URL https:

//epubs.siam.org/doi/10.
1137/22M1498310.

Zhong:2024:DHB

[ZW24] Haonan Zhong and Kaifa Wang. Dynamics on hepatitis B virus infection in vivo with interval delay. SIAM Journal on Applied Dynamical Systems, 23(1):205–235, January 2024. ISSN 1536-

Zhou:2005:RAD

[ZYO05] Shengfan Zhou, Fuqi Yin, and Zigen Ouyang. Random attractor for damped nonlinear wave equations with white noise. SIAM Journal on Applied Dynamical Systems, 4 (4):883–903, 2005. CODEN SJADAY. ISSN 1536-0040.

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        62309.

Zou:2009:SSR

[ZZ09] Wei Zou and Meng Zhan. Splay states in a ring of coupled oscillators: From local to global coupling. SIAM Journal on Applied Dynamical Systems, 8(3):1324–1340, ???? 2009. CODEN SJA-DAY. ISSN 1536-0040.

Zhang:2020:MMS

[ZZ20] Tailei Zhang and Xiao-Qiang Zhao. Mathematical modeling for schistosomiasis with seasonal influence: a case study in Hubei, China. SIAM Journal on Applied Dynamical Systems, 19(2):1438–1471, ???? 2020. CODEN SJA-DAY. ISSN 1536-0040.

Zhang:2018:ECL

[ZZQ18] Jinxing Zhang, Jiandong Zhu, and Chunjiang Qian. On equilibria and consensus of the Lohe model with identical oscillators. SIAM Journal on Applied Dynamical Systems, 17(2):1716–1741, ???? 2018.

REFERENCES 226 CODEN SJADAY. ISSN 1536-0040.